### v1

# Perturbation-based Analysis of Compositional Data

Abstract
Existing statistical methods for compositional data analysis are inadequate for many modern applications for two reasons. First, modern compositional datasets, for example in microbiome research, display traits such as high-dimensionality and sparsity that are poorly modelled with traditional approaches. Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward. In this work, we propose a framework based on hypothetical data perturbations that addresses both issues. Unlike existing methods for compositional data, we do not transform the data and instead use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects. These average perturbation effects, which can be employed in many applications, naturally account for confounding that biases frequently used marginal dependence analyses. We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques. We analyze the proposed estimators empirically on simulated data and demonstrate advantages over existing techniques on US census and microbiome data. For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
Keywords: compositional data, causality, perturbation, feature importance, diversity effects, nonparametrics
1 Introduction
Compositional data, that is, measurements consisting of parts of a whole, appear in many scientific disciplines, for example, rock compositions in geology, green house gases in the atmosphere, species compositions in ecology and racial or gender distributions in social sciences. The defining property of such data is that they only capture information relative to the unknown whole and hence do not contain any information on the absolute sizes. Depending on the application, the absolute information either might not be of interest (e.g., when analyzing racial or gender distributions) or might not be easily accessible with existing measuring techniques (e.g., in ecology). Mathematically, a $d$-coordinate composition can – without loss of information – be normalized such that it is represented as a point on the $(d-1)$-dimensional simplex
$\Delta^{d-1}:=\left\{z=(z^{1},\dots,z^{d})\in[0,1]^{d}\mid\textstyle\sum_{j=1}^{d}z^{j}=1\right\}.$
It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. Much work has gone into constructing new procedures for adjusting the correlation measure for compositional components, e.g., by transforming the data with log-ratio transforms (Aitchison, 1982). An instructive viewpoint is to see these approaches not as adjusting the correlation measure but rather as creating new target parameters (e.g., the correlation between log-ratios) that have meaningful interpretations in compositional data.
We adopt this point of view and develop a framework which uses hypothetical perturbations of the data to construct meaningful target parameters. Specifically, we consider a setting where we observe an outcome of interest $Y\in\mathbb{R}$ and an associated compositional predictor $Z\in\Delta^{d-1}$. Our goal is to construct functionals of the conditional distribution $Y$ given $Z$ that are useful when analyzing compositional data. To illustrate this, we consider two examples in which existing approaches are insufficient.
Example 1 (Effects of presence or absence of microbes).
Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease. Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances. Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ has been recorded. While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies. One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$ and then compare the distributions
$Y\,|\,Z^{j}=0\quad\text{and}\quad Y\,|\,Z^{j}\neq 0$
for example, with an independence test or by regressing $Y$ on $\mathbbm{1}_{\{Z^{j}=0\}}$. It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components. As an example, assume the disease indicator is generated by
$Y\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{Z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{Z^{2}=0\}}),$
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1. If $Z$ did not live on the simplex, we could adjust for the dependence in the remaining coordinates by considering conditional dependencies of the form $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$. However, due to the simplex constraint $Z^{-j}$ completely determines $Z^{j}$, so the conditional independence always holds. In the Table 1, we also provide the results from applying the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
Example 1
method estimate 95%-CI
OLS of $Y$ on $\mathbbm{1}_{\{Z^{1}=0\}}$ $-0.12$ $(-0.17,-0.06)$
$\mathrm{CKE}^{1}$ using rf $0.15$ $(0.10,0.21)$
Example 2
method estimate 95%-CI
OLS of Y on $-G(Z)$ $-1.57$ $(-2.18,-0.97)$
$\mathrm{CDI}_{\mathrm{Gini}}$ using rf $0.63$ $(0.14,1.12)$
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal effect estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. (Left) The estimated marginal effect is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. (Right) The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
Example 2 (Effects of increased diversity).
Consider the problem of determining the effect of an increase in diversity of $Z$ on the response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this, one would again like to adjust for the remaining dependence within $Z$ to avoid the confounding biases. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjust for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.2 where we estimate the effect of racial diversity on income on grouped real data. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
In this work, we propose a procedure that accounts for potential confounding within $Z$. The main idea is to introduce a new target parameter via a hypothetical perturbation of the data. More concretely, a perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a path $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. A target parameter that captures the desired adjusted effect is then given by
$\lambda_{\psi}:=\frac{\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,1)]\right]-\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,0)]\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point. Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
$\psi(z,\gamma)=z+\gamma\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}.$
We can then interpret $\psi(z,\epsilon)$ for a small $\epsilon>0$ as slightly increasing the diversity of the point $z$. With this notion of an increase in diversity, a target parameter that captures the desired adjusted effect is given by
$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,Z=\psi(Z,\gamma)\right]\big{|}_{\gamma=0}\right],$
which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity. A key strength of using perturbations to define target parameters is that it makes them easy to communicate, interpret and discuss with practitioners. From our experience, practitioners often have a good understanding of what perturbation they are interested in and formalizing them mathematically as above is straight-forward. Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities. To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section 2.3.
While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the conditional expectation $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which due the simplex structure is generally a complex functional. We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization. To make this more concrete, assume that the structural causal model (SCM) (e.g., Pearl, 2009) given in Figure 1 (left) generates the data (the causal terminology is used only to provide a more intuitive introduction). Given a perturbation $\psi$ that we are interested in, we now propose to construct a bijection $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ that maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure 1 (right)). Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which standard adjustment works, that is, using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the causal effect. Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$. We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
(a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
$Y$ $\leftarrow f(Z,\eta)$ $Y$$Z$$\psi$$f$
[…]
$Y$$L$$W$$\psi$$h$$m$$h$
$L$ $\leftarrow m(W,\xi)$ $Y$ $\leftarrow h(L,W,\varepsilon)$
Figure 1: (a) Data-generating SCM on the $Z$-scale. $Z\in\Delta^{d-1}$ is a compositional vector, $Y$ is a response variable and $\eta$ is exogenous noise. (b) Reparametrized SCM on the $(L,W)$-scale. $(L,W)=\phi(Z)$ are the reparametrized predictors, $Y$ is the same response as in (a) and $\xi$ and $\varepsilon$ are exogenous noise variables. The reparametrization needs to disentangle the perturbation effect on $Z$ in a way such that intervening on $Z$ by setting it to $\psi(Z,\gamma)$ corresponds to intervening on only $L$ (red double head arrows).
The remainder of the paper is structured as follows. In Section 1.1 we discuss existing methods for compositional data analysis and how our approach contributes to this research. Section 1.2 collects the mathematical notation used throughout this work. Our perturbation-based framework is introduced in Section 2. In Section 2.1, we provide a way to construct interpretable perturbations based on suitably choosing the endpoint, direction and speed of the perturbation. We also provide a general construction for the aforementioned $L$ and $W$ for a wide range of perturbations. We end the section by discussing how to adapt our proposals to general geometries on the simplex including those employed by many current methods in compositional data analysis. In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied. In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness. Our semiparametric estimators and perturbation effects are then illustrated on both simulated data and two real-world applications in Section 4. Finally, we present the theoretical guarantees of our estimates in Section 5.
1.1 Relation to existing compositional data analysis approaches
Compositional data analysis has a long history dating back to Pearson (1897). The most prevalent class of methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. By considering ratios instead of compositional coordinates it avoids the common pitfalls of spurious correlations that occur by comparing coordinates. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). Furthermore, some of the perturbation effects that we highlight as interpretable overlap with effects that can also be deduced from the log-ratio perspective. Most prominently, we consider a class of compositional feature influence measures for which a certain member (multiplicative speed, see Example 3) corresponds to the compositional feature influence which has been previously proposed by Huang et al. (2023). It also corresponds to one of the coefficients in a log-contrast model (Aitchison and Bacon-Shone, 1984) if the conditional mean lies in the log-contrast model class. Our perturbation-based framework, however, has a fundamentally different starting point and allows for the construction of perturbations that are not related to the log-ratio approach.
The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however we believe its indirect use of the data is restrictive. This is best illustrated by the so-called zero-problem that occurs because the log-ratio approach, in general, cannot be applied in settings with too many zeros. This problem has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Other approaches advocate the use of different transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022). While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches. In contrast to these methods, our perturbation-based framework works directly with the compositional data and defines the target parameter from this perspective. In the presence of zeros, perturbations that directly incorporate or omit zeros can be selected based on the application. Furthermore, we can construct perturbation effects that directly capture the effect of setting certain components to zero.
1.2 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator, with the additional convention that for $0\in\mathbb{R}^{d}$ we have $\mathsf{C}(0)=(1/d,\dots,1/d)=:z_{\mathrm{cen}}$.
2 Target parameters via perturbations
For a response $Y\in\mathbb{R}$ and a compositional predictor $Z\in\mathcal{Z}\subseteq\Delta^{d-1}$, we seek to summarize the effect of $Z$ on the expectation of $Y$ under a pre-specified change of $Z$. We specify this change via perturbations.
Definition 1 (perturbations).
Let $\mathcal{D}\subseteq\mathcal{Z}\times[0,\infty)$ be such that for all $z\in\mathcal{Z}$ it holds that $\{z\}\times\{0\}\in\mathcal{D}$. We call a function $\psi=(\psi^{1},\dots,\psi^{d}):\mathcal{D}\to\Delta^{d-1}$ a perturbation if it satisfies for all $z\in\mathcal{Z}$ that $\psi(z,0)=z$. For all $z\in\mathcal{Z}$ we define the $\gamma$-domain and endpoint of $\psi$ (if it exists) by
$\mathcal{I}_{\psi}(z):=\{\gamma\in[0,\infty)\mid(z,\gamma)\in\mathcal{D}\}\quad\text{and}\quad E_{\psi}(z):=\lim_{\gamma\to\sup\mathcal{I}_{\psi}(z)}\psi(z,\gamma),$
respectively. We furthermore say that
  1. (a)
    $\psi$ is a binary perturbation if for all $z\in\mathcal{Z}$ it holds that $\mathcal{I}_{\psi}(z)=\{0,1\}$, i.e. $\mathcal{D}=\mathcal{Z}\times\{0,1\}$, and
  2. (b)
    $\psi$ is a directional perturbation if for all $z\in\mathcal{Z}$ there exists $\epsilon>0$ with $[0,\epsilon)\subseteq\mathcal{I}_{\psi}(z)$ and $\omega_{\psi}(z):=\partial_{\gamma}\psi(z,\gamma)\big{|}_{\gamma=0}$ exists. We further define the direction and speed of the perturbation for all $z\in\mathcal{Z}$ by
    $v_{\psi}(z):=\mathbbm{1}_{\{\omega_{\psi}(z)=0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
    respectively.
Binary perturbations describe discrete changes that move each point to a specific endpoint, that is, they capture the difference between $\gamma=0$ (doing nothing) and $\gamma=1$ (applying the perturbation). An example of this is setting the first component of $Z$ to 00. Directional perturbations describe ‘local’ changes to $Z$, that is, differences in $\gamma=0$ (doing nothing) and $\gamma=\epsilon$ for some small $\epsilon>0$ (moving slightly along the perturbation path). We later focus on the derivative of the perturbation at $\gamma=0$ as this captures the rate of change of directional perturbations at the initial point. As we only intend to use this initial derivative, it suffices to consider perturbations of the form
$\psi(z,\gamma)=z+\gamma s_{\psi}(z)v_{\psi}(z).$ (1)
We will discuss how to construct perturbations with interpretable directions and speeds in Section 2.1.2. Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex. In this work, we only consider binary and directional perturbations but other classes of perturbations could be of interest too. Perturbations can also be defined on more general sample spaces than simplices and many of the perturbation-related concepts described here can be readily extended to these settings.
Using the two types of perturbations, we can define informative summary statistics that capture the effect of the perturbations on the response. We do this by using contrasts of the perturbation, similar to how causal effects are defined via contrasts of interventions in the causal inference literature. Formally, we define the following target parameters associated with each type of perturbation.
Definition 2 (average perturbation effects).
Let $\mathcal{Z}\subseteq\Delta^{d-1}$, let $(Y,Z)\in\mathbb{R}\times\mathcal{Z}$ be random variables and define $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$. Then, for a binary perturbation $\psi$, if $f$ is identified on $\{\psi(z^{\prime},1)\in\Delta^{d-1}\mid z^{\prime}\in\mathcal{Z}\}$ and $\mathbb{P}(Z\neq\psi(Z,1))>0$, we define the average (binary) perturbation effect by
$\lambda_{\psi}:=\frac{\mathbb{E}\left[f(\psi(Z,1))\right]-\mathbb{E}\left[Y\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex(see Section S1 for a precise definition) at $z$, we define the average (directional) perturbation effect by
$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}f(\psi(Z,\gamma))\big{|}_{\gamma=0}\right].$
We showcase different potential applications for both types of perturbation effect in the numerical experiments in Section 4. For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$. This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint. While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities. In particular, they are identifiable from the observational distribution as long as the conditions in Definition 2 are satisfied. We discuss the assumptions required for causal conclusions in Section 2.3.
Direct estimation of both target parameters relies on estimating the regression function $f:z\mapsto\mathbb{E}[Y\mid Z=z]$ and taking averages (see Section 3). An alternative estimation procedure for the effect of a binary perturbation $\lambda_{\psi}$ can be derived from noting that, if $L=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W=\psi(Z,1)$, then
$\mathbb{E}[Y\,|\,L=1,W]=\frac{\mathbb{E}[YL\,|\,W]}{\mathbb{E}[L\,|\,W]}=f(W).$
Therefore, we can write
$\lambda_{\psi}=\frac{\mathbb{E}[\mathbb{E}[Y\,|\,L=1,W]-Y]}{\mathbb{P}(L=0)},$ (2)
which is an expression more amenable to analysis by semiparametric theory and thus allows us to construct efficient estimators. If we are willing to assume that $f(\psi(Z,1))-f(Z)$ is constant when $L=0$ and equal to, say $\theta$, then
$\lambda_{\psi}=\frac{\mathbb{E}[\{f(W)-f(Z)\}(1-L)]}{\mathbb{P}(L=0)}=\theta$
and furthermore
$\mathbb{E}[Y\,|\,Z]=f(Z)=f(W)L+(f(Z)-f(W)+f(W))(1-L)=\theta L+\underbrace{f(W)-\theta}_{h(W)},$
so $\theta$ is the coefficient of $L$ in a partially linear model of $Y$ on $L$ and $W$.
It would be convenient to also phrase the average directional perturbation effect $\tau_{\psi}$ as a quantity that is similarly amenable to semiparametric analysis and permits us to impose a partially linear model. To motivate the upcoming condition, we consider what happens when $Z$ takes values in $\mathbb{R}^{d}$. In that case, we can write
$\partial_{\gamma}f(\psi(z,\gamma))\big{|}_{\gamma=0}=\nabla f(z)^{\top}\omega_{\psi}(z).$
We now want to isolate the change of $\psi$ in the direction $v_{\psi}$ into a real-valued variable $L$ while keeping the remainder of the ‘information’ in $Z$ in an additional variable $W$. For a reparametrization $\phi$ of $z$ into $(\ell,w)\in\mathbb{R}\times\mathcal{W}$, we define the reparametrized $f$ as $g:(\ell,w)\mapsto f(\phi^{-1}(\ell,w))$. Then, in order for the reparametrization $\phi$ to separate the ’information’ appropriately, we require
$\partial_{\ell}g(\ell,w)=\partial_{\ell}f(\phi^{-1}(\ell,w))=\nabla f(\phi^{-1}(\ell,w))^{\top}\partial_{\ell}\phi^{-1}(\ell,w)$
to equal $\partial_{\gamma}f(\psi(z,\gamma))\big{|}_{\gamma=0}$, that is,
$\omega_{\psi}(\phi^{-1}(\ell,w))=\partial_{\ell}\phi^{-1}(\ell,w).$
This derivation is only morally correct when $Z\in\Delta^{d-1}$ ($\nabla f$ is not defined in the usual sense for simplex-valued functions) and to make it formally correct one needs to apply techniques from differential geometry. The proof of the following result can be found in Section S2.1 of the supplementary material.
Proposition 1 (derivative-isolating).
Consider the setting of Definition 2. Let $\mathcal{W}$ be a set and $(\mathcal{L}_{w})_{w\in\mathcal{W}}$ a family of subsets of $\mathbb{R}$. Suppose that $\phi=(\phi^{L},\phi^{W}):\mathcal{Z}\to\{(\ell,w)\in\mathbb{R}\times\mathcal{W}\mid w\in\mathcal{W},\ell\in\mathcal{L}_{w}\}$ is a bijection and $\psi$ a directional perturbation such that for all $z\in\mathcal{Z}$ it holds that $\omega_{\psi}(z)\neq 0$. Suppose that for all $w\in\mathcal{W}$ and all $\ell\in\mathcal{L}_{w}$, we have
$\omega_{\psi}(\phi^{-1}(\ell,w))=\partial_{\ell}\phi^{-1}(\ell,w).$ (3)
Then, defining $g:(\ell,w)\mapsto f(\phi^{-1}(\ell,w))$ and $(L,W)=\phi(Z)$, the average perturbation effect satisfies
$\tau_{\psi}=\mathbb{E}\left[\partial_{\ell}g(\ell,W)\bigm{|}_{\ell=L}\right].$ (4)
We say that $\phi$ is derivative-isolating if it satisfies (3).
From this result it follows that if we are able to find a derivative-isolating reparametrization (we describe how to find such reparametrizations for certain perturbations in Section 2.1.2) the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993). It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (5)
Details on the different estimation approaches are discussed in Section 3.
Remark 1 (zero speeds).
While we do not allow $\omega_{\psi}(z)=0$ in the definition of derivative-isolating reparametrizations, it is clear from the definition of $\tau_{\psi}$ that such points do not contribute to the perturbation effect. We can therefore allow for such points if we appropriately adapt the estimates. In Section S3.2 of the supplementary material, we describe how to modify the estimation procedures introduced in Section 3 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (potentially geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from a summary statistic and assuming it satisfies a partially linear model as in (5) with $L$ being the summary statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
In the remainder of this section, we discuss interpretable simplex endpoints in Section 2.1.1, provide details on the selection of directions and speed together with the construction of derivative-isolating reparametrizations in Section 2.1.2 and finally in Section 2.1.3 discuss how one could construct directional perturbations that respect a different geometry than the Euclidean one. The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications. The more applied reader may jump directly to that section and skip the more technical discussion on perturbations here.
2.1.1 Endpoints in the simplex
As formally defined in Definition 1, an endpoint function $E:\Delta^{d-1}\rightarrow\Delta^{d-1}$, if it exists, determines where the perturbation ends when $\gamma$ is increased as much as possible. The simplest type of endpoint function on the simplex is a constant function, that is, there exists a fixed point $z_{*}\in\Delta^{d-1}$ such that for all $z\in\Delta^{d-1}$ it holds that $E(z)=z_{*}$. For example, we can choose the fixed endpoint $z_{*}=e_{j}$, which corresponds to a vertex on the simplex and hence a perturbation with this endpoint moves proportions from the other coordinate into the $j$th one. A further example, is to use the center of the simplex, i.e. $z_{*}=z_{\mathrm{cen}}$. As alluded to in the introduction, perturbations with $z_{\mathrm{cen}}$ as the endpoint can be seen as perturbing observations in the direction of equal distribution and hence capture a notion of diversity. For later reference, we call all perturbation $\psi$ with endpoint function $E_{\psi}:z\mapsto z_{\mathrm{cen}}$ centering perturbations.
A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex. An amalgamation is a well-studied operation (e.g., Aitchison, 2003) which consist of merging parts of the compositional components into others. Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
$\mathsf{A}_{A\to B}(z)^{A}:=0,\quad\mathsf{A}_{A\to B}(z)^{B}:=\mathsf{C}(z^{B})(\|z^{A}\|_{1}+\|z^{B}\|_{1}),\quad\mathsf{A}_{A\to B}(z)^{[d]\setminus(A\cup B)}:=z^{[d]\setminus(A\cup B)}.$
The $(A\to B)$-amalgamation has the effect of setting components in $A$ to 00, scaling up the components in $B$ such that the subcomposition in $B$ remains fixed (i.e., $\mathsf{C}(\mathsf{A}_{A\to B}(z)^{B})=\mathsf{C}(z^{B})$) and leaving the remaining components fixed. We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations. An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e. perturbations that push towards a vertex of the simplex. Amalgamations, however, do not need to have fixed endpoints but allow for more general endpoints that can also depend on the initial point. Several perturbations on $\Delta^{2}$ are visualized in Figure 2 below.
2.1.2 Directions, speeds and derivative-isolating reparametrizations
When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as
$\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$
An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
$v_{E}(z):=\frac{E(z)-z}{\|E(z)-z\|_{1}},$ (6)
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can be easily extended to settings where zero speeds are allowed. In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, for which a partially linear model as in (5) can be assumed, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
As a fundamental building block for the general construction below, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
$\psi(z,\gamma):=z+\gamma\frac{E(z)-z}{\|E(z)-z\|_{1}}.$
Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions. We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2. The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=-\ell w_{v}+w_{E}$, so
$\partial_{\ell}\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=w_{v}=\partial_{\gamma}\psi(\phi_{\mathrm{unit}}^{-1}(\ell,(w_{v},w_{E})),\gamma)\big{|}_{\gamma=0}=\omega_{\psi}(\phi_{\mathrm{unit}}^{-1}(\ell,w))$
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since these type of unit-speed perturbations may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic. The proof of the following result is given in Section S2.2 of the supplementary material.
Proposition 2.
For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$. Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$. Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(v(z),E(z))$ with $v$ the canonical endpoint direction defined in (6). Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
  1. (a)
    (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function. For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
    $-s_{w}(\delta)^{-1}=\partial_{\delta}t_{w}(\delta).$ (7)
    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$ and any other perturbation with the same speed and direction.
  2. (b)
    (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic. For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing. Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
    $s(z):=-\left(\partial_{\delta}t_{\phi^{W}(z)}(\delta)\big{|}_{\delta=\|E(z)-z\|_{1}}\right)^{-1},$
    and any other perturbation with the same speed and direction.
Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (7). Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint. Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized. We illustrate the use of Proposition 2 by constructing several directional perturbations.
Example 3 (Feature influence).
Suppose we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in\Delta^{d-1}$. We can investigate the effect of such a change by means of a $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{e_{j}\}$ that $v_{\psi}(z)=\frac{e_{j}-z}{\|e_{j}-z\|_{1}}$. While for some purposes it may be perfectly reasonable to use a constant speed $s_{\psi}\equiv 1$, in other settings this speed can be unintuitive. For example, if we imagine the observed $Z$ as being generated from an unobserved $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale by $Z:=\mathsf{C}(X)=\frac{1}{\|X\|_{1}}X$, it might be more natural to consider which speeds are induced on the simplex from hypothetical changes to $X^{j}$. A first attempt might be to consider the speed resulting from adding a constant $c$ to $X^{j}$, however, since
$\|\partial_{c}\mathsf{C}(X+ce_{j})\bigm{|}_{c=0}\|_{1}=2\tfrac{1}{\|X\|_{1}}\left(1-\tfrac{X^{j}}{\|X\|_{1}}\right),$
this speed is not scale-invariant for $X$, that is, the speed when considering the counts $X$ and $cX$ for $c>0$ may differ. This means that the speed on the simplex would depend on the unobserved quantity $\|X\|_{1}$ so the perturbation is ill-defined. Instead, we can consider the speed resulting from multiplying $X^{j}$ by $1+c$ (so that $c=0$ corresponds to no change) which yields a well-defined perturbation on the simplex since this operation is scale-invariant. More formally, letting $\odot$ denote the Hadamard (point-wise) product, the resulting speed is
$\|\partial_{c}\mathsf{C}(X\odot(1+ce_{j}))\bigm{|}_{c=0}\|_{1}=2\tfrac{X^{j}}{\|X\|_{1}}\left(1-\tfrac{X^{j}}{\|X\|_{1}}\right),$
which is scale-invariant and therefore meaningful to consider on the simplex. On the simplex, this corresponds to the speed
$s_{\psi}(z)=2z^{j}(1-z^{j})$
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a). Since for all $z\in\mathcal{Z}$ it holds that $E(z)^{j}=1$ and $v(z)^{j}=\frac{1}{2}$, we have that $s_{w}$ (defined in Proposition 2) does not depend on $w$ and is of the form
$s_{w}:\delta\mapsto\frac{(2-\delta)\delta}{2}.$
Hence, solving (7) amounts to solving
$-\frac{2}{(2-\delta)\delta}=\partial_{\delta}t_{w}(\delta)$
which has solutions of the form $t_{w}(\delta)=\log\left(\frac{2-\delta}{\delta}\right)+C$ for $C\in\mathbb{R}$. For simplicity, we choose $C=0$ and note that as $\|e_{j}-z\|_{1}=2(1-z^{j})$, we have
$t_{w}(\delta)=\log\left(\frac{z^{j}}{1-z^{j}}\right).$
We can therefore conclude (dropping the constant endpoint from the reparametrization) that
$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{z^{j}}{1-z^{j}}\right),v_{\psi}(z)\right)$
is a derivative-isolating reparametrization for the $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation $\psi$ with multiplicative speed.
Example 4 (Amalgamation influence).
We can extend the considerations of Example 3 to more general amalgamating perturbations. For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation. Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale. We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation so we suppose now that $A\cup B=[d]$. The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
$c\mapsto\mathsf{C}(X\odot(1+cu_{B})),$
which similarly corresponds to a well-defined perturbation on the simplex. Computing the speed, we obtain that
$\|\partial_{c}\mathsf{C}(X\odot(1+cu_{B}))\bigm{|}_{c=0}\|_{1}=\frac{2\|X^{A}\|_{1}\|X^{B}\|_{1}}{(\|X^{A}\|_{1}+\|X^{B}\|_{1})^{2}}.$
We can use this speed even when $A\cup B\neq[d]$ resulting in the general multiplicative speed
$s_{\psi}(z)=\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}.$ (8)
Using Proposition 2 (a) as in Example 3, we further obtain that
$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\right),(\mathsf{A}_{A\to B}(z),v_{\psi}(z))\right)$
is a derivative-isolating reparametrization for $\psi$ with multiplicative speed.
Example 5 (Diversity with Gini-speed).
Suppose we are interested in the effect of diversifying $z\in\Delta^{d-1}$. We can investigate such an effect by using a centering directional perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{z_{\mathrm{cen}}\}$ that $v_{\psi}(z)=\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}$. As every element of $\mathcal{W}$ as defined in Proposition 2 is of the form $(z_{\mathrm{cen}},u)$ for a direction $u$, we instead drop the endpoint from the set and use $\mathcal{W}=\{u\in\mathbb{R}^{d}\mid\|u\|_{1}=1,\sum_{j=1}^{d}u^{j}=0\}$ instead. In contrast to the Examples 3 and 4 it is not as obvious how to choose an interpretable speed. We therefore start from a summary statistic and apply Proposition 2 (b) to derive the corresponding speed and derivative-isolating reparametrization. For econometricians a natural choice of diversity measure might be the Gini coefficient. The Gini coefficient $G:\Delta^{d-1}\to[0,1]$ is defined for all $z\in\Delta^{d-1}$ by
$G(z):=\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|z^{j}-z^{k}|.$ (9)
Using $z=z_{\mathrm{cen}}-v_{\psi}(z)\|z_{\mathrm{cen}}-z\|_{1}$, we can express this as
$G(z)=\frac{\|z_{\mathrm{cen}}-z\|_{1}}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|.$
Hence, defining for all $w\in\mathcal{W}$ the strictly decreasing and differentiable function $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ ($\mathcal{I}_{w}$ is defined in Proposition 2) for all $\delta\in\mathcal{I}_{w}$ by
$t_{w}(\delta):=-\frac{\delta}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|w^{j}-w^{k}|,$
we have that $t_{v_{\psi}(z)}(\|z_{\mathrm{cen}}-z\|_{1})=-G(z)$. Applying Proposition 2 (b), we conclude that
$\phi=(\phi^{L},\phi^{W}):z\mapsto(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}})$
is a derivative-isolating reparametrization for perturbations with direction $v_{\psi}$ and speed given by
$s_{\psi}(z):=\frac{2d}{\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|}.$ (10)
Similar arguments could be applied to other popular measures of diversity such as the Shannon entropy or the Gini–Simpson index.
2.1.3 Perturbations arising from alternative simplex geometries
When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ by
$\mathrm{CLR}(z)^{j}:=\log(z^{j})-\frac{1}{d}\sum_{j=1}^{d}\log(z^{j}),\quad\forall j\in[d].$
This transformation maps the interior of $\Delta^{d-1}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$. The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
$\mathrm{CLR}^{-1}(x)^{j}:=\frac{\exp(x^{j})}{\sum_{j=1}^{d}\exp(x^{j})},\quad\forall j\in[d],$
or more simply as $\mathrm{CLR}^{-1}(x)=\mathsf{C}(\exp(x))$ where $\exp$ is applied componentwise. The inverse of the $\mathrm{CLR}$ transform can in fact be defined on all of $\mathbb{R}^{d}$ rather than just the range of the $\mathrm{CLR}$-function as is clear from the definition. The $\mathrm{CLR}$-transformation induces what is called the Aitchison geometry (Egozcue and Pawlowsky-Glahn, 2006) on the simplex. The Aitchison geometry induces a different notion of direction than the Euclidean notion considered above and, in particular, the direction of the line connecting $z$ to an endpoint may differ from the direction in (6) (see, for example, Figure 2). However, as $\mathrm{CLR}(z)$ is a bijection on the open simplex, we can just as well consider perturbations on $\mathbb{R}^{d}$ directly and use arguments similar to those in Proposition 2 but now for the perturbation on the log-ratio scale. We illustrate this procedure in the following two examples and discuss connections to the centering and amalgamating perturbations defined above.
Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations (the diversity perturbation in blue has Gini-speed (see (10)) and multiplicative speed (see (8)) for the amalgamating perturbations in red and green) and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$. The types of perturbation are distinguished by color. The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds. The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
Example 6 (Diversity in the Aitchison geometry).
Suppose we are interested in the effect of diversifying $z\in(\Delta^{d-1})^{\circ}$. As in Example 5, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space. This line transformed back to the simplex can be written as
$\{\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))\mid\gamma\in[0,1]\}.$
Therefore, computing the diversifying direction in the Aitchison geometry amounts to differentiating the expression above at $\gamma=0$ and normalizing. Applying the chain rule, this yields
$\partial_{\gamma}\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
which can be normalized to a unit $1$-norm vector. The direction of this derivative can be vastly different from the centering direction used in Example 5. To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale, $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale. It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
Example 7 (Feature influence in the Aitchison geometry).
Suppose, as in Example 3, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$. For $\mathrm{CLR}(z)\in\mathbb{R}^{d}$ we have a natural notion of increasing $z^{j}$ inherited from $\mathbb{R}^{d}$, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$. The derivative of the resulting perturbation on the simplex is for all $k\in[d]$ given by
$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathbbm{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
It is straight-forward to check that the speed and direction of this perturbation match exactly the multiplicative perturbation derived in Example 3. One can therefore view an additive log-ratio perturbation as a multiplicative speed perturbation.
2.2 Perturbation effects on the simplex
To make the discussion more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1. We motivate each perturbation effect with an example application to showcase potential use cases. A list summarizing all proposed perturbation effects together with their respective derivative-isolating reparametrizations that are used for efficient estimation is provided in Table 2. We encourage practitioners to develop further application-specific perturbations that optimally target the research question they are interested in.
2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.3, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then wants to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ the conditional independence
$Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$ (11)
is trivially satisfied and feature importance or influence measures based on (11) are meaningless. Intuitively, this stems from the fact that changing a single component always requires a change in at least some of the remaining components. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$. A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
In datasets with many zeros, which is common for example in microbiome data, it can also be interesting to focus only on the effect of zeros in individual components. This can be achieved by considering a binary perturbation $\psi$ given by $\psi(z,1)^{j}=0$ and $\psi(z,1)^{-j}=\mathsf{C}(z^{-j})$. We call the resulting average binary perturbation effect $\lambda_{\psi}$ the compositional knock-out effect for the $j$th feature, or $\mathrm{CKE}^{j}$ for short. The $\mathrm{CKE}^{j}$ allows for a straightforward analysis of zeros in compositional data, which has been a long-standing problem with many existing approaches.
2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (the centering perturbations defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 5 such that $L$ is equal to the negative Gini coefficient. The construction in Example 5 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.2, we use data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression.
We do not define a binary centering perturbation here as it requires a positive probability of observing $Z=z_{\mathrm{cen}}$, which is rarely the case.
2.2.3 Effects of amalgamations
As a third example we consider perturbation effects that capture the influence of general amalgamations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As above, we are again interested in effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example 4).
Analogously, we could be interested in the effect of removing particular activities (e.g. gambling) entirely and spending the time on desirable activities (e.g. sports) while fixing time spent on a third group of activities (e.g. eating, sleeping). We can capture this effect using a binary amalgamating perturbation $\psi$. We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
Effect of changes in Target Speed $s(z)$ Endpoint $E(z)$ $\phi^{L}(z)$
individual components $\mathrm{CFI}^{j}_{\mathrm{unit}}$ $1$ $e_{j}$ $-2(1-z^{j})$
$\mathrm{CFI}^{j}_{\mathrm{mult}}$ $2z^{j}(1-z^{j})$ $e_{j}$ $\log(\frac{z^{j}}{1-z^{j}})$
$\mathrm{CKE}^{j}$ binary $e_{j}$ $\mathds{1}_{\{z^{j}=0\}}$
diversity $\mathrm{CDI}_{\mathrm{unit}}$ $1$ $z_{\mathrm{cen}}$ $-\|z_{\mathrm{cen}}-z\|_{1}$
$\mathrm{CDI}_{\mathrm{Gini}}$ see (10) $z_{\mathrm{cen}}$ $-G(z)$, see (9)
amalgamations $\mathrm{CAI}^{A\to B}_{\mathrm{unit}}$ $1$ $\mathsf{A}_{A\to B}(z)$ $-2\|z^{A}\|_{1}$
$\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ $\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}$ $\mathsf{A}_{A\to B}(z)$ $\log\big{(}\tfrac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\big{)}$
$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
Table 2: List of perturbation effects considered in this work and what effects they capture. For each perturbation effect the corresponding speed, endpoint and $L$-term of the reparametrization used for efficient estimation is included. All reparametrizations use the canonical direction $v_{E}$ based on the endpoint, see (6). The $W$-term $\phi^{W}(z)$ of the reparametrization is $E(z)$ for the binary effects and $(v_{E}(z),E(z))$ for the directional effects.
2.3 Perturbations as causal quantities
Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions. To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference. Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data generating mechanism – often called interventions. In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data generating mechanism under hypothetical interventions. The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution. Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism. To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework (Pearl, 2009; Bongers et al., 2021). Similar considerations, however, also apply with the potential outcome framework.
As an example, consider the SCM over the variables $(H,Z,Y)\in\mathbb{R}^{q}\times\Delta^{d-1}\times\mathbb{R}$ given by
:$\mathcal{M}:\,\begin{cases}&H\leftarrow\varepsilon_{H}\\ &Z\leftarrow h(H,\varepsilon_{Z})\\ &Y\leftarrow g(H,Z)+\varepsilon_{Y},\end{cases}$
where $h:\mathbb{R}^{q}\times\mathbb{R}\rightarrow\Delta^{d-1}$ and $g:\mathbb{R}^{q}\times\Delta^{d-1}\rightarrow\mathbb{R}$ are measurable functions and $(\varepsilon_{H},\varepsilon_{Y},\varepsilon_{Z})$ are jointly independent noise terms with expectation zero. The SCM $\mathcal{M}$ generates a distribution over $(H,Y,Z)$ which is called the observational distribution. Intervening corresponds to changing individual structural assignments while keeping everything else fixed such that the modified SCM again induces a distribution – called the interventional distribution. The example SCM $\mathcal{M}$ is a commonly assumed causal structure in which the predictor $Z$ causally precedes the response but is confounded by $H$.
Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$. One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation (Pearl, 2009), this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$. For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components. Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions. Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors. More specifically, let $\psi$ be a perturbation as defined in Definition 1. Then, for all $\gamma\in[0,1]$, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$. In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$. Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself. The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
$\lambda_{\psi}^{\operatorname{causal}}:=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,1))\right]-\mathbb{E}\left[Y\right]\big{]}}{\mathbb{P}(Z\neq\psi(Z,1))}$
and the causal average directional perturbation effect
$\tau_{\psi}^{\operatorname{causal}}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,\gamma))\right]\big{|}_{\gamma=0}\right].$
Unlike their non-causal versions, the causal average perturbation effects may not be identifiable from the observational distribution. Therefore, to estimate them, we need to make additional causal assumptions. For example, we may assume the variable $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$. In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}\big{[}\mathbb{E}[Y\mid Z=\psi(Z,\gamma),H]\big{]},$ (12)
and hence is identifiable. The formula in (12) is a special case of the adjustment formula (e.g., Pearl, 2009), but other ways of identifying this effect also exist (e.g., propensity score matching). It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$. In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$. This is for example done in the numerical experiment in Section 4.2.
Finally, let us shortly discuss under which assumptions the causal and non-causal average perturbation effects coincide. Obviously, one sufficient condition is to assume that
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}[Y\mid Z=\psi(Z,\gamma)].$ (13)
Such a condition holds by definition whenever $Z$ is an exogenous variable (i.e., does not depend on other variables in its structural assignment) but can also hold more generally. A common assumption to ensure (13) is to assume that $Z$ and $Y$ are not confounded, that is, $Z$ and $Y$ have no common causes (this is known as ignorability in the potential outcome framework). In the SCM $\mathcal{M}$ this would for example mean that $H$ does not exist or either $g$ or $h$ do not depend on $H$. In many cases, this is not satisfied and one should instead carefully argue for which causal assumptions are reasonable and adjust the estimation accordingly, e.g., by conditioning on additional covariates as in (12).
3 Estimating perturbation effects
In this section, we describe how to estimate and test hypotheses about the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$. We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice. However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathbbm{1}_{\{}z=\psi(z,1)\}$ and $\phi^{W}:z\mapsto\psi(z,1)$. This allows us to transform our estimation problem from a compositional data problem where we observe $n$ i.i.d. copies of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ to a generic problem with $n$ observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$. We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient tests and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
$\tau_{\psi}=\mathbb{E}[\partial_{\ell}f(\ell,W)\bigm{|}_{\ell=L}],$
where $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$. A naive approach, which is known as plug-in estimation, is to estimate $\tau_{\psi}$ by first learning a differentiable estimate $\widehat{f}$ of $f$ using all $n$ observations and then computing
$\widetilde{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i}).$
While such an estimate can be consistent, it is rarely asymptotically Gaussian, which invalidates standard asymptotic inference. Using the corresponding plug-in estimate of the asymptotic variance, given by
$\widetilde{\sigma}^{2}:=\frac{1}{n}\sum_{i=1}^{n}(\partial_{\ell}\widehat{f}(L_{i},W_{i}))^{2}-\widetilde{\tau}_{\psi}^{2},$ (14)
hence does not provide valid inference (see Figure 3 (left)) as the confidence intervals can be biased or too narrow.
Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section 4.1 with $d=15$ and $n=1000$. The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter. The confidence intervals are ordered first by whether the true value is in the interval, then by the value of the estimate.
There are two main issues with this estimate; (a) repeated usage of the same data for both the estimation of $f$ and the computation of the target (double-dipping) and (b) estimation bias inherited from the bias in $\widehat{f}$. Issue (a) can be resolved by splitting the dataset into two parts of roughly equal size and using the first half to compute $\widehat{f}$ and the second part to compute $\widetilde{\tau}_{\psi}$. To increase finite-sample performance of this splitting procedure, one can swap the roles of the two halves and average the resulting estimates. This procedure is known as cross-fitting. Sometimes it is preferable to split in more than two folds to increase the sample size in the estimation of $\widehat{f}$ and to derandomize the resulting estimate. Issue (b) is more complicated to resolve and requires the machinery of semiparametric theory. One well-established approach is to start from the plug-in estimator and derive a bias-correction, known as a one-step correction, that can be added to the plug-in estimator and weakens the conditions under which the estimator is asymptotically Gaussian. The one-step correction can be found by computing the influence function of the target parameter, which can be thought of as analogous to a derivative when the target parameter is viewed as a function defined on a space of distributions. For more details on this general procedure, see e.g. Kennedy (2023). It can be shown that these bias-corrected estimators are asymptotically efficient, meaning that, at least asymptotically, they have the smallest possible variance of any estimator. Influence functions have already been computed for the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$ that we estimate here and we will describe the details of these estimates in the following sections. Figure 3 contains an illustration of the coverage of confidence intervals resulting from target parameter and variance estimates from the different methods mentioned above.
3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently powerful) regression methods.
3.2.1 Nonparametric estimation of the binary perturbation effect
Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}.$
The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities). The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of average predictive effects (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature. The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ is related to the augmented inverse propensity weighted (AIPW) estimator and is given by
$\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i},$
where $\widehat{f}$ is an estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\pi}$ is an estimator of the propensity score $\pi:w\mapsto\mathbb{P}(L=1\,|\,W=w)$. An efficient estimator for $\widehat{\lambda}_{\psi}$ is thus
$\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (15)
A full algorithm is given in Algorithm 2. It is worth noting that this estimator is sensitive to the estimation of the propensity score $\pi$ which can be difficult when the dimension of $W$ is large or the relationship between $L$ and $W$ is complicated (see Section S3.1 of the supplementary material for an illustration of this).
3.2.2 Nonparametric estimation of the directional perturbation effect
Consider a directional perturbation $\psi$ with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and the corresponding perturbation effect $\tau_{\psi}$. By the arguments in Section 2, setting $L:=\phi^{L}(Z)$ and $W:=\phi^{W}(Z)$, we have that
$\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}].$
This quantity is another well-known target of inference; the average partial effect (Newey and Stoker, 1993; Chernozhukov et al., 2022; Klyne and Shah, 2023). The one-step corrected estimator in this setting is given by
$\widehat{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i})-\widehat{\rho}(L_{i},W_{i})(Y_{i}-\widehat{f}(L_{i},W_{i}))$ (16)
where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$. A full algorithm for the estimation is given in Algorithm 1 below. Both the estimation of $f$ with computable derivative $\partial_{\ell}f$ and $\rho$ can be challenging and we show in Section 4 that the overall estimator (16) is sensitive to both estimates (see Section S3.1 of the supplementary material for an illustration of this).
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. A full description of our smoothing method is provided in Section S3.3 of the supplementary material. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1). Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplement.
3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions. One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (17)
With this assumption, $\theta$ is exactly equal to the perturbation effect. Efficient estimation of $\theta$ in a partially linear model is a well-studied problem (e.g., Robinson, 1988; Chernozhukov et al., 2018) and one estimator (sometimes referred to as the partialling out estimator) is given by
$\widehat{\theta}:=\frac{\sum_{i=1}^{n}\{Y_{i}-\widehat{g}(W_{i})\}\{L_{i}-\widehat{m}(W_{i})\}}{\sum_{i=1}^{n}\{L_{i}-\widehat{m}(W_{i})\}^{2}},$ (18)
where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$. A full algorithm is given in Algorithm 3. In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions. This comes at the expense of assuming that there is no interaction effect between $L$ and $W$ on $Y$ and that the effect of $L$ is linear on $Y$. Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above. In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S1 in Section S3.1.3 of the supplementary material for a proof of this fact).
4 Numerical experiments
We now present three numerical experiments illustrating the usage and performance of our methods. In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3. In Section 4.2 we analyze the effect of diversity on income based on grouped data from the 1994 US census. In Section 4.3 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.3 where we used the c-lasso package (Simpson et al., 2021).
4.1 Semiparametric estimators
In this section we investigate the performance of the semiparametric estimators outlined in Section 3 with $W$ as a simplex-valued random variable. More specifically, we compare the different approaches of estimating
$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}\quad\text{and}\quad\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}]$
from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (15) and (16). See Algorithms 1 and 2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
  • NPM_no_x: The same estimators as NPM but without cross-fitting.
  • NPM_oracle: The same estimators as NPM including cross-fitting but using the true nuisance functions for the one-step-correction (propensity score $\pi$ in the binary case and conditional score function $\rho$ in the directional case). This is an oracle as the true nuisance functions are unknown in practice.
  • PLM: The partially least model based estimators in (18) including cross-fitting. See Algorithm 3 with $K=2$ for the detailed estimators.
  • PLM_no_x: The same estimators as PLM but without cross-fitting.
  • plugin: The naive plugin versions of the NPM estimates. In the binary effect case, the estimator in (15) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (16) without the term involving $\widehat{\rho}$.
  • plugin_no_x: The same estimators as plugin but without cross-fitting.
The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plugin is estimated using an estimator akin to the one given in (14). Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{g}$ and $\widehat{m}$ and NPM and NPM_no_x additionally on the score estimate $\widehat{\rho}$. For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction. For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplement, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
We consider four data generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$. In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathbbm{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$. We then consider the following four generative models
  • Binary $L$ with partially linear $Y$:
    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=L+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
  • Binary $L$ with nonparametric $Y$:
    $L=U_{0}B(W_{1})+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
  • Continuous $L$ with partially linear $Y$:
    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
  • Continuous $L$ with nonparametric $Y$:
    $L=B(W_{1})+(1+B(W_{1}))\xi\quad\text{and}\quad Y=2B(W_{1})L+B(W_{1})+\varepsilon$
The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$). The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$. For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above. We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=15$ in Figure 4 while coverage rates for $d=3$ and $d=75$ are reported in Section S4.2 of the supplementary material.
Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedures over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$. PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting. NPM works in both settings but can be less robust for small sample sizes. Further, it can be seen that the plugin estimator and the estimators without cross-fitting do not lead to correct coverage in some of these settings.
In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated. We see that the plugin estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$. In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the tests, e.g. for $n=1000$ and binary $L$. In the cases where cross-fitting is not necessary, it does not seem to negatively impact the performance of the test.
4.2 Compositional confounding in relationship between income and race
In this section we provide an example based on real data that illustrates how failing to adjust for compositional confounding can lead to incorrect conclusions about effects of compositional variables. We consider the ‘Adult’ dataset available from the UCI machine learning repository (Becker and Kohavi, 1996). This dataset is based on an extract from the 1994 US Census database and consists of data on 48,842 individuals. The response variable in the data, compensation, is a binary variable where 00 and $1$ encode whether an individual makes below or above 50,000 USD per year, respectively. Additionally, we consider the demographic features age (numeric, between 17 and 90), education (categorical, with 16 categories), sex (binary, 00 encoding male and $1$ female) and race (categorical, with categories ’White’, ’Black’ and ’Other’).
We now use this individual-level data to construct community-level datasets by artificially grouping individuals together. This allows us to construct settings, where there is a known dependence between the racial diversity of the communities and the response variable. As we show below, this illustrates the potential pitfalls of ignoring compositional structure in the data and how the approach based on perturbations solves these issues. More specifically, we group all individual-level observations into groups of around $50$ individuals. We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group. Similarly, we create a predictor variable $X\in\mathbb{R}\times[0,1]\times\{1,\dots,16\}$ by averaging age and sex and selecting the majority class of the categorical variable education (breaking ties randomly). Finally, we create a compositional predictor $Z\in\Delta^{2}$ by computing the proportions of ‘White’, ‘Black’ and ‘Other’ individuals in each group. To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2). We construct these categories in two different ways:
  • Grouping by race and education: Groups in category 1 are sampled to be similarly diverse and with lower educational levels than the total population and category 2 groups are sampled to be more diverse and with higher educational levels than the total population. The groups in category 0 consist of groups constructed randomly from the remaining individuals.
  • Grouping by race and compensation: Groups in category 1 are sampled to be similarly diverse and with lower average compensation than the total population and category 2 groups are sampled to be more diverse and with higher average compensation than the total population. The groups in category 0 consist of groups constructed randomly from the remaining individuals.
We then sample groups from each category, using a weighted sampling approach described in detail in Section S4.3 of the supplementary material. This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation. For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 5. In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation).
Figure 5: Top: Racial compositions for both types of grouping in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$). Depending on the color the dependence between $Y$ and the Gini coefficient changes.
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use minus the Gini coefficient $L=-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 5, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3. The naive estimator that conditions either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded (e.g., educational levels may well be correlated with race and average compensation in actual communities).
These pitfalls can be avoided with a perturbation-based analysis. In this setting, we propose to apply $\mathrm{CDI}_{\mathrm{Gini}}$ with and without conditioning on the additional covariates $X$. The results are also shown in Table 3. As desired one gets significant positive effects with this method as it is adapted to the simplex. Interestingly, for $\mathrm{CDI}_{\mathrm{Gini}}$ that adjusts for $X$ the effect disappears for the grouping based on education. This is expected because the association between diversity and compensation was constructed based on education which is also part of $X$.
Method Grouping on education Grouping on compensation
Estimate $95\%$ CI Estimate $95\%$ CI
naive_diversity ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.082}}$ $(-0.144,-0.019)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.120}}$ $(-0.194,-0.046)$
[…]
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
Table 3: Table of estimates in the ‘Adult’ data experiment. Bold numbers indicate significance at a $5\%$ level, red indicates a negative effect, while green indicates a positive effect of diversity. The naive_diversity method indicates that either there is a significant negative or a highly uncertain effect of diversity on compensation despite the fact that the data is constructed with a positive effect of diversity. The CDI correctly identifies this effect.
4.3 Variable influence measures when predicting BMI from gut microbiome
In this section, we analyze the ‘American Gut’ (McDonald et al., 2018) dataset which contains a collection of microbiome measurements and metadata from over 10,000 human participants. We use the pre-processed version used in Bien et al. (2021) (available at https://github.com/jacobbien/trac-reproducible/tree/main/AmericanGut). Our focus is on the relationship between body mass index (BMI) and gut microbiome composition at the species-level and our primary goal is to determine which species are important for predicting BMI. We consider the subset of individuals from the United States with BMI between 15 and 40, height between 145 and 220 cm and age above 16 years. To reduce the dimensionality of the microbiome composition (i.e., the number of species), we perform an initial variable screening and only include species that appear in at least $10\%$ of the observations and where the mean abundance is greater than $10^{-4}$ (this is after scaling the absolute counts by the total number of reads for each observation). The resulting dataset consists of 4,581 observations of $\mathrm{BMI}$ measurements $Y\in\mathbb{R}$ and the relative abundances of $561$ microbial species $Z\in\Delta^{560}$.
We begin by evaluating, how sensitive the proposed perturbation effect estimates are with respect to the employed nuisance function estimators. To this end, we consider three different regression methods: random forest (rf), neural network (mlp) and kernelized support vector regression using an RBF kernel (svr). Each of the regressions is combined with a grid-search cross-validation to automatically tune the hyperparameters of each method. For each of the three regression methods and all the species we then compute $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ using the estimator based on a partially linear model and the $\mathrm{CKE}$ using the nonparametric estimator. As baselines, we additionally compute a nonparametric $R^{2}$ (Williamson et al., 2021) and permutation importance (Breiman, 2001). Figure 6 shows the Spearman correlation between the estimates from the different regression methods.
Figure 6: Spearman correlations of variable importance measures computed on each feature using different regressions. On the left, these are computed on all $561$ features and on the right they are computed on the subset of features where at least one of the two methods deems the effect significant at a $5\%$ level. The $R^{2}$ and permutation methods do not come with significance tests and are therefore omitted from the right plot.
While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not. The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator. For the permutation measure the inconsistency comes the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate. Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021). Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model. One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option. To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data. More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\mid Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint. For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression. As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species. In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
We now analyze different types of average perturbation effects and compare them with the commonly used log-contrast based analysis. We fit both a standard log-contrast regression and an $\ell^{1}$-penalized log-contrast regression and then use the coefficients as feature influences. Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex. We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures). To investigate how the log-contrast approach compares to the perturbation-based approach, we plot the most significant hits using $\mathrm{CFI}_{\mathrm{mult}}$ (with random forests as regression method) and a log-contrast regression model in the top row of Figure 7. These are comparable quantities since $\mathrm{CFI}_{\mathrm{mult}}^{j}$ is equal to the $j$th coefficient in a log-contrast model when it is correctly specified and no zeros are present. We omit plotting the $\ell^{1}$-penalized log-contrast estimates as these are broadly similar to the standard log-contrast estimates in this example. We also plot a marginal effect of $L=\log(Z^{j}/(1-Z^{j}))$ which corresponds to predicting $Y$ using OLS on $L$ and correcting for the presence of zeros as we mention in Remark 1. The results with corresponding $95\%$-confidence intervals are given in the top row of Figure 7; top $10$ most significant species sorted according to significance of $\mathrm{CFI}_{\mathrm{mult}}$ (left) and according to log-contrast coefficients (right).
Figure 7: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (top left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ (top right). The labels are colored by the size of the $p$-value. The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model. The bottom left plot shows the $p$-values from different methods for the three most significant species from each method. We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros. The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the top right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments. We see that the relative influences of the different species is highly contingent on the pseudocount.
While the effects overlap for some of the features, there are also significant differences. To further illustrate these differences, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the bottom left of Figure 7. From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species. This is despite the fact that the log-contrast model is supposed to model the continuous effect of each species. The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species. To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the bottom right of Figure 7. This further highlights the dependence of the log-contrast coefficients on the used pseudocount. The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach. In Section S4.4 of the supplementary material we provide an additional plot like the top row of Figure 7 which illustrates the significant hits of $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
5 Theoretical analysis of estimators
In this section we prove asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (17). These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere (e.g., Klyne and Shah, 2023; Chernozhukov et al., 2018). We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.
We start from a given perturbation $\psi$ and a corresponding reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ which reparametrizes $Z$ into $(L,W):=(\phi^{L}(Z),\phi^{W}(Z))$ and assume that $(L,W)$ satisfy (2) if $\psi$ is a binary perturbation and that $\phi$ is derivative-isolating if $\psi$ is a directional perturbation. Our goal is now to use the estimation procedures discussed in Section 3 to estimate the average perturbation effects (i.e., $\tau_{\psi}$ or $\lambda_{\psi}$) consistently and construct asymptotically valid confidence intervals for the estimates. To construct asymptotically valid confidence intervals, we estimate both the target parameter and its variance and then show that the resulting standardized estimators are asymptotically normal with a uniform convergence rate over the class of possible distributions of the data. Distributionally uniform results are desirable here as they ensure uniform coverage guarantees of the confidence intervals. The following result shows uniform asymptotic normality of each of the three estimation procedures.
Theorem 1 (Uniform asymptotic normality).
Let $\mathcal{P}$ denote a class of distributions of $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ and assume we are in one of the following settings.
  1. (a)
    Nonparametric directional effect estimation: Assumption 1 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\tau},\widehat{\sigma}^{2}_{\tau})$ from Algorithm 1.
  2. (b)
    Nonparametric binary effect estimation: Assumption 2 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\lambda},\widehat{\sigma}^{2}_{\lambda})$ from Algorithm 2.
  3. (c)
    Effect estimation using a partially linear model: Assumption 3 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\theta},\widehat{\sigma}^{2}_{\theta})$ from Algorithm 3.
Then, letting $\Phi$ denote the cumulative distribution function of the standard normal distribution, it holds that
$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widehat{\sigma}_{\beta}}(\widehat{\beta}-\beta_{P})\leq x\right)-\Phi\left(x\right)\right|=0.$
The result is proven in Section S5 of the supplementary material. In the following three sections, we go over each of the settings (a), (b) and (c) individually and provide the required assumptions and detailed algorithms for the target parameter and variance. For the remainder of this section, we omit $\psi$-subscripts for simplicity and use the following notation to help us express the uniformly asymptotic results rigorously. Let $(X_{P,n})_{n\in\mathbb{N},P\in\mathcal{P}}$ denote a family of sequences of random variables indexed by $\mathcal{P}$. For a nonnegative sequence $(a_{n})_{n\in\mathbb{N}}$ we write $X_{P,n}=o_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ it holds that
$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}\epsilon)=0,$
that is, $X_{P,n}/a_{n}$ converges to 00 in probability uniformly over $\mathcal{P}$. We also write $X_{P,n}=O_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ there exist $M_{\epsilon}>0$ and $N_{\epsilon}\in\mathbb{N}$ such that
$\sup_{n\geq N_{\epsilon}}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}M)<\epsilon,$
that is, $X_{P,n}/a_{n}$ is bounded in probability for large $n$. When we write conditional expectations given a fitted regression function $\widehat{f}$ we mean a conditional expectation given the random variables used to fit $\widehat{f}$ and any additional randomness involved in the fitting.
5.1 Nonparametric estimation for directional effects
We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem 1 (a). A high-level overview of this estimation was given in Section 3.2.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S3.4 in the supplementary material). We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm 1 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 1 Nonparametric estimation of $\tau$
1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, differentiable regression method $\widehat{f}$, score estimation method $\widehat{\rho}$ and nuisance function estimate $\widehat{\eta}$.
2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
3:For $k\in[K]$, split $I_{k}$ into $I_{k,0}$ and $I_{k,1}$ of (roughly) equal size and set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
6:     Fit nuisance function estimates $\widehat{\eta}_{k}$ using $I_{-k}$.
7:     for $r=0,1$ do
8:         Fit $\widehat{\rho}_{k,r}$ using $I_{k,(1-r)}$ and the nuisance function estimate $\widehat{\eta}_{k}$.
9:         Set $\widehat{\tau}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}$.
10:         Set $\widehat{\nu}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}\right)^{2}$.
11:return $\widehat{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\tau}_{k,r}$ and $\widehat{\sigma}^{2}_{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\nu}_{k,r}-\widehat{\tau}^{2}$.
To obtain the result in Theorem 1 (a), we require that the estimation methods $\widehat{f}$ and $\widehat{\rho}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 1.
Assume that for all $P\in\mathcal{P}$ and $w\in\mathcal{W}$, the conditional distribution of $L$ given $W=w$ has a density with respect to Lebesgue measure and denote it by $p_{P}(\cdot\,|\,w)$. For all $P\in\mathcal{P}$ define $f_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and assume that for almost all $(\ell,w)\in\phi(\mathcal{Z})$ both $p_{P}(\cdot\,|\,w)$ and $f_{P}(\cdot,w)$ are differentiable at $\ell$. Define $\tau_{P}:=\mathbb{E}_{P}[\partial_{\ell}f_{P}(\ell,W)\bigm{|}_{\ell=L}]$ and $\rho_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $\rho_{P}(\ell,w):=\partial_{\ell}\log(p_{P}(\ell\,|\,w))$.
Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{3n},L_{3n},W_{3n})$ denote $3n$ i.i.d. copies of $(Y,L,W)$ and define the training datasets $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$ and :=$\mathcal{B}:=\{(Y_{2n+1},L_{2n+1},W_{2n+1}),\dots,(Y_{3n},L_{3n},W_{3n})\}$. Let $\widehat{f}$ denote a differentiable estimator of $f_{P}$ based on $\mathcal{A}$, let $\widehat{\eta}$ denote an estimator of a nuisance function relevant for the estimation of $\rho_{P}$ based on $\mathcal{A}$ and let $\widehat{\rho}$ denote an estimator of $\rho_{P}$ based on $\mathcal{B}$ and using the nuisance function estimate $\widehat{\eta}$. Finally, assume there exist constants $\delta,c,C>0$ such that
  1. (a)
    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}[|Y|^{4+\delta}],\mathbb{E}_{P}[|\rho_{P}(L,W)|^{4+\delta}],\mathbb{E}_{P}[|\partial_{\ell}f_{P}(L,W)|^{2+\delta}],\mathrm{var}_{P}(Y\,|\,L,W)\}\leq C$
  2. (b)
    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,L,W),\mathbb{E}_{P}[\rho_{P}(L,W)^{2}]\}\geq c$
  3. (c)
    $\mathbb{E}_{P}\left[\left(\widehat{f}(L_{1},W_{1})-f_{P}(L_{1},W_{1})\right)^{4}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
  4. (d)
    $\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L_{1},W_{1})-\partial_{\ell}f_{P}(L_{1},W_{1})\right)^{2}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
  5. (e)
    $\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right)^{2}=o_{\mathcal{P}}(1)$
  6. (f)
    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
5.2 Nonparametric estimation for binary effects
We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem 1 (b). A high-level overview of this estimation was provided in Section 3.2.1. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$. We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm 2 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 2 Nonparametric estimation of $\lambda$
1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression method $\widehat{f}$ and propensity score estimator $\widehat{\pi}$.
2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
6:     Regress $L$ on $W$ using $I_{-k}$ yielding propensity score function $\widehat{\pi}_{k}$.
7:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}$.
8:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}\right)^{2}$.
9:Set $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}1-L_{i}$.
10:Set $\widehat{\kappa}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}$.
11:return $\widehat{\lambda}:=\widehat{p}^{-1}\widehat{\kappa}$ and $\widehat{\sigma}^{2}_{\lambda}:=\widehat{p}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}-\widehat{\kappa}^{2}\right)-\widehat{p}^{-3}\widehat{\kappa}^{2}(1-\widehat{p})$.
To obtain the result in Theorem 1 (b), we require that the estimation methods $\widehat{f}$ and $\widehat{\pi}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 2.
For all $P\in\mathcal{P}$ define $f_{P}:\{0,1\}\times\mathcal{W}\to\mathbb{R}$ and $\pi_{P}:\mathcal{W}\to[0,1]$ for all $(\ell,w)\in\{0,1\}\times\mathcal{W}$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and $\pi_{P}(w):=\mathbb{P}_{P}(L=1\,|\,W=w)$, respectively. Define $p_{P}:=\mathbb{P}_{P}(L=0)$ and $\lambda_{P}:=p_{P}^{-1}\left(\mathbb{E}_{P}[\mathbb{E}_{P}[Y\,|\,L=1,W]]-\mathbb{E}_{P}[Y]\right)$.
Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d. copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$. Let $\widehat{f}$ denote an estimator of $f_{P}$ based on $\mathcal{A}$ and let $\widehat{\pi}$ denote an estimator of $\pi_{P}$ also based on $\mathcal{A}$. Finally, assume there exist constants $\delta,c,C>0$ such that
  1. (a)
    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|Y|^{2+\delta}\right],\mathrm{var}_{P}(Y\,|\,W,L)\}\leq C$
  2. (b)
    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,W,L),p_{P},1-\pi_{P}(W),\pi_{P}(W),\widehat{\pi}(W)\}\geq c$
  3. (c)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{\pi}(W_{i})-\pi_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
  4. (d)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}(1,W_{i})-f_{P}(1,W_{i}))^{2}=o_{\mathcal{P}}(1)$
  5. (e)
    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}(W_{i})-\pi_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
5.3 Partially linear model for binary/directional effects
Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem 1 (c). A high-level overview of this estimation was provided in Section 3.2.3. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively. We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm 3 using a user-specified number of splits $K$ in the cross-validation.
Algorithm 3 Partially linear estimation of $\theta$ (where $\theta=\lambda$ or $\theta=\tau$)
1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression methods $\widehat{g}$ and $\widehat{m}$.
2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
4:for $k=1,\dots,K$ do
5:     Regress $Y$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{g}_{k}$.
6:     Regress $L$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{m}_{k}$.
7:     Set $\widehat{J}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}$.
8:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}$.
9:Set $\widehat{J}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{J}_{k}$
10:Set $\widehat{\theta}:=\widehat{J}^{-1}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}\right)$.
11:for $k=1,\dots,K$ do
12:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}-\widehat{\theta}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}\right)^{2}$
13:return $\widehat{\theta}$ and $\widehat{\sigma}^{2}_{\theta}:=\widehat{J}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}\right)$.
To obtain the result in Theorem 1 (c), we require that the estimation methods $\widehat{m}$ and $\widehat{g}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
Assumption 3.
Assume that for all $P\in\mathcal{P}$ there exist $\theta_{P}\in\mathbb{R}$ and $h_{P}:\mathcal{W}\to\mathbb{R}$ such that
$\mathbb{E}_{P}[Y\,|\,L,W]=\theta_{P}L+h_{P}(W).$
For all $P\in\mathcal{P}$ define $m_{P}:\mathcal{W}\to\mathbb{R}$ and $g_{P}:\mathcal{W}\to\mathbb{R}$ for all $w\in\mathcal{W}$ by $m(w):=\mathbb{E}_{P}[L\,|\,W=w]$ and $g_{P}(w):=\mathbb{E}_{P}[Y\,|\,W=w]=m_{P}(w)+\theta_{P}m_{P}(w)$, respectively. Define $\varepsilon:=Y-\mathbb{E}_{P}[Y\,|\,L,W]$ and $\xi:=Y-\mathbb{E}_{P}[L\,|\,W]$.
Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d. copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$. Let $\widehat{m}$ denote an estimator of $m_{P}$ based on $\mathcal{A}$ and let $\widehat{g}$ denote an estimator of $g_{P}$ also based on $\mathcal{A}$. Suppose that there exist constants $\delta,c,C>0$ such that
  1. (a)
    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|\varepsilon\xi|^{2+\delta}\right]\mathbb{E}_{P}\left[|\xi|^{4}\right],\mathrm{var}_{P}(Y\,|\,W),\mathrm{var}_{P}(L\,|\,W)\}\leq C$
  2. (b)
    for all $P\in\mathcal{P}:\quad$ $\min\{\mathbb{E}_{P}\left[\xi^{2}\right],\mathbb{E}_{P}\left[\varepsilon^{2}\xi^{2}\right]\}\geq c$
  3. (c)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{2}=o_{\mathcal{P}}(n^{-1/2})$
  4. (d)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{4}=o_{\mathcal{P}}(1)$
  5. (e)
    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{g}(W_{i})-g_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
  6. (f)
    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{m}(W_{i})-m_{P}(W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{g}(W_{i})-g_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
6 Discussion
We introduced perturbations on the simplex as a formalization of hypothetical changes to a compositional vector and used them to construct interpretable target parameters, which we call average perturbation effects. In our view, perturbations provide a natural way for practitioners to formalize research questions about non-standard changes to a system. We illustrated this, for example, by using perturbations to formalize the otherwise vague concept of an effect of diversity as well as to construct nonparametric feature influence measures for compositional coordinates. Unfortunately, many applied researchers working with compositional data, generally due to a lack of alternatives, still rely on marginal effects by analyzing dependence between a summary statistic and a response (e.g., by regressing a response on a diversity measure). As we demonstrated, such approaches can be misleading since they do not account for the confounding induced by the remaining variation in compositional vector. We hope that this work can draw attention to such pitfalls and improve the analysis of such data by providing a set of easily accessible alternative methods.
We additionally proposed efficient estimators for the average perturbation effects based on semiparametric theory and provided theoretical guarantees that enable us to perform valid inference. Most statistical methods for compositional data analysis to date are based on the parametric log-ratio approach introduced in the seminal paper Aitchison (1982). Our nonparametric approach goes beyond these established methods and allows practitioners to apply theoretically justified procedures to modern compositional data applications. Importantly, our target parameters and estimators circumvent the zero problem central to any log-ratio based analysis and naturally apply to high-dimensional data. This makes our proposed estimators applicable to a range of modern applications that previously had to rely on ad hoc methods or simply ignore the compositional constraints. An additional advantage of a perturbation-based analysis is that it easily facilitates a causal rather than observational interpretation, allowing practitioners to make standard causal inference assumptions to derive rigorous causal conclusions.
In this work we have focused on data constraints induced by the simplex. There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data. The idea of defining target parameters via perturbations easily extends to those settings. Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference(e.g., how to define causal variables). Nevertheless, for both extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory. We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
Acknowledgements
We thank Alexander Mangulad Christgau and Lucas Kook for helpful discussions. ARL and NP were supported by a research grant (0069071) from Novo Nordisk Fonden.
Appendix S1 Differentiability on the simplex
To define what we mean by differentiability on the simplex we use the theory of smooth manifolds [see Lee, 2012, for a comprehensive treatment of this theory]. The simplex $\Delta^{d-1}$ can be viewed as a smooth $(d-1)$-dimensional manifold with corners [Lee, 2012, p. 415], that is, the simplex locally behaves like a Euclidean half-space $\mathcal{H}^{d-1}:=\{x\in\mathbb{R}^{d-1}\mid\forall j\in[d-1]:x^{j}\geq 0\}$.
To apply the tools of differential geometry, we need to equip the simplex with a smooth atlas, that is, a set of mappings from subsets of the simplex to subsets of the half-space satisfying certain smoothness conditions. In particular, the mappings need to be homeomorphisms from subsets of $\Delta^{d-1}$ to relatively open subsets of $\mathcal{H}^{d-1}$ with the additional property that if $\varphi_{1}:\mathcal{Z}_{1}\to\mathcal{U}_{1}$ and $\varphi_{2}:\mathcal{Z}_{2}\to\mathcal{U}_{2}$ are both elements of the atlas, the composition $\varphi_{2}\circ\varphi_{1}^{-1}:\varphi_{1}(\mathcal{Z}_{1}\cap\mathcal{Z}_{2})\to\varphi_{2}(\mathcal{Z}_{1}\cap\mathcal{Z}_{2})$ is a diffeomorphism. A mapping is differentiable at a boundary point if there exists an extension of the mapping defined in a neighborhood of the boundary point that is differentiable at this point.
Definition S1 (the simplex as a smooth manifold).
Define $U:=\{x\in\mathcal{H}^{d-1}\mid\sum_{j=1}^{d-1}x^{j}<1\}$, which is relatively open in $\mathcal{H}^{d-1}$. For each $j\in[d]$, define $\mathcal{E}_{j}:=\{e_{1},\dots,e_{j-1},e_{j+1},\dots,e_{d}\}$, $\mathcal{Z}_{j}:=\Delta^{d-1}\setminus\mathcal{E}_{j}$ and $\varphi_{j}:\mathcal{Z}_{j}\to U$ by $\varphi_{j}(z):=(z^{1},\dots,z^{j-1},z^{j+1},\dots,z^{d})$. We call the set $\mathcal{A}:=\{(\mathcal{Z}_{j},\varphi_{j})\mid j\in[d]\}$ the canonical smooth atlas on $\Delta^{d-1}$ and each element of $\mathcal{A}$ is a canonical chart on $\Delta^{d-1}$.
For a canonical chart $(\mathcal{Z}_{j},\varphi_{j})$ the mapping $\varphi_{j}$ is indeed a homomorphism since both $\varphi_{j}$ and $\varphi_{j}^{-1}$ are continuous.
Definition S2 (differentiability on the simplex).
Let $\mathcal{Z}\subseteq\Delta^{d-1}$ and $f:\mathcal{Z}\to\mathbb{R}$. We say that $f$ is differentiable at $z\in\mathcal{Z}$ if there exists a canonical chart $(\mathcal{Z}_{j},\varphi_{j})$ with $z\in\mathcal{Z}_{j}$, an open set $O\subseteq\mathbb{R}^{d-1}$ with $\varphi_{j}(z)\in O$ and a function $g:O\to\mathbb{R}$ such that
  1. (a)
    $g$ is differentiable at $\varphi_{j}(z)$,
  2. (b)
    for all $u\in U\cap O$, it holds that $g(u)=f(\varphi_{j}^{-1}(u))$.
If $\varphi_{j}(z)$ is not a boundary point of $U$ in $\mathbb{R}^{d-1}$, then we can take $g=f\circ\varphi_{j}^{-1}$. It is easy to check that if $f:\Delta^{d-1}\to\mathbb{R}$ is differentiable at $z$ then for any $j\in[d]$ with $z\in\mathcal{Z}_{j}$ the function $f\circ\varphi_{j}^{-1}$ must be differentiable at $\varphi_{j}(z)$. With these tools from differential geometry in place, we will now prove Proposition 1.
Appendix S2 Proofs of results from Section 2
S2.1 Proof of Proposition 1
Proof.
It suffices to show that for all $z\in\mathcal{Z}$
$\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}=\partial_{\ell}g(\ell,\phi^{W}(z))\bigm{|}_{\ell=\phi^{L}(z)},$ (S1)
or equivalently for all $w\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$
$\partial_{\gamma}f(\psi(\phi^{-1}(\ell,w),\gamma))\bigm{|}_{\gamma=0}=\partial_{\ell^{\prime}}g(\ell^{\prime},w)\bigm{|}_{\ell^{\prime}=\ell}.$ (S2)
Fix $z\in\mathcal{Z}$, set $(\ell,w):=\phi(z)$ and define $\psi_{z}:\gamma\mapsto\psi(z,\gamma)$. Choose a canonical chart from the canonical smooth atlas $(\mathcal{Z}_{j},\varphi_{j})\in\mathcal{A}$ with $z\in\mathcal{Z}_{j}$ and note that
$f\circ\psi_{z}=\underbrace{f\circ\varphi_{j}^{-1}}_{\widetilde{f}}\circ\underbrace{\varphi_{j}\circ\psi_{z}}_{\widetilde{\psi}_{z}},$
where $\widetilde{f}:U\to\mathbb{R}$ and $\widetilde{\psi}_{z}:\mathcal{I}_{\psi}(z)\to U$ with $U$ defined in Definition S1. We can now apply the chain rule to compute the left-hand side of (S1):
$\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}=\nabla\widetilde{f}(z)^{\top}\partial_{\gamma}\widetilde{\psi}_{z}(\gamma)\bigm{|}_{\gamma=0}.$
Similarly,
$g=f\circ\phi^{-1}=\underbrace{f\circ\varphi_{j}^{-1}}_{\widetilde{f}}\circ\underbrace{\varphi_{j}\circ\phi^{-1}}_{\widetilde{\phi}^{-1}},$
where $\widetilde{\phi}^{-1}:\mathrm{Im}(\phi)\to U$. We can again apply the chain-rule to compute the right-hand side of (S2)
$\partial_{\ell^{\prime}}g(\ell^{\prime},w)\bigm{|}_{\ell^{\prime}=\ell}=\nabla\widetilde{f}(\widetilde{\phi}^{-1}(\ell,w))^{\top}\partial_{\ell^{\prime}}\widetilde{\phi}^{-1}(\ell^{\prime},w)\bigm{|}_{\ell^{\prime}=\ell}.$
Combining the above, we conclude that (S2) is satisfied if
$\partial_{\gamma}\widetilde{\psi}_{z}(\gamma)\bigm{|}_{\gamma=0}=\partial_{\ell^{\prime}}\widetilde{\phi}^{-1}(\ell^{\prime},w)\bigm{|}_{\ell^{\prime}=\ell}.$
This is true by Lee [2012, Proposition 3.24] under the assumption that $\phi$ is derivative-isolating, hence proving the result. We can therefore write
$\tau_{\psi}=\mathbb{E}[\partial_{\gamma}f(\psi(Z,\gamma))\bigm{|}_{\gamma=0}]=\mathbb{E}[\partial_{\ell}g(\ell,W)\bigm{|}_{\ell=L}]$
as desired.
S2.2 Proof of Proposition 2
Proof.
(a) For each $w\in\mathcal{W}$, the speed $s_{w}$ is a positive function, so by (7) $t_{w}$ must be strictly decreasing and hence invertible on $\mathcal{L}_{w}:=t_{w}(\mathcal{I}_{w})$. This lets us compute the inverse of $\phi$ for each $w=(w_{E},w_{v})\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$ as
$\phi^{-1}(\ell,w)=-t_{w}^{-1}(\ell)w_{v}+w_{E}$ (S3)
which shows that $\phi$ is indeed a bijection on the image of $\phi$. For all $w\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$ we have by (7) and the inverse function rule for derivatives that
$s(\phi^{-1}(\ell,w))=s_{w}(t_{w}^{-1}(\ell))=-\frac{1}{(\partial_{\delta}t_{w})(t_{w}^{-1}(\ell))}=-\partial_{\ell}t_{w}^{-1}(\ell).$
We conclude that (3) is satisfied and hence $\phi$ is derivative-isolating for $\psi$ since
$\partial_{\ell}\phi^{-1}(\ell,(w_{E},w_{v}))=-\partial_{\ell}t_{w}^{-1}(\ell)w_{v}.$
(b) By the same argument as in (a), we have that $\phi$ is invertible with inverse as computed in (S3). This reparametrization is derivative-isolating for perturbations with direction $v$ given by (6) and speed given by (reversing the computation in (a))
$s:z\mapsto-\partial_{\ell}t_{(E(z),v(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v(z))}(\|E(z)-z\|_{1}))^{-1}.$
Appendix S3 Estimation technicalities and algorithms
Section S3.1 provides an example that illustrates how the nuisance functions that need to be estimated for the nonparametric estimators in Section 3 can be more unstable and harder to estimate than those required for the estimators based on a partially linear model. Section S3.2 describes how to modify the estimation of $\tau_{\psi}$ using a derivative-isolating reparametrization in the presence of observations with $\omega_{\psi}(Z)=0$. Section S3.3 describes our procedure for the estimation of the derivative of a generic regression function. Section S3.4 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
S3.1 Robustness of nuisance function estimation for different estimators
A key motivation for applying the estimators based on a partially linear model on $Y$ is to avoid estimating nuisance functions that result in unstable estimators. We illustrate this point based on three examples below.
S3.1.1 Estimation of nuisance functions with continuous $L$
To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section 3.2.2) compared to $m$ and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
$L:=W^{1}+(1+\mathbbm{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
The out-of-sample predictions evaluated on a new dataset with $n=1000$ observations can be seen together with the true function in Figure S1.
Figure S1: Estimation of nuisance functions on the simulated data in Section S3.1.1. The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size. We see that the $m$ and $g$ estimates are reasonable estimates of the true functions, that roughly capture the behavior of the true function in an unbiased (conditioned on $W^{1}$) way. In contrast, the derivative and score estimates, are much worse at capturing the more complex behavior (with clear biases conditioned on $W^{1}$).
S3.1.2 Estimation of nuisance functions with binary $L$
To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section 3.2.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
$Y:=2LW^{1}+\varepsilon.$
The out-of-sample predictions evaluated on new dataset with $n=1000$ observations can be seen together with the true function in Figure S2.
Figure S2: Estimation of nuisance functions on the simulated data in Section S3.1.2. The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size. We see that the $m$, $g$ and $f$ estimates are reasonable estimates of the true functions whereas the inverse propensity (even when cutting off extreme values at $100$) is poorly estimated for $W^{1}$ close to zero.
S3.1.3 Partially linear model estimate when partial linearity is violated
The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified. The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting). We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section 4.1 which we do in Figure S3. Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
Proposition S1.
Let $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ denote random variables where $\max\{\mathbb{E}[L^{2}],\mathbb{E}[Y^{2}]\}<\infty$. Then,
$\theta:=\frac{\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]}{\mathbb{E}[\mathrm{var}(L\,|\,W)]}\quad\text{and}\quad h:w\mapsto\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$
are solutions to the minimization problem
$\min_{\theta\in\mathbb{R},h:\mathcal{W}\to\mathbb{R}}\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W)\right)^{2}\right].$ (S4)
Proof.
First, we note that instead of the optimization in (S4) we can also solve
$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[Y\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
since using $h:w\mapsto\widetilde{h}(w)+\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$ leads to a solution of (S4). Next, we expand the square in the objective function of the minimization problem (S5) and obtain that
$c(\theta,\widetilde{h})$ $=\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]\right)^{2}\right]+\theta^{2}\mathbb{E}\left[\mathrm{var}(L\,|\,W)\right]+\mathbb{E}[\widetilde{h}(W)^{2}]$
$-2\theta\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])].$
It is now clear that the optimal $\widetilde{h}$ is the zero function and by noting that
$\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])]=\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)],$
we obtain the desired result by minimizing the quadratic in $\theta$. ∎
Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section 4.1 with $d=15$. We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
S3.2 Computing directional perturbation effects in the presence of zero speeds
In Proposition 1 we only considered directional perturbations where $\omega_{\psi}(Z)\neq 0$ almost surely. We would, however, like to be able to use the derivative-isolating reparametrizations for estimation of the directional average perturbation effect $\tau_{\psi}$ even when this is not the case. To that end, we propose a modified estimation procedure that we outline in Algorithm S1. We propose to estimate $\tau_{\psi}$ on the subset of observations where $\omega_{\psi}(Z)\neq 0$ and then correct the resulting point and variance estimate appropriately.
Algorithm S1 Estimation of $\tau_{\psi}$ in the presence of observations with $\omega_{\psi}(Z)=0$.
1:Input: Data $(Y_{i},Z_{i})_{i\in[n]}$, directional perturbation $\psi$, derivative-isolating reparametrization $\phi$.
2:For $i\in[n]$, define $D_{i}:=\mathbbm{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
3:Define :=$\mathcal{I}:=\{i\in[n]\mid D_{i}=1\}$.
4:For $i\in\mathcal{I}$, define $(L_{i},W_{i})=\phi(Z_{i})$.
5:Compute estimate $\widehat{\tau}$ and asymptotic variance estimate $\widehat{\sigma}^{2}$ from either Algorithm 1 or 3.
6:Compute $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}$.
7:return estimate $\widetilde{\tau}:=\widehat{\tau}\widehat{p}$ and asymptotic variance estimate $\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$.
Theoretical guarantees for the proposed modification follow from repeating the arguments of the proof of Theorem 1 (a) and (c) to obtain the population quantities that determine the asymptotic behavior of $\widehat{\tau}$ and $\widehat{\sigma}^{2}$ and then applying Proposition S2 below.
Proposition S2.
Let $(D,S)\in\{0,1\}\times\mathbb{R}$ be a pair of random variables with distributions determined by $\mathcal{P}$ and let $(S_{n},D_{n})_{n\in\mathbb{N}}$ be an i.i.d. sequence of copies of $(S,D)$. Suppose that there exist $\delta,C,c>0$ such that $\mathrm{var}_{P}(DS)\geq c$ and $\mathbb{E}_{P}[D|S|^{2+\delta}]\leq C$. Suppose further that
$\widehat{\tau}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
and
$\widehat{\sigma}^{2}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
Then,
$\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i},\quad\widetilde{\tau}:=\widehat{\tau}\widehat{p},\quad\text{and}\quad\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$
satisfy
$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})\leq x\right)-\Phi(x)\right|=0.$
Proof.
We first note that
$\widetilde{\tau}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
since $\widehat{p}\leq 1$. Similarly, we have that
$\widetilde{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
Therefore, by several applications of Lemma S2, we have that
$\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(D_{i}S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1)$
and the result thus follows by Lemma S4 (a) by the bounds on the moments and variance of $(D,S)$. ∎
S3.3 Estimating differentiable regression functions
The following is a description of the smoothing method we apply in the nonparametric estimator of the average directional perturbation effect. The procedure uses random forest weights together with a local polynomial estimate to smooth a given regression function to produce a differentiable estimator and to compute its derivatives. Local polynomials are a prominent and natural method used to estimate derivatives [e.g., Masry and Fan, 1997, De Brabanter et al., 2013]. We are interested in estimating the partial derivative of a function $g:\mathbb{R}\times\mathcal{W}\rightarrow\mathbb{R}$. As in Klyne and Shah [2023] we start from a given arbitrary regression estimate $\widehat{g}$ of $g$ and apply the following local polynomial approach. Estimate, for all $i\in[n]$, polynomial coefficients $\beta\in\mathbb{R}^{3}$ such that the polynomial $p:\mathbb{R}\rightarrow\mathbb{R}$ defined for all $\ell\in\mathbb{R}$ by
$p(\ell)=\beta^{1}+\beta^{2}(\ell-L_{i})+\beta^{3}(\ell-L_{i})^{2}$
is a good local approximation of $\widehat{g}$ around $L_{i}$. Formally, for a given weight matrix $K\in\mathbb{R}^{n\times n}$, which encodes a closeness measure between sample points, we minimize for all $i\in[n]$ the weighted mean squared loss
$\hat{\beta}(i):=\operatorname{argmin}_{\beta\in\mathbb{R}^{3}}\sum_{k=1}^{n}\left(\widehat{g}(L_{k},W_{k})-\sum_{j=1}^{3}\beta^{j}(L_{k}-L_{i})^{j-1}\right)^{2}K_{i,\ell}.$ (S6)
Then, using the estimated coefficients $\hat{\beta}(i)$, we get a smoothed estimate $\widehat{g}_{s}$ which at each sample point $i\in[n]$ is given by
$\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)\quad\text{and}\quad\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i).$ (S7)
As weights $K$ we could, for example, use kernel weights, i.e., $K_{i,\ell}=k((X_{i}-X_{\ell})/\sigma)$ for a kernel function $k$ and a bandwidth $\sigma>0$. As we are considering a potentially high-dimensional setting, we instead propose to use random forest weights [Meinshausen, 2006, Athey et al., 2019]. The full procedure is detailed in Algorithm S2.
Algorithm S2 Local polynomial based smoothing with random forest weights
1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, fitted values $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$
2:Regress $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$ on $(L_{i},W_{i})_{i\in[n]}$ using a random forest yielding $\widehat{\mu}$.
3:Extract weight matrix $K$ from $\widehat{\mu}$.
4:for $i=1,\ldots,n$ do
5:     Set $\hat{\beta}(i)$ to equal the coefficient of a degree $2$ local polynomial fit in (S6).
6:     Set $\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)$.
7:     Set $\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i)$.
8:return smoothed predictions $(\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$ and derivatives $(\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$
S3.4 Location-scale modelling of the score function
The following is a description of the estimator of the conditional score based on a location-scale model proposed by Klyne and Shah [2023]. We apply this in our experiments to estimate conditional score functions and include it here for completeness. We are interested in the estimation of the conditional score of $L$ given $W$, where $L\in\mathbb{R}$ and $W\in\mathcal{W}$, that is, the derivative of the logarithm of the conditional density of $L$ given $W$. Conditional density estimation is difficult even in low dimensions. However, if we are willing to assume that there exist functions $m:\mathcal{W}\rightarrow\mathbb{R}$, $v:\mathcal{W}\rightarrow[0,\infty)$ and a random variable $\zeta\in\mathbb{R}$ that is independent of $W$ such that
$L=m(W)+v(W)^{1/2}\zeta,$
we can simplify estimation of the score function. The procedure is outlined in Algorithm S3 below. In all of our experiments, we use the univariate score estimation method given by Ng [1994] that is based on ideas from Cox [1985].
Algorithm S3 Location-scale score modelling estimation
1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, nuisance regression methods $\widehat{\eta}=(\widehat{m}$, $\widehat{v})$, univariate score estimator $\widehat{\rho}$.
2:Split indices $\{1,\dots,n\}$ into $2$ disjoint sets $I_{1},I_{2}$ of (roughly) equal size.
3:Regress $L$ on $W$ using $I_{1}$ yielding regression estimate $\widehat{m}$.
4:Regress $(L-\widehat{m}(w))^{2}$ on $W$ using $I_{1}$ yielding variance estimate $\widehat{v}$.
5:For $i\in I_{2}$, set $\widehat{\xi}_{i}:=\frac{L_{i}-\widehat{m}(W_{i})}{\widehat{v}(W_{i})^{1/2}}$.
6:Estimate a univariate score from $(\widehat{\xi}_{i})_{i\in I_{2}}$ yielding $\widehat{\rho}$.
7:return $\widehat{\rho}$.
Appendix S4 Details and additional results from numerical experiments
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.
S4.1 Simulations from Section 1
In this section we describe the data-generating processes and methods used in the numerical illustrations of Example 1 and Example 2. The results of are provided in Table 1.
S4.1.1 Example 1: Effects of absence or presence of microbes
Let $L\sim\mathrm{Bern}(1/2)$, $U^{W},U^{L}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
$Z:=(Z^{1},W^{2}(1-Z^{1}),(1-W^{2})(1-Z^{1}))\in\Delta^{2},$
where $W^{2}:=U^{W}(1-L)+U^{W}LB$ and $Z^{1}:=(1-L)U^{L}$. Furthermore, define $Y$ such that
$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathbbm{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathbbm{1}_{\{Z^{2}=0\}}\right).$
In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
$\mathrm{CKE}^{1}$ is computed using the $\lambda$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
S4.1.2 Example 2: Effects of increased diversity
Let $\Pi$ denote the orthogonal projection onto the orthogonal complement of $\mathrm{span}((1,1,1))$. Moreover, let $U\sim\mathcal{N}(0,\Pi\,\Pi^{\top})$ and $\xi,\varepsilon\sim\mathcal{N}(0,1)$ be independent random variables and define
$Z:=-DW+z_{\mathrm{cen}}\in\Delta^{2}\quad\text{and}\quad Y:=L+4W^{\top}e_{1}+\varepsilon,$
where $W:=\frac{U}{\|U\|_{1}}$, $D:=\mathrm{expit}(W^{\top}e_{1}+\xi)$ and $L:=-\frac{2}{9}D\sum_{j=1}^{3}\sum_{k=1}^{3}|W^{j}-W^{k}|$. Then, by construction $L$ is minus the Gini coefficient of $Z$. Hence, from the partially linear model of $Y$ it follows that the true effect of increasing the negative Gini coefficient is equal to $1$. In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
$\mathrm{CDI}_{\mathrm{Gini}}$ is computed using the $\tau$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
S4.2 Additional semiparametric estimation results
Here we provide additional results from the experiment of Section 4.1. Figure S4 and Figure S5 contain estimated coverage rates of the $95\%$ confidence intervals for the different methods and settings with $d=3$ and $d=75$, respectively. The patterns observed in Figure 4 of Section 4.1 are repeated in these figures.
Figure S4: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=3$. The patterns observed in Figure 4 of Section 4.1 are repeated here.
Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=75$. The patterns observed in Figure 4 of Section 4.1 are repeated here.
S4.3 Description of the construction of groups in Section 4.2
In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.2. On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
$(\widehat{\pi}_{R}(\mathrm{White}),\widehat{\pi}_{R}(\mathrm{Black}),\widehat{\pi}_{R}(\mathrm{Other}))$ $=(0.855,0.096,0.049),$
$(\widehat{\pi}_{E}(0),\widehat{\pi}_{E}(1))$ $=(0.454,0.546),$
$(\widehat{\pi}_{C}(0),\widehat{\pi}_{C}(1))$ $=(0.761,0.239),$
respectively. Now, for the two types of groupings we define target probabilities $p_{i}=(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ for each individual $i$ (see below). Based on these probabilities we randomly draw for each individual $i$ one of the three categories using the probabilities $p_{i}$. Finally, we randomly group individuals with the same categories into communities of approximately $50$ (as many communities of $50$ and the remaining with $49$) and aggregate the predictors and response as described in the main text. The probabilities $(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ are selected, in each of the two settings, as follows.
Grouping by race and education: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{E}^{1}(\texttt{education}_{i}),$
where $\texttt{race}_{i}$ and $\texttt{education}_{i}$ denote the race and education level of the $i$th individual, $p_{R}^{1}$ and $p_{E}^{1}$ are probability vectors satisfying
$p_{R}^{1}(\mathrm{White})\propto\frac{0.8}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{1}(\mathrm{Black})\propto\frac{0.02}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{1}(\mathrm{Other})\propto\frac{0.18}{\widehat{\pi}_{R}(\mathrm{Other})},$ (S8)
and
$p_{E}^{1}(0)\propto\frac{0.95}{\widehat{\pi}_{E}(0)},\quad p_{E}^{1}(1)\propto\frac{0.05}{\widehat{\pi}_{E}(1)}.$
That is, category $1$ is about as diverse ($(0.8,0.02,0.18)$ is about as diverse as $\widehat{\pi}_{R}$) and with lower education ($0.05<\widehat{\pi}_{E}(1)$). The probability $p_{i}^{2}$ of observation $i$ belonging to category $2$ is set to
$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{E}^{2}(\texttt{education}_{i}),$
where similarly $p_{R}^{2}$ and $p_{E}^{2}$ are probability vectors but now satisfying
$p_{R}^{2}(\mathrm{White})\propto\frac{0.65}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{2}(\mathrm{Black})\propto\frac{0.12}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{2}(\mathrm{Other})\propto\frac{0.23}{\widehat{\pi}_{R}(\mathrm{Other})}$ (S9)
and
$p_{E}^{2}(0)\propto\frac{0.3}{\widehat{\pi}_{E}(0)},\quad p_{E}^{2}(1)\propto\frac{0.7}{\widehat{\pi}_{E}(1)}.$
That is, category $2$ is more diverse ($(0.65,0.12,0.23)$ is more diverse than $\widehat{\pi}_{R}$) and with higher education ($0.7>\widehat{\pi}_{E}(1)$). Lastly, the probability $p_{i}^{0}$ for category 00 is set to $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
Grouping by race and compensation: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{C}^{1}(\texttt{compensation}_{i}),$
where $p_{R}^{1}$ is as in (S8) and $p_{C}^{1}$ is a probability vector satisfying
$p_{C}^{1}(0)\propto\frac{0.9}{\widehat{\pi}_{C}(0)},\quad p_{C}^{1}(1)\propto\frac{0.1}{\widehat{\pi}_{C}(1)}.$
That is, category $1$ is about as diverse and with lower compensation ($0.05<\widehat{\pi}_{C}(1)$). The probability $p_{i}^{2}$ for category $2$ is set to
$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{C}^{2}(\texttt{compensation}_{i}),$
where $p_{R}^{2}$ is as in (S9) and $p_{C}^{2}$ is a probability vector satisfying
$p_{C}^{2}(0)\propto\frac{0.7}{\widehat{\pi}_{C}(0)},\quad p_{C}^{2}(1)\propto\frac{0.3}{\widehat{\pi}_{C}(1)}.$
That is, category $2$ is more diverse and with higher compensation ($0.3>\widehat{\pi}_{C}(1)$). Lastly, we set the probability $p_{i}^{0}$ for category 00 is $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
S4.4 Additional plots from gut microbiome analysis in Section 4.3
In this section we provide an additional figure showing our results from the analysis of the gut microbiome data in Section 4.3. In Figure S6, we plot the most significant hits for $\mathrm{CFI}{\mathrm{unit}}$ and $\mathrm{CKE}$ both estimated using random forests together with $\mathrm{CFI}_{\mathrm{mult}}$ estimates. We plot the estimates scaled by the estimated standard deviation of $L$ to put the effects on the same scale (we do not correct the confidence intervals for the estimation of this standard deviation). Furthermore, we flip the sign of $\mathrm{CKE}$ to keep the interpretation similar to $\mathrm{CFI}$, that is, that we are looking at the effect of increasing $Z^{j}$.
Figure S6: Effect estimates scaled by the estimated standard deviation of $L$ for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{unit}}$ (left) and $\mathrm{CKE}$ (right). The labels are colored by the size of the $p$-value. For the $\mathrm{CFI}_{\mathrm{unit}}$ results, the first $9$ species remain below $0.05$ after Bonferroni-correction while this is only true for $2$ for the $\mathrm{CKE}$. There is little overlap between the selected species across the two methods.
Appendix S5 Proof of Theorem 1
S5.1 Proof of (a)
Proof.
We define for all $k\in[K]$ and $r\in\{0,1\}$, $N_{k,r}:=|I_{k,r}|$ and for all $i\in[n]$, $\varepsilon_{i}:=Y_{i}-f(L_{i},W_{i})$. We first argue that
$\widehat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\underbrace{\partial_{\ell}f_{P}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\varepsilon_{i}}_{S_{i}}+o_{\mathcal{P}}(n^{-1/2}).$ (S10)
To that end, we write for all $k\in[K]$ and $r\in\{0,1\}$
$\widehat{\tau}_{k,r}$ $=\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}S_{i}$
$+\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}}_{(\textup{I})}$
$-\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}\varepsilon_{i}}_{(\textup{II})}$
$+\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})\}+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}}_{(\textup{III})}.$
By the Cauchy–Schwarz inequality, Assumption 1 (f) and Lemma S3, we have
$|(\textup{I})|\leq\sqrt{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}=o_{\mathcal{P}}(n^{-1}).$
[…]
$\mathbb{P}_{P}(|(\textup{II})|\geq N_{k,r}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k,r}},\widehat{\rho}_{k,r})\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\mathrm{var}_{P}(Y_{i}\,|\,L_{i},W_{i}).$
Combining the above with Assumption 1 (a) and (e), we have
$\mathbb{P}_{P}(|(\textup{II})|\geq N_{k,r}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k,r}},\widehat{\rho}_{k,r})=o_{\mathcal{P}}(1)$
which implies (II$)=o_{\mathcal{P}}(N_{k}^{-1/2})( I$ as $\epsilon$ was arbitrary.
[…]
$\mathbb{E}_{P}\left[\left|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right|\,|\,\widehat{f}_{k}\right]$
$\leq\mathbb{E}_{P}[|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})|\,|\,\widehat{f}_{k}]+\mathbb{E}_{P}\left[\rho_{P}(L_{i},W_{i})^{2}\right]^{1/2}\mathbb{E}_{P}\left[\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}\,|\,\widehat{f}_{k}\right]^{1/2}.$
By the above, Assumption 1 (c), (d) and (a), Lemma S2 (a) and Jensen’s inequality, we have that
$\mathbb{E}_{P}\left[\left|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right|\,|\,\widehat{f}_{k}\right]$ $=O_{\mathcal{P}}(1),$
$\mathbb{E}_{P}\left[\left|\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right|\,|\,\widehat{f}_{k}\right]$ $=O_{\mathcal{P}}(1).$
This implies that for any $\zeta>0$, we can find $M>0$ and $n\in\mathbb{N}$ so that
$\Omega_{P}$ :=$:=\left\{\mathbb{E}_{P}\left[\left|\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right|\,|\,\widehat{f}_{k}\right]<M\right\}$
$\cap\left\{\mathbb{E}_{P}\left[\left|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right|\,|\,\widehat{f}_{k}\right]<M\right\}$
satisfies $\mathbb{P}_{P}(\Omega_{P})\geq 1-\zeta/2$ for all $P\in\mathcal{P}$. We therefore have for any $\epsilon>0$
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\{|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\}\cap\Omega_{P})+\zeta/2\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)+\zeta/2,$
hence it suffices to prove that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ to conclude that $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$. To that end, we first note that
$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}\right]=\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)\,|\,\widehat{f}_{k}\right].$
We will show that
$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}=h\right]=0$ (S11)
for $\mathbb{P}_{P}$-almost every $h$. If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=0$ then clearly the conditional expectation is 00. If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=1$, then we intend to apply Lemma S5 (integration by parts) with $g=(h-f_{P})$. Assumptions (a)-(c) of Lemma S5 are clear from the above and (d) follows from Lemma S6 and Assumption (a). We conclude that
$\mathbb{E}_{P}\left[\partial_{\ell}(h(L,W)-f_{P}(L,W))+\rho_{P}(L,W)(h(L,W)-f_{P}(L,W))\right]=0,$
and hence (S11) is satisfied. By conditional Chebyshev’s inequality (applicable since we have shown that summands are conditionally mean zero in (S11)), the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and conditional Cauchy–Schwarz inequality, we therefore have
$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\,|\,\widehat{f}_{k})$
$\qquad\leq\frac{1}{\epsilon^{2}N_{k,r}}\mathbb{E}_{P}[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})^{2}\,|\,\widehat{f}_{k}]$
$\qquad=\frac{1}{\epsilon^{2}}\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)^{2}\,|\,\widehat{f}_{k}\right]$
$\qquad\leq\frac{2}{\epsilon^{2}}\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})\right)^{2}\,|\,\widehat{f}_{k}\right]$
$\qquad+\frac{2}{\epsilon^{2}}\mathbb{E}_{P}\left[|\rho_{P}(L_{i},W_{i})|^{4}\right]^{1/2}\mathbb{E}_{P}\left[\left|\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right|^{4}\,|\,\widehat{f}_{k}\right]^{1/2}.$
Assumption 1 (c), (d) and (a) and Lemma S3 thus yield that
$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
which implies that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ so $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ by our earlier argument. The above combined with Lemma S1 proves (S10).
Let $\delta^{\prime}:=\delta/2$. We now prove that for all $i\in[n]$ and $P\in\mathcal{P}$, we have
$\mathbb{E}_{P}\left[|S_{i}|^{2+\delta^{\prime}}\right]\leq 2^{1+\delta^{\prime}}\left(C^{(2+\delta^{\prime})/(2+\delta)}+C2^{2+\delta^{\prime}}\right)=:C^{\prime}.$ (S12)
By using that $|a+b|^{1+\delta}\leq 2^{\delta}(|a|^{2+\delta}+|b|^{2+\delta})$ for $\delta>0$ (sometimes known as the $c_{r}$-inequality), we have
$|S_{i}|^{2+\delta^{\prime}}\leq 2^{1+\delta^{\prime}}\left(\left|\partial_{\ell}f_{P}(L_{i},W_{i})\right|^{2+\delta^{\prime}}+\left|\rho_{P}(L_{i},W_{i})\right|^{2+\delta}\left|\varepsilon_{i}\right|^{2+\delta}\right),$
and
$|\varepsilon_{i}|^{4+\delta}\leq 2^{3+\delta}\left(|Y_{i}|^{4+\delta}+|\mathbb{E}_{P}[Y_{i}\,|\,L_{i},W_{i}]|^{4+\delta}\right).$
Therefore, using Assumption 1 (a), the Cauchy–Schwarz inequality and Jensen’s inequality, we have
$\mathbb{E}_{P}\left[|S_{i}|^{2+\delta^{\prime}}\right]$ $\leq 2^{1+\delta^{\prime}}\left(C^{(2+\delta^{\prime})/(2+\delta)}+\left(\mathbb{E}_{P}\left[\left|\rho_{P}(L_{i},W_{i})\right|^{4+\delta}\right]\mathbb{E}_{P}\left[\left|\varepsilon_{i}\right|^{4+\delta}\right]\right)^{1/2}\right)$
$\leq 2^{1+\delta^{\prime}}\left(C^{(2+\delta^{\prime})/(2+\delta)}+\left(C^{2}(2^{4+\delta})\right)^{1/2}\right)=C^{\prime}$
as desired.
We now show that
$\widehat{\sigma}^{2}_{\tau}=\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$ (S13)
Note that $\frac{1}{n}\sum_{i=1}^{n}S_{i}=O_{\mathcal{P}}(1)$ by (S12) and Markov’s inequality, hence by (S10) and Lemma S3, we have that
$\widehat{\tau}^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
We have thus shown (S13) if we can show that
$\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\nu}_{k,r}=\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}+o_{\mathcal{P}}(1).$
To that end using the same decomposition as for $\widehat{\tau}_{k,r}$, we write for $k\in[K]$, $r\in\{0,1\}$, $i\in I_{k,r}$
$\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k,r}(L_{i},W_{i})\}=\underbrace{\partial_{\ell}f_{P}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\varepsilon_{i}}_{S_{i}}$
[…]
so that
$\widehat{\nu}_{k,r}=\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}S_{i}^{2}+\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}R_{i}^{2}+\frac{2}{N_{k,r}}\sum_{i\in I_{k,r}}S_{i}R_{i}$
Since $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ (which can be seen by the Cauchy–Schwarz inequality), we have
$\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}R_{i}^{2}$ $\leq 3\biggl{(}\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}_{\widetilde{(\textup{I})}}$
$+\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\varepsilon_{i}^{2}}_{\widetilde{(\textup{II})}}$
$+\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\left(\{\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})\}+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)^{2}}_{\widetilde{(\textup{III})}}\biggr{)}.$
By using that $\sum_{i}a_{i}^{2}b_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}b_{i}^{2}\right)$, we have
$\widetilde{(\textup{I})}\leq\left(\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\right)\left(\frac{1}{N_{k}}\sum_{i\in I_{k,r}}\left(\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right)^{2}\right).$
[…]
$\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}R_{i}^{2}=o_{\mathcal{P}}(1).$
By the Cauchy–Schwarz inequality, we have that
$\frac{2}{N_{k,r}}\left|\sum_{i\in I_{k,r}}S_{i}R_{i}\right|\leq 2\left(\frac{1}{N_{k,r}^{2}}\sum_{i\in I_{k,r}}R_{i}^{2}\sum_{i\in I_{k,r}}S_{i}^{2}\right)^{1/2}.$
Since $1/N_{k,r}\sum_{i\in I_{k,r}}S_{i}^{2}=O_{\mathcal{P}}(1)$ by (S12) and Markov’s inequality and $\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}R_{i}^{2}=o_{\mathcal{P}}(1)$ by the above, we conclude by Lemma S3 that
$\frac{2}{N_{k,r}}\sum_{i\in I_{k,r}}S_{i}R_{i}=o_{\mathcal{P}}(1)$
and Lemma S1 concludes the proof of (S13).
Since for all $i\in[n]$,
$\mathbb{E}_{P}[S_{i}]=\mathbb{E}_{P}[\partial_{\ell}f_{P}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\varepsilon_{i}]=\tau_{P},$
and (S12) implies that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\mathrm{CFI}^{j})=O_{\mathcal{P}}(1)$, we have by Lemma S3, (S10) and (S13) that
$\frac{\sqrt{n}}{\widehat{\sigma}_{\tau}}(\widehat{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1).$
The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption 1 (a) and (b), since
$\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)\right)^{2}\right]$
$=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
S5.2 Proof of (b)
Proof.
We define for all $k\in[K]$, $N_{k}:=|I_{k}|$ and for all $i\in[n]$ $\varepsilon_{i}:=Y_{i}-f_{P}(L_{i},W_{i})$. We first argue that
$\widehat{\kappa}=\frac{1}{n}\sum_{i=1}^{n}\underbrace{f_{P}(1,W_{i})-Y_{i}+\frac{\varepsilon_{i}}{\pi_{P}(W_{i})}L_{i}}_{S_{i}}+o_{\mathcal{P}}(n^{-1/2}).$ (S14)
To that end, we write for $k\in[K]$
$\widehat{\kappa}_{k}$ $=\frac{1}{N_{k}}\sum_{i\in I_{k}}S_{i}$
$-\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)\left(\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right)L_{i}}_{(\textup{I})}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)\varepsilon_{i}L_{i}}_{(\textup{II})}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(1,W_{i})-f_{P}(1,W_{i})\right)\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)}_{(\textup{III})}.$
Applying the triangle and Cauchy–Schwarz inequality, followed by Assumption 2 (b) and (e) leads to
$|(\textup{I})|$ $\leq\sqrt{\frac{1}{N_{k}^{2}}\sum_{i\in I_{k}}\frac{\{\pi_{P}(W_{i})-\widehat{\pi}_{k}(W_{i})\}^{2}}{\widehat{\pi}_{k}(W_{i})^{2}\pi_{P}(W_{i})^{2}}\sum_{i\in I_{k}}\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}$
$\leq\frac{1}{\zeta^{2}}\sqrt{\frac{1}{N_{k}^{2}}\sum_{i\in I_{k}}\{\pi_{P}(W_{i})-\widehat{\pi}_{k}(W_{i})\}^{2}\sum_{i\in I_{k}}\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}=o_{\mathcal{P}}(N_{k}^{-1/2}).$
[…]
$\mathbb{P}_{P}(|(\textup{II})|\geq N_{k}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k}},\widehat{\pi}_{k})$ $\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)^{2}\mathrm{var}_{P}(Y_{i}\,|\,W_{i},L_{i})$
$=\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\frac{\{\pi_{P}(W_{i})-\widehat{\pi}_{k}(W_{i})\}^{2}}{\widehat{\pi}_{k}(W_{i})^{2}\pi_{P}(W_{i})^{2}}\mathrm{var}_{P}(Y_{i}\,|\,W_{i},L_{i}).$
Combining Assumption 2 (b), (c) and (a), we conclude that
$\mathbb{P}_{P}(|(\textup{II})|\geq N_{k}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k}},\widehat{\pi}_{k})=o_{\mathcal{P}}(1)$
which implies that (II$)=o_{\mathcal{P}}(N_{k}^{-1/2})( I$ as $\epsilon$ was arbitrary.
[…]
$\mathbb{P}(|(\textup{III})|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{f}_{k})$ $\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\mathbb{E}_{P}\left[\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)^{2}\,\Bigm{|}\,W_{i}\right]\left(\widehat{f}_{k}(1,W_{i})-f_{P}(1,W_{i})\right)^{2}.$
By Assumption 2 (b), we have
$\mathbb{E}_{P}\left[\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)^{2}\,\Bigm{|}\,W_{i}\right]=\frac{1}{\pi_{P}(W_{i})}-1\leq\frac{1-c}{c},$
hence, by Assumption 2 (d),
$\mathbb{P}_{P}(|(\textup{III})|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{f}_{k})=o_{\mathcal{P}}(1)$
which implies that $(\textup{III})=o_{\mathcal{P}}(N_{k}^{-1/2})$ as $\epsilon$ was arbitrary. The above combined with Lemma S1 proves (S14).
We now prove that for $i\in[n]$
$\mathbb{E}_{P}\left[|S_{i}|^{2+\delta}\right]\leq 3^{1+\delta}C\left(1+c^{-(2+\delta)}(1+2^{2+\delta})\right)=:C^{\prime}.$ (S15)
We have by repeated applications of Jensen’s inequality that
$|S_{i}|^{2+\delta}\leq 3^{1+\delta}\left(|f_{P}(1,W_{i})|^{2+\delta}+|Y_{i}|^{2+\delta}+\frac{|\varepsilon_{i}|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\right),$
by Assumption 2 (b),
$|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathbbm{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
and
$|\varepsilon_{i}|^{2+\delta}\leq 2^{1+\delta}\left(|Y_{i}|^{2+\delta}+|\mathbb{E}_{P}[Y_{i}\,|\,W_{i},L_{i}]|^{2+\delta}\right).$
Therefore, using Assumption 2 (a)
$\mathbb{E}_{P}\left[|S_{i}|^{2+\delta}\right]\leq 3^{1+\delta}\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\right]\left(1+c^{-(2+\delta)}(1+2^{1+\delta})\right)\leq C^{\prime}$
as desired.
Define $\kappa_{P}:=\mathbb{E}_{P}[S_{i}]$ and $\sigma^{2}_{\kappa_{P}}:=\mathrm{var}_{P}(S_{i})$. We now show that
$\frac{1}{K}\sum_{j=1}^{k}\widehat{\nu}_{k}-\widehat{\kappa}^{2}=\sigma^{2}_{\kappa_{P}}+o_{\mathcal{P}}(1).$ (S16)
By (S14) and Chebyshev’s inequality, we have that
$\widehat{\kappa}=\kappa_{P}+o_{\mathcal{P}}(1),$
thus by (S15) and Lemma S3, we have
$\widehat{\kappa}^{2}=\kappa_{P}^{2}+o_{\mathcal{P}}(1).$
We have therefore shown (S16) if we can show that
$\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}=\mathbb{E}_{P}[S_{i}^{2}]+o_{\mathcal{P}}(1).$ (S17)
To that end using the same composition as for $\widehat{\kappa}_{k}$, we write for $k\in[K]$, $i\in I_{k}$
$\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}=\underbrace{f_{P}(1,W_{i})-Y_{i}+\frac{\varepsilon_{i}}{\pi_{P}(W_{i})}L_{i}}_{S_{i}}$
[…]
so that
$\widehat{\nu}_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}S_{i}^{2}+\frac{1}{N_{k}}\sum_{i\in I_{k}}R_{i}^{2}+\frac{2}{N_{k}}\sum_{i\in I_{k}}S_{i}R_{i}.$
We have
$\frac{1}{N_{k}}\sum_{i\in I_{k}}R_{i}^{2}$ $\leq 3\biggl{(}\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)^{2}\left(\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right)^{2}}_{\widetilde{(\textup{I})}}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)^{2}\varepsilon_{i}^{2}}_{\widetilde{(\textup{II})}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)^{2}\left(\widehat{f}_{k}(1,W_{i})-f_{P}(1,W_{i})\right)^{2}}_{\widetilde{(\textup{III})}}\biggr{)}.$
By using that $\sum_{i}a_{i}^{2}b_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}b_{i}^{2}\right)$, we have
$\widetilde{(\textup{I})}\leq\left(\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\frac{1}{\widehat{\pi}_{k}(W_{i})}-\frac{1}{\pi_{P}(W_{i})}\right)^{2}\right)\left(\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right)^{2}\right).$
[…]
$\frac{1}{N_{k}}\sum_{i\in I_{k}}R_{i}^{2}=o_{\mathcal{P}}(1).$
By the Cauchy–Schwarz inequality, we have that
$\frac{2}{N_{k}}\left|\sum_{i\in I_{k}}S_{i}R_{i}\right|\leq 2\left(\frac{1}{N_{k}^{2}}\sum_{i\in I_{k}}R_{i}^{2}\sum_{i\in I_{k}}S_{i}^{2}\right)^{1/2}.$
Since $1/N_{k}\sum_{i\in I_{k}}S_{i}^{2}=O_{\mathcal{P}}(1)$ by (S15) and Markov’s inequality and $\frac{1}{N_{k}}\sum_{i\in I_{k}}R_{i}^{2}=o_{\mathcal{P}}(1)$ by the above, we conclude by Lemma S3 that
$\frac{2}{N_{k}}\sum_{i\in I_{k}}S_{i}R_{i}=o_{\mathcal{P}}(1).$
Applying Lemma S1, we conclude that
$\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}=\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}+o_{\mathcal{P}}(1)$
so (S17) now holds by Chebyshev’s inequality and thus also (S16) follows.
Applying Chebyshev’s inequality once more, we see that
$\widehat{p}=p_{P}+o_{\mathcal{P}}(1).$ (S18)
Using this, Assumption 2 (b), (S15) and (S16), we have by Lemma S2 that
$\widehat{\sigma}^{2}_{\lambda}=\underbrace{\frac{\sigma^{2}_{\kappa_{P}}}{p_{P}^{2}}-\frac{\kappa_{P}^{2}(1-p_{P})}{p_{P}^{3}}}_{\sigma^{2}_{\lambda_{P}}}+o_{\mathcal{P}}(1).$ (S19)
We note that by the Cauchy–Schwarz inequality
$\kappa_{P}^{2}=E_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
thus by Assumption 2 (b)
$p_{P}\sigma^{2}_{\kappa_{P}}-\kappa_{P}^{2}(1-p_{P})=p_{P}\mathbb{E}_{P}[S_{i}^{2}]-\kappa_{P}^{2}\geq p_{P}\mathbb{E}_{P}\left[(Y-f_{P}(L,W))^{2}\left(\frac{(L-\pi(W))^{2}}{\pi(W)^{2}}\right)\right]\geq 1-c>0.$ (S20)
This in turn implies that $\sigma^{2}_{\lambda_{P}}\geq c^{-3}(1-c)$ so, by repeated applications of Lemma S2 using (S14), (S16), (S18) and (S12), we have
$\frac{\sqrt{n}}{\widehat{\sigma}_{\lambda}}(\widehat{\lambda}-\lambda_{P})=\frac{\sqrt{n}}{p_{P}\widehat{p}\widehat{\sigma}_{\lambda}}(\widehat{\kappa}p_{P}-\widehat{p}\kappa_{P})=\frac{\sqrt{n}}{p_{P}^{2}\sigma_{\lambda}}\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}p_{P}-(1-L_{i})\kappa_{P}\right)+o_{\mathcal{P}}(1).$
Since
$\mathrm{var}_{P}(S_{i}p_{P}-(1-L_{i})\kappa_{P})=\sigma^{2}_{\kappa_{P}}p_{P}^{2}-\kappa_{P}(1-p_{P})p_{P}=\sigma^{2}_{\lambda_{P}}p_{P}^{4},$
and $\sigma^{2}_{\lambda_{P}}p_{P}^{4}\geq c(1-c)$ by (S20), the result now follows by Lemma S4 (a) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
S5.3 Proof of (c)
Proof.
We set $\mu_{P}:=\mathbb{E}_{P}\left[\xi^{2}\right]$ and define for all $k\in[K]$, $N_{k}:=|I_{k}|$.
We first argue that
$\widehat{J}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}+o_{\mathcal{P}}(n^{-1/2}).$ (S21)
To that end, for all $k\in[K]$, we note that
$\widehat{J}_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{2}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{(\textup{I})}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}\xi_{i}}_{(\textup{II})}.$
By Assumption 3 (c), we have $(\textup{I})=o_{\mathcal{P}}(N_{k}^{-1/2})$. Since the summands of the third term are conditionally i.i.d. and mean zero given $\widehat{m}_{k}$ and $(W_{i})_{i\in I_{k}}$, we have by the conditional version of Chebyshev’s inequality that
$\mathbb{P}_{P}\left(\left|(\textup{II})\right|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{m}_{k}\right)\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}\mathrm{var}_{P}(L_{i}\,|\,W_{i}).$
Combining the above with Assumption 3 (a) and (c), we conclude that
$\mathbb{P}_{P}\left(\left|(\textup{II})\right|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{m}_{k}\right)=o_{\mathcal{P}}(N_{k}^{-1/2})=o_{\mathcal{P}}(1)$
which implies that (II$)=o_{\mathcal{P}}(N_{k}^{-1/2})( I$ as $\epsilon$ was arbitrary. Combining the above with Lemma S1, we see that (S21) holds.
We now argue that
$\widehat{\theta}-\theta_{P}=\frac{1}{n}\sum_{i=1}^{n}\frac{\varepsilon_{i}\xi_{i}}{\mu_{P}}+o_{\mathcal{P}}(n^{-1/2}).$ (S22)
To do so, we first note that
$\sup_{P\in\mathcal{P}}|\theta_{P}|=\sup_{P\in\mathcal{P}}\left|\frac{\mathbb{E}_{P}[\varepsilon\xi]}{\mathbb{E}_{P}\left[\xi^{2}\right]}\right|\leq\sup_{P\in\mathcal{P}}\frac{\mathbb{E}_{P}[|\varepsilon\xi|]}{\mu_{P}}\leq\frac{C^{1/(2+\delta)}}{c}$ (S23)
by Assumption 3 (a) and (b). Thus, for all $k\in[K]$, we note that, using the fact that $Y-g_{P}(W)=\varepsilon+\theta_{P}\xi$ and (S21),
$\widehat{\kappa}_{k}-\theta_{P}\widehat{J}$ $=\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}\xi_{i}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{(\textup{III})}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}}_{(\textup{IV})}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}}_{(\textup{V})}+o_{\mathcal{P}}(N_{k}^{-1/2}).$
By the Cauchy–Schwarz inequality and Assumption 3 (f), we have
$|(\textup{III})|\leq\sqrt{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}=o_{\mathcal{P}}(N_{k}^{-1/2}).$
[…]
$\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}-\theta_{P}\widehat{J}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}+o_{\mathcal{P}}(n^{-1/2}).$ (S24)
Next, combining (S21), Chebyshev’s inequality and Assumption 3 (a) yields that
$\widehat{J}=\mu_{P}+o_{\mathcal{P}}(1),$
hence Assumption 3 (b) and Lemma S2 (d) yields
$\widehat{J}^{-1}=\mu_{P}^{-1}+o_{\mathcal{P}}(1).$ (S25)
Combining (S25) with (S24) yields that (S22) is satisfied.
Defining $\nu_{P}:=\mathbb{E}_{P}[\varepsilon^{2}\xi^{2}]$, we now argue that
$\widehat{\sigma}^{2}_{\theta}=\frac{\nu_{P}}{\mu_{P}^{2}}+o_{\mathcal{P}}(1).$ (S26)
To that end, for all $k\in[K]$, we have
$\widehat{\nu}_{k}$ $=\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{Y_{i}-\widehat{g}_{k}(W_{i})\}^{2}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}}_{\widehat{\nu}_{k}^{(1)}}+\widehat{\theta}^{2}\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{4}}_{\widehat{\nu}_{k}^{(2)}}$ (S27)
$-2\widehat{\theta}\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{3}}_{\widehat{\nu}_{k}^{(3)}}.$
Dealing first with $\widehat{\nu}_{k}^{(1)}$, we decompose further and write
$\widehat{\nu}_{k}^{(1)}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}^{2}\xi_{i}^{2}+\frac{1}{N_{k}}\sum_{i\in I_{k}}\theta_{P}^{2}\xi_{i}^{4}+2\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\theta_{P}\varepsilon_{i}\xi_{i}^{3}}_{\theta_{P}\cdot\widetilde{(\textup{I})}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{II})}_{m}}$
$+2\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{III})}_{m}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}\xi_{i}^{2}}_{\widetilde{(\textup{II})}_{g}}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{IV})}}+2\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{V})}_{g}}$
$+2\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}^{2}}_{\widetilde{(\textup{III})}_{g}}$
$+2\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{V})}_{m}}$
$+4\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{VI})}}.$
Note first that
$\mathbb{E}_{P}\left[\varepsilon\xi^{3}\right]=\mathbb{E}_{P}\left[\mathbb{E}_{P}\left[\varepsilon\,|\,L,W\right]\xi^{3}\right]=0.$
Further, the Cauchy–Schwarz inequality yields
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[\left|\varepsilon\xi^{3}\right|^{1+\delta/6}\right]\leq\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[\left|\varepsilon\xi\right|^{1+\delta/2}\xi^{2}\right]\leq\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[\left|\varepsilon\xi\right|^{2+\delta}\right]\mathbb{E}_{P}\left[\xi^{4}\right]\leq C^{2},$
[…]
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}\left(\left|\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi^{2}\right|\geq M\right)\leq\sup_{P\in\mathcal{P}}\frac{\theta_{P}^{2}\mathbb{E}_{P}\left[\xi^{4}\right]+\mathbb{E}_{P}\left[\varepsilon^{2}\xi^{2}\right]}{M}\leq\frac{1}{M}\left(\frac{C^{(4+\delta)/(2+\delta)}}{c^{2}}+C^{2/(2+\delta)}\right),$
hence
$\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi_{i}^{2}=O_{\mathcal{P}}(1).$ (S28)
By the Cauchy–Schwarz inequality, (S28), the fact that $\widetilde{(\textup{II})}_{m}=o_{\mathcal{P}}(1)$ and Lemmas S2 and S3, we have
$|\widetilde{(\textup{III})}_{m}|\leq\left(\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi^{2}\right)^{1/2}\left(\widetilde{(\textup{II})}_{m}\right)^{1/2}=o_{\mathcal{P}}(1).$
An analogous argument can be used to show that $\widetilde{(\textup{III})}_{g}=o_{\mathcal{P}}(1)$. Using $\sum_{i}a_{i}^{2}a_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}a_{i}^{2}\right)$ and Assumption 3 (f), we have
$\widetilde{(\textup{IV})}\leq\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}=o_{\mathcal{P}}(1).$
By the Cauchy–Schwarz inequality, our results above and Lemmas S2 and S3, we have
$|\widetilde{(\textup{V})}_{g}|\leq\left(\widetilde{(\textup{II})}_{g}\right)^{1/2}\left(\widetilde{(\textup{IV})}\right)^{1/2}=o_{\mathcal{P}}(1).$
An analogous argument can be applied to show that $\widetilde{(\textup{V})}_{m}=o_{\mathcal{P}}(1)$. Finally, again by similar arguments as for the previous terms
$|\widetilde{(\textup{VI})}|\leq\left(\widetilde{(\textup{II})}_{g}\right)^{1/2}\left(\widetilde{(\textup{II})}_{m}\right)^{1/2}=o_{\mathcal{P}}(1).$
Putting things together, we conclude that
$\widehat{\nu}_{k}^{(1)}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}^{2}\xi_{i}^{2}+\frac{1}{N_{k}}\sum_{i\in I_{k}}\theta_{P}^{2}\xi_{i}^{4}+o_{\mathcal{P}}(1).$ (S29)
Turning now to $\widehat{\nu}_{k}^{(2)}$, we decompose further and write
$\widehat{\nu}_{k}^{(2)}$ $=\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{4}+4\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{3}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{VII})}_{m}}+6\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{2}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{VIII})}}$
$+4\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{3}}_{\widetilde{(\textup{IX})}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{4}}_{\widetilde{(\textup{X})}}$
Note that by similar arguments as used in (S28), we can show that $\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{4}=O_{\mathcal{P}}(1)$. By the Cauchy–Schwarz inequality and Lemmas S2 and S3, we have
$|\widetilde{(\textup{VII})}_{m}|\leq\left(\frac{1}{N_{k}}\sum_{i=1}^{n}\xi_{i}^{4}\right)^{1/2}\left((\textup{II})\right)^{1/2}=o_{\mathcal{P}}(1),$
[…]
$|\widetilde{(\textup{IX})}|\leq\left((\textup{II})\right)^{1/2}\left(\widetilde{(\textup{IX})}\right)^{1/2}=o_{\mathcal{P}}(1).$
In summary, we have that
$\widehat{\nu}_{k}^{(2)}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{4}+o_{\mathcal{P}}(1).$ (S30)
To deal with $\widehat{\nu}_{k}^{(3)}$, we decompose further and write
$\frac{1}{2}\widehat{\nu}_{k}^{(3)}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\theta_{P}\xi_{i}^{4}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}\xi_{i}^{3}}_{\widetilde{(\textup{I})}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}^{3}}_{\widetilde{(\textup{VII})}_{g}}$
$+3\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\xi_{i}^{2}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{XI})}}+3\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}^{2}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{\widetilde{(\textup{XII})}}$
$+3\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{XIII})}}+3\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{\widetilde{(\textup{XIV})}}$
$+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{3}}_{\widetilde{(\textup{XV})}}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{3}}_{\widetilde{(\textup{XVI})}}$
We have already argued that $\widetilde{(\textup{I})}=o_{\mathcal{P}}(1)$ above and using similar arguments as for $\widetilde{(\textup{VII})}_{m}$, we see that $\widetilde{(\textup{VII})}_{g}=o_{\mathcal{P}}(1)$. For the remaining terms, we use the Cauchy–Schwarz inequality and the previous results as follows
$|\widetilde{(\textup{XI})}|\leq\left(\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi^{2}\right)^{1/2}\left((\textup{II})\right)^{1/2}=o_{\mathcal{P}}(1),$
$|\widetilde{(\textup{XII})}|\leq\left(\widetilde{(\textup{II})}_{g}\right)^{1/2}\left((\textup{II})\right)^{1/2}=o_{\mathcal{P}}(1),$
$|\widetilde{(\textup{XIII})}|\leq\left(\frac{1}{N_{k}}\sum_{i\in I_{k}}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}^{2}\xi^{2}\right)^{1/2}\left(\widetilde{(\textup{X})}\right)^{1/2}=o_{\mathcal{P}}(1),$
$|\widetilde{(\textup{XIV})}|\leq\left(\widetilde{(\textup{II})}_{g}\right)^{1/2}\left(\widetilde{(\textup{X})}\right)^{1/2}=o_{\mathcal{P}}(1),$
$|\widetilde{(\textup{XV})}|\leq\left(\widetilde{(\textup{II})}_{m}\right)^{1/2}\left(\widetilde{(\textup{X})}\right)^{1/2}=o_{\mathcal{P}}(1),$
$|\widetilde{(\textup{XVI})}|\leq\left(\widetilde{(\textup{IV})}\right)^{1/2}\left(\widetilde{(\textup{X})}\right)^{1/2}=o_{\mathcal{P}}(1).$
Putting things together, we have that
$\widehat{\nu}_{k}^{(3)}=2\frac{1}{N_{k}}\sum_{i\in I_{k}}\theta_{P}\xi_{i}^{4}+o_{\mathcal{P}}(1).$ (S31)
Combining (S23) and (S22), we have that
$\widehat{\theta}=\theta_{P}+o_{\mathcal{P}}(1)$
thus by (S28), (S27), (S29), (S30) and (S31), we have
$\widehat{\nu}_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}^{2}\xi_{i}^{2}+o_{\mathcal{P}}(1),$
and hence Lemma S1 yields that
$\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\xi_{i}^{2}+o_{\mathcal{P}}(1).$
Applying Chebyshev’s inequality, we obtain
$\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}=\nu_{P}+o_{\mathcal{P}}(1),$
so by Lemma S2, (S25) and Assumption 3 (a) and (b), (S26) holds.
Finally, combining (S22), (S26) with Lemmas S3 and S2 using Assumption 3 (a) and (b), we have that
$\frac{\sqrt{n}}{\widehat{\sigma}_{\theta}}(\widehat{\theta}-\theta_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}}{\nu_{P}^{1/2}}+o_{\mathcal{P}}(1).$
The conclusion now follows from Assumption 3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
S5.4 Auxiliary results
This is a collection of minor or known results that we have included for completeness.
Lemma S1.
Let $X$ be a random variable with distribution determined by $\mathcal{P}$ and let $(X_{n})_{n\in\mathbb{N}}$ be an i.i.d. sequence of copies of $X$ with $\sup_{P\in\mathcal{P}}\mathbb{E}_{P}[|X|]\leq C<\infty$. Let $K\in\mathbb{N}$, let $I_{1},\dots,I_{K}$ be a partition of $[n]$ and for all $k\in[K]$ define $N_{k}:=|I_{k}|$ and assume that $\lfloor n/K\rfloor\leq N_{k}\leq\lceil n/K\rceil$. Suppose that there exist random variables $Y_{1},\dots,Y_{k}$ and $\delta\in[0,1)$ such that for all $k\in[K]$ it holds that
$Y_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}X_{i}+o_{\mathcal{P}}(N_{k}^{-\delta}).$
Then
$\frac{1}{K}\sum_{k=1}^{K}Y_{k}=\frac{1}{n}\sum_{i=1}^{n}X_{i}+o_{\mathcal{P}}(n^{-\delta}).$
Proof.
Without loss of generality we will assume that $\lceil n/K\rceil=N_{1}\geq\dots\geq N_{K}=\lfloor n/K\rfloor$, since otherwise we may relabel the partition so that this is satisfied.
By assumption, we can write
$Y_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}X_{i}+R_{k}$
with $R_{k}=o_{\mathcal{P}}(N_{k}^{-\delta})$ hence
$\frac{1}{K}\sum_{k=1}^{K}Y_{k}=\frac{1}{n}\sum_{i=1}^{n}X_{i}+\underbrace{\sum_{k=1}^{K}\left(\frac{1}{KN_{k}}-\frac{1}{n}\right)\sum_{i\in I_{k}}X_{i}}_{(\textup{I})}+\underbrace{\frac{1}{K}\sum_{k=1}^{K}R_{k}}_{(\textup{II})},$
and it suffices to show that $(\textup{I})=o_{\mathcal{P}}(n^{-\delta})$ and (II$)=o_{\mathcal{P}}(n^{-\delta})( I$.
[…]
$|(\textup{I})|\leq\sum_{k=1}^{K}\left|\frac{1}{KN_{k}}-\frac{1}{n}\right|\sum_{i\in I_{k}}|X_{i}|\leq\max_{k\in[K]}\left|\frac{n}{KN_{k}}-1\right|\frac{1}{n}\sum_{i=1}^{n}|X_{i}|.$
By Markov’s inequality and the assumption on $\mathbb{E}[|X|]$, we have for any $\epsilon>0$ that
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{I})|\geq n^{-\delta}\epsilon)\leq\frac{C}{\epsilon}\max_{k\in[K]}|n-KN_{k}|\frac{n^{\delta}}{KN_{k}}.$
By definition of $N_{k}$, we have
$\frac{n}{K}-1\leq N_{k}\leq n/K+1\Longrightarrow n-K\leq KN_{k}\leq n+K\Longrightarrow|n-N_{k}K|\leq K,$
thus, for $n\geq K+1$,
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{I})|\geq n^{-\delta}\epsilon)\leq\frac{CK}{\epsilon}\frac{n^{\delta}}{n-K}.$
As the upper bound goes to 00 as $n\to\infty$ by our assumption on $\delta$, we have $(\textup{I})=o_{\mathcal{P}}(n^{-\delta})$ as desired.
[…]
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}\left(\left|(\textup{II})\right|\geq n^{-\delta}\epsilon\right)\leq\sum_{k=1}^{K}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|R_{k}|\geq n^{-\delta}\epsilon)\leq\sum_{k=1}^{K}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|R_{k}|\geq(2K)^{-\delta}N_{k}^{-\delta}\epsilon).$
We conclude that (II$)=o_{\mathcal{P}}(n^{-\delta})( I$ by the assumption that $R_{k}=o_{\mathcal{P}}(N_{k}^{-\delta})$. ∎
Lemma S2.
Let $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ be sequences of real-valued random variables and let $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ be sequences of non-negative real numbers.
  1. (a)
    If $X_{n}=o_{\mathcal{P}}(a_{n})$ then $X_{n}=O_{\mathcal{P}}(a_{n})$.
  2. (b)
    If $X_{n}=o_{\mathcal{P}}(a_{n})$ and $Y_{n}=O_{\mathcal{P}}(b_{n})$ then $X_{n}Y_{n}=o_{\mathcal{P}}(a_{n}b_{n})$.
  3. (c)
    If $X_{n}=o_{\mathcal{P}}(a_{n})$ and $Y_{n}=O_{\mathcal{P}}(a_{n})$ then $X_{n}+Y_{n}=O_{\mathcal{P}}(a_{n})$.
  4. (d)
    If $X_{n}=Y_{n}+o_{\mathcal{P}}(1)$ and $Y_{n}^{-1}=O_{\mathcal{P}}(1)$, then $X_{n}^{-1}=Y_{n}^{-1}+o_{\mathcal{P}}(1)$.
Proof.
  1. (a)
    Given $\epsilon>0$, we can set $M_{\epsilon}=1$ and, since $X_{n}=o_{\mathcal{P}}(a_{n})$, we have the existence of $N_{\epsilon}$ such that
    $\sup_{n\geq N_{\epsilon}}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}|\geq a_{n})<\epsilon.$
  2. (b)
    Given $\epsilon>0$, we have for any $M>0$ that
    $\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}Y_{n}|\geq a_{n}b_{n}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}|\geq a_{n}\epsilon/M)+\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|Y_{n}|\geq b_{n}M).$
    Choosing $N$ and $M$ large enough, by assumption on $X_{n}$ and $Y_{n}$, we can make each of the two terms smaller than $\epsilon/2$ hence proving the desired result.
  3. (c)
    For any $M>0$, we have that
    $\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}+Y_{n}|\geq a_{n}M)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{n}|\geq a_{n}M/2)+\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|Y_{n}|\geq a_{n}M/2).$
    Since $X_{n}=o_{\mathcal{P}}(a_{n})$, the first term goes to 00 for any $M>0$ and since $Y_{n}=O_{\mathcal{P}}(a_{n})$, we can make the second term arbitrarily small by choosing $M$ large enough. This proves the desired result.
  4. (d)
    Let $R_{n}:=X_{n}-Y_{n}$ so that $R_{n}=o_{\mathcal{P}}(1)$. We have
    $X_{n}^{-1}-Y_{n}^{-1}=R_{n}Y_{n}^{-1}(Y_{n}+R_{n})^{-1},$
    so we are done by the previous claims if we can show that $(Y_{n}+R_{n})^{-1}=O_{\mathcal{P}}(1)$. For any $M,K,\zeta>0$ we have
    $\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|Y_{n}+R_{n}|^{-1}\geq M)$ $\leq\underbrace{\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|Y_{n}+R_{n}|^{-1}\geq M,|R_{n}|<\zeta,|Y_{n}|^{-1}<K)}_{(\textup{I})}$
    $+\underbrace{\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|Y_{n}|^{-1}\geq K)}_{(\textup{II})}+\underbrace{\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|R_{n}|\geq\zeta)}_{(\textup{III})}.$
    Given $\epsilon>0$, we can make (II$)<\epsilon/2( I$ by choosing $K$ sufficiently large. Letting $\zeta:=K^{-1}/2$, we can make $(\textup{III})<\epsilon/2$ for $n$ large enough. Letting $M:=K^{-1}/2$, $(\textup{I})=0$, since
    $|Y_{n}+R_{n}|\geq|Y_{n}|-|R_{n}|<K^{-1}/2\Longrightarrow|Y_{n}+R_{n}|^{-1}<K^{-1}/2=M.$
    As $\epsilon$ was arbitrary, this proves the desired result.
Lemma S3.
Let $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ be sequences of random variables taking values in the closed set $\mathcal{X}\subseteq\mathbb{R}^{d}$. Suppose that $X_{n}=Y_{n}+o_{\mathcal{P}}(1)$ and let $h:\mathcal{X}\rightarrow\mathbb{R}$ be a continuous function. Suppose that at least one of the following conditions hold:
  1. (a)
    $h$ is uniformly continuous,
  2. (b)
    $Y_{n}=O_{\mathcal{P}}(1)$.
Then $h(X_{n})=h(Y_{n})+o_{\mathcal{P}}(1)$.
Proof.
Let $\epsilon>0$ be given. We need to show that
$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|h(X_{n})-h(Y_{n})|>\epsilon)=0.$
If $h$ is uniformly continuous, then we can find $\delta>0$ such that $|h(x)-h(y)|\leq\epsilon$ whenever $\|x-y\|\leq\delta$. Thus,
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|h(X_{n})-h(Y_{n})|>\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\|X_{n}-Y_{n}\|>\delta)\rightarrow 0$
as $n\to\infty$. On the other hand, suppose now that $Y_{n}=O_{\mathcal{P}}(1)$ and let $M>0$. Since $h$ is continuous, it is uniformly continuous on $B(0,M):=\{x\in\mathcal{X}:\|x\|\leq M\}$, so we can choose $\delta>0$ such that $|h(x)-h(y)|\leq\epsilon$ whenever $x,y\in B(0,M)$ satisfy $\|x-y\|\leq\delta$. Hence, for $M>\delta$,
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|h(X_{n})-h(Y_{n})|>\epsilon)$ $\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\|X_{n}-Y_{n}\|>\delta)$
$\hskip 28.45274pt+\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\|X_{n}\|\vee\|Y_{n}\|>M,\|X_{n}-Y_{n}\|\leq\delta)$
$\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\|X_{n}-Y_{n}\|>\delta)+\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\|Y_{n}\|>M-\delta)\rightarrow 0$
as $n,M\to\infty$. ∎
Lemma S4.
Let $(X_{n})_{n\in\mathbb{N}}$ be a sequence of i.i.d. copies of a random variable $X$ taking values in $\mathbb{R}$. Suppose there exist constants $c,C,\delta>0$ such that for all $P\in\mathcal{P}$ $\mathrm{var}_{P}(X)\geq c$ and $\mathbb{E}_{P}\left[|X|^{2+\delta}\right]\leq C$. Then
  1. (a)
    $T:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
  2. (b)
    $\widetilde{T}:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\mathrm{var}_{P}(X)\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
Proof.
Defining $\mu_{P}:=\mathbb{E}_{P}[X]$, we have that
$\frac{1}{n}\sum_{i=1}^{n}X_{i}=\mu_{P}+o_{\mathcal{P}}(1)$
by the uniform law of large numbers in Shah and Peters [2020, Lemma 19] and the assumption that $\mathbb{E}_{P}\left[|X|^{2+\delta}\right]\leq C$. The same argument reveals that, defining $\nu_{P}:=\mathbb{E}_{P}\left[X^{2}\right]$,
$\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}=\nu_{P}+o_{\mathcal{P}}(1).$
We conclude by Lemma S2 and Lemma S3 that
$\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}=\sigma+o_{\mathcal{P}}(1).$
For any $M>0$ and $P\in\mathcal{P}$, by Markov’s inequality, Jensen’s inequality and the fact that $\mathbb{E}_{P}\left[|X|^{2+\delta}\right]\leq C$,
$\mathbb{P}_{P}\left(\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])\right|\geq M\right)\leq\frac{1}{M^{2}}\mathbb{E}_{P}\left(X^{2}\right)\leq\frac{1}{M^{2}}C^{2/(2+\delta)},$
hence the numerator of $T$ is $O_{\mathcal{P}}(1)$. Combining the facts above, by Lemma S2 since $\sigma\geq c^{1/2}$, we have that
$T=\frac{1}{\sqrt{n}\sigma}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])+o_{\mathcal{P}}(1).$
The result now follows by a uniform CLT and version of Slutsky’s theorem in Shah and Peters [2020, Lemmas 18 and 20].
Similar arguments can be applied to show the result for $\widetilde{T}$. ∎
Lemma S5.
Let $(L,W)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-2}$ and let $P_{W}$ denote the distribution of $W$. Suppose that the conditional distribution of $L$ given $W=w$ is absolutely continuous with respect to Lebesgue measure with conditional density $p(\ell\,|\,w)$. Let $g:\mathbb{R}\times\Delta^{d-2}\to\mathbb{R}$ denote a function. Assume that for $P_{W}$-almost every $w\in\Delta^{d-2}$:
  1. (a)
    $\partial_{\ell}g(\ell,w)$ exists,
  2. (b)
    $\partial_{\ell}p(\ell\,|\,w)$ exists,
  3. (c)
    $\mathbb{E}\left[|\partial_{\ell}g(L,W)+\rho(L,W)g(L,W)|\right]<\infty$, where $\rho(\ell,w):=\partial_{\ell}\log(p(\ell\,|\,w))$,
  4. (d)
    $\lim_{t\to\infty}g(t,w)p(t\,|\,w)-g(-t,w)p(-t\,|\,w)=0$.
Then
$\mathbb{E}\left[\partial_{\ell}g(L,W)+\rho(L,W)g(L,W)\,|\,W\right]=0,$
and thus integration by parts is valid, that is,
$\mathbb{E}\left[\partial_{\ell}g(L,W)\right]=-\mathbb{E}\left[\rho(L,W)g(L,W)\right].$
Proof.
The following is an adaptation of Klyne and Shah [2023, Proposition 1] (the primary change is the slight rephrasing of assumption (d)). We prove the result by showing that for $P_{W}$-almost all $w\in\Delta^{d-2}$, we have
$\mathbb{E}\left[\partial_{\ell}g(L,W)+\rho(L,W)g(L,W)\,|\,W=w\right]=0.$
Note that assumption (c) and Fubini’s theorem [Schilling, 2017, Corollary 14.9] implies that
$\ell\mapsto\partial_{\ell}g(\ell,w)p(\ell\,|\,w)+\rho(\ell,w)g(\ell,w)p(\ell\,|\,w)=\partial_{\ell}(g(\ell,w)p(\ell\,|\,w))$
is integrable with respect to Lebesgue measure for $P_{W}$ almost all $w$. Hence,
$\mathbb{E}\left[\partial_{\ell}g(L,W)+\rho(L,W)g(L,W)\,|\,W=w\right]=\int\partial_{\ell}(g(\ell,w)p(\ell\,|\,w))\,\mathrm{d}\ell$
$=\lim_{t\to\infty}\int_{-t}^{t}\partial_{\ell}(g(\ell,w)p(\ell\,|\,w))\,\mathrm{d}\ell=\lim_{t\to\infty}g(t,w)p(t\,|\,w)-g(-t,w)p(-t\,|\,w)=0,$
by dominated convergence [Schilling, 2017, Theorem 12.2], the fundamental theorem of calculus and (iv). This proves the desired result. ∎
Lemma S6.
Let $\mu$ denote the Lebesgue measure on $\mathbb{R}$, let $f:\mathbb{R}\to\mathbb{R}$ be differentiable with derivative $f^{\prime}$ and suppose that
$\max\left(\int|f|\,\mathrm{d}\mu,\int|f^{\prime}|\,\mathrm{d}\mu\right)<\infty.$
Then $\lim_{|x|\to\infty}f(x)=0$.
Proof.
Since $f^{\prime}$ is the derivative of $f$ by Rudin [1986, Theorem 7.21], we have for all $x\in\mathbb{R}$ and $y\geq x$ that
$|f(x)|=\left|f(y)-\int_{x}^{y}f^{\prime}(t)\,\mathrm{d}\mu(t)\right|\leq|f(y)|+\int_{x}^{y}|f^{\prime}(t)|\,\mathrm{d}\mu(t).$
Multiplying both sides of the inequality by $\mathbbm{1}_{[x,x+1]}(y)$, using monotonicity of integrals and Fubini’s theorem, we see that
$|f(x)|$ $=\int_{x}^{x+1}|f(x)|\,\mathrm{d}\mu(y)\leq\int_{x}^{x+1}|f(y)|\,\mathrm{d}\mu(y)+\int_{x}^{x+1}\int_{x}^{y}|f^{\prime}(t)|\,\mathrm{d}\mu(t)\,\mathrm{d}\mu(y)$
$\leq\int_{x}^{\infty}|f(y)|\,\mathrm{d}\mu(y)+\int_{x}^{\infty}\int_{x}^{x+1}|f^{\prime}(t)|\,\mathrm{d}\mu(t)\,\mathrm{d}\mu(y)=\int_{x}^{\infty}|f(y)|\,\mathrm{d}\mu(y)+\int_{x}^{\infty}|f^{\prime}(t)|\,\mathrm{d}\mu(t).$
By integrability of $f$ and $f^{\prime}$ the right-hand side must go to 00 as $x\to\infty$ which proves that $\lim_{x\to\infty}|f(x)|=0$. An analogous argument can be applied to show that $\lim_{x\to-\infty}|f(x)|=0$. ∎
Ljf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
-Perturbation-based Analysis of Compositional Data
+Perturbation-based Effect Measures for Compositional Data
 Abstract
-Existing statistical methods for compositional data analysis are inadequate for many modern applications for two reasons.
-First, modern compositional datasets, for example in microbiome research, display traits such as high-dimensionality and sparsity that are poorly modelled with traditional approaches.
+Existing effect measures for compositional features are inadequate for many modern applications for two reasons.
+First, modern datasets with compositional covariates, for example in microbiome research, display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches.
 Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward.
 In this work, we propose a framework based on hypothetical data perturbations that addresses both issues.
-Unlike existing methods for compositional data, we do not transform the data and instead use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects.
-These average perturbation effects, which can be employed in many applications, naturally account for confounding that biases frequently used marginal dependence analyses.
+Unlike many existing effect measures for compositional features, we do not define our effects based on a parametric model or a transformation of the data.
+Instead, we use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects.
+These effects naturally account for confounding that biases frequently used marginal dependence analyses.
 We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques.
-We analyze the proposed estimators empirically on simulated data and demonstrate advantages over existing techniques on US census and microbiome data.
+We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data.
 For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
-Keywords: compositional data, causality, perturbation, feature importance, diversity effects, nonparametrics
+Keywords: compositional data, effect estimation, causality, perturbation, feature importance, diversity effects, nonparametrics, mixture experiments
 1 Introduction
 Compositional data, that is, measurements consisting of parts of a whole, appear in many scientific disciplines, for example, rock compositions in geology, green house gases in the atmosphere, species compositions in ecology and racial or gender distributions in social sciences.
-The defining property of such data is that they only capture information relative to the unknown whole and hence do not contain any information on the absolute sizes.
+The defining property of such data is that they only capture information relative to the whole.
 Depending on the application, the absolute information either might not be of interest (e.g., when analyzing racial or gender distributions) or might not be easily accessible with existing measuring techniques (e.g., in ecology).
@@
 $\Delta^{d-1}:=\left\{z=(z^{1},\dots,z^{d})\in[0,1]^{d}\mid\textstyle\sum_{j=1}^{d}z^{j}=1\right\}.$
-It is well-known that many conventional statistical procedures become invalid when applied to compositional data.
-For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint.
-Much work has gone into constructing new procedures for adjusting the correlation measure for compositional components, e.g., by transforming the data with log-ratio transforms (Aitchison, 1982).
-An instructive viewpoint is to see these approaches not as adjusting the correlation measure but rather as creating new target parameters (e.g., the correlation between log-ratios) that have meaningful interpretations in compositional data.
-We adopt this point of view and develop a framework which uses hypothetical perturbations of the data to construct meaningful target parameters.
-Specifically, we consider a setting where we observe an outcome of interest $Y\in\mathbb{R}$ and an associated compositional predictor $Z\in\Delta^{d-1}$.
-Our goal is to construct functionals of the conditional distribution $Y$ given $Z$ that are useful when analyzing compositional data.
-To illustrate this, we consider two examples in which existing approaches are insufficient.
+Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982), which we discuss in more detail in Section 1.1.
+A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates.
+More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically.
+However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
+Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures.
+Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984).
+Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model.
+Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications.
+First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g.
+the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}$ to the simplex.
+In Section 4.3 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature.
+Second, effects based on directional perturbations summarize continuous changes in specific directions.
+This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$.
+For example, we can define the effect of ‘diversifying’ $Z$, i.e.
+moving closer to the center of $\Delta^{d-1}$.
+Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity.
+However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.2).
+In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
+Our proposal is intimately related to the growing literature on targeted learning and semiparametric inference (Chernozhukov et al., 2018; Berk et al., 2021; Vansteelandt and Dukes, 2022).
+An important aspect of our work is, however, to clarify the need for tailored target estimands in the presence of compositional structure in the covariates.
+Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Example 1 and Section 4.3).
+Beyond defining average perturbation effects, we also provide several algorithm-agnostic procedures to estimate the average perturbation effects.
+These procedures are essentially well-known semiparametric learning algorithms applied after appropriately transforming the compositional covariates.
+While these algorithms require us to estimate nuisance functions relating to the distribution of $(Y,Z)$, which are particularly delicate for compositional covariates, they only require these estimates to be accurate and not interpretable.
+This allows us to apply arbitrary regression methods as long as they are sufficiently predictive.
+Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$.
+To motivate this, consider Euclidean covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$.
+Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function.
+Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the rate of change in the $j$th coordinate, or more formally,
+$\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$
+We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$.
+Then, for any $x\in\mathbb{R}^{d}$, we have
+$\partial_{x_{j}}f(x)=\lim_{\gamma\to 0}\frac{f(\psi(x,\gamma))-f(\psi(x,0))}{\gamma}=\partial_{\gamma}f(\psi(x,\gamma))\bigm{|}_{\gamma=0}.$
+Therefore, the coefficient $\beta^{j}$ also corresponds to the initial rate of change (initial referring to $\gamma=0$) of $f$ along the perturbation $\psi$.
+Similar observations are possible regarding other interpretable effect measures for Euclidean covariates and under more general nonlinear models.
+Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined.
+The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$.
+We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures.
+The following two examples provide toy applications that illustrate the pitfall and short-comings of existing approaches.
 Example 1 (Effects of presence or absence of microbes).
@@
 Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances.
-Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ has been recorded.
+Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ ($1$ encoding disease) has been recorded.
 While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies.
-One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$ and then compare the distributions
-$Y\,|\,Z^{j}=0\quad\text{and}\quad Y\,|\,Z^{j}\neq 0$
-for example, with an independence test or by regressing $Y$ on $\mathbbm{1}_{\{Z^{j}=0\}}$.
+One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$, define the treatment variable $T:=\mathds{1}_{\{Z^{j}=0\}}$ corresponding to the presence or absence of microbe $j$ and then estimate a marginal treatment effect given by
+$\mathrm{MTE}:=\mathbb{E}[Y\,|\,T=1]-\mathbb{E}[Y\,|\,T=0]=\mathbb{P}(Y=1\,|\,T=1)-\mathbb{P}(Y=1\,|\,T=0),$
+where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$.
 It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components.
-As an example, assume the disease indicator is generated by
-$Y\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{Z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{Z^{2}=0\}}),$
+[^1] This is true when $Z$ is Euclidean, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$.
+To illustrate this, assume the disease indicator given the composition has conditional distribution for all $z\in\Delta^{d-1}$ given by
+$Y\,|\,Z=z\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}),$
 that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease.
 In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1.
-If $Z$ did not live on the simplex, we could adjust for the dependence in the remaining coordinates by considering conditional dependencies of the form $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$.
-However, due to the simplex constraint $Z^{-j}$ completely determines $Z^{j}$, so the conditional independence always holds.
-In the Table 1, we also provide the results from applying the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
+If $Z$ was Euclidean, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
+$h:z^{-j}\mapsto\mathbb{E}[Y\,|\,T=1,Z^{-j}=z^{-j}]-\mathbb{E}[Y\,|\,T=0,Z^{-j}=z^{-j}]$
+and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$.
+[^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE).
+Estimating this quantity is a well-studied problem, e.g.
+using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018).
+Unfortunately, neither of these estimators are well-defined when $Z$ is compositional as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}T\,|\,Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$.
+In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
 Example 1
 method estimate 95%-CI
-OLS of $Y$ on $\mathbbm{1}_{\{Z^{1}=0\}}$ $-0.12$ $(-0.17,-0.06)$
-$\mathrm{CKE}^{1}$ using rf $0.15$ $(0.10,0.21)$
+$\mathrm{MTE}$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.12}}$ $(-0.17,-0.06)$
+$\mathrm{CKE}^{1}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.15}}$ $(0.10,0.21)$
 Example 2
 method estimate 95%-CI
-OLS of Y on $-G(Z)$ $-1.57$ $(-2.18,-0.97)$
-$\mathrm{CDI}_{\mathrm{Gini}}$ using rf $0.63$ $(0.14,1.12)$
+OLS of $Y$ on $-G(Z)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-1.57}}$ $(-2.18,-0.97)$
+$\mathrm{CDI}_{\mathrm{Gini}}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.63}}$ $(0.14,1.12)$
 Table 1: Numerical illustration of Examples 1 and 2.
-The first row in each table corresponds to the marginal effect estimate based on a correctly specified linear model.
+The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model.
 The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators.
-(Left) The estimated marginal effect is negative, even though the true effect is positive.
+Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect.
+(Left) The estimated MTE is negative, even though the true effect is positive.
 The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign.
@@
 Example 2 (Effects of increased diversity).
-Consider the problem of determining the effect of an increase in diversity of $Z$ on the response $Y$.
+Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$.
 Such a question is of immediate interest in many applied fields, e.g.
@@
 In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material).
-To avoid this, one would again like to adjust for the remaining dependence within $Z$ to avoid the confounding biases.
+To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$.
 However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable.
-The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjust for the confounding as shown in Table 1.
-A more in-depth example in which the marginal approach is misleading is given in Section 4.2 where we estimate the effect of racial diversity on income on grouped real data.
+The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1.
+A more in-depth example in which the marginal approach is misleading is given in Section 4.2.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data.
+A real application is given in Section 4.2.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York.
 Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data.
 Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
-In this work, we propose a procedure that accounts for potential confounding within $Z$.
-The main idea is to introduce a new target parameter via a hypothetical perturbation of the data.
-More concretely, a perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a path $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed.
-For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
-A target parameter that captures the desired adjusted effect is then given by
-$\lambda_{\psi}:=\frac{\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,1)]\right]-\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,0)]\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
+We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here.
+A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed.
+For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
+By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$.
+The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual Euclidean approach outlined in Example 1.
+By marginalizing over $Z$ the perturbation-based approach leads to the average effect
+$\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$
+In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
+$\lambda_{\psi}:=\frac{\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
+$\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$.
 In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point.
@@
 We can then interpret $\psi(z,\epsilon)$ for a small $\epsilon>0$ as slightly increasing the diversity of the point $z$.
-With this notion of an increase in diversity, a target parameter that captures the desired adjusted effect is given by
-$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,Z=\psi(Z,\gamma)\right]\big{|}_{\gamma=0}\right],$
+With this notion of an increase in diversity, we propose to measure the effect of increasing diversity by
+$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}f(\psi(Z,\gamma))\big{|}_{\gamma=0}\right],$
 which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity.
@@
 To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section 2.3.
-While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the conditional expectation $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which due the simplex structure is generally a complex functional.
+While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional.
 We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization.
@@
 (a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
-$Y$ $\leftarrow f(Z,\eta)$ $Y$$Z$$\psi$$f$
-[…]
+$Y$ $\leftarrow f(Z,\eta)$
+$Y$$Z$$\psi$$f$
+bijection $\phi$$\phi(Z)=(L,W)$$Z=\phi^{-1}(L,W)$
 $Y$$L$$W$$\psi$$h$$m$$h$
-$L$ $\leftarrow m(W,\xi)$ $Y$ $\leftarrow h(L,W,\varepsilon)$
+$L$ $\leftarrow m(W,\xi)$
+$Y$ $\leftarrow h(L,W,\varepsilon)$
 Figure 1: (a) Data-generating SCM on the $Z$-scale.
@@
 In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied.
+We provide a short discussion on how to connect the average perturbation effects to causal quantities in Section 2.3.
 In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness.
-Our semiparametric estimators and perturbation effects are then illustrated on both simulated data and two real-world applications in Section 4.
+Our semiparametric estimators and perturbation effects are then illustrated on both simulated and semi-synthetic data and applied to two real-world applications in Section 4.
 Finally, we present the theoretical guarantees of our estimates in Section 5.
-1.1 Relation to existing compositional data analysis approaches
-Compositional data analysis has a long history dating back to Pearson (1897).
-The most prevalent class of methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach.
-In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space.
-By considering ratios instead of compositional coordinates it avoids the common pitfalls of spurious correlations that occur by comparing coordinates.
+1.1 Related literature
+This paper addresses problems from several different fields of statistics.
+Our proposed perturbation-based framework is perhaps most closely related to the literature on causal inference, semiparametric inference and targeted learning (Bickel et al., 1993; Tsiatis, 2006; van der Laan and Rose, 2011).
+Spirtes and Scheines (2004) discuss some of the problems of modelling causal effects when certain variables are deterministic functions of other variables.
+This discussion was later continued by Woodward (2015) and new causal models have been developed to address this problem (Beckers et al., 2023).
+We contribute to these developments by providing a novel framework for causal effect estimation when some covariates are compositional (see Section 2.3).
+An important aspect of causal effect estimation more generally is the need for an interpretable summary measure of the effect.
+Historically, such interpretable measures originated from parametric models that were easy to interpret.
+Much work has since gone into the interpretation of and uncertainty quantification for such parameters when the model is misspecified.
+Perhaps one of the most impactful such analyses was given by White (1980) who provided an estimation procedure for the coefficients in a linear model that did not require the model to be correctly specified.
+To what extent parameters based on linear models are useful and interpretable when the model is misspecified remains an active research topic particularly in econometrics (Aronow and Samii, 2016; Graham and de Xavier Pinto, 2022; Słoczyński, 2022).
+More recently, interpretable parameters based on partially linear models have been proposed and analyzed extensively (Chernozhukov et al., 2018; Vansteelandt and Dukes, 2022).
+Despite these and similar parameters depending on potentially complicated nuisance functions, it is often possible to estimate them efficiently under relatively mild conditions, as we also do here, by utilizing semiparametric theory (Kennedy, 2023).
+Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002).
+While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting.
+Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials.
+Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks.
+Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g.
+so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$.
+It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material).
+Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary.
+Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients.
+Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021).
+One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al.
+(2023) where this was first shown).
+Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022).
+Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis.
+It is well-known that many conventional statistical procedures become invalid when applied to compositional data.
+For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint.
+The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach.
+In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space.
 Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3).
-Furthermore, some of the perturbation effects that we highlight as interpretable overlap with effects that can also be deduced from the log-ratio perspective.
-Most prominently, we consider a class of compositional feature influence measures for which a certain member (multiplicative speed, see Example 3) corresponds to the compositional feature influence which has been previously proposed by Huang et al.
-(2023).
-It also corresponds to one of the coefficients in a log-contrast model (Aitchison and Bacon-Shone, 1984) if the conditional mean lies in the log-contrast model class.
-Our perturbation-based framework, however, has a fundamentally different starting point and allows for the construction of perturbations that are not related to the log-ratio approach.
-The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however we believe its indirect use of the data is restrictive.
-This is best illustrated by the so-called zero-problem that occurs because the log-ratio approach, in general, cannot be applied in settings with too many zeros.
-This problem has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach.
-Other approaches advocate the use of different transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022).
-While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
+The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex.
+This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach.
 Greenacre et al.
 (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches.
-In contrast to these methods, our perturbation-based framework works directly with the compositional data and defines the target parameter from this perspective.
-In the presence of zeros, perturbations that directly incorporate or omit zeros can be selected based on the application.
-Furthermore, we can construct perturbation effects that directly capture the effect of setting certain components to zero.
+While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
+There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021).
+There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022).
 1.2 Notation
@@
 We further define the direction and speed of the perturbation for all $z\in\mathcal{Z}$ by
-    $v_{\psi}(z):=\mathbbm{1}_{\{\omega_{\psi}(z)=0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
+    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
     respectively.
@@
 $\lambda_{\psi}:=\frac{\mathbb{E}\left[f(\psi(Z,1))\right]-\mathbb{E}\left[Y\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
-Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex(see Section S1 for a precise definition) at $z$, we define the average (directional) perturbation effect by
+Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 of the supplementary material for a precise definition) at $z$, we define the average (directional) perturbation effect by
 $\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}f(\psi(Z,\gamma))\big{|}_{\gamma=0}\right].$
@@
 This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint.
+For instance, the denominator makes $\lambda_{\psi}$ invariant to modifications of the distribution of $Z$ that only changes $\mathbb{P}(Z\neq\psi(Z,1))$, i.e.
+the probability that $Z$ is not already perturbed.
 While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities.
@@
 Direct estimation of both target parameters relies on estimating the regression function $f:z\mapsto\mathbb{E}[Y\mid Z=z]$ and taking averages (see Section 3).
-An alternative estimation procedure for the effect of a binary perturbation $\lambda_{\psi}$ can be derived from noting that, if $L=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W=\psi(Z,1)$, then
-$\mathbb{E}[Y\,|\,L=1,W]=\frac{\mathbb{E}[YL\,|\,W]}{\mathbb{E}[L\,|\,W]}=f(W).$
+An alternative estimation procedure for the effect of a binary perturbation $\lambda_{\psi}$ can be derived from noting that, if $L=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W=\psi(Z,1)$, then
+$\mathbb{E}[Y\,|\,L=1,W]=\frac{\mathbb{E}[YL\,|\,W]}{\mathbb{E}[L\,|\,W]}=\frac{\mathbb{E}[f(Z)L\,|\,W]}{\mathbb{E}[L\,|\,W]}=f(W).$
 Therefore, we can write
@@
 The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation.
-We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from a summary statistic and assuming it satisfies a partially linear model as in (5) with $L$ being the summary statistic.
+We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic.
 For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
@@
 The simplest type of endpoint function on the simplex is a constant function, that is, there exists a fixed point $z_{*}\in\Delta^{d-1}$ such that for all $z\in\Delta^{d-1}$ it holds that $E(z)=z_{*}$.
-For example, we can choose the fixed endpoint $z_{*}=e_{j}$, which corresponds to a vertex on the simplex and hence a perturbation with this endpoint moves proportions from the other coordinate into the $j$th one.
+For example, we can choose the fixed endpoint $z_{*}=e_{j}$, which corresponds to a vertex on the simplex and hence a perturbation with this endpoint moves proportions from the other coordinates into the $j$th one.
 A further example, is to use the center of the simplex, i.e.
@@
 As alluded to in the introduction, perturbations with $z_{\mathrm{cen}}$ as the endpoint can be seen as perturbing observations in the direction of equal distribution and hence capture a notion of diversity.
-For later reference, we call all perturbation $\psi$ with endpoint function $E_{\psi}:z\mapsto z_{\mathrm{cen}}$ centering perturbations.
+For later reference, we call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto z_{\mathrm{cen}}$ centering perturbations.
 A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex.
-An amalgamation is a well-studied operation (e.g., Aitchison, 2003) which consist of merging parts of the compositional components into others.
+An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consist of merging parts of the compositional components into others.
 Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
@@
 that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint.
-Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can be easily extended to settings where zero speeds are allowed.
+Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed.
 In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction.
 Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect.
-We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, for which a partially linear model as in (5) can be assumed, and then constructing a corresponding speed.
+We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed.
 In both cases, the construction of a derivative-isolating reparametrization is similar.
@@
 Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions.
-We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2.
-The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=-\ell w_{v}+w_{E}$, so
+We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2.
+The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=\ell w_{v}+w_{E}$, so
 $\partial_{\ell}\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=w_{v}=\partial_{\gamma}\psi(\phi_{\mathrm{unit}}^{-1}(\ell,(w_{v},w_{E})),\gamma)\big{|}_{\gamma=0}=\omega_{\psi}(\phi_{\mathrm{unit}}^{-1}(\ell,w))$
 and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$.
-Since these type of unit-speed perturbations may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic.
+Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic.
 The proof of the following result is given in Section S2.2 of the supplementary material.
@@
 Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$.
-Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(v(z),E(z))$ with $v$ the canonical endpoint direction defined in (6).
+Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=-\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(E(z),v_{E}(z))$ with $v_{E}$ defined in (6).
 Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
@@
     (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function.
-For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
+For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
     $-s_{w}(\delta)^{-1}=\partial_{\delta}t_{w}(\delta).$ (7)
-    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$ and any other perturbation with the same speed and direction.
+    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$ and any other perturbation with the same speed and direction.
   2. (b)
     (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic.
-For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing.
-Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
+For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing.
+Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
     $s(z):=-\left(\partial_{\delta}t_{\phi^{W}(z)}(\delta)\big{|}_{\delta=\|E(z)-z\|_{1}}\right)^{-1},$
@@
 For example, if we imagine the observed $Z$ as being generated from an unobserved $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale by $Z:=\mathsf{C}(X)=\frac{1}{\|X\|_{1}}X$, it might be more natural to consider which speeds are induced on the simplex from hypothetical changes to $X^{j}$.
-A first attempt might be to consider the speed resulting from adding a constant $c$ to $X^{j}$, however, since
+A first attempt might be to consider the speed resulting from adding a constant $c$ to $X^{j}$.
+However, since
 $\|\partial_{c}\mathsf{C}(X+ce_{j})\bigm{|}_{c=0}\|_{1}=2\tfrac{1}{\|X\|_{1}}\left(1-\tfrac{X^{j}}{\|X\|_{1}}\right),$
 this speed is not scale-invariant for $X$, that is, the speed when considering the counts $X$ and $cX$ for $c>0$ may differ.
-This means that the speed on the simplex would depend on the unobserved quantity $\|X\|_{1}$ so the perturbation is ill-defined.
+This means that the speed on the simplex would depend on the unobserved quantity $\|X\|_{1}$, so the perturbation is ill-defined.
 Instead, we can consider the speed resulting from multiplying $X^{j}$ by $1+c$ (so that $c=0$ corresponds to no change) which yields a well-defined perturbation on the simplex since this operation is scale-invariant.
@@
 To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a).
-Since for all $z\in\mathcal{Z}$ it holds that $E(z)^{j}=1$ and $v(z)^{j}=\frac{1}{2}$, we have that $s_{w}$ (defined in Proposition 2) does not depend on $w$ and is of the form
-$s_{w}:\delta\mapsto\frac{(2-\delta)\delta}{2}.$
+Since for all $w\in\mathcal{W}$ it holds that $w_{E}^{j}=1$ (component $w_{E}$ corresponds to the endpoint $E(z)$) and $w_{v}^{j}=\frac{1}{2}$ (component $w_{v}$ corresponds to the direction $v_{E}(z)$ and we used that $v_{E}(z)^{j}=\frac{1-z^{j}}{\|e_{j}-z^{j}\|_{1}}=\frac{1-z^{j}}{2(1-z^{j})}=\frac{1}{2}$), we have that $\phi_{\mathrm{unit}}^{-1}(\ell,w)^{j}=\frac{\ell}{2}+1$.
+Therefore,
+$s_{w}:\delta\mapsto s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))=\frac{(2-\delta)\delta}{2}.$
 Hence, solving (7) amounts to solving
@@
 which has solutions of the form $t_{w}(\delta)=\log\left(\frac{2-\delta}{\delta}\right)+C$ for $C\in\mathbb{R}$. For simplicity, we choose $C=0$ and note that as $\|e_{j}-z\|_{1}=2(1-z^{j})$, we have
-$t_{w}(\delta)=\log\left(\frac{z^{j}}{1-z^{j}}\right).$
+$\phi^{W}(z)=(e_{1},v_{E}(z))\quad\text{and}\quad\phi^{L}(z)=t_{\phi^{W}(z)}(\|e_{j}-z\|_{1})=\log\left(\frac{z^{j}}{1-z^{j}}\right)$
 We can therefore conclude (dropping the constant endpoint from the reparametrization) that
@@
 Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale.
-We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation so we suppose now that $A\cup B=[d]$.
+We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation, so we suppose now that $A\cup B=[d]$.
 The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
@@
 we have that $t_{v_{\psi}(z)}(\|z_{\mathrm{cen}}-z\|_{1})=-G(z)$. Applying Proposition 2 (b), we conclude that
-$\phi=(\phi^{L},\phi^{W}):z\mapsto(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}})$
+$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
 is a derivative-isolating reparametrization for perturbations with direction $v_{\psi}$ and speed given by
@@
 2.1.3 Perturbations arising from alternative simplex geometries
-When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
+When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
 While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions.
@@
 The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data.
-A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ by
+A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
 $\mathrm{CLR}(z)^{j}:=\log(z^{j})-\frac{1}{d}\sum_{j=1}^{d}\log(z^{j}),\quad\forall j\in[d].$
-This transformation maps the interior of $\Delta^{d-1}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$.
+This transformation maps $(\Delta^{d-1})^{\circ}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$.
 The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
@@
 The direction of this derivative can be vastly different from the centering direction used in Example 5.
-To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale, $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale.
+To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale (normalizing with the Euclidean norm as is natural here), $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale.
 It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
@@
 The derivative of the resulting perturbation on the simplex is for all $k\in[d]$ given by
-$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathbbm{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
+$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
 It is straight-forward to check that the speed and direction of this perturbation match exactly the multiplicative perturbation derived in Example 3.
@@
 For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured.
-We then wants to determine for which microbes a high abundance is associated with a high BMI.
+We then want to determine for which microbes a high abundance is associated with a high BMI.
 The simplex constraint implies that for all coordinates $j$ the conditional independence
@@
 Huang et al.
-(2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$.
+(2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material).
 A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$.
@@
 The construction in Example 5 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics.
-In Section 4.2, we use data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression.
+In Section 4.2.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression.
+In Section 4.2.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
 We do not define a binary centering perturbation here as it requires a positive probability of observing $Z=z_{\mathrm{cen}}$, which is rarely the case.
@@
 Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness.
-As above, we are again interested in effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
+As above, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
 However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally.
@@
 To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference.
-Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data generating mechanism – often called interventions.
-In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data generating mechanism under hypothetical interventions.
+Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions.
+In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions.
 The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution.
@@
 More specifically, let $\psi$ be a perturbation as defined in Definition 1.
-Then, for all $\gamma\in[0,1]$, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$.
+Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$.
 In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$.
@@
 Therefore, to estimate them, we need to make additional causal assumptions.
-For example, we may assume the variable $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$.
+For example, we may assume the variables $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$.
 In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
@@
 In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$.
-This is for example done in the numerical experiment in Section 4.2.
+This is for example done in the numerical experiments in Section 4.2.
 Finally, let us shortly discuss under which assumptions the causal and non-causal average perturbation effects coincide. Obviously, one sufficient condition is to assume that
@@
 We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice.
-However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathbbm{1}_{\{}z=\psi(z,1)\}$ and $\phi^{W}:z\mapsto\psi(z,1)$.
-This allows us to transform our estimation problem from a compositional data problem where we observe $n$ i.i.d.
-copies of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ to a generic problem with $n$ observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$.
+However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathds{1}_{\{z=\psi(z,1)\}}$ and $\phi^{W}:z\mapsto\psi(z,1)$.
+This allows us to reparametrize our estimation problem from a compositional data problem with $n$ i.i.d.
+observations of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ into a generic problem with $n$ i.i.d.
+observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$.
 We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
@@
 It can be shown that these bias-corrected estimators are asymptotically efficient, meaning that, at least asymptotically, they have the smallest possible variance of any estimator.
-Influence functions have already been computed for the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$ that we estimate here and we will describe the details of these estimates in the following sections.
+Influence functions have already been computed for the reparametrized target parameters $\tau_{\psi}$ and $\lambda_{\psi}$ that we estimate here and we will describe the details of these estimates in the following sections.
 Figure 3 contains an illustration of the coverage of confidence intervals resulting from target parameter and variance estimates from the different methods mentioned above.
@@
 We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here.
-All the presented estimators involve regressions that can be performed by arbitrary (sufficiently powerful) regression methods.
+All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
 3.2.1 Nonparametric estimation of the binary perturbation effect
-Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
+Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
 $\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}.$
 The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities).
-The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of average predictive effects (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature.
-The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ is related to the augmented inverse propensity weighted (AIPW) estimator and is given by
+The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature.
+The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
 $\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i},$
@@
 This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1).
-Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplement.
+Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplementary material.
 3.2.3 Partially linear estimation of perturbation effects
-While the estimators in the previous sections make only mild assumption on the data generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions.
+While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions.
 One way to avoid these complicated estimations is by additionally assuming a partially linear model.
@@
 In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3.
-In Section 4.2 we analyze the effect of diversity on income based on grouped data from the 1994 US census.
+In Section 4.2.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census.
+In Section 4.2.2, we analyze the effect of racial diversity on grades in New York schools.
 In Section 4.3 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data.
@@
   • PLM_no_x: The same estimators as PLM but without cross-fitting.
-  • plugin: The naive plugin versions of the NPM estimates.
+  • plugin: The naive plug-in versions of the NPM estimates.
 In the binary effect case, the estimator in (15) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (16) without the term involving $\widehat{\rho}$.
   • plugin_no_x: The same estimators as plugin but without cross-fitting.
-The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plugin is estimated using an estimator akin to the one given in (14).
-Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{g}$ and $\widehat{m}$ and NPM and NPM_no_x additionally on the score estimate $\widehat{\rho}$.
+The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (14).
+Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$.
 For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction.
-For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplement, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
-We consider four data generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
-In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathbbm{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
+For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplementary material, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
+We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
+In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
 We then consider the following four generative models
@@
   • Binary $L$ with nonparametric $Y$:
-    $L=U_{0}B(W_{1})+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
+    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
   • Continuous $L$ with partially linear $Y$:
@@
 We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=15$ in Figure 4 while coverage rates for $d=3$ and $d=75$ are reported in Section S4.2 of the supplementary material.
-Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedures over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$.
+Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$.
 PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting.
 NPM works in both settings but can be less robust for small sample sizes.
-Further, it can be seen that the plugin estimator and the estimators without cross-fitting do not lead to correct coverage in some of these settings.
+Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
 In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated.
-We see that the plugin estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
-In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the tests, e.g.
+We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
+In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g.
 for $n=1000$ and binary $L$.
-In the cases where cross-fitting is not necessary, it does not seem to negatively impact the performance of the test.
-4.2 Compositional confounding in relationship between income and race
-In this section we provide an example based on real data that illustrates how failing to adjust for compositional confounding can lead to incorrect conclusions about effects of compositional variables.
+In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
+4.2 Estimating effects of diversity on socio-economic outcomes
+4.2.1 Effects of diversity on income in semi-synthetic US census data
+In this section we provide an example based on semi-synthetic data that illustrates how failing to adjust for compositional confounding can lead to incorrect conclusions about effects of compositional variables.
 We consider the ‘Adult’ dataset available from the UCI machine learning repository (Becker and Kohavi, 1996).
@@
 Additionally, we consider the demographic features age (numeric, between 17 and 90), education (categorical, with 16 categories), sex (binary, 00 encoding male and $1$ female) and race (categorical, with categories ’White’, ’Black’ and ’Other’).
-We now use this individual-level data to construct community-level datasets by artificially grouping individuals together.
-This allows us to construct settings, where there is a known dependence between the racial diversity of the communities and the response variable.
-As we show below, this illustrates the potential pitfalls of ignoring compositional structure in the data and how the approach based on perturbations solves these issues.
-More specifically, we group all individual-level observations into groups of around $50$ individuals.
+We now use this raw individual-level data to construct community-level datasets by artificially grouping individuals together.
+The grouping is deliberately constructed in a way such that diversity affects income and this relationship is confounded by the remaining variation of the racial composition.
+This semi-synthetic example is only intended as an illustration.
+A genuine analysis of this data would require knowledge of the true community structure, which is not available in the data extract we are using here.
+We emphasize that even if the individual-level data is available, one may still be interested in effects of the racial composition of the community an individual lives in on that individual’s income.
+For such an analysis, one can add the racial composition of an individual’s community as a variable and then perform the same analysis as here but now on the individual-level data while adjusting for individual-level confounding factors.
+Importantly, even for such an individual-level analysis, the simplex confounding can bias the results in the same way as we illustrate here.
+We provide an example of this bias occurring on the individual-level effect in Section S4.4 of the supplementary material.
+We group all individual-level observations into groups of around $50$ individuals.
 We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group.
@@
 Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
-Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$).
+Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
 Depending on the color the dependence between $Y$ and the Gini coefficient changes.
@@
 This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption.
-The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3.
-The naive estimator that conditions either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect.
+The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3 where random forests are used for nuisance function estimation when including additional covariates.
+The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect.
 Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure.
@@
 In this setting, we propose to apply $\mathrm{CDI}_{\mathrm{Gini}}$ with and without conditioning on the additional covariates $X$.
-The results are also shown in Table 3.
+The results are also shown in Table 3 again using random forests for nuisance function estimation.
 As desired one gets significant positive effects with this method as it is adapted to the simplex.
@@
 Estimate $95\%$ CI Estimate $95\%$ CI
-naive_diversity ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.082}}$ $(-0.144,-0.019)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.120}}$ $(-0.194,-0.046)$
+naive_diversity ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.082}}$ $(-0.144,-0.019)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.120}}$ $(-0.194,-0.046)$
 […]
@@
 […]
-$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
-$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
+$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
 Table 3: Table of estimates in the ‘Adult’ data experiment.
-Bold numbers indicate significance at a $5\%$ level, red indicates a negative effect, while green indicates a positive effect of diversity.
+Bold red numbers indicate a significant negative effect of diversity at a $5\%$ level, while bold green indicates a significant positive effect.
 The naive_diversity method indicates that either there is a significant negative or a highly uncertain effect of diversity on compensation despite the fact that the data is constructed with a positive effect of diversity.
 The CDI correctly identifies this effect.
+4.2.2 Effects of diversity on grades in New York schools
+In this section we analyze data from a 2018 Kaggle competition containing data on schools in New York (PASSNYC, 2018).
+The data (collected in 2016) contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S4.5 of the supplementary material).
+We let $X$ denote all background variables except racial composition which we denote by $Z$.
+Race is given as whole percentages of ‘Asian’, ‘Black’, ‘Hispanic’ and ‘White’ students which we rescale so that $Z\in\Delta^{3}$.
+We are interested in the effect of diversity, again measured by the Gini coefficient, on the average grade of the schools (either ELA or math).
+We repeat the analysis in Section 4.2.1 on this real data and the results can be seen in Table 4.
+The naive_diversity method indicates that there is a significant effect of diversity on grades even when controlling for additional covariates $X$.
+In contrast, the CDI estimates smaller positive effects for ELA and effects that cannot be distinguished from 00 for math.
+This discrepancy between the two methods indicates that there is a significant amount of compositional confounding in this data, which invalidates the marginal analysis.
+Method Effect on ELA grades Effect on math grades
+Estimate $95\%$ CI Estimate $95\%$ CI
+naive_diversity ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{1.079}}$ $(0.934,1.223)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{1.231}}$ $(1.040,1.422)$
+[…]
+$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.345}}$ $(0.105,0.585)$ $0.098$ $(-0.175,0.371)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.163}}$ $(0.002,0.323)$ $0.047$ $(-0.167,0.261)$
+Table 4: Table of estimates in the New York schools data experiment.
+Bold green numbers indicate a significant positive effect of diversity at a $5\%$ level.
+The naive_diversity method indicates that there is a significant effect of diversity on grades even when controlling for additional covariates $X$.
+The CDI (with or without controlling for $X$) generally estimates smaller positive effects or even effects that cannot be distinguished from 00, indicating a significant amount of compositional confounding in this data.
 4.3 Variable influence measures when predicting BMI from gut microbiome
@@
 The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator.
-For the permutation measure the inconsistency comes the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate.
+For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate.
 Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021).
@@
 To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data.
-More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\mid Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint.
+More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint.
 For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression.
@@
 6 Discussion
-We introduced perturbations on the simplex as a formalization of hypothetical changes to a compositional vector and used them to construct interpretable target parameters, which we call average perturbation effects.
+We introduced perturbations on the simplex as a formalization of hypothetical changes to a compositional vector and used them to construct interpretable compositional effect measures, which we call average perturbation effects.
 In our view, perturbations provide a natural way for practitioners to formalize research questions about non-standard changes to a system.
@@
 Unfortunately, many applied researchers working with compositional data, generally due to a lack of alternatives, still rely on marginal effects by analyzing dependence between a summary statistic and a response (e.g., by regressing a response on a diversity measure).
-As we demonstrated, such approaches can be misleading since they do not account for the confounding induced by the remaining variation in compositional vector.
+As we demonstrated, such approaches can be misleading since they do not account for the confounding induced by the remaining variation in the compositional vector.
 We hope that this work can draw attention to such pitfalls and improve the analysis of such data by providing a set of easily accessible alternative methods.
 We additionally proposed efficient estimators for the average perturbation effects based on semiparametric theory and provided theoretical guarantees that enable us to perform valid inference.
-Most statistical methods for compositional data analysis to date are based on the parametric log-ratio approach introduced in the seminal paper Aitchison (1982).
+Most statistical methods for compositional data analysis to date are based on the parametric log-ratio approach introduced in the seminal paper Aitchison (1982) or other parametric and transformation-based approaches.
 Our nonparametric approach goes beyond these established methods and allows practitioners to apply theoretically justified procedures to modern compositional data applications.
@@
 There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data.
-The idea of defining target parameters via perturbations easily extends to those settings.
+The idea of defining effect measures via perturbations easily extends to those settings.
 Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference (e.g., how to define causal variables).
@@
 Acknowledgements
-We thank Alexander Mangulad Christgau and Lucas Kook for helpful discussions. ARL and NP were supported by a research grant (0069071) from Novo Nordisk Fonden.
+We thank Alexander Mangulad Christgau and Lucas Kook for helpful discussions.
+We also thank Viktor Skjold Lindegaard for spotting some minor errors in an earlier version of the manuscript.
+ARL and NP were supported by a research grant (0069071) from Novo Nordisk Fonden.
 Appendix S1 Differentiability on the simplex
@@
 (b) By the same argument as in (a), we have that $\phi$ is invertible with inverse as computed in (S3).
-This reparametrization is derivative-isolating for perturbations with direction $v$ given by (6) and speed given by (reversing the computation in (a))
-$s:z\mapsto-\partial_{\ell}t_{(E(z),v(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v(z))}(\|E(z)-z\|_{1}))^{-1}.$
+This reparametrization is derivative-isolating for perturbations with direction $v_{E}$ given by (6) and speed given by (reversing the computation in (a))
+$s:z\mapsto-\partial_{\ell}t_{(E(z),v_{E}(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v_{E}(z))}(\|E(z)-z\|_{1}))^{-1}.$
 Appendix S3 Estimation technicalities and algorithms
@@
 We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
-$L:=W^{1}+(1+\mathbbm{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
+$L:=W^{1}+(1+\mathds{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
 The out-of-sample predictions evaluated on a new dataset with $n=1000$ observations can be seen together with the true function in Figure S1.
@@
 First, we note that instead of the optimization in (S4) we can also solve
-$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[Y\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
+$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[L\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
 since using $h:w\mapsto\widetilde{h}(w)+\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$ leads to a solution of (S4).
@@
 1:Input: Data $(Y_{i},Z_{i})_{i\in[n]}$, directional perturbation $\psi$, derivative-isolating reparametrization $\phi$.
-2:For $i\in[n]$, define $D_{i}:=\mathbbm{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
+2:For $i\in[n]$, define $D_{i}:=\mathds{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
 3:Define :=$\mathcal{I}:=\{i\in[n]\mid D_{i}=1\}$.
@@
 Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators.
-Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.
+Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.1.
+Section S4.5 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.2.2.
 S4.1 Simulations from Section 1
@@
 where $W^{2}:=U^{W}(1-L)+U^{W}LB$ and $Z^{1}:=(1-L)U^{L}$. Furthermore, define $Y$ such that
-$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathbbm{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathbbm{1}_{\{Z^{2}=0\}}\right).$
+$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\right).$
 In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
@@
 The patterns observed in Figure 4 of Section 4.1 are repeated here.
-S4.3 Description of the construction of groups in Section 4.2
-In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.2.
+S4.3 Description of the construction of groups in Section 4.2.1
+In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.2.1.
 On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
@@
 That is, category $2$ is more diverse and with higher compensation ($0.3>\widehat{\pi}_{C}(1)$). Lastly, we set the probability $p_{i}^{0}$ for category 00 is $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
-S4.4 Additional plots from gut microbiome analysis in Section 4.3
+S4.4 Semi-synthetic individual-level experiment using groups from Section 4.2.1
+In this section we consider a variant of the experiment performed in Section 4.2.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups.
+More specifically, we generate the groups as before and estimate the probability that $Y=1$ within each group and use this probability to generate new $Y$s for each observation.
+For each individual, we thus have a response $Y$, a race $R$, a ‘contextual’ variable $Z$ that gives the racial distribution of each group and additional covariates $X$.
+By construction of the groups and sampling of the $Y$, we know that there is a positive effect of diversity on the income.
+We now apply the naive_diversity and $\mathrm{CDI}_{\mathrm{Gini}}$ methods as in Section 4.2.1 except we always control for the race of inviduals and omit the unstable estimators that condition on $Z$.
+The results can be seen in Table 5.
+Method Grouping on compensation
+Estimate $95\%$ CI
+[…]
+[…]
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.459}}$ $(0.247,0.671)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.487}}$ $(0.282,0.692)$
+Table 5: Table of estimates in the individual-level ‘Adult’ data experiment.
+Bold red numbers indicate a significant negative effect of diversity at a 5% level, while bold green indicates a significant positive effect.
+The naive diversity method indicates that there is a significant negative effect of diversity on compensation, even when controlling for individual race ($R$) and additional covariates ($X$), despite the fact that the data is constructed with a positive effect of diversity.
+The CDI correctly identifies that the effect is positive.
+We see that, as in Section 4.2.1, the naive_diversity method incorrectly estimates a negative effect of diversity on income while $\mathrm{CDI}_{\mathrm{Gini}}$ correctly identifies that the effect is positive.
+S4.5 Description of additional control variables used in Section 4.2.2
+The additional background variables in the New York schools data used in Section 4.2.2 are described below. Unless otherwise mentioned, no preprocessing was used.
+School income estimate
+The school income estimates were grouped into four categories; one for missing values and three more based on quantiles of the observed values.
+School grade range (grade referring to years of school)
+The range of grades for the schools were grouped into three categories; one for ‘high grades’ where the lowest grade was greater or equal to 6th, one for ‘low grades’ where the highest grade was less than or equal to 4th and the remaining schools were put in a ‘mixed’ category.
+Student attendance rate
+This number was computed as the total number of days attended by all students / total number of days on register for all students.
+% chronically absent students
+The percentage of students missing 10% of school days or at least 18 days per year in a 180-day school year.
+Economic Need Index
+Computed as the percentage of students in temporary housing plus 0.5 times the percentage of students that are eligible for Health Reimbursement Arrangements plus 0.5 times the percentage of students eligible for free lunch.
+Community school
+An indicator of whether the school is a community school.
+% English language learners
+Percentage of students who are learning English for the first time.
+Rigorous Instruction %
+School rating for ‘how well the curriculum and instruction engage students, build critical-thinking skills, and are aligned to the Common Core’.
+Collaborative Teachers %
+School rating for ‘how well teachers participate in opportunities to develop, grow, and contribute to the continuous improvement of the school community’.
+Supportive Environment %
+School rating for ‘how well the school establishes a culture where students feel safe, challenged to grow, and supported to meet high expectations’.
+Effective School Leadership %
+School rating for ‘how well school leadership inspires the school community with a clear instructional vision and effectively distributes leadership to realize this vision’.
+Strong Family-Community Ties %
+School rating for ‘how well the school forms effective partnerships with families to improve the school’.
+Trust %
+School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
+S4.6 Additional plots from gut microbiome analysis in Section 4.3
 In this section we provide an additional figure showing our results from the analysis of the gut microbiome data in Section 4.3.
@@
 satisfies $\mathbb{P}_{P}(\Omega_{P})\geq 1-\zeta/2$ for all $P\in\mathcal{P}$. We therefore have for any $\epsilon>0$
-$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\{|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\}\cap\Omega_{P})+\zeta/2\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)+\zeta/2,$
-hence it suffices to prove that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ to conclude that $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$.
+$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\{|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\}\cap\Omega_{P})+\zeta/2\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)+\zeta/2,$
+hence it suffices to prove that $\mathds{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ to conclude that $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$.
 To that end, we first note that
-$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}\right]=\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)\,|\,\widehat{f}_{k}\right].$
+$\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}\right]=\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)\,|\,\widehat{f}_{k}\right].$
 We will show that
-$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}=h\right]=0$ (S11)
+$\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}=h\right]=0$ (S11)
 for $\mathbb{P}_{P}$-almost every $h$.
-If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=0$ then clearly the conditional expectation is 00.
-If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=1$, then we intend to apply Lemma S5 (integration by parts) with $g=(h-f_{P})$.
-Assumptions (a)-(c) of Lemma S5 are clear from the above and (d) follows from Lemma S6 and Assumption (a).
+If $h$ is such that $\mathds{1}_{\Omega_{P}}=0$ then clearly the conditional expectation is 00.
+If $h$ is such that $\mathds{1}_{\Omega_{P}}=1$, then we intend to apply Lemma S5 (integration by parts) with $g=(h-f_{P})$.
+Assumptions (a)-(c) of Lemma S5 are clear from the above.
+Assumption (d) of Lemma S5 also holds, since then both $\mathbb{E}[|h(L,W)-f_{P}(L,W)|\,|\,W]$ and $\mathbb{E}[|\partial_{\ell}(h(L,W)-f_{P}(L,W))+\rho_{P}(L,W)(h(L,W)-f_{P}(L,W))|\,|\,W]$ are almost surely finite when $\mathds{1}_{\Omega_{P}}=1$, so Lemma S6 yields that the assumption is satisfied.
 We conclude that
@@
 By conditional Chebyshev’s inequality (applicable since we have shown that summands are conditionally mean zero in (S11)), the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and conditional Cauchy–Schwarz inequality, we therefore have
-$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\,|\,\widehat{f}_{k})$
-$\qquad\leq\frac{1}{\epsilon^{2}N_{k,r}}\mathbb{E}_{P}[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})^{2}\,|\,\widehat{f}_{k}]$
+$\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\,|\,\widehat{f}_{k})$
+$\qquad\leq\frac{1}{\epsilon^{2}N_{k,r}}\mathbb{E}_{P}[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})^{2}\,|\,\widehat{f}_{k}]$
 $\qquad=\frac{1}{\epsilon^{2}}\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)^{2}\,|\,\widehat{f}_{k}\right]$
@@
 Assumption 1 (c), (d) and (a) and Lemma S3 thus yield that
-$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
-which implies that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ so $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ by our earlier argument.
+$\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
+which implies that $\mathds{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ so $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ by our earlier argument.
 The above combined with Lemma S1 proves (S10).
@@
 $\mathbb{E}_{P}[S_{i}]=\mathbb{E}_{P}[\partial_{\ell}f_{P}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\varepsilon_{i}]=\tau_{P},$
-and (S12) implies that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\mathrm{CFI}^{j})=O_{\mathcal{P}}(1)$, we have by Lemma S3, (S10) and (S13) that
+and (S12) implies that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\tau_{P})=O_{\mathcal{P}}(1)$, we have by Lemma S3, (S10) and (S13) that
 $\frac{\sqrt{n}}{\widehat{\sigma}_{\tau}}(\widehat{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1).$
 The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption 1 (a) and (b), since
-$\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)\right)^{2}\right]$
-$=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
+$\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}f_{P}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\partial_{\ell}f_{P}(L,W)\right]^{2}$
+$\geq\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
 S5.2 Proof of (b)
@@
 by Assumption 2 (b),
-$|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathbbm{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
+$|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathds{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
 and
@@
 We note that by the Cauchy–Schwarz inequality
-$\kappa_{P}^{2}=E_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
+$\kappa_{P}^{2}=\mathbb{E}_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
 thus by Assumption 2 (b)
@@
   1. (a)
-    $T:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
+    $T:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}}\quad\text{satisfies}\quad\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{t\in\mathbb{R}}|\mathbb{P}_{P}(T\leq t)-\Phi(t)|=0$
   2. (b)
-    $\widetilde{T}:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\mathrm{var}_{P}(X)\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
+    $T:=\widetilde{T}:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\mathrm{var}_{P}(X)\right)^{1/2}}\quad\text{satisfies}\quad\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{t\in\mathbb{R}}|\mathbb{P}_{P}(\widetilde{T}\leq t)-\Phi(t)|=0$
 Proof.
@@
 Lemma S5.
-Let $(L,W)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-2}$ and let $P_{W}$ denote the distribution of $W$.
-Suppose that the conditional distribution of $L$ given $W=w$ is absolutely continuous with respect to Lebesgue measure with conditional density $p(\ell\,|\,w)$.
-Let $g:\mathbb{R}\times\Delta^{d-2}\to\mathbb{R}$ denote a function.
-Assume that for $P_{W}$-almost every $w\in\Delta^{d-2}$:
+[…]
   1. (a)
@@
 The following is an adaptation of Klyne and Shah [2023, Proposition 1] (the primary change is the slight rephrasing of assumption (d)).
-We prove the result by showing that for $P_{W}$-almost all $w\in\Delta^{d-2}$, we have
+We prove the result by showing that for $P_{W}$-almost all $w\in\mathcal{W}$, we have
 $\mathbb{E}\left[\partial_{\ell}g(L,W)+\rho(L,W)g(L,W)\,|\,W=w\right]=0.$
@@
 $|f(x)|=\left|f(y)-\int_{x}^{y}f^{\prime}(t)\,\mathrm{d}\mu(t)\right|\leq|f(y)|+\int_{x}^{y}|f^{\prime}(t)|\,\mathrm{d}\mu(t).$
-Multiplying both sides of the inequality by $\mathbbm{1}_{[x,x+1]}(y)$, using monotonicity of integrals and Fubini’s theorem, we see that
+Multiplying both sides of the inequality by $\mathds{1}_{[x,x+1]}(y)$, using monotonicity of integrals and Fubini’s theorem, we see that
 $|f(x)|$ $=\int_{x}^{x+1}|f(x)|\,\mathrm{d}\mu(y)\leq\int_{x}^{x+1}|f(y)|\,\mathrm{d}\mu(y)+\int_{x}^{x+1}\int_{x}^{y}|f^{\prime}(t)|\,\mathrm{d}\mu(t)\,\mathrm{d}\mu(y)$
@@
 ∎
-Ljf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+Appendix S6 Miscellaneous results
+The following result provides a relationship between the $\mathrm{CFI}^{j}$ and coefficients in a well-specified log-contrast regression model.
+The result has been previously proven in Huang et al.
+[2023].
+Proposition S3.
+Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$.
+Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}\log(z^{j})+\mu$.
+Then, for all $j\in[d]$, we have
+$\mathrm{CFI}^{j}_{\mathrm{mult}}=\beta^{j}.$
+Proof.
+The perturbation corresponding to $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is given by
+$\psi(z,\gamma)=z+\gamma\frac{2z^{j}(1-z^{j})}{\|e_{j}-z\|_{1}}(e_{j}-z)=z+\gamma z^{j}(e_{j}-z).$
+We thus have
+$f(\psi(z,\gamma))=\sum_{k=1}^{d}\beta^{k}\log(z^{k}+\gamma z^{j}(\mathds{1}_{\{k=j\}}-z^{k}))+\mu$
+and therefore
+$\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}=\sum_{k=1}^{d}\frac{\beta^{k}z^{j}(\mathds{1}_{\{k=j\}}-z^{k})}{z^{k}}=\beta^{j}-z^{j}\sum_{k=1}^{d}\beta^{k}=\beta^{j}.$
+We conclude that
+$\mathrm{CFI}^{j}_{\mathrm{mult}}=\tau_{\psi}=\mathbb{E}\left[\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}\right]=\beta^{j}.$
+The next result shows that the effects proposed in Cox [1971] are special cases of average perturbation effects in a parametric model.
+Proposition S4.
+Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $z_{*}\in\Delta^{d-1}$.
+Let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ and suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}z_{*}^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}z^{j}+\mu$.
+For $k\in[d]$, define the perturbation $\psi_{k}$ such that for all components $j\in[d]$ it holds
+$\psi_{k}^{j}(z,\gamma)=\begin{cases}z^{j}+\gamma\quad&\text{if }j=k\\ z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\quad&\text{else}.\end{cases}$
+Then
+$\tau_{\psi_{k}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+Proof.
+We have
+$f(\psi_{k}(z,\gamma))=\beta^{k}(z^{k}+\gamma)+\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\left(z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\right)+\mu$
+and therefore
+$\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}=\beta^{k}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\beta^{k}\left(1+\frac{z^{k}_{*}}{1-z^{k}_{*}}\right)-\sum\limits_{j=1}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+We conclude that
+$\tau_{\psi_{k}}=\mathbb{E}\left[\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}\right]=\frac{\beta^{k}}{1-z^{k}_{*}}.$

### v3

@@
 $\Delta^{d-1}:=\left\{z=(z^{1},\dots,z^{d})\in[0,1]^{d}\mid\textstyle\sum_{j=1}^{d}z^{j}=1\right\}.$
-Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982), which we discuss in more detail in Section 1.1.
+Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982).
+We briefly summarize the developments of these fields in Section 1.1.
 A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates.
@@
 First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g.
-the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}$ to the simplex.
+the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex.
 In Section 4.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature.
@@
 Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$.
-To motivate this, consider Euclidean covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$.
+To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$.
 Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function.
-Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the rate of change in the $j$th coordinate, or more formally,
+Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally,
 $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$
@@
 Therefore, the coefficient $\beta^{j}$ also corresponds to the initial rate of change (initial referring to $\gamma=0$) of $f$ along the perturbation $\psi$.
-Similar observations are possible regarding other interpretable effect measures for Euclidean covariates and under more general nonlinear models.
+Similar observations are possible regarding other interpretable effect measures for unconstrained covariates and under more general nonlinear models.
 Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined.
@@
 It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components.
-[^1] This is true when $Z$ is Euclidean, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$.
-To illustrate this, assume the disease indicator given the composition has conditional distribution for all $z\in\Delta^{d-1}$ given by
-$Y\,|\,Z=z\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}),$
-that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease.
-In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1.
-If $Z$ was Euclidean, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
+[^1] This is true when $Z\in\mathbb{R}^{d}$ is unconstrained, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$.
+To illustrate this, assume the disease indicator $Y$ given the composition $Z$ has conditional distribution for all $z\in\Delta^{d-1}$ given by a Bernoulli distribution with
+$\mathbb{P}(Y=1\,|\,Z=z)=\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}$
+that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ reduces the probability of having the disease.
+In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1 (details are provided in Section S4.1 of the supplementary material).
+If $Z\in\mathbb{R}^{d}$ was unconstrained, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
 $h:z^{-j}\mapsto\mathbb{E}[Y\,|\,T=1,Z^{-j}=z^{-j}]-\mathbb{E}[Y\,|\,T=0,Z^{-j}=z^{-j}]$
@@
 using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018).
-Unfortunately, neither of these estimators are well-defined when $Z$ is compositional as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}T\,|\,Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$.
+Unfortunately, neither of the estimators is well defined when $Z\in\Delta^{d-1}$ as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y$ is conditionally independent of $T$ given $Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$.
 In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
@@
 Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect.
-(Left) The estimated MTE is negative, even though the true effect is positive.
+Left: The estimated MTE is negative, even though the true effect is positive.
 The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign.
-(Right) The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction.
+Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction.
 The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect.
@@
 A real application is given in Section 4.1.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York.
-Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data.
+Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained data.
 Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
@@
 By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$.
-The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual Euclidean approach outlined in Example 1.
+The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach outlined in Example 1.
 By marginalizing over $Z$ the perturbation-based approach leads to the average effect
@@
 (2023) where this was first shown).
-Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022).
+Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022; Li et al., 2023; Bhaduri et al., 2025).
 Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis.
@@
 Greenacre et al.
-(2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches.
+(2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches which are further highlighted in Scealy and Welsh (2014).
 While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
 There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021).
-There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022).
+There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022; Kurisu et al., 2024).
 1.2 Notation
@@
 For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$.
-Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator, with the additional convention that for $0\in\mathbb{R}^{d}$ we have $\mathsf{C}(0)=(1/d,\dots,1/d)=:z_{\mathrm{cen}}$.
+Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
 2 Target parameters via perturbations
@@
 We further define the direction and speed of the perturbation for all $z\in\mathcal{Z}$ by
-    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
+    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{and}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
     respectively.
@@
 The proof of the following result can be found in Section S2.1 of the supplementary material.
-Proposition 1 (derivative-isolating).
+Proposition 1 (Isolating derivatives).
 Consider the setting of Definition 2.
@@
 For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1).
-For a fixed endpoint there is a canonical (potentially geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint.
+For a fixed endpoint there is a canonical (geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint.
 The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation.
@@
 The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications.
-The more applied reader may jump directly to that section and skip the more technical discussion on perturbations here.
+Section 2.2 may be particularly interesting for the reader who is more interested in applications of our framework.
 2.1.1 Endpoints in the simplex
@@
 A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex.
-An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consist of merging parts of the compositional components into others.
-Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
+An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consists of adding together the values of groups of components of $z$ resulting in a lower dimensional composition.
+Formally, for disjoint $A,B\subseteq[d]$, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
 $\mathsf{A}_{A\to B}(z)^{A}:=0,\quad\mathsf{A}_{A\to B}(z)^{B}:=\mathsf{C}(z^{B})(\|z^{A}\|_{1}+\|z^{B}\|_{1}),\quad\mathsf{A}_{A\to B}(z)^{[d]\setminus(A\cup B)}:=z^{[d]\setminus(A\cup B)}.$
@@
 We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations.
+We give an example of a scientific question involving a general amalgamation in Section 2.2.3.
 An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e.
@@
 The proof of the following result is given in Section S2.2 of the supplementary material.
-Proposition 2.
+Proposition 2 (Finding derivative-isolating reparametrizations).
 For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$.
@@
 We illustrate this procedure in the following two examples and discuss connections to the centering and amalgamating perturbations defined above.
-Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations (the diversity perturbation in blue has Gini-speed (see (10)) and multiplicative speed (see (8)) for the amalgamating perturbations in red and green) and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$.
+Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$.
 The types of perturbation are distinguished by color.
+The alternative speed for the diversity perturbation in blue is Gini-speed (see (10)) and for the amalgamating perturbations in red and green the speed is multiplicative speed (see (8)).
 The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds.
@@
 As in Example 5, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space.
-This line transformed back to the simplex can be written as
-$\{\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))\mid\gamma\in[0,1]\}.$
-Therefore, computing the diversifying direction in the Aitchison geometry amounts to differentiating the expression above at $\gamma=0$ and normalizing. Applying the chain rule, this yields
-$\partial_{\gamma}\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
-which can be normalized to a unit $1$-norm vector.
+Using the Aitchison geometry, this defines a perturbation via
+$\psi_{\mathrm{Ait}}:(z,\gamma)\mapsto\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z)).$
+We can compute $\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)\bigm{|}_{\gamma=0}$ to find a perturbation $\psi$ of the form (1) for which $\tau_{\psi}=\tau_{\psi_{\mathrm{Ait}}}$.
+Applying the chain rule, we obtain
+$\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
+which can be decomposed into a speed and direction as in Definition 1 to obtain $\psi$.
 The direction of this derivative can be vastly different from the centering direction used in Example 5.
-To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale (normalizing with the Euclidean norm as is natural here), $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale.
-It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
+To obtain a derivative-isolating reparametrization, since $\psi_{\mathrm{Ait}}$ shrinks towards 00 in CLR-space, it is straightforward to show that $\phi^{L}:z\mapsto-\|\mathrm{CLR}(z)\|_{2}$ and $\phi^{W}:z\mapsto\mathrm{CLR}(z)/\|\mathrm{CLR}(z)\|_{2}$ are derivative-isolating reparametrizations.
 Example 7 (Feature influence in the Aitchison geometry).
 Suppose, as in Example 3, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$.
-For $\mathrm{CLR}(z)\in\mathbb{R}^{d}$ we have a natural notion of increasing $z^{j}$ inherited from $\mathbb{R}^{d}$, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$.
-The derivative of the resulting perturbation on the simplex is for all $k\in[d]$ given by
-$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
+There is a natural notion of increasing $z^{j}$ inherited from CLR-space, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$ and transforming back to the simplex.
+The resulting perturbation is given by
+$\psi:(z,\gamma)\mapsto\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j}).$
+The derivative of the resulting perturbation is for all $k\in[d]$ given by
+$\partial_{\gamma}\psi(z,\gamma)^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
 It is straight-forward to check that the speed and direction of this perturbation match exactly the multiplicative perturbation derived in Example 3.
@@
 For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured.
-We then want to determine for which microbes a high abundance is associated with a high BMI.
-The simplex constraint implies that for all coordinates $j$ the conditional independence
-$Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$ (11)
-is trivially satisfied and feature importance or influence measures based on (11) are meaningless.
-Intuitively, this stems from the fact that changing a single component always requires a change in at least some of the remaining components.
+We then want to determine for which microbes a high abundance is associated with a high BMI.The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
+Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$.
 We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally.
@@
 This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023).
+The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest.
 For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
@@
 hence does not provide valid inference (see Figure 3 (left)) as the confidence intervals can be biased or too narrow.
-Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section 4.1 with $d=15$ and $n=1000$.
+Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S4.2 of the supplementary material with $d=15$ and $n=1000$.
 The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter.
@@
 4 Numerical experiments
-We now present three numerical experiments illustrating the usage and performance of our methods.
-In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3.
+We now present numerical experiments illustrating the usage and performance of our methods.
 In Section 4.1.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census.
@@
 In Section 4.2 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data.
+In addition to the experiments included here, Section S4.2 of the supplementary material contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
 The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa.
 For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.2 where we used the c-lasso package (Simpson et al., 2021).
-4.1 Semiparametric estimators
-In this section we investigate the performance of the semiparametric estimators outlined in Section 3 with $W$ as a simplex-valued random variable.
-More specifically, we compare the different approaches of estimating
-$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}\quad\text{and}\quad\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}]$
-from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
-  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (15) and (16).
-See Algorithms 1 and 2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
-  • NPM_no_x: The same estimators as NPM but without cross-fitting.
-  • NPM_oracle: The same estimators as NPM including cross-fitting but using the true nuisance functions for the one-step-correction (propensity score $\pi$ in the binary case and conditional score function $\rho$ in the directional case).
-This is an oracle as the true nuisance functions are unknown in practice.
-  • PLM: The partially least model based estimators in (18) including cross-fitting. See Algorithm 3 with $K=2$ for the detailed estimators.
-  • PLM_no_x: The same estimators as PLM but without cross-fitting.
-  • plugin: The naive plug-in versions of the NPM estimates.
-In the binary effect case, the estimator in (15) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (16) without the term involving $\widehat{\rho}$.
-  • plugin_no_x: The same estimators as plugin but without cross-fitting.
-The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (14).
-Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$.
-For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction.
-For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplementary material, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
-We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
-In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
-We then consider the following four generative models
-  • Binary $L$ with partially linear $Y$:
-    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=L+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
-  • Binary $L$ with nonparametric $Y$:
-    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
-  • Continuous $L$ with partially linear $Y$:
-    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
-  • Continuous $L$ with nonparametric $Y$:
-    $L=B(W_{1})+(1+B(W_{1}))\xi\quad\text{and}\quad Y=2B(W_{1})L+B(W_{1})+\varepsilon$
-The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$).
-The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$.
-For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above.
-We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=15$ in Figure 4 while coverage rates for $d=3$ and $d=75$ are reported in Section S4.2 of the supplementary material.
-Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$.
-PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting.
-NPM works in both settings but can be less robust for small sample sizes.
-Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
-In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated.
-We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
-In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g.
-for $n=1000$ and binary $L$.
-In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
 4.1 Estimating effects of diversity on socio-economic outcomes
@@
 This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation.
+For the data where we have grouped on education the racial composition for the resulting groups is illustrated in the left plot of Figure 4.
+We expect that – by construction – there is a positive effect of increased diversity on compensation because education is positively associated with compensation.
+This can be seen in the right plot of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
+The same occurs when grouping on compensation and the corresponding figure can be found in Section S4.4 of the supplementary material.
 For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 4.
 In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation).
-Figure 5: Top: Racial compositions for both types of grouping in a ternary plot.
+This can be seen in the lower half of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
+Figure 4: Summary of the semi-synthetic ‘Adult’ dataset when grouping on education.
+Left: Racial compositions in a ternary plot.
 Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
-Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
+Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
 Depending on the color the dependence between $Y$ and the Gini coefficient changes.
+The Gini coefficient represents a similarity rather than a diversity measure.
 Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g.
@@
 Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option.
-To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data.
+To this end we consider a standard approach for estimating effects of coordinates for unconstrained data.
 More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint.
@@
 The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model.
-The bottom left plot shows the $p$-values from different methods for the three most significant species from each method.
+The bottom left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability).
 We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros.
@@
 In this work we have focused on data constraints induced by the simplex.
-There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data.
+There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained data.
 The idea of defining effect measures via perturbations easily extends to those settings.
 Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference (e.g., how to define causal variables).
-Nevertheless, for both extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory.
+Furthermore, there may be settings where the perturbation of interest depends on additional covariates which then changes the resulting target of inference.
+Nevertheless, for all of the above extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory.
 We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
@@
 The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting).
-We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section 4.1 which we do in Figure S3.
+We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section S4.2 which we do in Figure S3.
 Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
@@
 we obtain the desired result by minimizing the quadratic in $\theta$. ∎
-Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section 4.1 with $d=15$.
+Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section S4.2 with $d=15$.
 We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
@@
 Section S4.1 provides the details behind the experiments in Table 1 in Section 1.
-Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators.
+Section S4.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
 Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1.
+Section S4.4 includes an additional figure showing the semi-synthetic dataset used in Section 4.1.1.
 Section S4.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2.
@@
 S4.1.1 Example 1: Effects of absence or presence of microbes
-Let $L\sim\mathrm{Bern}(1/2)$, $U^{W},U^{L}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
+Let $L\sim\mathrm{Bern}(1/2)$, $U^{1},U^{2}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
 $Z:=(Z^{1},W^{2}(1-Z^{1}),(1-W^{2})(1-Z^{1}))\in\Delta^{2},$
-where $W^{2}:=U^{W}(1-L)+U^{W}LB$ and $Z^{1}:=(1-L)U^{L}$. Furthermore, define $Y$ such that
-$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\right).$
+where $W^{2}:=U^{1}(1-L)+U^{1}LB$ and $Z^{1}:=(1-L)U^{2}$. Furthermore, define $Y$ such that $Y$ is Bernoulli distributed with
+$\mathbb{P}(Y=1\,|\,Z)=\sim\mathrm{Bern}\biggl{(}\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\biggr{)}.$
 In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
@@
 $\mathrm{CDI}_{\mathrm{Gini}}$ is computed using the $\tau$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
-S4.2 Additional semiparametric estimation results
-Here we provide additional results from the experiment of Section 4.1.
-Figure S4 and Figure S5 contain estimated coverage rates of the $95\%$ confidence intervals for the different methods and settings with $d=3$ and $d=75$, respectively.
-The patterns observed in Figure 4 of Section 4.1 are repeated in these figures.
-Figure S4: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=3$.
-The patterns observed in Figure 4 of Section 4.1 are repeated here.
-Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=75$.
-The patterns observed in Figure 4 of Section 4.1 are repeated here.
+S4.2 Semiparametric estimators
+In this section we investigate the performance of the semiparametric estimators outlined in Section 3 with $W$ as a simplex-valued random variable.
+More specifically, we compare the different approaches of estimating
+$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}\quad\text{and}\quad\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}]$
+from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
+  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (14) and (15).
+See Algorithms 1 and 2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
+  • NPM_no_x: The same estimators as NPM but without cross-fitting.
+  • NPM_oracle: The same estimators as NPM including cross-fitting but using the true nuisance functions for the one-step-correction (propensity score $\pi$ in the binary case and conditional score function $\rho$ in the directional case).
+This is an oracle as the true nuisance functions are unknown in practice.
+  • PLM: The partially least model based estimators in (17) including cross-fitting. See Algorithm 3 with $K=2$ for the detailed estimators.
+  • PLM_no_x: The same estimators as PLM but without cross-fitting.
+  • plugin: The naive plug-in versions of the NPM estimates.
+In the binary effect case, the estimator in (14) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (15) without the term involving $\widehat{\rho}$.
+  • plugin_no_x: The same estimators as plugin but without cross-fitting.
+The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (13).
+Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$.
+For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction.
+For the score estimator we use the location-scale based estimator described in Section S3.4, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
+We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
+In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
+We then consider the following four generative models
+  • Binary $L$ with partially linear $Y$:
+    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=L+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
+  • Binary $L$ with nonparametric $Y$:
+    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
+  • Continuous $L$ with partially linear $Y$:
+    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
+  • Continuous $L$ with nonparametric $Y$:
+    $L=B(W_{1})+(1+B(W_{1}))\xi\quad\text{and}\quad Y=2B(W_{1})L+B(W_{1})+\varepsilon$
+The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$).
+The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$.
+For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above.
+We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=3$ in Figure S4, for $d=15$ in Figure S5 and for $d=75$ in Figure S6.
+Figure S4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section S4.2 with $d=3$.
+PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting.
+NPM works in both settings but can be less robust for small sample sizes.
+Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
+Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=15$.
+The patterns observed in Figure S4 are repeated here.
+Figure S6: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=75$.
+The patterns observed in Figure S4 are repeated here.
+In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated.
+We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
+In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g.
+for $n=1000$ and binary $L$.
+In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
 S4.3 Description of the construction of groups in Section 4.1.1
@@
 That is, category $2$ is more diverse and with higher compensation ($0.3>\widehat{\pi}_{C}(1)$). Lastly, we set the probability $p_{i}^{0}$ for category 00 is $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
-S4.4 Semi-synthetic individual-level experiment using groups from Section 4.2.1
+S4.4 Additional figure for the ‘Adult’ data experiment in Section 4.1.1
+Figure S7 complements Figure 4 in Section S4.4 by summarizing the semi-synthetic ‘Adult’ dataset when grouping on compensation.
+Figure S7: Summary of the semi-synthetic ‘Adult’ dataset when grouping on compensation.
+Left: Racial compositions in a ternary plot.
+Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
+Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
+Depending on the color the dependence between $Y$ and the Gini coefficient changes.
+Note that the Gini coefficient represents a similarity rather than a diversity measure.
+S4.5 Semi-synthetic individual-level experiment using groups from Section 4.1.1
 In this section we consider a variant of the experiment performed in Section 4.1.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups.
@@
 We see that, as in Section 4.1.1, the naive_diversity method incorrectly estimates a negative effect of diversity on income while $\mathrm{CDI}_{\mathrm{Gini}}$ correctly identifies that the effect is positive.
-S4.5 Description of additional control variables used in Section 4.2.2
+S4.6 Description of additional control variables used in Section 4.1.2
 The additional background variables in the New York schools data used in Section 4.1.2 are described below. Unless otherwise mentioned, no preprocessing was used.
@@
 School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
-S4.6 Additional plots from gut microbiome analysis in Section 4.3
+S4.7 Additional plots from gut microbiome analysis in Section 4.2
 In this section we provide an additional figure showing our results from the analysis of the gut microbiome data in Section 4.2.

### v4

@@
 Abstract
-Existing effect measures for compositional features are inadequate for many modern applications for two reasons.
-First, modern datasets with compositional covariates, for example in microbiome research, display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches.
-Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward.
-In this work, we propose a framework based on hypothetical data perturbations that addresses both issues.
-Unlike many existing effect measures for compositional features, we do not define our effects based on a parametric model or a transformation of the data.
-Instead, we use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects.
+Existing effect measures for compositional features are inadequate for many modern applications, for example, in microbiome research, since they display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches.
+Further, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward.
+We propose a framework based on hypothetical data perturbations which defines interpretable statistical functionals on the compositions themselves, which we call average perturbation effects.
 These effects naturally account for confounding that biases frequently used marginal dependence analyses.
@@
 We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data.
-For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
 Keywords: compositional data, effect estimation, causality, perturbation, feature importance, diversity effects, nonparametrics, mixture experiments
@@
 Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982).
-We briefly summarize the developments of these fields in Section 1.1.
+We summarize the historical development of these fields in Section S2.1 of the supplementary material.
 A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates.
@@
 Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity.
-However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.1).
+However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Section 4.1).
 In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
 An important aspect of our work is, however, to clarify the need for tailored target estimands in the presence of compositional structure in the covariates.
-Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Example 1 and Section 4.2).
+Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Section 4.2).
 Beyond defining average perturbation effects, we also provide several algorithm-agnostic procedures to estimate the average perturbation effects.
@@
 To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$.
-Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function.
-Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally,
-$\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$
-We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$.
+Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function of $x$.
+Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$.
 Then, for any $x\in\mathbb{R}^{d}$, we have
@@
 We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures.
-The following two examples provide toy applications that illustrate the pitfall and short-comings of existing approaches.
-Example 1 (Effects of presence or absence of microbes).
-Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease.
-Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances.
-Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ ($1$ encoding disease) has been recorded.
-While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies.
-One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$, define the treatment variable $T:=\mathds{1}_{\{Z^{j}=0\}}$ corresponding to the presence or absence of microbe $j$ and then estimate a marginal treatment effect given by
-$\mathrm{MTE}:=\mathbb{E}[Y\,|\,T=1]-\mathbb{E}[Y\,|\,T=0]=\mathbb{P}(Y=1\,|\,T=1)-\mathbb{P}(Y=1\,|\,T=0),$
-where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$.
-It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components.
-[^1] This is true when $Z\in\mathbb{R}^{d}$ is unconstrained, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$.
-To illustrate this, assume the disease indicator $Y$ given the composition $Z$ has conditional distribution for all $z\in\Delta^{d-1}$ given by a Bernoulli distribution with
-$\mathbb{P}(Y=1\,|\,Z=z)=\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}$
-that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ reduces the probability of having the disease.
-In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1 (details are provided in Section S4.1 of the supplementary material).
-If $Z\in\mathbb{R}^{d}$ was unconstrained, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
-$h:z^{-j}\mapsto\mathbb{E}[Y\,|\,T=1,Z^{-j}=z^{-j}]-\mathbb{E}[Y\,|\,T=0,Z^{-j}=z^{-j}]$
-and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$.
-[^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE).
-Estimating this quantity is a well-studied problem, e.g.
-using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018).
-Unfortunately, neither of the estimators is well defined when $Z\in\Delta^{d-1}$ as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y$ is conditionally independent of $T$ given $Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$.
-In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
-Example 1
-method estimate 95%-CI
-$\mathrm{MTE}$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.12}}$ $(-0.17,-0.06)$
-$\mathrm{CKE}^{1}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.15}}$ $(0.10,0.21)$
-Example 2
-method estimate 95%-CI
-OLS of $Y$ on $-G(Z)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-1.57}}$ $(-2.18,-0.97)$
-$\mathrm{CDI}_{\mathrm{Gini}}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.63}}$ $(0.14,1.12)$
-Table 1: Numerical illustration of Examples 1 and 2.
-The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model.
-The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators.
-Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect.
-Left: The estimated MTE is negative, even though the true effect is positive.
-The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign.
-Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction.
-The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect.
-Details on both experiments are provided in Section S4.1 of the supplementary material.
-Example 2 (Effects of increased diversity).
-Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$.
-Such a question is of immediate interest in many applied fields, e.g.
-when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012).
-The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$.
-As in Example 1 this type of marginal dependence analysis is potentially misleading in practice.
-In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material).
-To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$.
-However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable.
-The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1.
-A more in-depth example in which the marginal approach is misleading is given in Section 4.1.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data.
-A real application is given in Section 4.1.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York.
-Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained data.
-Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
-We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here.
-A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed.
-For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
+In Section S2.2 of the supplementary material we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
+A perturbation, formally defined in Section 2, is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed.
+For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
 By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$.
-The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach outlined in Example 1.
-By marginalizing over $Z$ the perturbation-based approach leads to the average effect
-$\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$
-In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
+The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach.
+By marginalizing over $Z$ the perturbation-based approach leads to the average effect $\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$ In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
 $\lambda_{\psi}:=\frac{\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
-$\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$.
-In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point.
+The quantity $\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$.
+An increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point.
 Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
@@
 Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities.
-To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section 2.3.
+To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3 of the supplementary material.
 While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional.
 We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization.
-To make this more concrete, assume that the structural causal model (SCM) (e.g., Pearl, 2009) given in Figure 1 (left) generates the data (the causal terminology is used only to provide a more intuitive introduction).
-Given a perturbation $\psi$ that we are interested in, we now propose to construct a bijection $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ that maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure 1 (right)).
-Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which standard adjustment works, that is, using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the causal effect.
+Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the desired effect.
 Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$.
 We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
-(a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
-$Y$ $\leftarrow f(Z,\eta)$
-$Y$$Z$$\psi$$f$
-bijection $\phi$$\phi(Z)=(L,W)$$Z=\phi^{-1}(L,W)$
-$Y$$L$$W$$\psi$$h$$m$$h$
-$L$ $\leftarrow m(W,\xi)$
-$Y$ $\leftarrow h(L,W,\varepsilon)$
-Figure 1: (a) Data-generating SCM on the $Z$-scale.
-$Z\in\Delta^{d-1}$ is a compositional vector, $Y$ is a response variable and $\eta$ is exogenous noise.
-(b) Reparametrized SCM on the $(L,W)$-scale.
-$(L,W)=\phi(Z)$ are the reparametrized predictors, $Y$ is the same response as in (a) and $\xi$ and $\varepsilon$ are exogenous noise variables.
-The reparametrization needs to disentangle the perturbation effect on $Z$ in a way such that intervening on $Z$ by setting it to $\psi(Z,\gamma)$ corresponds to intervening on only $L$ (red double head arrows).
-The remainder of the paper is structured as follows.
-In Section 1.1 we discuss existing methods for compositional data analysis and how our approach contributes to this research.
-Section 1.2 collects the mathematical notation used throughout this work.
-Our perturbation-based framework is introduced in Section 2.
-In Section 2.1, we provide a way to construct interpretable perturbations based on suitably choosing the endpoint, direction and speed of the perturbation.
-We also provide a general construction for the aforementioned $L$ and $W$ for a wide range of perturbations.
-We end the section by discussing how to adapt our proposals to general geometries on the simplex including those employed by many current methods in compositional data analysis.
-In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied.
-We provide a short discussion on how to connect the average perturbation effects to causal quantities in Section 2.3.
-In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness.
-Our semiparametric estimators and perturbation effects are then illustrated on both simulated and semi-synthetic data and applied to two real-world applications in Section 4.
-Finally, we present the theoretical guarantees of our estimates in Section 5.
-1.1 Related literature
-This paper addresses problems from several different fields of statistics.
-Our proposed perturbation-based framework is perhaps most closely related to the literature on causal inference, semiparametric inference and targeted learning (Bickel et al., 1993; Tsiatis, 2006; van der Laan and Rose, 2011).
-Spirtes and Scheines (2004) discuss some of the problems of modelling causal effects when certain variables are deterministic functions of other variables.
-This discussion was later continued by Woodward (2015) and new causal models have been developed to address this problem (Beckers et al., 2023).
-We contribute to these developments by providing a novel framework for causal effect estimation when some covariates are compositional (see Section 2.3).
-An important aspect of causal effect estimation more generally is the need for an interpretable summary measure of the effect.
-Historically, such interpretable measures originated from parametric models that were easy to interpret.
-Much work has since gone into the interpretation of and uncertainty quantification for such parameters when the model is misspecified.
-Perhaps one of the most impactful such analyses was given by White (1980) who provided an estimation procedure for the coefficients in a linear model that did not require the model to be correctly specified.
-To what extent parameters based on linear models are useful and interpretable when the model is misspecified remains an active research topic particularly in econometrics (Aronow and Samii, 2016; Graham and de Xavier Pinto, 2022; Słoczyński, 2022).
-More recently, interpretable parameters based on partially linear models have been proposed and analyzed extensively (Chernozhukov et al., 2018; Vansteelandt and Dukes, 2022).
-Despite these and similar parameters depending on potentially complicated nuisance functions, it is often possible to estimate them efficiently under relatively mild conditions, as we also do here, by utilizing semiparametric theory (Kennedy, 2023).
-Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002).
-While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting.
-Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials.
-Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks.
-Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g.
-so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$.
-It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material).
-Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary.
-Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients.
-Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021).
-One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al.
-(2023) where this was first shown).
-Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022; Li et al., 2023; Bhaduri et al., 2025).
-Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis.
-It is well-known that many conventional statistical procedures become invalid when applied to compositional data.
-For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint.
-The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach.
-In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space.
-Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3).
-The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex.
-This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach.
-Greenacre et al.
-(2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches which are further highlighted in Scealy and Welsh (2014).
-While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
-There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021).
-There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022; Kurisu et al., 2024).
+The paper is structured as follows.
+Our perturbation-based framework is formally introduced and developed in Section 2.
+We describe how to construct perturbations in Section 2.1 and provide a table of pre-defined simplex effects along with examples of their use in Section 2.2.
+In Section 3 we give an overview of the semiparametric techniques used to estimate our proposed effects.
+In Section 4 we provide a semi-synthetic and two real world applications of our proposal.
+The supplementary material contains technical details on theory and algorithms along with additional examples and experiments.
 1.1 Notation
@@
 Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
+All references to sections and results beginning with ‘S’ point to the supplementary material.
 2 Target parameters via perturbations
@@
 $\psi(z,\gamma)=z+\gamma s_{\psi}(z)v_{\psi}(z).$ (1)
-We will discuss how to construct perturbations with interpretable directions and speeds in Section 2.1.2.
+We construct interpretable perturbations in Section 2.1.2.
 Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex.
-In this work, we only consider binary and directional perturbations but other classes of perturbations could be of interest too.
-Perturbations can also be defined on more general sample spaces than simplices and many of the perturbation-related concepts described here can be readily extended to these settings.
+The concept of a perturbation can be defined in greater generality than we do here but we stick to the simplex for simplicity.
 Using the two types of perturbations, we can define informative summary statistics that capture the effect of the perturbations on the response.
@@
 Let $\mathcal{Z}\subseteq\Delta^{d-1}$, let $(Y,Z)\in\mathbb{R}\times\mathcal{Z}$ be random variables and define $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$.
-Then, for a binary perturbation $\psi$, if $f$ is identified on $\{\psi(z^{\prime},1)\in\Delta^{d-1}\mid z^{\prime}\in\mathcal{Z}\}$ and $\mathbb{P}(Z\neq\psi(Z,1))>0$, we define the average (binary) perturbation effect by
+Then, for a binary perturbation $\psi$, if $\mathbb{P}(Z\neq\psi(Z,1))>0$ and $z\mapsto\psi(z,1)$ preserves null sets of $Z$, we define the average (binary) perturbation effect by
 $\lambda_{\psi}:=\frac{\mathbb{E}\left[f(\psi(Z,1))\right]-\mathbb{E}\left[Y\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
-Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 of the supplementary material for a precise definition) at $z$, we define the average (directional) perturbation effect by
+Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 for a precise definition) at $z$, we define the average (directional) perturbation effect by
 $\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}f(\psi(Z,\gamma))\big{|}_{\gamma=0}\right].$
-We showcase different potential applications for both types of perturbation effect in the numerical experiments in Section 4.
+The conditions we impose on the perturbation and distribution of $Z$ are only to ensure that the parameters are well-defined (see Section S2.3).
+We showcase different potential applications for both types of perturbation effect in the examples in Section S2.2 and the numerical experiments in Section 4.
 For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$.
@@
 While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities.
-In particular, they are identifiable from the observational distribution as long as the conditions in Definition 2 are satisfied.
-We discuss the assumptions required for causal conclusions in Section 2.3.
+We discuss the assumptions required for causal conclusions in Section S3.
 Direct estimation of both target parameters relies on estimating the regression function $f:z\mapsto\mathbb{E}[Y\mid Z=z]$ and taking averages (see Section 3).
@@
 which is an expression more amenable to analysis by semiparametric theory and thus allows us to construct efficient estimators.
-If we are willing to assume that $f(\psi(Z,1))-f(Z)$ is constant when $L=0$ and equal to, say $\theta$, then
+If we are willing to assume that $f(\psi(Z,1))-f(Z)$ is constant when $L=0$ and equal to $\theta$, then
 $\lambda_{\psi}=\frac{\mathbb{E}[\{f(W)-f(Z)\}(1-L)]}{\mathbb{P}(L=0)}=\theta$
@@
 This derivation is only morally correct when $Z\in\Delta^{d-1}$ ($\nabla f$ is not defined in the usual sense for simplex-valued functions) and to make it formally correct one needs to apply techniques from differential geometry.
-The proof of the following result can be found in Section S2.1 of the supplementary material.
+The proof of the following result can be found in Section S2.4.
 Proposition 1 (Isolating derivatives).
@@
 We say that $\phi$ is derivative-isolating if it satisfies (3).
-From this result it follows that if we are able to find a derivative-isolating reparametrization (we describe how to find such reparametrizations for certain perturbations in Section 2.1.2) the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993).
+From this result it follows that if we are able to find a derivative-isolating reparametrization the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993).
 It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
@@
 We can therefore allow for such points if we appropriately adapt the estimates.
-In Section S3.2 of the supplementary material, we describe how to modify the estimation procedures introduced in Section 3 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
+In Section S4.4, we describe how to modify the estimation procedures introduced in Section 3 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
 2.1 Perturbations on the simplex
@@
 We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic.
-For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
-In the remainder of this section, we discuss interpretable simplex endpoints in Section 2.1.1, provide details on the selection of directions and speed together with the construction of derivative-isolating reparametrizations in Section 2.1.2 and finally in Section 2.1.3 discuss how one could construct directional perturbations that respect a different geometry than the Euclidean one.
-The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications.
-Section 2.2 may be particularly interesting for the reader who is more interested in applications of our framework.
+For both approaches, we provide – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
 2.1.1 Endpoints in the simplex
@@
 We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations.
-We give an example of a scientific question involving a general amalgamation in Section 2.2.3.
-An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e.
+We discuss general amalgamating perturbations in Section S2.6.
+Here, we focus on an important class of amalgamating perturbations: those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e.
 perturbations that push towards a vertex of the simplex.
-Amalgamations, however, do not need to have fixed endpoints but allow for more general endpoints that can also depend on the initial point.
-Several perturbations on $\Delta^{2}$ are visualized in Figure 2 below.
 2.1.2 Directions, speeds and derivative-isolating reparametrizations
-When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as
-$\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$
-An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
+When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as $\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$ An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
 $v_{E}(z):=\frac{E(z)-z}{\|E(z)-z\|_{1}},$ (6)
@@
 Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed.
-In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction.
+In Section S2.7 we describe how working in other geometries can lead to different notions of direction.
 Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect.
@@
 In both cases, the construction of a derivative-isolating reparametrization is similar.
-As a fundamental building block for the general construction below, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
+As a fundamental building block for the general construction, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
 $\psi(z,\gamma):=z+\gamma\frac{E(z)-z}{\|E(z)-z\|_{1}}.$
 Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions.
-We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2.
+We therefore consider the reparametrization defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$.
 The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=\ell w_{v}+w_{E}$, so
@@
 and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$.
-Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic.
-The proof of the following result is given in Section S2.2 of the supplementary material.
-Proposition 2 (Finding derivative-isolating reparametrizations).
-For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$.
-Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$.
-Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=-\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(E(z),v_{E}(z))$ with $v_{E}$ defined in (6).
-Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
-  1. (a)
-    (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function.
-For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
-    $-s_{w}(\delta)^{-1}=\partial_{\delta}t_{w}(\delta).$ (7)
-    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$ and any other perturbation with the same speed and direction.
-  2. (b)
-    (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic.
-For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing.
-Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
-    $s(z):=-\left(\partial_{\delta}t_{\phi^{W}(z)}(\delta)\big{|}_{\delta=\|E(z)-z\|_{1}}\right)^{-1},$
-    and any other perturbation with the same speed and direction.
-Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (7).
-Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint.
-Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized.
-We illustrate the use of Proposition 2 by constructing several directional perturbations.
+Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction in Section S2.5 to start from either a given interpretable speed or a summary statistic.
+We illustrate both options in the following two examples.
 Example 1 (Feature influence).
@@
 $s_{\psi}(z)=2z^{j}(1-z^{j})$
-which we call multiplicative speed.
-To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a).
-Since for all $w\in\mathcal{W}$ it holds that $w_{E}^{j}=1$ (component $w_{E}$ corresponds to the endpoint $E(z)$) and $w_{v}^{j}=\frac{1}{2}$ (component $w_{v}$ corresponds to the direction $v_{E}(z)$ and we used that $v_{E}(z)^{j}=\frac{1-z^{j}}{\|e_{j}-z^{j}\|_{1}}=\frac{1-z^{j}}{2(1-z^{j})}=\frac{1}{2}$), we have that $\phi_{\mathrm{unit}}^{-1}(\ell,w)^{j}=\frac{\ell}{2}+1$.
-Therefore,
-$s_{w}:\delta\mapsto s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))=\frac{(2-\delta)\delta}{2}.$
-Hence, solving (7) amounts to solving
-$-\frac{2}{(2-\delta)\delta}=\partial_{\delta}t_{w}(\delta)$
-which has solutions of the form $t_{w}(\delta)=\log\left(\frac{2-\delta}{\delta}\right)+C$ for $C\in\mathbb{R}$. For simplicity, we choose $C=0$ and note that as $\|e_{j}-z\|_{1}=2(1-z^{j})$, we have
-$\phi^{W}(z)=(e_{1},v_{E}(z))\quad\text{and}\quad\phi^{L}(z)=t_{\phi^{W}(z)}(\|e_{j}-z\|_{1})=\log\left(\frac{z^{j}}{1-z^{j}}\right)$
-We can therefore conclude (dropping the constant endpoint from the reparametrization) that
+which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
 $\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{z^{j}}{1-z^{j}}\right),v_{\psi}(z)\right)$
 is a derivative-isolating reparametrization for the $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation $\psi$ with multiplicative speed.
-Example 4 (Amalgamation influence).
-We can extend the considerations of Example 3 to more general amalgamating perturbations.
-For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation.
-Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale.
-We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation, so we suppose now that $A\cup B=[d]$.
-The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
-$c\mapsto\mathsf{C}(X\odot(1+cu_{B})),$
-which similarly corresponds to a well-defined perturbation on the simplex. Computing the speed, we obtain that
-$\|\partial_{c}\mathsf{C}(X\odot(1+cu_{B}))\bigm{|}_{c=0}\|_{1}=\frac{2\|X^{A}\|_{1}\|X^{B}\|_{1}}{(\|X^{A}\|_{1}+\|X^{B}\|_{1})^{2}}.$
-We can use this speed even when $A\cup B\neq[d]$ resulting in the general multiplicative speed
-$s_{\psi}(z)=\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}.$ (8)
-Using Proposition 2 (a) as in Example 3, we further obtain that
-$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\right),(\mathsf{A}_{A\to B}(z),v_{\psi}(z))\right)$
-is a derivative-isolating reparametrization for $\psi$ with multiplicative speed.
 Example 2 (Diversity with Gini-speed).
@@
 We can investigate such an effect by using a centering directional perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{z_{\mathrm{cen}}\}$ that $v_{\psi}(z)=\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}$.
-As every element of $\mathcal{W}$ as defined in Proposition 2 is of the form $(z_{\mathrm{cen}},u)$ for a direction $u$, we instead drop the endpoint from the set and use $\mathcal{W}=\{u\in\mathbb{R}^{d}\mid\|u\|_{1}=1,\sum_{j=1}^{d}u^{j}=0\}$ instead.
-In contrast to the Examples 3 and 4 it is not as obvious how to choose an interpretable speed.
-We therefore start from a summary statistic and apply Proposition 2 (b) to derive the corresponding speed and derivative-isolating reparametrization.
-For econometricians a natural choice of diversity measure might be the Gini coefficient.
+In contrast to Example 1 it is not as obvious how to choose an interpretable speed.
+For econometricians a natural choice of similarity measure might be the Gini coefficient.
 The Gini coefficient $G:\Delta^{d-1}\to[0,1]$ is defined for all $z\in\Delta^{d-1}$ by
-$G(z):=\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|z^{j}-z^{k}|.$ (9)
-Using $z=z_{\mathrm{cen}}-v_{\psi}(z)\|z_{\mathrm{cen}}-z\|_{1}$, we can express this as
-$G(z)=\frac{\|z_{\mathrm{cen}}-z\|_{1}}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|.$
-Hence, defining for all $w\in\mathcal{W}$ the strictly decreasing and differentiable function $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ ($\mathcal{I}_{w}$ is defined in Proposition 2) for all $\delta\in\mathcal{I}_{w}$ by
-$t_{w}(\delta):=-\frac{\delta}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|w^{j}-w^{k}|,$
-we have that $t_{v_{\psi}(z)}(\|z_{\mathrm{cen}}-z\|_{1})=-G(z)$. Applying Proposition 2 (b), we conclude that
-$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
+$G(z):=\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|z^{j}-z^{k}|,$ (7)
+whence we define a diversity measure by $D(z):=1-G(z)$. Using Proposition S1 (b) (see Example S4 for details), we obtain that
+$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(1-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
 is a derivative-isolating reparametrization for perturbations with direction $v_{\psi}$ and speed given by
@@
 Similar arguments could be applied to other popular measures of diversity such as the Shannon entropy or the Gini–Simpson index.
-2.1.3 Perturbations arising from alternative simplex geometries
-When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
-While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions.
-We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature.
-The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data.
-A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
-$\mathrm{CLR}(z)^{j}:=\log(z^{j})-\frac{1}{d}\sum_{j=1}^{d}\log(z^{j}),\quad\forall j\in[d].$
-This transformation maps $(\Delta^{d-1})^{\circ}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$.
-The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
-$\mathrm{CLR}^{-1}(x)^{j}:=\frac{\exp(x^{j})}{\sum_{j=1}^{d}\exp(x^{j})},\quad\forall j\in[d],$
-or more simply as $\mathrm{CLR}^{-1}(x)=\mathsf{C}(\exp(x))$ where $\exp$ is applied componentwise.
-The inverse of the $\mathrm{CLR}$ transform can in fact be defined on all of $\mathbb{R}^{d}$ rather than just the range of the $\mathrm{CLR}$-function as is clear from the definition.
-The $\mathrm{CLR}$-transformation induces what is called the Aitchison geometry (Egozcue and Pawlowsky-Glahn, 2006) on the simplex.
-The Aitchison geometry induces a different notion of direction than the Euclidean notion considered above and, in particular, the direction of the line connecting $z$ to an endpoint may differ from the direction in (6) (see, for example, Figure 2).
-However, as $\mathrm{CLR}(z)$ is a bijection on the open simplex, we can just as well consider perturbations on $\mathbb{R}^{d}$ directly and use arguments similar to those in Proposition 2 but now for the perturbation on the log-ratio scale.
-We illustrate this procedure in the following two examples and discuss connections to the centering and amalgamating perturbations defined above.
-Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$.
-The types of perturbation are distinguished by color.
-The alternative speed for the diversity perturbation in blue is Gini-speed (see (10)) and for the amalgamating perturbations in red and green the speed is multiplicative speed (see (8)).
-The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds.
-The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
-Example 6 (Diversity in the Aitchison geometry).
-Suppose we are interested in the effect of diversifying $z\in(\Delta^{d-1})^{\circ}$.
-As in Example 5, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space.
-Using the Aitchison geometry, this defines a perturbation via
-$\psi_{\mathrm{Ait}}:(z,\gamma)\mapsto\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z)).$
-We can compute $\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)\bigm{|}_{\gamma=0}$ to find a perturbation $\psi$ of the form (1) for which $\tau_{\psi}=\tau_{\psi_{\mathrm{Ait}}}$.
-Applying the chain rule, we obtain
-$\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
-which can be decomposed into a speed and direction as in Definition 1 to obtain $\psi$.
-The direction of this derivative can be vastly different from the centering direction used in Example 5.
-To obtain a derivative-isolating reparametrization, since $\psi_{\mathrm{Ait}}$ shrinks towards 00 in CLR-space, it is straightforward to show that $\phi^{L}:z\mapsto-\|\mathrm{CLR}(z)\|_{2}$ and $\phi^{W}:z\mapsto\mathrm{CLR}(z)/\|\mathrm{CLR}(z)\|_{2}$ are derivative-isolating reparametrizations.
-Example 7 (Feature influence in the Aitchison geometry).
-Suppose, as in Example 3, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$.
-There is a natural notion of increasing $z^{j}$ inherited from CLR-space, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$ and transforming back to the simplex.
-The resulting perturbation is given by
-$\psi:(z,\gamma)\mapsto\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j}).$
-The derivative of the resulting perturbation is for all $k\in[d]$ given by
-$\partial_{\gamma}\psi(z,\gamma)^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
-It is straight-forward to check that the speed and direction of this perturbation match exactly the multiplicative perturbation derived in Example 3.
-One can therefore view an additive log-ratio perturbation as a multiplicative speed perturbation.
+Figure S1 contains an illustration of the different perturbations we have considered in this work.
 2.2 Perturbation effects on the simplex
-To make the discussion more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1.
+To make our proposal more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1.
 We motivate each perturbation effect with an example application to showcase potential use cases.
-A list summarizing all proposed perturbation effects together with their respective derivative-isolating reparametrizations that are used for efficient estimation is provided in Table 2.
-We encourage practitioners to develop further application-specific perturbations that optimally target the research question they are interested in.
+In Section S2.6 we describe an additional class of perturbation effects based on general amalgamating perturbations.
 2.2.1 Effects of changes in individual components
@@
 For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured.
-We then want to determine for which microbes a high abundance is associated with a high BMI.The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
+We then want to determine for which microbes a high abundance is associated with a high BMI.
+The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
 Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$.
@@
 We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$.
-To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds).
+We already considered these perturbations in Example 1 where we examined two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed.
 $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al.
@@
 Huang et al.
-(2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material).
+(2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S2.8).
 A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$.
@@
 As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response.
-Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (the centering perturbations defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$).
-We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 5 such that $L$ is equal to the negative Gini coefficient.
+Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$).
+We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 2 such that $L$ is equal to one minus the Gini coefficient.
 The construction in Example 2 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics.
@@
 We do not define a binary centering perturbation here as it requires a positive probability of observing $Z=z_{\mathrm{cen}}$, which is rarely the case.
-2.2.3 Effects of amalgamations
-As a third example we consider perturbation effects that capture the influence of general amalgamations.
-As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness).
-Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness.
-As above, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
-However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally.
-For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic.
-To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports.
-Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$.
-We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short.
-As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example 4).
-Analogously, we could be interested in the effect of removing particular activities (e.g.
-gambling) entirely and spending the time on desirable activities (e.g.
-sports) while fixing time spent on a third group of activities (e.g.
-eating, sleeping).
-We can capture this effect using a binary amalgamating perturbation $\psi$.
-We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
 Effect of changes in Target Speed $s(z)$ Endpoint $E(z)$ $\phi^{L}(z)$
@@
 diversity $\mathrm{CDI}_{\mathrm{unit}}$ $1$ $z_{\mathrm{cen}}$ $-\|z_{\mathrm{cen}}-z\|_{1}$
-$\mathrm{CDI}_{\mathrm{Gini}}$ see (10) $z_{\mathrm{cen}}$ $-G(z)$, see (9)
-amalgamations $\mathrm{CAI}^{A\to B}_{\mathrm{unit}}$ $1$ $\mathsf{A}_{A\to B}(z)$ $-2\|z^{A}\|_{1}$
-$\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ $\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}$ $\mathsf{A}_{A\to B}(z)$ $\log\big{(}\tfrac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\big{)}$
-$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
+$\mathrm{CDI}_{\mathrm{Gini}}$ see (8) $z_{\mathrm{cen}}$ $1-G(z)$, see (7)
 Table 1: List of perturbation effects considered in this work and what effects they capture.
@@
 The $W$-term $\phi^{W}(z)$ of the reparametrization is $E(z)$ for the binary effects and $(v_{E}(z),E(z))$ for the directional effects.
-2.3 Perturbations as causal quantities
-Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions.
-To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference.
-Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions.
-In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions.
-The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution.
-Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism.
-To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework (Pearl, 2009; Bongers et al., 2021).
-Similar considerations, however, also apply with the potential outcome framework.
-As an example, consider the SCM over the variables $(H,Z,Y)\in\mathbb{R}^{q}\times\Delta^{d-1}\times\mathbb{R}$ given by
-:$\mathcal{M}:\,\begin{cases}&H\leftarrow\varepsilon_{H}\\ &Z\leftarrow h(H,\varepsilon_{Z})\\ &Y\leftarrow g(H,Z)+\varepsilon_{Y},\end{cases}$
-where $h:\mathbb{R}^{q}\times\mathbb{R}\rightarrow\Delta^{d-1}$ and $g:\mathbb{R}^{q}\times\Delta^{d-1}\rightarrow\mathbb{R}$ are measurable functions and $(\varepsilon_{H},\varepsilon_{Y},\varepsilon_{Z})$ are jointly independent noise terms with expectation zero.
-The SCM $\mathcal{M}$ generates a distribution over $(H,Y,Z)$ which is called the observational distribution.
-Intervening corresponds to changing individual structural assignments while keeping everything else fixed such that the modified SCM again induces a distribution – called the interventional distribution.
-The example SCM $\mathcal{M}$ is a commonly assumed causal structure in which the predictor $Z$ causally precedes the response but is confounded by $H$.
-Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$.
-One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation (Pearl, 2009), this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$.
-For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components.
-Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions.
-Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors.
-More specifically, let $\psi$ be a perturbation as defined in Definition 1.
-Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$.
-In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$.
-Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself.
-The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
-$\lambda_{\psi}^{\operatorname{causal}}:=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,1))\right]-\mathbb{E}\left[Y\right]\big{]}}{\mathbb{P}(Z\neq\psi(Z,1))}$
-and the causal average directional perturbation effect
-$\tau_{\psi}^{\operatorname{causal}}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,\gamma))\right]\big{|}_{\gamma=0}\right].$
-Unlike their non-causal versions, the causal average perturbation effects may not be identifiable from the observational distribution.
-Therefore, to estimate them, we need to make additional causal assumptions.
-For example, we may assume the variables $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$.
-In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
-$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}\big{[}\mathbb{E}[Y\mid Z=\psi(Z,\gamma),H]\big{]},$ (11)
-and hence is identifiable.
-The formula in (11) is a special case of the adjustment formula (e.g., Pearl, 2009), but other ways of identifying this effect also exist (e.g., propensity score matching).
-It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$.
-In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$.
-This is for example done in the numerical experiments in Section 4.1.
-Finally, let us shortly discuss under which assumptions the causal and non-causal average perturbation effects coincide. Obviously, one sufficient condition is to assume that
-$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}[Y\mid Z=\psi(Z,\gamma)].$ (12)
-Such a condition holds by definition whenever $Z$ is an exogenous variable (i.e., does not depend on other variables in its structural assignment) but can also hold more generally.
-A common assumption to ensure (12) is to assume that $Z$ and $Y$ are not confounded, that is, $Z$ and $Y$ have no common causes (this is known as ignorability in the potential outcome framework).
-In the SCM $\mathcal{M}$ this would for example mean that $H$ does not exist or either $g$ or $h$ do not depend on $H$.
-In many cases, this is not satisfied and one should instead carefully argue for which causal assumptions are reasonable and adjust the estimation accordingly, e.g., by conditioning on additional covariates as in (11).
 3 Estimating perturbation effects
-In this section, we describe how to estimate and test hypotheses about the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$.
+In this section, we describe how to estimate and construct confidence intervals for the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$.
 We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice.
@@
 observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$.
-We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
+We present an overview of the different estimators in this section and provide further details of the algorithms and theory in Section S4.1.
 3.1 Efficient estimation using semiparametric theory
-We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient tests and how to remedy this.
+We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient confidence intervals and how to remedy this.
 This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023).
@@
 hence does not provide valid inference (see Figure 1 (left)) as the confidence intervals can be biased or too narrow.
-Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S4.2 of the supplementary material with $d=15$ and $n=1000$.
+Figure 1: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S5.2 with $d=15$ and $n=1000$.
 The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter.
@@
 In this section, we present our proposals for estimating perturbation effects.
-We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here.
+Further details and full algorithms can be found in Section S4.1.
 All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
@@
 $\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (10)
-A full algorithm is given in Algorithm 2.
-It is worth noting that this estimator is sensitive to the estimation of the propensity score $\pi$ which can be difficult when the dimension of $W$ is large or the relationship between $L$ and $W$ is complicated (see Section S3.1 of the supplementary material for an illustration of this).
 3.2.2 Nonparametric estimation of the directional perturbation effect
@@
 where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$.
-A full algorithm for the estimation is given in Algorithm 1 below.
-Both the estimation of $f$ with computable derivative $\partial_{\ell}f$ and $\rho$ can be challenging and we show in Section 4 that the overall estimator (15) is sensitive to both estimates (see Section S3.1 of the supplementary material for an illustration of this).
 For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate).
@@
 Here, we use an alternative smoothing method which uses random forests and local polynomials.
-A full description of our smoothing method is provided in Section S3.3 of the supplementary material.
 Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model.
-This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1).
-Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplementary material.
+This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1).
+Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.5 and S4.6, respectively.
 3.2.3 Partially linear estimation of perturbation effects
-While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions.
+While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.3 for an illustration of this sensitivity).
 One way to avoid these complicated estimations is by additionally assuming a partially linear model.
@@
 where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$.
-A full algorithm is given in Algorithm 3.
 In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions.
@@
 Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above.
-In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S1 in Section S3.1.3 of the supplementary material for a proof of this fact).
+In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S4 in Section S4.3.3 for a proof).
 4 Numerical experiments
 We now present numerical experiments illustrating the usage and performance of our methods.
-In Section 4.1.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census.
-In Section 4.1.2, we analyze the effect of racial diversity on grades in New York schools.
-In Section 4.2 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data.
-In addition to the experiments included here, Section S4.2 of the supplementary material contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
+In addition to the experiments included here, Section S5.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
 The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa.
@@
 Importantly, even for such an individual-level analysis, the simplex confounding can bias the results in the same way as we illustrate here.
-We provide an example of this bias occurring on the individual-level effect in Section S4.5 of the supplementary material.
+We provide an example of this bias occurring on the individual-level effect in Section S5.5.
 We group all individual-level observations into groups of around $50$ individuals.
@@
 To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2).
-We construct these categories in two different ways:
-  • Grouping by race and education: Groups in category 1 are sampled to be similarly diverse and with lower educational levels than the total population and category 2 groups are sampled to be more diverse and with higher educational levels than the total population.
+We construct these categories based on either education or compensation as follows: Groups in category 1 are sampled to be similarly diverse and with lower educational (compensation) levels than the total population and category 2 groups are sampled to be more diverse and with higher educational (compensation) levels than the total population.
 The groups in category 0 consist of groups constructed randomly from the remaining individuals.
-  • Grouping by race and compensation: Groups in category 1 are sampled to be similarly diverse and with lower average compensation than the total population and category 2 groups are sampled to be more diverse and with higher average compensation than the total population.
-The groups in category 0 consist of groups constructed randomly from the remaining individuals.
-We then sample groups from each category, using a weighted sampling approach described in detail in Section S4.3 of the supplementary material.
+We then sample groups from each category, using a weighted sampling approach described in detail in Section S5.3.
 This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation.
-For the data where we have grouped on education the racial composition for the resulting groups is illustrated in the left plot of Figure 4.
 We expect that – by construction – there is a positive effect of increased diversity on compensation because education is positively associated with compensation.
-This can be seen in the right plot of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
-The same occurs when grouping on compensation and the corresponding figure can be found in Section S4.4 of the supplementary material.
-For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 4.
-In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation).
-This can be seen in the lower half of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
-Figure 4: Summary of the semi-synthetic ‘Adult’ dataset when grouping on education.
-Left: Racial compositions in a ternary plot.
-Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
-Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
-Depending on the color the dependence between $Y$ and the Gini coefficient changes.
-The Gini coefficient represents a similarity rather than a diversity measure.
+We visualize the distribution of the data and this confounding in Section S5.4.
 Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g.
@@
 (2018)) with the perturbation-based approach.
-The naive approach summarizes racial diversity in a single measure, here we use minus the Gini coefficient $L=-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 5, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$.
+The naive approach summarizes racial diversity in a single measure, here we use one minus the Gini coefficient $L=1-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 2, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$.
 We call the method that regresses $Y$ on $L$ linearly naive_diversity.
 This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption.
-The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3 where random forests are used for nuisance function estimation when including additional covariates.
+The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 3.2.3) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates.
 The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect.
@@
 This is also visible in the large resulting confidence intervals.
-While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded (e.g., educational levels may well be correlated with race and average compensation in actual communities).
+While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded e.g., educational levels may well be correlated with race and average compensation in actual communities.
 These pitfalls can be avoided with a perturbation-based analysis.
@@
 $\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
-$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.074$ $(-0.046,0.194)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
 Table 2: Table of estimates in the ‘Adult’ data experiment.
@@
 In this section we analyze data from a 2018 Kaggle competition containing data on schools in New York (PASSNYC, 2018).
-The data (collected in 2016) contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S4.6 of the supplementary material).
+The data contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S5.6).
 We let $X$ denote all background variables except racial composition which we denote by $Z$.
 Race is given as whole percentages of ‘Asian’, ‘Black’, ‘Hispanic’ and ‘White’ students which we rescale so that $Z\in\Delta^{3}$.
-We are interested in the effect of diversity, again measured by the Gini coefficient, on the average grade of the schools (either ELA or math).
+We are interested in the effect of diversity, again measured by one minus the Gini coefficient, on the average grade of the schools (either ELA or math).
 We repeat the analysis in Section 4.1.1 on this real data and the results can be seen in Table 3.
@@
 […]
-$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.345}}$ $(0.105,0.585)$ $0.098$ $(-0.175,0.371)$
-$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.163}}$ $(0.002,0.323)$ $0.047$ $(-0.167,0.261)$
+$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.348}}$ $(0.108,0.589)$ $0.101$ $(-0.171,0.373)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.161}}$ $(0.002,0.320)$ $0.042$ $(-0.171,0.255)$
 Table 3: Table of estimates in the New York schools data experiment.
@@
 We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures).
-To investigate how the log-contrast approach compares to the perturbation-based approach, we plot the most significant hits using $\mathrm{CFI}_{\mathrm{mult}}$ (with random forests as regression method) and a log-contrast regression model in the top row of Figure 6.
-These are comparable quantities since $\mathrm{CFI}_{\mathrm{mult}}^{j}$ is equal to the $j$th coefficient in a log-contrast model when it is correctly specified and no zeros are present.
-We omit plotting the $\ell^{1}$-penalized log-contrast estimates as these are broadly similar to the standard log-contrast estimates in this example.
-We also plot a marginal effect of $L=\log(Z^{j}/(1-Z^{j}))$ which corresponds to predicting $Y$ using OLS on $L$ and correcting for the presence of zeros as we mention in Remark 1.
-The results with corresponding $95\%$-confidence intervals are given in the top row of Figure 6; top $10$ most significant species sorted according to significance of $\mathrm{CFI}_{\mathrm{mult}}$ (left) and according to log-contrast coefficients (right).
-Figure 6: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (top left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ (top right).
-The labels are colored by the size of the $p$-value.
-The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model.
-The bottom left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability).
-We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros.
-The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the top right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments.
-We see that the relative influences of the different species is highly contingent on the pseudocount.
-While the effects overlap for some of the features, there are also significant differences.
-To further illustrate these differences, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the bottom left of Figure 6.
+To illustrate the difference between the approaches, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the left of Figure 3.
 From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species.
@@
 The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species.
-To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the bottom right of Figure 6.
+To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the right of Figure 3.
 This further highlights the dependence of the log-contrast coefficients on the used pseudocount.
 The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach.
-In Section S4.7 of the supplementary material we provide an additional plot like the top row of Figure 6 which illustrates the significant hits of $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
-5 Theoretical analysis of estimators
-In this section we prove asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (16).
-These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere (e.g., Klyne and Shah, 2023; Chernozhukov et al., 2018).
-We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.
-We start from a given perturbation $\psi$ and a corresponding reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ which reparametrizes $Z$ into $(L,W):=(\phi^{L}(Z),\phi^{W}(Z))$ and assume that $(L,W)$ satisfy (2) if $\psi$ is a binary perturbation and that $\phi$ is derivative-isolating if $\psi$ is a directional perturbation.
-Our goal is now to use the estimation procedures discussed in Section 3 to estimate the average perturbation effects (i.e., $\tau_{\psi}$ or $\lambda_{\psi}$) consistently and construct asymptotically valid confidence intervals for the estimates.
-To construct asymptotically valid confidence intervals, we estimate both the target parameter and its variance and then show that the resulting standardized estimators are asymptotically normal with a uniform convergence rate over the class of possible distributions of the data.
-Distributionally uniform results are desirable here as they ensure uniform coverage guarantees of the confidence intervals.
-The following result shows uniform asymptotic normality of each of the three estimation procedures.
-Theorem 1 (Uniform asymptotic normality).
-Let $\mathcal{P}$ denote a class of distributions of $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ and assume we are in one of the following settings.
-  1. (a)
-    Nonparametric directional effect estimation: Assumption 1 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\tau},\widehat{\sigma}^{2}_{\tau})$ from Algorithm 1.
-  2. (b)
-    Nonparametric binary effect estimation: Assumption 2 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\lambda},\widehat{\sigma}^{2}_{\lambda})$ from Algorithm 2.
-  3. (c)
-    Effect estimation using a partially linear model: Assumption 3 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\theta},\widehat{\sigma}^{2}_{\theta})$ from Algorithm 3.
-Then, letting $\Phi$ denote the cumulative distribution function of the standard normal distribution, it holds that
-$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widehat{\sigma}_{\beta}}(\widehat{\beta}-\beta_{P})\leq x\right)-\Phi\left(x\right)\right|=0.$
-The result is proven in Section S5 of the supplementary material.
-In the following three sections, we go over each of the settings (a), (b) and (c) individually and provide the required assumptions and detailed algorithms for the target parameter and variance.
-For the remainder of this section, we omit $\psi$-subscripts for simplicity and use the following notation to help us express the uniformly asymptotic results rigorously.
-Let $(X_{P,n})_{n\in\mathbb{N},P\in\mathcal{P}}$ denote a family of sequences of random variables indexed by $\mathcal{P}$.
-For a nonnegative sequence $(a_{n})_{n\in\mathbb{N}}$ we write $X_{P,n}=o_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ it holds that
-$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}\epsilon)=0,$
-that is, $X_{P,n}/a_{n}$ converges to 00 in probability uniformly over $\mathcal{P}$.
-We also write $X_{P,n}=O_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ there exist $M_{\epsilon}>0$ and $N_{\epsilon}\in\mathbb{N}$ such that
-$\sup_{n\geq N_{\epsilon}}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}M)<\epsilon,$
-that is, $X_{P,n}/a_{n}$ is bounded in probability for large $n$.
-When we write conditional expectations given a fitted regression function $\widehat{f}$ we mean a conditional expectation given the random variables used to fit $\widehat{f}$ and any additional randomness involved in the fitting.
-5.1 Nonparametric estimation for directional effects
-We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem 1 (a).
-A high-level overview of this estimation was given in Section 3.2.2.
-More formally, assume we have a dataset of $n$ i.i.d.
-copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S3.4 in the supplementary material).
-We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm 1 using a user-specified number of splits $K$ in the cross-validation.
-Algorithm 1 Nonparametric estimation of $\tau$
-1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, differentiable regression method $\widehat{f}$, score estimation method $\widehat{\rho}$ and nuisance function estimate $\widehat{\eta}$.
-2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
-3:For $k\in[K]$, split $I_{k}$ into $I_{k,0}$ and $I_{k,1}$ of (roughly) equal size and set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
-4:for $k=1,\dots,K$ do
-5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
-6:     Fit nuisance function estimates $\widehat{\eta}_{k}$ using $I_{-k}$.
-7:     for $r=0,1$ do
-8:         Fit $\widehat{\rho}_{k,r}$ using $I_{k,(1-r)}$ and the nuisance function estimate $\widehat{\eta}_{k}$.
-9:         Set $\widehat{\tau}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}$.
-10:         Set $\widehat{\nu}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}\right)^{2}$.
-11:return $\widehat{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\tau}_{k,r}$ and $\widehat{\sigma}^{2}_{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\nu}_{k,r}-\widehat{\tau}^{2}$.
-To obtain the result in Theorem 1 (a), we require that the estimation methods $\widehat{f}$ and $\widehat{\rho}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
-Assumption 1.
-Assume that for all $P\in\mathcal{P}$ and $w\in\mathcal{W}$, the conditional distribution of $L$ given $W=w$ has a density with respect to Lebesgue measure and denote it by $p_{P}(\cdot\,|\,w)$.
-For all $P\in\mathcal{P}$ define $f_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and assume that for almost all $(\ell,w)\in\phi(\mathcal{Z})$ both $p_{P}(\cdot\,|\,w)$ and $f_{P}(\cdot,w)$ are differentiable at $\ell$.
-Define $\tau_{P}:=\mathbb{E}_{P}[\partial_{\ell}f_{P}(\ell,W)\bigm{|}_{\ell=L}]$ and $\rho_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $\rho_{P}(\ell,w):=\partial_{\ell}\log(p_{P}(\ell\,|\,w))$.
-Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{3n},L_{3n},W_{3n})$ denote $3n$ i.i.d.
-copies of $(Y,L,W)$ and define the training datasets $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$ and :=$\mathcal{B}:=\{(Y_{2n+1},L_{2n+1},W_{2n+1}),\dots,(Y_{3n},L_{3n},W_{3n})\}$.
-Let $\widehat{f}$ denote a differentiable estimator of $f_{P}$ based on $\mathcal{A}$, let $\widehat{\eta}$ denote an estimator of a nuisance function relevant for the estimation of $\rho_{P}$ based on $\mathcal{A}$ and let $\widehat{\rho}$ denote an estimator of $\rho_{P}$ based on $\mathcal{B}$ and using the nuisance function estimate $\widehat{\eta}$.
-Finally, assume there exist constants $\delta,c,C>0$ such that
-  1. (a)
-    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}[|Y|^{4+\delta}],\mathbb{E}_{P}[|\rho_{P}(L,W)|^{4+\delta}],\mathbb{E}_{P}[|\partial_{\ell}f_{P}(L,W)|^{2+\delta}],\mathrm{var}_{P}(Y\,|\,L,W)\}\leq C$
-  2. (b)
-    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,L,W),\mathbb{E}_{P}[\rho_{P}(L,W)^{2}]\}\geq c$
-  3. (c)
-    $\mathbb{E}_{P}\left[\left(\widehat{f}(L_{1},W_{1})-f_{P}(L_{1},W_{1})\right)^{4}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
-  4. (d)
-    $\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L_{1},W_{1})-\partial_{\ell}f_{P}(L_{1},W_{1})\right)^{2}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
-  5. (e)
-    $\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right)^{2}=o_{\mathcal{P}}(1)$
-  6. (f)
-    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
-5.2 Nonparametric estimation for binary effects
-We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem 1 (b).
-A high-level overview of this estimation was provided in Section 3.2.1.
-More formally, assume we have a dataset of $n$ i.i.d.
-copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$.
-We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm 2 using a user-specified number of splits $K$ in the cross-validation.
-Algorithm 2 Nonparametric estimation of $\lambda$
-1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression method $\widehat{f}$ and propensity score estimator $\widehat{\pi}$.
-2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
-3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
-4:for $k=1,\dots,K$ do
-5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
-6:     Regress $L$ on $W$ using $I_{-k}$ yielding propensity score function $\widehat{\pi}_{k}$.
-7:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}$.
-8:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}\right)^{2}$.
-9:Set $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}1-L_{i}$.
-10:Set $\widehat{\kappa}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}$.
-11:return $\widehat{\lambda}:=\widehat{p}^{-1}\widehat{\kappa}$ and $\widehat{\sigma}^{2}_{\lambda}:=\widehat{p}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}-\widehat{\kappa}^{2}\right)-\widehat{p}^{-3}\widehat{\kappa}^{2}(1-\widehat{p})$.
-To obtain the result in Theorem 1 (b), we require that the estimation methods $\widehat{f}$ and $\widehat{\pi}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
-Assumption 2.
-For all $P\in\mathcal{P}$ define $f_{P}:\{0,1\}\times\mathcal{W}\to\mathbb{R}$ and $\pi_{P}:\mathcal{W}\to[0,1]$ for all $(\ell,w)\in\{0,1\}\times\mathcal{W}$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and $\pi_{P}(w):=\mathbb{P}_{P}(L=1\,|\,W=w)$, respectively.
-Define $p_{P}:=\mathbb{P}_{P}(L=0)$ and $\lambda_{P}:=p_{P}^{-1}\left(\mathbb{E}_{P}[\mathbb{E}_{P}[Y\,|\,L=1,W]]-\mathbb{E}_{P}[Y]\right)$.
-Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d.
-copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$.
-Let $\widehat{f}$ denote an estimator of $f_{P}$ based on $\mathcal{A}$ and let $\widehat{\pi}$ denote an estimator of $\pi_{P}$ also based on $\mathcal{A}$.
-Finally, assume there exist constants $\delta,c,C>0$ such that
-  1. (a)
-    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|Y|^{2+\delta}\right],\mathrm{var}_{P}(Y\,|\,W,L)\}\leq C$
-  2. (b)
-    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,W,L),p_{P},1-\pi_{P}(W),\pi_{P}(W),\widehat{\pi}(W)\}\geq c$
-  3. (c)
-    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{\pi}(W_{i})-\pi_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
-  4. (d)
-    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}(1,W_{i})-f_{P}(1,W_{i}))^{2}=o_{\mathcal{P}}(1)$
-  5. (e)
-    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}(W_{i})-\pi_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
-5.3 Partially linear model for binary/directional effects
-Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem 1 (c).
-A high-level overview of this estimation was provided in Section 3.2.3.
-More formally, assume we have a dataset of $n$ i.i.d.
-copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively.
-We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm 3 using a user-specified number of splits $K$ in the cross-validation.
-Algorithm 3 Partially linear estimation of $\theta$ (where $\theta=\lambda$ or $\theta=\tau$)
-1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression methods $\widehat{g}$ and $\widehat{m}$.
-2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
-3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
-4:for $k=1,\dots,K$ do
-5:     Regress $Y$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{g}_{k}$.
-6:     Regress $L$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{m}_{k}$.
-7:     Set $\widehat{J}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}$.
-8:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}$.
-9:Set $\widehat{J}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{J}_{k}$
-10:Set $\widehat{\theta}:=\widehat{J}^{-1}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}\right)$.
-11:for $k=1,\dots,K$ do
-12:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}-\widehat{\theta}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}\right)^{2}$
-13:return $\widehat{\theta}$ and $\widehat{\sigma}^{2}_{\theta}:=\widehat{J}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}\right)$.
-To obtain the result in Theorem 1 (c), we require that the estimation methods $\widehat{m}$ and $\widehat{g}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
-Assumption 3.
-Assume that for all $P\in\mathcal{P}$ there exist $\theta_{P}\in\mathbb{R}$ and $h_{P}:\mathcal{W}\to\mathbb{R}$ such that
-$\mathbb{E}_{P}[Y\,|\,L,W]=\theta_{P}L+h_{P}(W).$
-For all $P\in\mathcal{P}$ define $m_{P}:\mathcal{W}\to\mathbb{R}$ and $g_{P}:\mathcal{W}\to\mathbb{R}$ for all $w\in\mathcal{W}$ by $m(w):=\mathbb{E}_{P}[L\,|\,W=w]$ and $g_{P}(w):=\mathbb{E}_{P}[Y\,|\,W=w]=m_{P}(w)+\theta_{P}m_{P}(w)$, respectively.
-Define $\varepsilon:=Y-\mathbb{E}_{P}[Y\,|\,L,W]$ and $\xi:=Y-\mathbb{E}_{P}[L\,|\,W]$.
-Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d.
-copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$.
-Let $\widehat{m}$ denote an estimator of $m_{P}$ based on $\mathcal{A}$ and let $\widehat{g}$ denote an estimator of $g_{P}$ also based on $\mathcal{A}$.
-Suppose that there exist constants $\delta,c,C>0$ such that
-  1. (a)
-    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|\varepsilon\xi|^{2+\delta}\right]\mathbb{E}_{P}\left[|\xi|^{4}\right],\mathrm{var}_{P}(Y\,|\,W),\mathrm{var}_{P}(L\,|\,W)\}\leq C$
-  2. (b)
-    for all $P\in\mathcal{P}:\quad$ $\min\{\mathbb{E}_{P}\left[\xi^{2}\right],\mathbb{E}_{P}\left[\varepsilon^{2}\xi^{2}\right]\}\geq c$
-  3. (c)
-    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{2}=o_{\mathcal{P}}(n^{-1/2})$
-  4. (d)
-    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{4}=o_{\mathcal{P}}(1)$
-  5. (e)
-    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{g}(W_{i})-g_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
-  6. (f)
-    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{m}(W_{i})-m_{P}(W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{g}(W_{i})-g_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
+In Section S5.7 we provide additional plots which compare the significant hits of $\mathrm{CFI}_{\mathrm{mult}}$, log-contrast estimates and a naive marginal analysis and illustrate the differences between $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
+Figure 3: The left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability).
+We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros.
+The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments.
+We see that the relative influences of the different species is highly contingent on the pseudocount.
 5 Discussion
@@
 Appendix S1 Differentiability on the simplex
-To define what we mean by differentiability on the simplex we use the theory of smooth manifolds [see Lee, 2012, for a comprehensive treatment of this theory].
+In this section we provide the background needed to define differentiability of a function defined on the simplex, which we use in Section 2, by means of the theory of smooth manifolds [see Lee, 2012, for a comprehensive treatment of this theory].
 The simplex $\Delta^{d-1}$ can be viewed as a smooth $(d-1)$-dimensional manifold with corners [Lee, 2012, p.
@@
 It is easy to check that if $f:\Delta^{d-1}\to\mathbb{R}$ is differentiable at $z$ then for any $j\in[d]$ with $z\in\mathcal{Z}_{j}$ the function $f\circ\varphi_{j}^{-1}$ must be differentiable at $\varphi_{j}(z)$.
-With these tools from differential geometry in place, we will now prove Proposition 1.
-Appendix S2 Proofs of results from Section 2
-S2.1 Proof of Proposition 1
+We will use these tools to prove results about perturbations below.
+Appendix S2 Perturbations: background, theory and examples
+In this section we provide additional details on perturbations and perturbation effects.
+Section S2.1 provides background on the history of effect measures and compositional data.
+Section S2.2 contains two examples to motivate why applying standard methods to compositional predictors can be misleading.
+Section S2.3 discusses well-definedness of the perturbation effects defined in Definition 2.
+Section S2.4 contains a proof of Proposition 1.
+Section S2.5 contains a general result on the construction of a derivative-isolating reparametrization and examples of the use of the result.
+Section S2.6 includes a generalization of the ‘individual component’ effect measures that we presented in Section 2.
+Section S2.7 investigates how perturbations can be derived starting from the Aitchison geometry and how such perturbations compare to the ones we construct in Section 2.
+Section S2.8 contains two results relating our average perturbation effects to well-known classical effect measure.
+S2.1 Historical review of effect measures for compositional data
+In this section we provide historical context for our proposal by summarizing relevant developments from several different statistical fields related to our proposal.
+Our proposed perturbation-based framework is perhaps most closely related to the literature on causal inference, semiparametric inference and targeted learning [Bickel et al., 1993, Tsiatis, 2006, van der Laan and Rose, 2011].
+Spirtes and Scheines [2004] discuss some of the problems of modelling causal effects when certain variables are deterministic functions of other variables.
+This discussion was later continued by Woodward [2015] and new causal models have been developed to address this problem [Beckers et al., 2023].
+We contribute to these developments by providing a novel framework for causal effect estimation when some covariates are compositional (see Section S3).
+An important aspect of causal effect estimation more generally is the need for an interpretable summary measure of the effect.
+Historically, such interpretable measures originated from parametric models that were easy to interpret.
+Much work has since gone into the interpretation of and uncertainty quantification for such parameters when the model is misspecified.
+Perhaps one of the most impactful such analyses was given by White [1980] who provided an estimation procedure for the coefficients in a linear model that did not require the model to be correctly specified.
+To what extent parameters based on linear models are useful and interpretable when the model is misspecified remains an active research topic particularly in econometrics [Aronow and Samii, 2016, Graham and de Xavier Pinto, 2022, Słoczyński, 2022].
+More recently, interpretable parameters based on partially linear models have been proposed and analyzed extensively [Chernozhukov et al., 2018, Vansteelandt and Dukes, 2022].
+Despite these and similar parameters depending on potentially complicated nuisance functions, it is often possible to estimate them efficiently under relatively mild conditions, as we also do here, by utilizing semiparametric theory [Kennedy, 2023].
+Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ [Claringbold, 1955, Cornell, 2002].
+While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting.
+Scheffé [1958] pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials.
+Becker [1968] critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks.
+Cox [1971] argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g.
+so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$.
+It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S3 in Section S2.8 of the supplementary material).
+Draper and John [1977] developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary.
+Aitchison and Bacon-Shone [1984] introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients.
+Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ [Lin et al., 2014, Simpson et al., 2021].
+One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 1), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S2 in Section S2.8 of the supplementary material or Huang et al.
+[2023] where this was first shown).
+Defining practical predictive models for compositional data remains an active topic of research [Xiong et al., 2015, Fiksel et al., 2022, Li et al., 2023, Bhaduri et al., 2025].
+Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis.
+It is well-known that many conventional statistical procedures become invalid when applied to compositional data.
+For example, already Pearson [1897] drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint.
+The most prevalent class of general methods is based on the seminal paper by Aitchison [1982] which introduces the log-ratio approach.
+In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms [Aitchison, 1982, 1986, Egozcue et al., 2003] and then performing the statistical analysis in the transformed space.
+Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section S2.7).
+The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex.
+This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes [e.g., Martín-Fernández et al., 2000, 2003, 2015, Palarea-Albaladejo and Martín-Fernández, 2008, Lubbe et al., 2021] that can be used before applying the log-ratio approach.
+Greenacre et al.
+[2023] provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches which are further highlighted in Scealy and Welsh [2014].
+While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
+There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations [Greenacre, 2009] or spherical transformations [Scealy and Welsh, 2011, Li and Ahn, 2022], or by appealing to more general notions of object-oriented data analysis [Marron and Dryden, 2021].
+There has also been much recent work on manifold-valued data where compositional data is a special case [Dai and Müller, 2018, Dai et al., 2021, Petersen et al., 2022, Kurisu et al., 2024].
+S2.2 Illustrative examples
+In this section we provide two helpful examples meant to motivate the need to account for the compositional nature of covariates. The examples are meant to supplement the development in Section 1.
+Example S1 (Effects of presence or absence of microbes).
+Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease.
+Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances.
+Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ ($1$ encoding disease) has been recorded.
+While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies.
+One possible approach [e.g., Hu et al., 2021] would be to consider a fixed microbe $j$, define the treatment variable $T:=\mathds{1}_{\{Z^{j}=0\}}$ corresponding to the presence or absence of microbe $j$ and then estimate a marginal treatment effect given by
+$\mathrm{MTE}:=\mathbb{E}[Y\,|\,T=1]-\mathbb{E}[Y\,|\,T=0]=\mathbb{P}(Y=1\,|\,T=1)-\mathbb{P}(Y=1\,|\,T=0),$
+where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$.
+It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components.
+[^1] This is true when $Z\in\mathbb{R}^{d}$ is unconstrained, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$.
+To illustrate this, assume the disease indicator $Y$ given the composition $Z$ has conditional distribution for all $z\in\Delta^{d-1}$ given by a Bernoulli distribution with
+$\mathbb{P}(Y=1\,|\,Z=z)=\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}$
+that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ reduces the probability of having the disease.
+In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table S1 (details are provided in Section S5.1 of the supplementary material).
+If $Z\in\mathbb{R}^{d}$ was unconstrained, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
+$h:z^{-j}\mapsto\mathbb{E}[Y\,|\,T=1,Z^{-j}=z^{-j}]-\mathbb{E}[Y\,|\,T=0,Z^{-j}=z^{-j}]$
+and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$.
+[^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE).
+Estimating this quantity is a well-studied problem, e.g.
+using the AIPW estimator [Robins et al., 1994] or using a partially linear model [Chernozhukov et al., 2018].
+Unfortunately, neither of the estimators is well defined when $Z\in\Delta^{d-1}$ as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y$ is conditionally independent of $T$ given $Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$.
+In the Table S1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
+Example S1
+method estimate 95%-CI
+$\mathrm{MTE}$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.12}}$ $(-0.17,-0.06)$
+$\mathrm{CKE}^{1}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.15}}$ $(0.10,0.21)$
+Example S2
+method estimate 95%-CI
+OLS of $Y$ on $-G(Z)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-1.57}}$ $(-2.18,-0.97)$
+$\mathrm{CDI}_{\mathrm{Gini}}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.63}}$ $(0.14,1.12)$
+Table S1: Numerical illustration of Examples S1 and S2.
+The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model.
+The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators.
+Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect.
+Left: The estimated MTE is negative, even though the true effect is positive.
+The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign.
+Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction.
+The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect.
+Details on both experiments are provided in Section S5.1 of the supplementary material.
+Example S2 (Effects of increased diversity).
+Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$.
+Such a question is of immediate interest in many applied fields, e.g.
+when investigating whether more racially diverse university programs improve educational outcomes [Antonio et al., 2004] or whether increased income equality boosts economic output [Shin, 2012].
+The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$.
+As in Example S1 this type of marginal dependence analysis is potentially misleading in practice.
+In Table S1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S5.1 of the supplementary material).
+To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$.
+However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable.
+The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table S1.
+A more in-depth example in which the marginal approach is misleading is given in Section 4.1.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data.
+A real application is given in Section 4.1.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York.
+Unlike in Example S1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained data.
+Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
+S2.3 Well-definedness of perturbation effects
+In this section, we argue that the perturbation effects we define in Definition 2 are well-defined.
+This is immediate for the directional effect and similarly it is clear why we need $\mathbb{P}(Z\neq\psi(Z,1))>0$ for the binary effect.
+To see why we need that $z\mapsto\psi(z,1)$ preserves null sets, we can argue as follows.
+If $z\mapsto\psi(z,1)$ does not preserve null sets, there exists a set $A\subseteq\mathcal{Z}$ such that $\mathbb{P}(Z\in A)=0$ but $\mathbb{P}(\psi(Z,1)\in A)>0$.
+Let $f_{1}$ be any version of $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ and let $f_{2}:z\mapsto f_{1}(z)+\mathds{1}_{A}(z)$.
+Since $A$ is a null set for the distribution of $Z$, $f_{2}(Z)$ is a valid version of the conditional expectation of $Y$ given $Z$.
+However,
+$\mathbb{E}[f_{1}(\psi(Z,1))]\neq\mathbb{E}[f_{1}(\psi(Z,1))]+\mathbb{P}(\psi(Z,1)\in A)=\mathbb{E}[f_{2}(\psi(Z,1))],$
+so it is necessary for $z\mapsto\psi(z,1)$ to preserve null sets for the binary perturbation effect to be well-defined.
+On the contrary, if $z\mapsto\psi(z,1)$ preserves null sets, then for any two version $f_{1}$ and $f_{2}$ of $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ we have that $\mathbb{P}(f_{1}(Z)=f_{2}(Z))=1$ and thus also $\mathbb{P}(f_{1}(\psi(Z,1))=f_{2}(\psi(Z,1)))=1$.
+This immediately yields that $\mathbb{E}[f_{1}(\psi(Z,1))]=\mathbb{E}[f_{2}(\psi(Z,1))]$, so the binary perturbation effect is well-defined.
+S2.4 Proof of Proposition 1
+In this section we prove Proposition 1 using the tools from Section S1.
 Proof.
@@
 as desired.
-S2.2 Proof of Proposition 2
+S2.5 Finding derivative-isolating reparametrizations
+In this section we provide a general method for finding derivative-isolating reparametrizations based on the development in Section 2.1.2. Our main result is as follows.
+Proposition S1 (Finding derivative-isolating reparametrizations).
+For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$.
+Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$.
+Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=-\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(E(z),v_{E}(z))$ with $v_{E}$ defined in (6).
+Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
+  1. (a)
+    (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function.
+For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
+    $-s_{w}(\delta)^{-1}=\partial_{\delta}t_{w}(\delta).$ (S3)
+    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$ and any other perturbation with the same speed and direction.
+  2. (b)
+    (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic.
+For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing.
+Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
+    $s(z):=-\left(\partial_{\delta}t_{\phi^{W}(z)}(\delta)\big{|}_{\delta=\|E(z)-z\|_{1}}\right)^{-1},$
+    and any other perturbation with the same speed and direction.
 Proof.
-(a) For each $w\in\mathcal{W}$, the speed $s_{w}$ is a positive function, so by (7) $t_{w}$ must be strictly decreasing and hence invertible on $\mathcal{L}_{w}:=t_{w}(\mathcal{I}_{w})$.
+(a) For each $w\in\mathcal{W}$, the speed $s_{w}$ is a positive function, so by (S3) $t_{w}$ must be strictly decreasing and hence invertible on $\mathcal{L}_{w}:=t_{w}(\mathcal{I}_{w})$.
 This lets us compute the inverse of $\phi$ for each $w=(w_{E},w_{v})\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$ as
 $\phi^{-1}(\ell,w)=-t_{w}^{-1}(\ell)w_{v}+w_{E}$ (S4)
-which shows that $\phi$ is indeed a bijection on the image of $\phi$. For all $w\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$ we have by (7) and the inverse function rule for derivatives that
+which shows that $\phi$ is indeed a bijection on the image of $\phi$. For all $w\in\mathcal{W}$ and $\ell\in\mathcal{L}_{w}$ we have by (S3) and the inverse function rule for derivatives that
 $s(\phi^{-1}(\ell,w))=s_{w}(t_{w}^{-1}(\ell))=-\frac{1}{(\partial_{\delta}t_{w})(t_{w}^{-1}(\ell))}=-\partial_{\ell}t_{w}^{-1}(\ell).$
@@
 $s:z\mapsto-\partial_{\ell}t_{(E(z),v_{E}(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v_{E}(z))}(\|E(z)-z\|_{1}))^{-1}.$
-Appendix S3 Estimation technicalities and algorithms
-Section S3.1 provides an example that illustrates how the nuisance functions that need to be estimated for the nonparametric estimators in Section 3 can be more unstable and harder to estimate than those required for the estimators based on a partially linear model.
+Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (S3).
+Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint.
+Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized.
+In the following two examples, we give the missing details from Examples 1 and 2 from Section 2.1.2.
+Example S3.
+Recall the setting of Example 1.
+To find a derivative-isolating reparametrization which corresponds to the speed $s_{\psi}(z)=2z^{j}(1-z^{j})$, we apply Proposition S1 (a).
+Since for all $w\in\mathcal{W}$ it holds that $w_{E}^{j}=1$ (component $w_{E}$ corresponds to the endpoint $E(z)$) and $w_{v}^{j}=\frac{1}{2}$ (component $w_{v}$ corresponds to the direction $v_{E}(z)$ and we used that $v_{E}(z)^{j}=\frac{1-z^{j}}{\|e_{j}-z^{j}\|_{1}}=\frac{1-z^{j}}{2(1-z^{j})}=\frac{1}{2}$), we have that $\phi_{\mathrm{unit}}^{-1}(\ell,w)^{j}=\frac{\ell}{2}+1$.
+Therefore,
+$s_{w}:\delta\mapsto s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))=\frac{(2-\delta)\delta}{2}.$
+Hence, solving (S3) amounts to solving
+$-\frac{2}{(2-\delta)\delta}=\partial_{\delta}t_{w}(\delta)$
+which has solutions of the form $t_{w}(\delta)=\log\left(\frac{2-\delta}{\delta}\right)+C$ for $C\in\mathbb{R}$. For simplicity, we choose $C=0$ and note that as $\|e_{j}-z\|_{1}=2(1-z^{j})$, we have
+$\phi^{W}(z)=(e_{1},v_{E}(z))\quad\text{and}\quad\phi^{L}(z)=t_{\phi^{W}(z)}(\|e_{j}-z\|_{1})=\log\left(\frac{z^{j}}{1-z^{j}}\right)$
+We can therefore conclude (dropping the constant endpoint from the reparametrization) that
+$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{z^{j}}{1-z^{j}}\right),v_{\psi}(z)\right)$
+is a derivative-isolating reparametrization for the $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation $\psi$ with multiplicative speed.
+Example S4.
+Recall the setting of Example 2. Using that $z=z_{\mathrm{cen}}-v_{\psi}(z)\|z_{\mathrm{cen}}-z\|_{1}$, we can express
+$G(z)=\frac{\|z_{\mathrm{cen}}-z\|_{1}}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|.$
+Hence, defining for all $w\in\mathcal{W}$ the strictly decreasing and differentiable function $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ ($\mathcal{I}_{w}$ is defined in Proposition S1) for all $\delta\in\mathcal{I}_{w}$ by
+$t_{w}(\delta):=1-\frac{\delta}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|w^{j}-w^{k}|,$
+we have that $t_{v_{\psi}(z)}(\|z_{\mathrm{cen}}-z\|_{1})=1-G(z)$. Applying Proposition S1 (b), we obtain that
+$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(1-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
+is a derivative-isolating reparametrization for perturbations with direction $v_{\psi}$ and speed given by
+$s_{\psi}(z):=\frac{2d}{\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|}.$
+S2.6 Amalgamations
+In this section we consider general amalgamating perturbations as defined in Section 2.1.1 rather than only those which push towards a vertex of the simplex, i.e., those we describe in Section 2.2.1.
+We first generalize Example 1.
+Example S5 (Amalgamation influence).
+For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation.
+Suppose as in Example 1 that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale.
+We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation, so we suppose now that $A\cup B=[d]$.
+The multiplicative operation described in Example 1 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
+$c\mapsto\mathsf{C}(X\odot(1+cu_{B})),$
+which similarly corresponds to a well-defined perturbation on the simplex. Computing the speed, we obtain that
+$\|\partial_{c}\mathsf{C}(X\odot(1+cu_{B}))\bigm{|}_{c=0}\|_{1}=\frac{2\|X^{A}\|_{1}\|X^{B}\|_{1}}{(\|X^{A}\|_{1}+\|X^{B}\|_{1})^{2}}.$
+We can use this speed even when $A\cup B\neq[d]$ resulting in the general multiplicative speed
+$s_{\psi}(z)=\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}.$ (S5)
+Using Proposition S1 (a) as in Example 1, we further obtain that
+$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\right),(\mathsf{A}_{A\to B}(z),v_{\psi}(z))\right)$
+is a derivative-isolating reparametrization for $\psi$ with multiplicative speed.
+With this generalization, we can consider perturbation effects for general amalgmating perturbations.
+As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness).
+Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness.
+As in Section 2.2.1, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
+However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally.
+For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic.
+To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports.
+Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$.
+We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short.
+As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example S5).
+Analogously, we could be interested in the effect of removing particular activities (e.g.
+gambling) entirely and spending the time on desirable activities (e.g.
+sports) while fixing time spent on a third group of activities (e.g.
+eating, sleeping).
+We can capture this effect using a binary amalgamating perturbation $\psi$.
+We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
+We summarize these target parameters in Table S2 which complements Table 1.
+amalgamations $\mathrm{CAI}^{A\to B}_{\mathrm{unit}}$ $1$ $\mathsf{A}_{A\to B}(z)$ $-2\|z^{A}\|_{1}$
+$\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ $\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}$ $\mathsf{A}_{A\to B}(z)$ $\log\big{(}\tfrac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\big{)}$
+$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
+Table S2: List of amalgamating perturbation effects considered in this work and what effects they capture.
+For each perturbation effect the corresponding speed, endpoint and $L$-term of the reparametrization used for efficient estimation is included.
+All reparametrizations use the canonical direction $v_{E}$ based on the endpoint, see (6).
+The $W$-term $\phi^{W}(z)$ of the reparametrization is $E(z)$ for the binary effects and $(v_{E}(z),E(z))$ for the directional effects.
+S2.7 Perturbations arising from alternative simplex geometries
+In this section we describe an alternative way to construct perturbations than the one considered in Sections 2.1.1 and 2.1.2.
+When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
+While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions.
+We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature.
+The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison [1982] who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data.
+A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
+$\mathrm{CLR}(z)^{j}:=\log(z^{j})-\frac{1}{d}\sum_{j=1}^{d}\log(z^{j}),\quad\forall j\in[d].$
+This transformation maps $(\Delta^{d-1})^{\circ}$ to the subspace of $\mathbb{R}^{d}$ given by $\{x\in\mathbb{R}^{d}\mid\sum_{j=1}^{d}x^{j}=0\}$.
+The inverse CLR is defined for all $x\in\operatorname{Im}(\mathrm{CLR})$ by
+$\mathrm{CLR}^{-1}(x)^{j}:=\frac{\exp(x^{j})}{\sum_{j=1}^{d}\exp(x^{j})},\quad\forall j\in[d],$
+or more simply as $\mathrm{CLR}^{-1}(x)=\mathsf{C}(\exp(x))$ where $\exp$ is applied componentwise.
+The inverse of the $\mathrm{CLR}$ transform can in fact be defined on all of $\mathbb{R}^{d}$ rather than just the range of the $\mathrm{CLR}$-function as is clear from the definition.
+The $\mathrm{CLR}$-transformation induces what is called the Aitchison geometry [Egozcue and Pawlowsky-Glahn, 2006] on the simplex.
+The Aitchison geometry induces a different notion of direction than the Euclidean notion considered above and, in particular, the direction of the line connecting $z$ to an endpoint may differ from the direction in (6) (see, for example, Figure S1).
+However, as $\mathrm{CLR}(z)$ is a bijection on the open simplex, we can just as well consider perturbations on $\mathbb{R}^{d}$ directly and use arguments similar to those in Proposition S1 but now for the perturbation on the log-ratio scale.
+We illustrate this procedure in the following two examples and discuss connections to the centering and amalgamating perturbations defined in Examples 1 and 2.
+Example S6 (Diversity in the Aitchison geometry).
+Suppose we are interested in the effect of diversifying $z\in(\Delta^{d-1})^{\circ}$.
+As in Example 2, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space.
+Using the Aitchison geometry, this defines a perturbation via
+$\psi_{\mathrm{Ait}}:(z,\gamma)\mapsto\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z)).$
+We can compute $\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)\bigm{|}_{\gamma=0}$ to find a perturbation $\psi$ of the form (1) for which $\tau_{\psi}=\tau_{\psi_{\mathrm{Ait}}}$.
+Applying the chain rule, we obtain
+$\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
+which can be decomposed into a speed and direction as in Definition 1 to obtain $\psi$.
+The direction of this derivative can be vastly different from the centering direction used in Example 2.
+To obtain a derivative-isolating reparametrization, since $\psi_{\mathrm{Ait}}$ shrinks towards 00 in CLR-space, it is straightforward to show that $\phi^{L}:z\mapsto-\|\mathrm{CLR}(z)\|_{2}$ and $\phi^{W}:z\mapsto\mathrm{CLR}(z)/\|\mathrm{CLR}(z)\|_{2}$ are derivative-isolating reparametrizations.
+Example S7 (Feature influence in the Aitchison geometry).
+Suppose, as in Example 1, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$.
+There is a natural notion of increasing $z^{j}$ inherited from CLR-space, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$ and transforming back to the simplex.
+The resulting perturbation is given by
+$\psi:(z,\gamma)\mapsto\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j}).$
+The derivative of the resulting perturbation is for all $k\in[d]$ given by
+$\partial_{\gamma}\psi(z,\gamma)^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
+It is straight-forward to check that the speed and direction of this perturbation match exactly the multiplicative perturbation derived in Example 1.
+One can therefore view an additive log-ratio perturbation as a multiplicative speed perturbation.
+We summarize the types of perturbations considered in this work in Figure S1.
+Figure S1: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$.
+The types of perturbation are distinguished by color.
+The alternative speed for the diversity perturbation in blue is Gini-speed (see (8)) and for the amalgamating perturbations in red and green the speed is multiplicative speed (see (S5)).
+The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds.
+The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
+S2.8 Perturbation effects and classical effect measures
+In this section we present two results which connect our perturbation effects to classical effect measures in the literature.
+The first result provides a relationship between the $\mathrm{CFI}^{j}$ and coefficients in a well-specified log-contrast regression model.
+The result has previously been proven in Huang et al.
+[2023].
+Proposition S2.
+Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$.
+Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}\log(z^{j})+\mu$.
+Then, for all $j\in[d]$, we have
+$\mathrm{CFI}^{j}_{\mathrm{mult}}=\beta^{j}.$
+Proof.
+The perturbation corresponding to $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is given by
+$\psi(z,\gamma)=z+\gamma\frac{2z^{j}(1-z^{j})}{\|e_{j}-z\|_{1}}(e_{j}-z)=z+\gamma z^{j}(e_{j}-z).$
+We thus have
+$f(\psi(z,\gamma))=\sum_{k=1}^{d}\beta^{k}\log(z^{k}+\gamma z^{j}(\mathds{1}_{\{k=j\}}-z^{k}))+\mu$
+and therefore
+$\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}=\sum_{k=1}^{d}\frac{\beta^{k}z^{j}(\mathds{1}_{\{k=j\}}-z^{k})}{z^{k}}=\beta^{j}-z^{j}\sum_{k=1}^{d}\beta^{k}=\beta^{j}.$
+We conclude that
+$\mathrm{CFI}^{j}_{\mathrm{mult}}=\tau_{\psi}=\mathbb{E}\left[\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}\right]=\beta^{j}.$
+The second result shows that the effects proposed in Cox [1971] are special cases of average perturbation effects in a parametric model.
+Proposition S3.
+Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $z_{*}\in\Delta^{d-1}$.
+Let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ and suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}z_{*}^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}z^{j}+\mu$.
+For $k\in[d]$, define the perturbation $\psi_{k}$ such that for all components $j\in[d]$ it holds
+$\psi_{k}^{j}(z,\gamma)=\begin{cases}z^{j}+\gamma\quad&\text{if }j=k\\ z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\quad&\text{else}.\end{cases}$
+Then
+$\tau_{\psi_{k}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+Proof.
+We have
+$f(\psi_{k}(z,\gamma))=\beta^{k}(z^{k}+\gamma)+\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\left(z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\right)+\mu$
+and therefore
+$\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}=\beta^{k}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\beta^{k}\left(1+\frac{z^{k}_{*}}{1-z^{k}_{*}}\right)-\sum\limits_{j=1}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+We conclude that
+$\tau_{\psi_{k}}=\mathbb{E}\left[\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}\right]=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+Appendix S3 Perturbations as causal quantities
+Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions.
+To make this distinction clear, we discuss here how our perturbation framework can be used for causal inference.
+Causal models [Pearl, 2009, Rubin, 2005] provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions.
+In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions.
+The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution.
+Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism.
+To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework [Pearl, 2009, Bongers et al., 2021].
+Similar considerations, however, also apply with the potential outcome framework.
+As a first example, assume that the SCM given in Figure S2 (left) generates the data.
+Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 2.1.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
+(a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
+$Y$ $\leftarrow f(Z,\eta)$
+$Y$$Z$$\psi$$f$
+bijection $\phi$$\phi(Z)=(L,W)$$Z=\phi^{-1}(L,W)$
+$Y$$L$$W$$\psi$$h$$m$$h$
+$L$ $\leftarrow m(W,\xi)$
+$Y$ $\leftarrow h(L,W,\varepsilon)$
+Figure S2: (a) Data-generating SCM on the $Z$-scale.
+$Z\in\Delta^{d-1}$ is a compositional vector, $Y$ is a response variable and $\eta$ is exogenous noise.
+(b) Reparametrized SCM on the $(L,W)$-scale.
+$(L,W)=\phi(Z)$ are the reparametrized predictors, $Y$ is the same response as in (a) and $\xi$ and $\varepsilon$ are exogenous noise variables.
+The reparametrization needs to disentangle the perturbation effect on $Z$ in a way such that intervening on $Z$ by setting it to $\psi(Z,\gamma)$ corresponds to intervening on only $L$ (red double head arrows).
+As a second example, consider the SCM over the variables $(H,Z,Y)\in\mathbb{R}^{q}\times\Delta^{d-1}\times\mathbb{R}$ given by
+:$\mathcal{M}:\,\begin{cases}&H\leftarrow\varepsilon_{H}\\ &Z\leftarrow h(H,\varepsilon_{Z})\\ &Y\leftarrow g(H,Z)+\varepsilon_{Y},\end{cases}$
+where $h:\mathbb{R}^{q}\times\mathbb{R}\rightarrow\Delta^{d-1}$ and $g:\mathbb{R}^{q}\times\Delta^{d-1}\rightarrow\mathbb{R}$ are measurable functions and $(\varepsilon_{H},\varepsilon_{Y},\varepsilon_{Z})$ are jointly independent noise terms with expectation zero.
+The SCM $\mathcal{M}$ generates a distribution over $(H,Y,Z)$ which is called the observational distribution.
+Intervening corresponds to changing individual structural assignments while keeping everything else fixed such that the modified SCM again induces a distribution – called the interventional distribution.
+The example SCM $\mathcal{M}$ is a commonly assumed causal structure in which the predictor $Z$ causally precedes the response but is confounded by $H$.
+Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$.
+One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation [Pearl, 2009], this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$.
+For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components.
+Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions.
+Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors.
+More specifically, let $\psi$ be a perturbation as defined in Definition 1.
+Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$.
+In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$.
+Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself.
+The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
+$\lambda_{\psi}^{\operatorname{causal}}:=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,1))\right]-\mathbb{E}\left[Y\right]\big{]}}{\mathbb{P}(Z\neq\psi(Z,1))}$
+and the causal average directional perturbation effect
+$\tau_{\psi}^{\operatorname{causal}}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,\operatorname{do}(Z=\psi(Z,\gamma))\right]\big{|}_{\gamma=0}\right].$
+Unlike their non-causal versions, the causal average perturbation effects may not be identifiable from the observational distribution.
+Therefore, to estimate them, we need to make additional causal assumptions.
+For example, we may assume the variables $(H,Y,Z)$ were generated by an SCM with the same causal structure as $\mathcal{M}$.
+In that case, if $H$ is observed, the expectation of $Y$ under perturbation induced interventions on $Z$ can be expressed as
+$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}\big{[}\mathbb{E}[Y\mid Z=\psi(Z,\gamma),H]\big{]},$ (S6)
+and hence is identifiable.
+The formula in (S6) is a special case of the adjustment formula [e.g., Pearl, 2009], but other ways of identifying this effect also exist (e.g., propensity score matching).
+It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$.
+In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$.
+This is for example done in the numerical experiments in Section 4.1.
+There are several possible assumptions which permit us to conclude that the causal and non-causal perturbation effect coincide. Obviously, one sufficient condition is to assume that
+$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}[Y\mid Z=\psi(Z,\gamma)].$ (S7)
+Such a condition holds by definition whenever $Z$ is an exogenous variable (i.e., does not depend on other variables in its structural assignment) but can also hold more generally.
+A common assumption to ensure (S7) is to assume that $Z$ and $Y$ are not confounded, that is, $Z$ and $Y$ have no common causes (this is known as ignorability in the potential outcome framework).
+In the SCM $\mathcal{M}$ this would for example mean that $H$ does not exist or either $g$ or $h$ do not depend on $H$.
+In many cases, this is not satisfied and one should instead carefully argue for which causal assumptions are reasonable and adjust the estimation accordingly, e.g., by conditioning on additional covariates as in (S6).
+Appendix S4 Inference and algorithms for perturbation effects
+In this section we provide details on the theory and algorithms for the estimation of perturbation effects.
+Section S4.1 contains a summary of the theoretical results and algorithms for the main estimation procedures.
+Section S4.2 contains a proof of our main theoretical result.
+Section S4.3 investigates the robustness of the different estimators.
 Section S4.4 describes how to modify the estimation of $\tau_{\psi}$ using a derivative-isolating reparametrization in the presence of observations with $\omega_{\psi}(Z)=0$.
@@
 Section S4.6 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
-S3.1 Robustness of nuisance function estimation for different estimators
-A key motivation for applying the estimators based on a partially linear model on $Y$ is to avoid estimating nuisance functions that result in unstable estimators.
-We illustrate this point based on three examples below.
-S3.1.1 Estimation of nuisance functions with continuous $L$
-To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section 3.2.2) compared to $m$ and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
-We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
-$L:=W^{1}+(1+\mathds{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
-The out-of-sample predictions evaluated on a new dataset with $n=1000$ observations can be seen together with the true function in Figure S1.
-Figure S1: Estimation of nuisance functions on the simulated data in Section S3.1.1.
-The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size.
-We see that the $m$ and $g$ estimates are reasonable estimates of the true functions, that roughly capture the behavior of the true function in an unbiased (conditioned on $W^{1}$) way.
-In contrast, the derivative and score estimates, are much worse at capturing the more complex behavior (with clear biases conditioned on $W^{1}$).
-S3.1.2 Estimation of nuisance functions with binary $L$
-To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section 3.2.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
-We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
-$Y:=2LW^{1}+\varepsilon.$
-The out-of-sample predictions evaluated on new dataset with $n=1000$ observations can be seen together with the true function in Figure S2.
-Figure S2: Estimation of nuisance functions on the simulated data in Section S3.1.2.
-The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size.
-We see that the $m$, $g$ and $f$ estimates are reasonable estimates of the true functions whereas the inverse propensity (even when cutting off extreme values at $100$) is poorly estimated for $W^{1}$ close to zero.
-S3.1.3 Partially linear model estimate when partial linearity is violated
-The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified.
-The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting).
-We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section S4.2 which we do in Figure S3.
-Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
-Proposition S1.
-Let $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ denote random variables where $\max\{\mathbb{E}[L^{2}],\mathbb{E}[Y^{2}]\}<\infty$. Then,
-$\theta:=\frac{\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]}{\mathbb{E}[\mathrm{var}(L\,|\,W)]}\quad\text{and}\quad h:w\mapsto\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$
-are solutions to the minimization problem
-$\min_{\theta\in\mathbb{R},h:\mathcal{W}\to\mathbb{R}}\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W)\right)^{2}\right].$ (S4)
-Proof.
-First, we note that instead of the optimization in (S4) we can also solve
-$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[L\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
-since using $h:w\mapsto\widetilde{h}(w)+\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$ leads to a solution of (S4).
-Next, we expand the square in the objective function of the minimization problem (S5) and obtain that
-$c(\theta,\widetilde{h})$ $=\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]\right)^{2}\right]+\theta^{2}\mathbb{E}\left[\mathrm{var}(L\,|\,W)\right]+\mathbb{E}[\widetilde{h}(W)^{2}]$
-$-2\theta\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])].$
-It is now clear that the optimal $\widetilde{h}$ is the zero function and by noting that
-$\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])]=\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)],$
-we obtain the desired result by minimizing the quadratic in $\theta$. ∎
-Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section S4.2 with $d=15$.
-We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
-S3.2 Computing directional perturbation effects in the presence of zero speeds
-In Proposition 1 we only considered directional perturbations where $\omega_{\psi}(Z)\neq 0$ almost surely.
-We would, however, like to be able to use the derivative-isolating reparametrizations for estimation of the directional average perturbation effect $\tau_{\psi}$ even when this is not the case.
-To that end, we propose a modified estimation procedure that we outline in Algorithm S1.
-We propose to estimate $\tau_{\psi}$ on the subset of observations where $\omega_{\psi}(Z)\neq 0$ and then correct the resulting point and variance estimate appropriately.
-Algorithm S1 Estimation of $\tau_{\psi}$ in the presence of observations with $\omega_{\psi}(Z)=0$.
-1:Input: Data $(Y_{i},Z_{i})_{i\in[n]}$, directional perturbation $\psi$, derivative-isolating reparametrization $\phi$.
-2:For $i\in[n]$, define $D_{i}:=\mathds{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
-3:Define :=$\mathcal{I}:=\{i\in[n]\mid D_{i}=1\}$.
-4:For $i\in\mathcal{I}$, define $(L_{i},W_{i})=\phi(Z_{i})$.
-5:Compute estimate $\widehat{\tau}$ and asymptotic variance estimate $\widehat{\sigma}^{2}$ from either Algorithm 1 or 3.
-6:Compute $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}$.
-7:return estimate $\widetilde{\tau}:=\widehat{\tau}\widehat{p}$ and asymptotic variance estimate $\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$.
-Theoretical guarantees for the proposed modification follow from repeating the arguments of the proof of Theorem 1 (a) and (c) to obtain the population quantities that determine the asymptotic behavior of $\widehat{\tau}$ and $\widehat{\sigma}^{2}$ and then applying Proposition S2 below.
-Proposition S2.
-Let $(D,S)\in\{0,1\}\times\mathbb{R}$ be a pair of random variables with distributions determined by $\mathcal{P}$ and let $(S_{n},D_{n})_{n\in\mathbb{N}}$ be an i.i.d.
-sequence of copies of $(S,D)$.
-Suppose that there exist $\delta,C,c>0$ such that $\mathrm{var}_{P}(DS)\geq c$ and $\mathbb{E}_{P}[D|S|^{2+\delta}]\leq C$.
-Suppose further that
-$\widehat{\tau}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
-and
-$\widehat{\sigma}^{2}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
-Then,
-$\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i},\quad\widetilde{\tau}:=\widehat{\tau}\widehat{p},\quad\text{and}\quad\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$
-satisfy
-$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})\leq x\right)-\Phi(x)\right|=0.$
-Proof.
-We first note that
-$\widetilde{\tau}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
-since $\widehat{p}\leq 1$. Similarly, we have that
-$\widetilde{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
-Therefore, by several applications of Lemma S2, we have that
-$\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(D_{i}S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1)$
-and the result thus follows by Lemma S4 (a) by the bounds on the moments and variance of $(D,S)$. ∎
-S3.3 Estimating differentiable regression functions
-The following is a description of the smoothing method we apply in the nonparametric estimator of the average directional perturbation effect.
-The procedure uses random forest weights together with a local polynomial estimate to smooth a given regression function to produce a differentiable estimator and to compute its derivatives.
-Local polynomials are a prominent and natural method used to estimate derivatives [e.g., Masry and Fan, 1997, De Brabanter et al., 2013].
-We are interested in estimating the partial derivative of a function $g:\mathbb{R}\times\mathcal{W}\rightarrow\mathbb{R}$.
-As in Klyne and Shah [2023] we start from a given arbitrary regression estimate $\widehat{g}$ of $g$ and apply the following local polynomial approach.
-Estimate, for all $i\in[n]$, polynomial coefficients $\beta\in\mathbb{R}^{3}$ such that the polynomial $p:\mathbb{R}\rightarrow\mathbb{R}$ defined for all $\ell\in\mathbb{R}$ by
-$p(\ell)=\beta^{1}+\beta^{2}(\ell-L_{i})+\beta^{3}(\ell-L_{i})^{2}$
-is a good local approximation of $\widehat{g}$ around $L_{i}$.
-Formally, for a given weight matrix $K\in\mathbb{R}^{n\times n}$, which encodes a closeness measure between sample points, we minimize for all $i\in[n]$ the weighted mean squared loss
-$\hat{\beta}(i):=\operatorname{argmin}_{\beta\in\mathbb{R}^{3}}\sum_{k=1}^{n}\left(\widehat{g}(L_{k},W_{k})-\sum_{j=1}^{3}\beta^{j}(L_{k}-L_{i})^{j-1}\right)^{2}K_{i,\ell}.$ (S6)
-Then, using the estimated coefficients $\hat{\beta}(i)$, we get a smoothed estimate $\widehat{g}_{s}$ which at each sample point $i\in[n]$ is given by
-$\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)\quad\text{and}\quad\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i).$ (S7)
-As weights $K$ we could, for example, use kernel weights, i.e., $K_{i,\ell}=k((X_{i}-X_{\ell})/\sigma)$ for a kernel function $k$ and a bandwidth $\sigma>0$.
-As we are considering a potentially high-dimensional setting, we instead propose to use random forest weights [Meinshausen, 2006, Athey et al., 2019].
-The full procedure is detailed in Algorithm S2.
-Algorithm S2 Local polynomial based smoothing with random forest weights
-1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, fitted values $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$
-2:Regress $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$ on $(L_{i},W_{i})_{i\in[n]}$ using a random forest yielding $\widehat{\mu}$.
-3:Extract weight matrix $K$ from $\widehat{\mu}$.
-4:for $i=1,\ldots,n$ do
-5:     Set $\hat{\beta}(i)$ to equal the coefficient of a degree $2$ local polynomial fit in (S6).
-6:     Set $\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)$.
-7:     Set $\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i)$.
-8:return smoothed predictions $(\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$ and derivatives $(\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$
-S3.4 Location-scale modelling of the score function
-The following is a description of the estimator of the conditional score based on a location-scale model proposed by Klyne and Shah [2023].
-We apply this in our experiments to estimate conditional score functions and include it here for completeness.
-We are interested in the estimation of the conditional score of $L$ given $W$, where $L\in\mathbb{R}$ and $W\in\mathcal{W}$, that is, the derivative of the logarithm of the conditional density of $L$ given $W$.
-Conditional density estimation is difficult even in low dimensions.
-However, if we are willing to assume that there exist functions $m:\mathcal{W}\rightarrow\mathbb{R}$, $v:\mathcal{W}\rightarrow[0,\infty)$ and a random variable $\zeta\in\mathbb{R}$ that is independent of $W$ such that
-$L=m(W)+v(W)^{1/2}\zeta,$
-we can simplify estimation of the score function.
-The procedure is outlined in Algorithm S3 below.
-In all of our experiments, we use the univariate score estimation method given by Ng [1994] that is based on ideas from Cox [1985].
-Algorithm S3 Location-scale score modelling estimation
-1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, nuisance regression methods $\widehat{\eta}=(\widehat{m}$, $\widehat{v})$, univariate score estimator $\widehat{\rho}$.
-2:Split indices $\{1,\dots,n\}$ into $2$ disjoint sets $I_{1},I_{2}$ of (roughly) equal size.
-3:Regress $L$ on $W$ using $I_{1}$ yielding regression estimate $\widehat{m}$.
-4:Regress $(L-\widehat{m}(w))^{2}$ on $W$ using $I_{1}$ yielding variance estimate $\widehat{v}$.
-5:For $i\in I_{2}$, set $\widehat{\xi}_{i}:=\frac{L_{i}-\widehat{m}(W_{i})}{\widehat{v}(W_{i})^{1/2}}$.
-6:Estimate a univariate score from $(\widehat{\xi}_{i})_{i\in I_{2}}$ yielding $\widehat{\rho}$.
-7:return $\widehat{\rho}$.
-Appendix S4 Details and additional results from numerical experiments
-Section S4.1 provides the details behind the experiments in Table 1 in Section 1.
-Section S4.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
-Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1.
-Section S4.4 includes an additional figure showing the semi-synthetic dataset used in Section 4.1.1.
-Section S4.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2.
-S4.1 Simulations from Section 1
-In this section we describe the data-generating processes and methods used in the numerical illustrations of Example 1 and Example 2. The results of are provided in Table 1.
-S4.1.1 Example 1: Effects of absence or presence of microbes
-Let $L\sim\mathrm{Bern}(1/2)$, $U^{1},U^{2}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
-$Z:=(Z^{1},W^{2}(1-Z^{1}),(1-W^{2})(1-Z^{1}))\in\Delta^{2},$
-where $W^{2}:=U^{1}(1-L)+U^{1}LB$ and $Z^{1}:=(1-L)U^{2}$. Furthermore, define $Y$ such that $Y$ is Bernoulli distributed with
-$\mathbb{P}(Y=1\,|\,Z)=\sim\mathrm{Bern}\biggl{(}\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\biggr{)}.$
-In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
-$\mathrm{CKE}^{1}$ is computed using the $\lambda$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
-S4.1.2 Example 2: Effects of increased diversity
-Let $\Pi$ denote the orthogonal projection onto the orthogonal complement of $\mathrm{span}((1,1,1))$.
-Moreover, let $U\sim\mathcal{N}(0,\Pi\,\Pi^{\top})$ and $\xi,\varepsilon\sim\mathcal{N}(0,1)$ be independent random variables and define
-$Z:=-DW+z_{\mathrm{cen}}\in\Delta^{2}\quad\text{and}\quad Y:=L+4W^{\top}e_{1}+\varepsilon,$
-where $W:=\frac{U}{\|U\|_{1}}$, $D:=\mathrm{expit}(W^{\top}e_{1}+\xi)$ and $L:=-\frac{2}{9}D\sum_{j=1}^{3}\sum_{k=1}^{3}|W^{j}-W^{k}|$.
-Then, by construction $L$ is minus the Gini coefficient of $Z$.
-Hence, from the partially linear model of $Y$ it follows that the true effect of increasing the negative Gini coefficient is equal to $1$.
-In the experiment, we generate $n=1000$ i.i.d.
-copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
-$\mathrm{CDI}_{\mathrm{Gini}}$ is computed using the $\tau$-estimator based on a partially linear model for $Y$ (Algorithm 3) with $10$-fold cross-fitting and random forests for all regressions.
-S4.2 Semiparametric estimators
-In this section we investigate the performance of the semiparametric estimators outlined in Section 3 with $W$ as a simplex-valued random variable.
-More specifically, we compare the different approaches of estimating
-$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}\quad\text{and}\quad\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}]$
-from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
-  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (14) and (15).
-See Algorithms 1 and 2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
-  • NPM_no_x: The same estimators as NPM but without cross-fitting.
-  • NPM_oracle: The same estimators as NPM including cross-fitting but using the true nuisance functions for the one-step-correction (propensity score $\pi$ in the binary case and conditional score function $\rho$ in the directional case).
-This is an oracle as the true nuisance functions are unknown in practice.
-  • PLM: The partially least model based estimators in (17) including cross-fitting. See Algorithm 3 with $K=2$ for the detailed estimators.
-  • PLM_no_x: The same estimators as PLM but without cross-fitting.
-  • plugin: The naive plug-in versions of the NPM estimates.
-In the binary effect case, the estimator in (14) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (15) without the term involving $\widehat{\rho}$.
-  • plugin_no_x: The same estimators as plugin but without cross-fitting.
-The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (13).
-Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$.
-For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction.
-For the score estimator we use the location-scale based estimator described in Section S3.4, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
-We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
-In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
-We then consider the following four generative models
-  • Binary $L$ with partially linear $Y$:
-    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=L+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
-  • Binary $L$ with nonparametric $Y$:
-    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
-  • Continuous $L$ with partially linear $Y$:
-    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
-  • Continuous $L$ with nonparametric $Y$:
-    $L=B(W_{1})+(1+B(W_{1}))\xi\quad\text{and}\quad Y=2B(W_{1})L+B(W_{1})+\varepsilon$
-The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$).
-The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$.
-For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above.
-We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=3$ in Figure S4, for $d=15$ in Figure S5 and for $d=75$ in Figure S6.
-Figure S4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section S4.2 with $d=3$.
-PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting.
-NPM works in both settings but can be less robust for small sample sizes.
-Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
-Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=15$.
-The patterns observed in Figure S4 are repeated here.
-Figure S6: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=75$.
-The patterns observed in Figure S4 are repeated here.
-In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated.
-We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
-In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g.
-for $n=1000$ and binary $L$.
-In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
-S4.3 Description of the construction of groups in Section 4.1.1
-In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.1.1.
-On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
-$(\widehat{\pi}_{R}(\mathrm{White}),\widehat{\pi}_{R}(\mathrm{Black}),\widehat{\pi}_{R}(\mathrm{Other}))$ $=(0.855,0.096,0.049),$
-$(\widehat{\pi}_{E}(0),\widehat{\pi}_{E}(1))$ $=(0.454,0.546),$
-$(\widehat{\pi}_{C}(0),\widehat{\pi}_{C}(1))$ $=(0.761,0.239),$
-respectively.
-Now, for the two types of groupings we define target probabilities $p_{i}=(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ for each individual $i$ (see below).
-Based on these probabilities we randomly draw for each individual $i$ one of the three categories using the probabilities $p_{i}$.
-Finally, we randomly group individuals with the same categories into communities of approximately $50$ (as many communities of $50$ and the remaining with $49$) and aggregate the predictors and response as described in the main text.
-The probabilities $(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ are selected, in each of the two settings, as follows.
-Grouping by race and education: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
-$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{E}^{1}(\texttt{education}_{i}),$
-where $\texttt{race}_{i}$ and $\texttt{education}_{i}$ denote the race and education level of the $i$th individual, $p_{R}^{1}$ and $p_{E}^{1}$ are probability vectors satisfying
-$p_{R}^{1}(\mathrm{White})\propto\frac{0.8}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{1}(\mathrm{Black})\propto\frac{0.02}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{1}(\mathrm{Other})\propto\frac{0.18}{\widehat{\pi}_{R}(\mathrm{Other})},$ (S8)
-and
-$p_{E}^{1}(0)\propto\frac{0.95}{\widehat{\pi}_{E}(0)},\quad p_{E}^{1}(1)\propto\frac{0.05}{\widehat{\pi}_{E}(1)}.$
-That is, category $1$ is about as diverse ($(0.8,0.02,0.18)$ is about as diverse as $\widehat{\pi}_{R}$) and with lower education ($0.05<\widehat{\pi}_{E}(1)$).
-The probability $p_{i}^{2}$ of observation $i$ belonging to category $2$ is set to
-$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{E}^{2}(\texttt{education}_{i}),$
-where similarly $p_{R}^{2}$ and $p_{E}^{2}$ are probability vectors but now satisfying
-$p_{R}^{2}(\mathrm{White})\propto\frac{0.65}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{2}(\mathrm{Black})\propto\frac{0.12}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{2}(\mathrm{Other})\propto\frac{0.23}{\widehat{\pi}_{R}(\mathrm{Other})}$ (S9)
-and
-$p_{E}^{2}(0)\propto\frac{0.3}{\widehat{\pi}_{E}(0)},\quad p_{E}^{2}(1)\propto\frac{0.7}{\widehat{\pi}_{E}(1)}.$
-That is, category $2$ is more diverse ($(0.65,0.12,0.23)$ is more diverse than $\widehat{\pi}_{R}$) and with higher education ($0.7>\widehat{\pi}_{E}(1)$).
-Lastly, the probability $p_{i}^{0}$ for category 00 is set to $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
-Grouping by race and compensation: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
-$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{C}^{1}(\texttt{compensation}_{i}),$
-where $p_{R}^{1}$ is as in (S8) and $p_{C}^{1}$ is a probability vector satisfying
-$p_{C}^{1}(0)\propto\frac{0.9}{\widehat{\pi}_{C}(0)},\quad p_{C}^{1}(1)\propto\frac{0.1}{\widehat{\pi}_{C}(1)}.$
-That is, category $1$ is about as diverse and with lower compensation ($0.05<\widehat{\pi}_{C}(1)$). The probability $p_{i}^{2}$ for category $2$ is set to
-$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{C}^{2}(\texttt{compensation}_{i}),$
-where $p_{R}^{2}$ is as in (S9) and $p_{C}^{2}$ is a probability vector satisfying
-$p_{C}^{2}(0)\propto\frac{0.7}{\widehat{\pi}_{C}(0)},\quad p_{C}^{2}(1)\propto\frac{0.3}{\widehat{\pi}_{C}(1)}.$
-That is, category $2$ is more diverse and with higher compensation ($0.3>\widehat{\pi}_{C}(1)$). Lastly, we set the probability $p_{i}^{0}$ for category 00 is $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
-S4.4 Additional figure for the ‘Adult’ data experiment in Section 4.1.1
-Figure S7 complements Figure 4 in Section S4.4 by summarizing the semi-synthetic ‘Adult’ dataset when grouping on compensation.
-Figure S7: Summary of the semi-synthetic ‘Adult’ dataset when grouping on compensation.
-Left: Racial compositions in a ternary plot.
-Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
-Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
-Depending on the color the dependence between $Y$ and the Gini coefficient changes.
-Note that the Gini coefficient represents a similarity rather than a diversity measure.
-S4.5 Semi-synthetic individual-level experiment using groups from Section 4.1.1
-In this section we consider a variant of the experiment performed in Section 4.1.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups.
-More specifically, we generate the groups as before and estimate the probability that $Y=1$ within each group and use this probability to generate new $Y$s for each observation.
-For each individual, we thus have a response $Y$, a race $R$, a ‘contextual’ variable $Z$ that gives the racial distribution of each group and additional covariates $X$.
-By construction of the groups and sampling of the $Y$, we know that there is a positive effect of diversity on the income.
-We now apply the naive_diversity and $\mathrm{CDI}_{\mathrm{Gini}}$ methods as in Section 4.1.1 except we always control for the race of inviduals and omit the unstable estimators that condition on $Z$.
-The results can be seen in Table 5.
-Method Grouping on compensation
-Estimate $95\%$ CI
-[…]
-[…]
-$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.459}}$ $(0.247,0.671)$
-$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.487}}$ $(0.282,0.692)$
-Table 5: Table of estimates in the individual-level ‘Adult’ data experiment.
-Bold red numbers indicate a significant negative effect of diversity at a 5% level, while bold green indicates a significant positive effect.
-The naive diversity method indicates that there is a significant negative effect of diversity on compensation, even when controlling for individual race ($R$) and additional covariates ($X$), despite the fact that the data is constructed with a positive effect of diversity.
-The CDI correctly identifies that the effect is positive.
-We see that, as in Section 4.1.1, the naive_diversity method incorrectly estimates a negative effect of diversity on income while $\mathrm{CDI}_{\mathrm{Gini}}$ correctly identifies that the effect is positive.
-S4.6 Description of additional control variables used in Section 4.1.2
-The additional background variables in the New York schools data used in Section 4.1.2 are described below. Unless otherwise mentioned, no preprocessing was used.
-School income estimate
-The school income estimates were grouped into four categories; one for missing values and three more based on quantiles of the observed values.
-School grade range (grade referring to years of school)
-The range of grades for the schools were grouped into three categories; one for ‘high grades’ where the lowest grade was greater or equal to 6th, one for ‘low grades’ where the highest grade was less than or equal to 4th and the remaining schools were put in a ‘mixed’ category.
-Student attendance rate
-This number was computed as the total number of days attended by all students / total number of days on register for all students.
-% chronically absent students
-The percentage of students missing 10% of school days or at least 18 days per year in a 180-day school year.
-Economic Need Index
-Computed as the percentage of students in temporary housing plus 0.5 times the percentage of students that are eligible for Health Reimbursement Arrangements plus 0.5 times the percentage of students eligible for free lunch.
-Community school
-An indicator of whether the school is a community school.
-% English language learners
-Percentage of students who are learning English for the first time.
-Rigorous Instruction %
-School rating for ‘how well the curriculum and instruction engage students, build critical-thinking skills, and are aligned to the Common Core’.
-Collaborative Teachers %
-School rating for ‘how well teachers participate in opportunities to develop, grow, and contribute to the continuous improvement of the school community’.
-Supportive Environment %
-School rating for ‘how well the school establishes a culture where students feel safe, challenged to grow, and supported to meet high expectations’.
-Effective School Leadership %
-School rating for ‘how well school leadership inspires the school community with a clear instructional vision and effectively distributes leadership to realize this vision’.
-Strong Family-Community Ties %
-School rating for ‘how well the school forms effective partnerships with families to improve the school’.
-Trust %
-School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
-S4.7 Additional plots from gut microbiome analysis in Section 4.2
-In this section we provide an additional figure showing our results from the analysis of the gut microbiome data in Section 4.2.
-In Figure S8, we plot the most significant hits for $\mathrm{CFI}{\mathrm{unit}}$ and $\mathrm{CKE}$ both estimated using random forests together with $\mathrm{CFI}_{\mathrm{mult}}$ estimates.
-We plot the estimates scaled by the estimated standard deviation of $L$ to put the effects on the same scale (we do not correct the confidence intervals for the estimation of this standard deviation).
-Furthermore, we flip the sign of $\mathrm{CKE}$ to keep the interpretation similar to $\mathrm{CFI}$, that is, that we are looking at the effect of increasing $Z^{j}$.
-Figure S8: Effect estimates scaled by the estimated standard deviation of $L$ for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{unit}}$ (left) and $\mathrm{CKE}$ (right).
-The labels are colored by the size of the $p$-value.
-For the $\mathrm{CFI}_{\mathrm{unit}}$ results, the first $9$ species remain below $0.05$ after Bonferroni-correction while this is only true for $2$ for the $\mathrm{CKE}$.
-There is little overlap between the selected species across the two methods.
-Appendix S5 Proof of Theorem 1
-S5.1 Proof of (a)
+S4.1 Theoretical analysis of estimators
+In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (12).
+These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere [e.g., Klyne and Shah, 2023, Chernozhukov et al., 2018].
+We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.
+We start from a given perturbation $\psi$ and a corresponding reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ which reparametrizes $Z$ into $(L,W):=(\phi^{L}(Z),\phi^{W}(Z))$ and assume that $(L,W)$ satisfy (2) if $\psi$ is a binary perturbation and that $\phi$ is derivative-isolating if $\psi$ is a directional perturbation.
+Our goal is now to use the estimation procedures discussed in Section 3 to estimate the average perturbation effects (i.e., $\tau_{\psi}$ or $\lambda_{\psi}$) consistently and construct asymptotically valid confidence intervals for the estimates.
+To construct asymptotically valid confidence intervals, we estimate both the target parameter and its asymptotic variance and then show that the resulting standardized estimators are asymptotically standard normal with a uniform convergence rate over the class of possible distributions of the data.
+Distributionally uniform results are desirable here as they ensure uniform coverage guarantees of the confidence intervals.
+The following result shows uniform asymptotic normality of each of the three estimation procedures.
+Theorem S1 (Uniform asymptotic normality).
+Let $\mathcal{P}$ denote a class of distributions of $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ and assume we are in one of the following settings.
+  1. (a)
+    Nonparametric directional effect estimation: Assumption S1 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\tau},\widehat{\sigma}^{2}_{\tau})$ from Algorithm S1.
+  2. (b)
+    Nonparametric binary effect estimation: Assumption S2 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\lambda},\widehat{\sigma}^{2}_{\lambda})$ from Algorithm S2.
+  3. (c)
+    Effect estimation using a partially linear model: Assumption S3 and $(\widehat{\beta},\widehat{\sigma}^{2}_{\beta}):=(\widehat{\theta},\widehat{\sigma}^{2}_{\theta})$ from Algorithm S3.
+Then, letting $\Phi$ denote the cumulative distribution function of the standard normal distribution, it holds that
+$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widehat{\sigma}_{\beta}}(\widehat{\beta}-\beta_{P})\leq x\right)-\Phi\left(x\right)\right|=0.$
+The result is proven in Section S4.2 below.
+In the following three sections, we go over each of the settings (a), (b) and (c) individually and provide the required assumptions and detailed algorithms for the target parameter and variance.
+For the remainder of this section, we omit $\psi$-subscripts for simplicity and use the following notation to help us express the uniformly asymptotic results rigorously.
+Let $(X_{P,n})_{n\in\mathbb{N},P\in\mathcal{P}}$ denote a family of sequences of random variables indexed by $\mathcal{P}$.
+For a nonnegative sequence $(a_{n})_{n\in\mathbb{N}}$ we write $X_{P,n}=o_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ it holds that
+$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}\epsilon)=0,$
+that is, $X_{P,n}/a_{n}$ converges to 00 in probability uniformly over $\mathcal{P}$.
+We also write $X_{P,n}=O_{\mathcal{P}}(a_{n})$ if for all $\epsilon>0$ there exist $M_{\epsilon}>0$ and $N_{\epsilon}\in\mathbb{N}$ such that
+$\sup_{n\geq N_{\epsilon}}\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|X_{P,n}|\geq a_{n}M)<\epsilon,$
+that is, $X_{P,n}/a_{n}$ is bounded in probability for large $n$.
+When we write conditional expectations given a fitted regression function $\widehat{f}$ we mean a conditional expectation given the random variables used to fit $\widehat{f}$ and any additional randomness involved in the fitting.
+S4.1.1 Nonparametric estimation for directional effects
+We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem S1 (a).
+A high-level overview of this estimation was given in Section 3.2.2.
+More formally, assume we have a dataset of $n$ i.i.d.
+copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S4.6).
+We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm S1 using a user-specified number of splits $K$ in the cross-validation.
+Algorithm S1 Nonparametric estimation of $\tau$
+1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, differentiable regression method $\widehat{f}$, score estimation method $\widehat{\rho}$ and nuisance function estimate $\widehat{\eta}$.
+2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
+3:For $k\in[K]$, split $I_{k}$ into $I_{k,0}$ and $I_{k,1}$ of (roughly) equal size and set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
+4:for $k=1,\dots,K$ do
+5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
+6:     Fit nuisance function estimates $\widehat{\eta}_{k}$ using $I_{-k}$.
+7:     for $r=0,1$ do
+8:         Fit $\widehat{\rho}_{k,r}$ using $I_{k,(1-r)}$ and the nuisance function estimate $\widehat{\eta}_{k}$.
+9:         Set $\widehat{\tau}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}$.
+10:         Set $\widehat{\nu}_{k,r}:=|I_{k,r}|^{-1}\sum_{i\in I_{k,r}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\widehat{\rho}_{k,r}(L_{i},W_{i})\{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})\}\right)^{2}$.
+11:return $\widehat{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\tau}_{k,r}$ and $\widehat{\sigma}^{2}_{\tau}:=\frac{1}{2K}\sum_{k=1}^{K}\sum_{r=0}^{1}\widehat{\nu}_{k,r}-\widehat{\tau}^{2}$.
+To obtain the result in Theorem S1 (a), we require that the estimation methods $\widehat{f}$ and $\widehat{\rho}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
+Assumption S1.
+Assume that for all $P\in\mathcal{P}$ and $w\in\mathcal{W}$, the conditional distribution of $L$ given $W=w$ has a density with respect to Lebesgue measure and denote it by $p_{P}(\cdot\,|\,w)$.
+For all $P\in\mathcal{P}$ define $f_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and assume that for almost all $(\ell,w)\in\phi(\mathcal{Z})$ both $p_{P}(\cdot\,|\,w)$ and $f_{P}(\cdot,w)$ are differentiable at $\ell$.
+Define $\tau_{P}:=\mathbb{E}_{P}[\partial_{\ell}f_{P}(\ell,W)\bigm{|}_{\ell=L}]$ and $\rho_{P}:\phi(\mathcal{Z})\to\mathbb{R}$ for all $(\ell,w)\in\phi(\mathcal{Z})$ by $\rho_{P}(\ell,w):=\partial_{\ell}\log(p_{P}(\ell\,|\,w))$.
+Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{3n},L_{3n},W_{3n})$ denote $3n$ i.i.d.
+copies of $(Y,L,W)$ and define the training datasets $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$ and :=$\mathcal{B}:=\{(Y_{2n+1},L_{2n+1},W_{2n+1})$, $\dots$, $(Y_{3n},L_{3n},W_{3n})\}$.
+Let $\widehat{f}$ denote a differentiable estimator of $f_{P}$ based on $\mathcal{A}$, let $\widehat{\eta}$ denote an estimator of a nuisance function relevant for the estimation of $\rho_{P}$ based on $\mathcal{A}$ and let $\widehat{\rho}$ denote an estimator of $\rho_{P}$ based on $\mathcal{B}$ and using the nuisance function estimate $\widehat{\eta}$.
+Finally, assume there exist constants $\delta,c,C>0$ such that
+  1. (a)
+    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}[|Y|^{4+\delta}],\mathbb{E}_{P}[|\rho_{P}(L,W)|^{4+\delta}],\mathbb{E}_{P}[|\partial_{\ell}f_{P}(L,W)|^{2+\delta}],\mathrm{var}_{P}(Y\,|\,L,W)\}\leq C$
+  2. (b)
+    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,L,W),\mathbb{E}_{P}[\rho_{P}(L,W)^{2}]\}\geq c$
+  3. (c)
+    $\mathbb{E}_{P}\left[\left(\widehat{f}(L_{1},W_{1})-f_{P}(L_{1},W_{1})\right)^{4}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
+  4. (d)
+    $\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L_{1},W_{1})-\partial_{\ell}f_{P}(L_{1},W_{1})\right)^{2}\,|\,\,\widehat{f}\,\right]=o_{\mathcal{P}}(1)$
+  5. (e)
+    $\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right)^{2}=o_{\mathcal{P}}(1)$
+  6. (f)
+    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
+S4.1.2 Nonparametric estimation for binary effects
+We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem S1 (b).
+A high-level overview of this estimation was provided in Section 3.2.1.
+More formally, assume we have a dataset of $n$ i.i.d.
+copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$.
+We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm S2 using a user-specified number of splits $K$ in the cross-validation.
+Algorithm S2 Nonparametric estimation of $\lambda$
+1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression method $\widehat{f}$ and propensity score estimator $\widehat{\pi}$.
+2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
+3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
+4:for $k=1,\dots,K$ do
+5:     Regress $Y$ on $(L,W)$ using $I_{-k}$ yielding regression estimate $\widehat{f}_{k}$.
+6:     Regress $L$ on $W$ using $I_{-k}$ yielding propensity score function $\widehat{\pi}_{k}$.
+7:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}$.
+8:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}_{k}(L_{i},W_{i})}{\widehat{\pi}_{k}(W_{i})}L_{i}\right)^{2}$.
+9:Set $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}1-L_{i}$.
+10:Set $\widehat{\kappa}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}$.
+11:return $\widehat{\lambda}:=\widehat{p}^{-1}\widehat{\kappa}$ and $\widehat{\sigma}^{2}_{\lambda}:=\widehat{p}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}-\widehat{\kappa}^{2}\right)-\widehat{p}^{-3}\widehat{\kappa}^{2}(1-\widehat{p})$.
+To obtain the result in Theorem S1 (b), we require that the estimation methods $\widehat{f}$ and $\widehat{\pi}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
+Assumption S2.
+For all $P\in\mathcal{P}$ define $f_{P}:\{0,1\}\times\mathcal{W}\to\mathbb{R}$ and $\pi_{P}:\mathcal{W}\to[0,1]$ for all $(\ell,w)\in\{0,1\}\times\mathcal{W}$ by $f_{P}(\ell,w):=\mathbb{E}_{P}[Y\,|\,L=\ell,W=w]$ and $\pi_{P}(w):=\mathbb{P}_{P}(L=1\,|\,W=w)$, respectively.
+Define $p_{P}:=\mathbb{P}_{P}(L=0)$ and $\lambda_{P}:=p_{P}^{-1}\left(\mathbb{E}_{P}[\mathbb{E}_{P}[Y\,|\,L=1,W]]-\mathbb{E}_{P}[Y]\right)$.
+Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d.
+copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$.
+Let $\widehat{f}$ denote an estimator of $f_{P}$ based on $\mathcal{A}$ and let $\widehat{\pi}$ denote an estimator of $\pi_{P}$ also based on $\mathcal{A}$.
+Finally, assume there exist constants $\delta,c,C>0$ such that
+  1. (a)
+    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|Y|^{2+\delta}\right],\mathrm{var}_{P}(Y\,|\,W,L)\}\leq C$
+  2. (b)
+    for all $P\in\mathcal{P}:\quad$ $\min\{\mathrm{var}_{P}(Y\,|\,W,L),p_{P},1-\pi_{P}(W),\pi_{P}(W),\widehat{\pi}(W)\}\geq c$
+  3. (c)
+    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{\pi}(W_{i})-\pi_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
+  4. (d)
+    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}(1,W_{i})-f_{P}(1,W_{i}))^{2}=o_{\mathcal{P}}(1)$
+  5. (e)
+    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}(W_{i})-\pi_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
+S4.1.3 Partially linear model for binary/directional effects
+Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem S1 (c).
+A high-level overview of this estimation was provided in Section 3.2.3.
+More formally, assume we have a dataset of $n$ i.i.d.
+copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively.
+We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm S3 using a user-specified number of splits $K$ in the cross-validation.
+Algorithm S3 Partially linear estimation of $\theta$ (where $\theta=\lambda$ or $\theta=\tau$)
+1:Input: Data $(Y_{i},L_{i},W_{i})_{i\in[n]}$, number of folds $K$, regression methods $\widehat{g}$ and $\widehat{m}$.
+2:Split indices $\{1,\dots,n\}$ into $K$ disjoint sets $I_{1},\dots,I_{K}$ of (roughly) equal size.
+3:For $k\in[K]$, set $I_{-k}:=\cup_{k^{\prime}\neq k}I_{k^{\prime}}$.
+4:for $k=1,\dots,K$ do
+5:     Regress $Y$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{g}_{k}$.
+6:     Regress $L$ on $W$ using $I_{-k}$ yielding regression estimate $\widehat{m}_{k}$.
+7:     Set $\widehat{J}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}$.
+8:     Set $\widehat{\kappa}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}$.
+9:Set $\widehat{J}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{J}_{k}$
+10:Set $\widehat{\theta}:=\widehat{J}^{-1}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}\right)$.
+11:for $k=1,\dots,K$ do
+12:     Set $\widehat{\nu}_{k}:=|I_{k}|^{-1}\sum_{i\in I_{k}}\left(\{Y_{i}-\widehat{g}_{k}(W_{i})\}\{L_{i}-\widehat{m}_{k}(W_{i})\}-\widehat{\theta}\{L_{i}-\widehat{m}_{k}(W_{i})\}^{2}\right)^{2}$
+13:return $\widehat{\theta}$ and $\widehat{\sigma}^{2}_{\theta}:=\widehat{J}^{-2}\left(\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}\right)$.
+To obtain the result in Theorem S1 (c), we require that the estimation methods $\widehat{m}$ and $\widehat{g}$ converge sufficiently fast and that the class of distributions $\mathcal{P}$ satisfies several regularity conditions.
+Assumption S3.
+Assume that for all $P\in\mathcal{P}$ there exist $\theta_{P}\in\mathbb{R}$ and $h_{P}:\mathcal{W}\to\mathbb{R}$ such that
+$\mathbb{E}_{P}[Y\,|\,L,W]=\theta_{P}L+h_{P}(W).$
+For all $P\in\mathcal{P}$ define $m_{P}:\mathcal{W}\to\mathbb{R}$ and $g_{P}:\mathcal{W}\to\mathbb{R}$ for all $w\in\mathcal{W}$ by $m(w):=\mathbb{E}_{P}[L\,|\,W=w]$ and $g_{P}(w):=\mathbb{E}_{P}[Y\,|\,W=w]=m_{P}(w)+\theta_{P}m_{P}(w)$, respectively.
+Define $\varepsilon:=Y-\mathbb{E}_{P}[Y\,|\,L,W]$ and $\xi:=Y-\mathbb{E}_{P}[L\,|\,W]$.
+Let $(Y_{1},L_{1},W_{1}),\dots,(Y_{2n},L_{2n},W_{2n})$ denote $2n$ i.i.d.
+copies of $(Y,L,W)$ and define the training dataset $\mathcal{A}:=\{(Y_{n+1},L_{n+1},W_{n+1}),\dots,(Y_{2n},L_{2n},W_{2n})\}$.
+Let $\widehat{m}$ denote an estimator of $m_{P}$ based on $\mathcal{A}$ and let $\widehat{g}$ denote an estimator of $g_{P}$ also based on $\mathcal{A}$.
+Suppose that there exist constants $\delta,c,C>0$ such that
+  1. (a)
+    for all $P\in\mathcal{P}:\quad$ $\max\{\mathbb{E}_{P}\left[|\varepsilon\xi|^{2+\delta}\right]\mathbb{E}_{P}\left[|\xi|^{4}\right],\mathrm{var}_{P}(Y\,|\,W),\mathrm{var}_{P}(L\,|\,W)\}\leq C$
+  2. (b)
+    for all $P\in\mathcal{P}:\quad$ $\min\{\mathbb{E}_{P}\left[\xi^{2}\right],\mathbb{E}_{P}\left[\varepsilon^{2}\xi^{2}\right]\}\geq c$
+  3. (c)
+    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{2}=o_{\mathcal{P}}(n^{-1/2})$
+  4. (d)
+    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{m}(W_{i})-m_{P}(W_{i}))^{4}=o_{\mathcal{P}}(1)$
+  5. (e)
+    $\frac{1}{n}\sum_{i=1}^{n}(\widehat{g}(W_{i})-g_{P}(W_{i}))^{2}=o_{\mathcal{P}}(1)$
+  6. (f)
+    $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{m}(W_{i})-m_{P}(W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{g}(W_{i})-g_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
+S4.2 Proof of Theorem S1
+In this section, we prove Theorem S1 and some additional auxiliary results.
+S4.2.1 Proof of (a)
 Proof.
@@
 $+\underbrace{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})\}+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}}_{(\textup{III})}.$
-By the Cauchy–Schwarz inequality, Assumption 1 (f) and Lemma S3, we have
+By the Cauchy–Schwarz inequality, Assumption S1 (f) and Lemma S3, we have
 $|(\textup{I})|\leq\sqrt{\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\frac{1}{N_{k,r}}\sum_{i\in I_{k,r}}\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}=o_{\mathcal{P}}(n^{-1}).$
@@
 $\mathbb{P}_{P}(|(\textup{II})|\geq N_{k,r}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k,r}},\widehat{\rho}_{k,r})\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k,r}}\{\widehat{\rho}_{k,r}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\}^{2}\mathrm{var}_{P}(Y_{i}\,|\,L_{i},W_{i}).$
-Combining the above with Assumption 1 (a) and (e), we have
+Combining the above with Assumption S1 (a) and (e), we have
 $\mathbb{P}_{P}(|(\textup{II})|\geq N_{k,r}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k,r}},\widehat{\rho}_{k,r})=o_{\mathcal{P}}(1)$
@@
 $\leq\mathbb{E}_{P}[|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})|\,|\,\widehat{f}_{k}]+\mathbb{E}_{P}\left[\rho_{P}(L_{i},W_{i})^{2}\right]^{1/2}\mathbb{E}_{P}\left[\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}\,|\,\widehat{f}_{k}\right]^{1/2}.$
-By the above, Assumption 1 (c), (d) and (a), Lemma S2 (a) and Jensen’s inequality, we have that
+By the above, Assumption S1 (c), (d) and (a), Lemma S2 (a) and Jensen’s inequality, we have that
 $\mathbb{E}_{P}\left[\left|\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right|\,|\,\widehat{f}_{k}\right]$ $=O_{\mathcal{P}}(1),$
@@
 $\qquad+\frac{2}{\epsilon^{2}}\mathbb{E}_{P}\left[|\rho_{P}(L_{i},W_{i})|^{4}\right]^{1/2}\mathbb{E}_{P}\left[\left|\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right|^{4}\,|\,\widehat{f}_{k}\right]^{1/2}.$
-Assumption 1 (c), (d) and (a) and Lemma S3 thus yield that
+Assumption S1 (c), (d) and (a) and Lemma S3 thus yield that
 $\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
@@
 $|\varepsilon_{i}|^{4+\delta}\leq 2^{3+\delta}\left(|Y_{i}|^{4+\delta}+|\mathbb{E}_{P}[Y_{i}\,|\,L_{i},W_{i}]|^{4+\delta}\right).$
-Therefore, using Assumption 1 (a), the Cauchy–Schwarz inequality and Jensen’s inequality, we have
+Therefore, using Assumption S1 (a), the Cauchy–Schwarz inequality and Jensen’s inequality, we have
 $\mathbb{E}_{P}\left[|S_{i}|^{2+\delta^{\prime}}\right]$ $\leq 2^{1+\delta^{\prime}}\left(C^{(2+\delta^{\prime})/(2+\delta)}+\left(\mathbb{E}_{P}\left[\left|\rho_{P}(L_{i},W_{i})\right|^{4+\delta}\right]\mathbb{E}_{P}\left[\left|\varepsilon_{i}\right|^{4+\delta}\right]\right)^{1/2}\right)$
@@
 $\frac{\sqrt{n}}{\widehat{\sigma}_{\tau}}(\widehat{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1).$
-The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption 1 (a) and (b), since
+The conclusion now follows from a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20] and Lemma S4 (a), which we can apply by Assumption S1 (a) and (b), since
 $\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}f_{P}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\partial_{\ell}f_{P}(L,W)\right]^{2}$
 $\geq\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
-S5.2 Proof of (b)
+S4.2.2 Proof of (b)
 Proof.
@@
 $+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\left(\widehat{f}_{k}(1,W_{i})-f_{P}(1,W_{i})\right)\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)}_{(\textup{III})}.$
-Applying the triangle and Cauchy–Schwarz inequality, followed by Assumption 2 (b) and (e) leads to
+Applying the triangle and Cauchy–Schwarz inequality, followed by Assumption S2 (b) and (e) leads to
 $|(\textup{I})|$ $\leq\sqrt{\frac{1}{N_{k}^{2}}\sum_{i\in I_{k}}\frac{\{\pi_{P}(W_{i})-\widehat{\pi}_{k}(W_{i})\}^{2}}{\widehat{\pi}_{k}(W_{i})^{2}\pi_{P}(W_{i})^{2}}\sum_{i\in I_{k}}\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}^{2}}$
@@
 $=\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\frac{\{\pi_{P}(W_{i})-\widehat{\pi}_{k}(W_{i})\}^{2}}{\widehat{\pi}_{k}(W_{i})^{2}\pi_{P}(W_{i})^{2}}\mathrm{var}_{P}(Y_{i}\,|\,W_{i},L_{i}).$
-Combining Assumption 2 (b), (c) and (a), we conclude that
+Combining Assumption S2 (b), (c) and (a), we conclude that
 $\mathbb{P}_{P}(|(\textup{II})|\geq N_{k}^{-1/2}\epsilon\,|\,(L_{i},W_{i})_{i\in I_{k}},\widehat{\pi}_{k})=o_{\mathcal{P}}(1)$
@@
 $\mathbb{P}(|(\textup{III})|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{f}_{k})$ $\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\mathbb{E}_{P}\left[\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)^{2}\,\Bigm{|}\,W_{i}\right]\left(\widehat{f}_{k}(1,W_{i})-f_{P}(1,W_{i})\right)^{2}.$
-By Assumption 2 (b), we have
+By Assumption S2 (b), we have
 $\mathbb{E}_{P}\left[\left(1-\frac{L_{i}}{\pi_{P}(W_{i})}\right)^{2}\,\Bigm{|}\,W_{i}\right]=\frac{1}{\pi_{P}(W_{i})}-1\leq\frac{1-c}{c},$
-hence, by Assumption 2 (d),
+hence, by Assumption S2 (d),
 $\mathbb{P}_{P}(|(\textup{III})|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{f}_{k})=o_{\mathcal{P}}(1)$
@@
 $|S_{i}|^{2+\delta}\leq 3^{1+\delta}\left(|f_{P}(1,W_{i})|^{2+\delta}+|Y_{i}|^{2+\delta}+\frac{|\varepsilon_{i}|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\right),$
-by Assumption 2 (b),
+by Assumption S2 (b),
 $|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathds{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
@@
 $|\varepsilon_{i}|^{2+\delta}\leq 2^{1+\delta}\left(|Y_{i}|^{2+\delta}+|\mathbb{E}_{P}[Y_{i}\,|\,W_{i},L_{i}]|^{2+\delta}\right).$
-Therefore, using Assumption 2 (a)
+Therefore, using Assumption S2 (a)
 $\mathbb{E}_{P}\left[|S_{i}|^{2+\delta}\right]\leq 3^{1+\delta}\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\right]\left(1+c^{-(2+\delta)}(1+2^{1+\delta})\right)\leq C^{\prime}$
@@
 $\widehat{p}=p_{P}+o_{\mathcal{P}}(1).$ (S16)
-Using this, Assumption 2 (b), (S15) and (S16), we have by Lemma S2 that
+Using this, Assumption S2 (b), (S13) and (S14), we have by Lemma S2 that
 $\widehat{\sigma}^{2}_{\lambda}=\underbrace{\frac{\sigma^{2}_{\kappa_{P}}}{p_{P}^{2}}-\frac{\kappa_{P}^{2}(1-p_{P})}{p_{P}^{3}}}_{\sigma^{2}_{\lambda_{P}}}+o_{\mathcal{P}}(1).$ (S17)
@@
 $\kappa_{P}^{2}=\mathbb{E}_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
-thus by Assumption 2 (b)
+thus by Assumption S2 (b)
 $p_{P}\sigma^{2}_{\kappa_{P}}-\kappa_{P}^{2}(1-p_{P})=p_{P}\mathbb{E}_{P}[S_{i}^{2}]-\kappa_{P}^{2}\geq p_{P}\mathbb{E}_{P}\left[(Y-f_{P}(L,W))^{2}\left(\frac{(L-\pi(W))^{2}}{\pi(W)^{2}}\right)\right]\geq 1-c>0.$ (S18)
@@
 and $\sigma^{2}_{\lambda_{P}}p_{P}^{4}\geq c(1-c)$ by (S18), the result now follows by Lemma S4 (a) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
-S5.3 Proof of (c)
+S4.2.3 Proof of (c)
 Proof.
@@
 $\widehat{J}_{k}=\frac{1}{N_{k}}\sum_{i\in I_{k}}\xi_{i}^{2}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}_{(\textup{I})}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}\xi_{i}}_{(\textup{II})}.$
-By Assumption 3 (c), we have $(\textup{I})=o_{\mathcal{P}}(N_{k}^{-1/2})$.
+By Assumption S3 (c), we have $(\textup{I})=o_{\mathcal{P}}(N_{k}^{-1/2})$.
 Since the summands of the third term are conditionally i.i.d.
@@
 $\mathbb{P}_{P}\left(\left|(\textup{II})\right|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{m}_{k}\right)\leq\frac{1}{N_{k}\epsilon^{2}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}\mathrm{var}_{P}(L_{i}\,|\,W_{i}).$
-Combining the above with Assumption 3 (a) and (c), we conclude that
+Combining the above with Assumption S3 (a) and (c), we conclude that
 $\mathbb{P}_{P}\left(\left|(\textup{II})\right|\geq N_{k}^{-1/2}\epsilon\,|\,(W_{i})_{i\in I_{k}},\widehat{m}_{k}\right)=o_{\mathcal{P}}(N_{k}^{-1/2})=o_{\mathcal{P}}(1)$
@@
 $\sup_{P\in\mathcal{P}}|\theta_{P}|=\sup_{P\in\mathcal{P}}\left|\frac{\mathbb{E}_{P}[\varepsilon\xi]}{\mathbb{E}_{P}\left[\xi^{2}\right]}\right|\leq\sup_{P\in\mathcal{P}}\frac{\mathbb{E}_{P}[|\varepsilon\xi|]}{\mu_{P}}\leq\frac{C^{1/(2+\delta)}}{c}$ (S21)
-by Assumption 3 (a) and (b). Thus, for all $k\in[K]$, we note that, using the fact that $Y-g_{P}(W)=\varepsilon+\theta_{P}\xi$ and (S21),
+by Assumption S3 (a) and (b). Thus, for all $k\in[K]$, we note that, using the fact that $Y-g_{P}(W)=\varepsilon+\theta_{P}\xi$ and (S19),
 $\widehat{\kappa}_{k}-\theta_{P}\widehat{J}$ $=\frac{1}{N_{k}}\sum_{i\in I_{k}}\varepsilon_{i}\xi_{i}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}}_{(\textup{III})}$
 $+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}\xi_{i}}_{(\textup{IV})}+\underbrace{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}\{\theta_{P}\xi_{i}+\varepsilon_{i}\}}_{(\textup{V})}+o_{\mathcal{P}}(N_{k}^{-1/2}).$
-By the Cauchy–Schwarz inequality and Assumption 3 (f), we have
+By the Cauchy–Schwarz inequality and Assumption S3 (f), we have
 $|(\textup{III})|\leq\sqrt{\frac{1}{N_{k}}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}}=o_{\mathcal{P}}(N_{k}^{-1/2}).$
@@
 $\frac{1}{K}\sum_{k=1}^{K}\widehat{\kappa}_{k}-\theta_{P}\widehat{J}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}+o_{\mathcal{P}}(n^{-1/2}).$ (S22)
-Next, combining (S21), Chebyshev’s inequality and Assumption 3 (a) yields that
+Next, combining (S19), Chebyshev’s inequality and Assumption S3 (a) yields that
 $\widehat{J}=\mu_{P}+o_{\mathcal{P}}(1),$
-hence Assumption 3 (b) and Lemma S2 (d) yields
+hence Assumption S3 (b) and Lemma S2 (d) yields
 $\widehat{J}^{-1}=\mu_{P}^{-1}+o_{\mathcal{P}}(1).$ (S23)
@@
 An analogous argument can be used to show that $\widetilde{(\textup{III})}_{g}=o_{\mathcal{P}}(1)$.
-Using $\sum_{i}a_{i}^{2}a_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}a_{i}^{2}\right)$ and Assumption 3 (f), we have
+Using $\sum_{i}a_{i}^{2}a_{i}^{2}\leq\left(\sum_{i}a_{i}^{2}\right)\left(\sum_{i}a_{i}^{2}\right)$ and Assumption S3 (f), we have
 $\widetilde{(\textup{IV})}\leq\frac{1}{N_{k}}\sum_{i\in I_{k}}\{m_{P}(W_{i})-\widehat{m}_{k}(W_{i})\}^{2}\sum_{i\in I_{k}}\{g_{P}(W_{i})-\widehat{g}_{k}(W_{i})\}^{2}=o_{\mathcal{P}}(1).$
@@
 $\frac{1}{K}\sum_{k=1}^{K}\widehat{\nu}_{k}=\nu_{P}+o_{\mathcal{P}}(1),$
-so by Lemma S2, (S25) and Assumption 3 (a) and (b), (S26) holds.
-Finally, combining (S22), (S26) with Lemmas S3 and S2 using Assumption 3 (a) and (b), we have that
+so by Lemma S2, (S23) and Assumption S3 (a) and (b), (S24) holds.
+Finally, combining (S20), (S24) with Lemmas S3 and S2 using Assumption S3 (a) and (b), we have that
 $\frac{\sqrt{n}}{\widehat{\sigma}_{\theta}}(\widehat{\theta}-\theta_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}}{\nu_{P}^{1/2}}+o_{\mathcal{P}}(1).$
-The conclusion now follows from Assumption 3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
-S5.4 Auxiliary results
-This is a collection of minor or known results that we have included for completeness.
+The conclusion now follows from Assumption S3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
+S4.2.4 Auxiliary results
 Lemma S1.
@@
 ∎
-Appendix S6 Miscellaneous results
-The following result provides a relationship between the $\mathrm{CFI}^{j}$ and coefficients in a well-specified log-contrast regression model.
-The result has been previously proven in Huang et al.
-[2023].
+S4.3 Robustness of nuisance function estimation for different estimators
+A key motivation for applying the estimators based on a partially linear model on $Y$ is to avoid estimating nuisance functions that result in unstable estimators.
+We illustrate this point based on three examples below.
+S4.3.1 Estimation of nuisance functions with continuous $L$
+To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section 3.2.2) compared to $m$ and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
+We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
+$L:=W^{1}+(1+\mathds{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
+The out-of-sample predictions evaluated on a new dataset with $n=1000$ observations can be seen together with the true function in Figure S3.
+Figure S3: Estimation of nuisance functions on the simulated data in Section S4.3.1.
+The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size.
+We see that the $m$ and $g$ estimates are reasonable estimates of the true functions, that roughly capture the behavior of the true function in an unbiased (conditioned on $W^{1}$) way.
+In contrast, the derivative and score estimates, are much worse at capturing the more complex behavior (with clear biases conditioned on $W^{1}$).
+S4.3.2 Estimation of nuisance functions with binary $L$
+To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section 3.2.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
+We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
+$Y:=2LW^{1}+\varepsilon.$
+The out-of-sample predictions evaluated on new dataset with $n=1000$ observations can be seen together with the true function in Figure S4.
+Figure S4: Estimation of nuisance functions on the simulated data in Section S4.3.2.
+The estimates are computed on a dataset of size $n=1000$ and the estimates (red) and true functions (blue) are evaluated on a new dataset of the same size.
+We see that the $m$, $g$ and $f$ estimates are reasonable estimates of the true functions whereas the inverse propensity (even when cutting off extreme values at $100$) is poorly estimated for $W^{1}$ close to zero.
+S4.3.3 Partially linear model estimate when partial linearity is violated
+The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified.
+The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting).
+We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section S5.2 which we do in Figure S5.
+Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
 Proposition S4.
-Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$.
-Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}\log(z^{j})+\mu$.
-Then, for all $j\in[d]$, we have
-$\mathrm{CFI}^{j}_{\mathrm{mult}}=\beta^{j}.$
+Let $(Y,L,W)\in\mathbb{R}\times\mathbb{R}\times\mathcal{W}$ denote random variables where $\max\{\mathbb{E}[L^{2}],\mathbb{E}[Y^{2}]\}<\infty$. Then,
+$\theta:=\frac{\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]}{\mathbb{E}[\mathrm{var}(L\,|\,W)]}\quad\text{and}\quad h:w\mapsto\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$
+are solutions to the minimization problem
+$\min_{\theta\in\mathbb{R},h:\mathcal{W}\to\mathbb{R}}\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W)\right)^{2}\right].$ (S30)
 Proof.
-The perturbation corresponding to $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is given by
-$\psi(z,\gamma)=z+\gamma\frac{2z^{j}(1-z^{j})}{\|e_{j}-z\|_{1}}(e_{j}-z)=z+\gamma z^{j}(e_{j}-z).$
-We thus have
-$f(\psi(z,\gamma))=\sum_{k=1}^{d}\beta^{k}\log(z^{k}+\gamma z^{j}(\mathds{1}_{\{k=j\}}-z^{k}))+\mu$
-and therefore
-$\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}=\sum_{k=1}^{d}\frac{\beta^{k}z^{j}(\mathds{1}_{\{k=j\}}-z^{k})}{z^{k}}=\beta^{j}-z^{j}\sum_{k=1}^{d}\beta^{k}=\beta^{j}.$
-We conclude that
-$\mathrm{CFI}^{j}_{\mathrm{mult}}=\tau_{\psi}=\mathbb{E}\left[\partial_{\gamma}f(\psi(z,\gamma))\bigm{|}_{\gamma=0}\right]=\beta^{j}.$
-The next result shows that the effects proposed in Cox [1971] are special cases of average perturbation effects in a parametric model.
+First, we note that instead of the optimization in (S30) we can also solve
+$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[L\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S31)
+since using $h:w\mapsto\widetilde{h}(w)+\mathbb{E}[Y\,|\,W=w]-\theta\mathbb{E}[L\,|\,W=w]$ leads to a solution of (S30).
+Next, we expand the square in the objective function of the minimization problem (S31) and obtain that
+$c(\theta,\widetilde{h})$ $=\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]\right)^{2}\right]+\theta^{2}\mathbb{E}\left[\mathrm{var}(L\,|\,W)\right]+\mathbb{E}[\widetilde{h}(W)^{2}]$
+$-2\theta\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])].$
+It is now clear that the optimal $\widetilde{h}$ is the zero function and by noting that
+$\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W])(L-\mathbb{E}[L\,|\,W])]=\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)],$
+we obtain the desired result by minimizing the quadratic in $\theta$. ∎
+Figure S5: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section S5.2 with $d=15$.
+We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
+S4.4 Computing directional perturbation effects in the presence of zero speeds
+In Proposition 1 we only considered directional perturbations where $\omega_{\psi}(Z)\neq 0$ almost surely.
+We would, however, like to be able to use the derivative-isolating reparametrizations for estimation of the directional average perturbation effect $\tau_{\psi}$ even when this is not the case.
+To that end, we propose a modified estimation procedure that we outline in Algorithm S4.
+We propose to estimate $\tau_{\psi}$ on the subset of observations where $\omega_{\psi}(Z)\neq 0$ and then correct the resulting point and variance estimate appropriately.
+Algorithm S4 Estimation of $\tau_{\psi}$ in the presence of observations with $\omega_{\psi}(Z)=0$.
+1:Input: Data $(Y_{i},Z_{i})_{i\in[n]}$, directional perturbation $\psi$, derivative-isolating reparametrization $\phi$.
+2:For $i\in[n]$, define $D_{i}:=\mathds{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
+3:Define :=$\mathcal{I}:=\{i\in[n]\mid D_{i}=1\}$.
+4:For $i\in\mathcal{I}$, define $(L_{i},W_{i})=\phi(Z_{i})$.
+5:Compute estimate $\widehat{\tau}$ and asymptotic variance estimate $\widehat{\sigma}^{2}$ from either Algorithm S1 or S3.
+6:Compute $\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}$.
+7:return estimate $\widetilde{\tau}:=\widehat{\tau}\widehat{p}$ and asymptotic variance estimate $\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$.
+Theoretical guarantees for the proposed modification follow from repeating the arguments of the proof of Theorem S1 (a) and (c) to obtain the population quantities that determine the asymptotic behavior of $\widehat{\tau}$ and $\widehat{\sigma}^{2}$ and then applying Proposition S5 below.
 Proposition S5.
-Let $(Y,Z)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-1}$ and let $z_{*}\in\Delta^{d-1}$.
-Let $f:\Delta^{d-1}\to\mathbb{R}$ be given by $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ and suppose that there exists $\beta=(\beta^{1},\dots,\beta^{j})\in\mathbb{R}^{d}$ with $\sum_{j=1}^{d}\beta^{j}z_{*}^{j}=0$ and $\mu\in\mathbb{R}$ such that for all $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$ we have $f(z)=\sum_{j=1}^{d}\beta^{j}z^{j}+\mu$.
-For $k\in[d]$, define the perturbation $\psi_{k}$ such that for all components $j\in[d]$ it holds
-$\psi_{k}^{j}(z,\gamma)=\begin{cases}z^{j}+\gamma\quad&\text{if }j=k\\ z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\quad&\text{else}.\end{cases}$
-Then
-$\tau_{\psi_{k}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+Let $(D,S)\in\{0,1\}\times\mathbb{R}$ be a pair of random variables with distributions determined by $\mathcal{P}$ and let $(S_{n},D_{n})_{n\in\mathbb{N}}$ be an i.i.d.
+sequence of copies of $(S,D)$.
+Suppose that there exist $\delta,C,c>0$ such that $\mathrm{var}_{P}(DS)\geq c$ and $\mathbb{E}_{P}[D|S|^{2+\delta}]\leq C$.
+Suppose further that
+$\widehat{\tau}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
+and
+$\widehat{\sigma}^{2}=\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{\sum_{i=1}^{n}D_{i}}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
+Then,
+$\widehat{p}:=\frac{1}{n}\sum_{i=1}^{n}D_{i},\quad\widetilde{\tau}:=\widehat{\tau}\widehat{p},\quad\text{and}\quad\widetilde{\sigma}^{2}:=\widehat{p}\widehat{\sigma}^{2}+\widehat{\tau}^{2}\widehat{p}(1-\widehat{p})$
+satisfy
+$\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|\mathbb{P}_{P}\left(\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})\leq x\right)-\Phi(x)\right|=0.$
 Proof.
-We have
-$f(\psi_{k}(z,\gamma))=\beta^{k}(z^{k}+\gamma)+\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\left(z^{j}-\gamma\frac{z^{j}_{*}}{1-z^{k}_{*}}\right)+\mu$
-and therefore
-$\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}=\beta^{k}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\beta^{k}\left(1+\frac{z^{k}_{*}}{1-z^{k}_{*}}\right)-\sum\limits_{j=1}^{d}\beta^{j}\frac{z^{j}_{*}}{1-z^{k}_{*}}=\frac{\beta^{k}}{1-z^{k}_{*}}.$
-We conclude that
-$\tau_{\psi_{k}}=\mathbb{E}\left[\partial_{\gamma}f(\psi_{k}(z,\gamma))\bigm{|}_{\gamma=0}\right]=\frac{\beta^{k}}{1-z^{k}_{*}}.$
+We first note that
+$\widetilde{\tau}:=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}+o_{\mathcal{P}}(n^{-1/2})$
+since $\widehat{p}\leq 1$. Similarly, we have that
+$\widetilde{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}+o_{\mathcal{P}}(1).$
+Therefore, by several applications of Lemma S2, we have that
+$\frac{\sqrt{n}}{\widetilde{\sigma}}(\widetilde{\tau}-\tau_{P})=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(D_{i}S_{i}-\tau_{P})}{\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}D_{i}S_{i}\right)^{2}\right)^{1/2}}+o_{\mathcal{P}}(1)$
+and the result thus follows by Lemma S4 (a) by the bounds on the moments and variance of $(D,S)$. ∎
+S4.5 Estimating differentiable regression functions
+The following is a description of the smoothing method we apply in the nonparametric estimator of the average directional perturbation effect.
+The procedure uses random forest weights together with a local polynomial estimate to smooth a given regression function to produce a differentiable estimator and to compute its derivatives.
+Local polynomials are a prominent and natural method used to estimate derivatives [e.g., Masry and Fan, 1997, De Brabanter et al., 2013].
+We are interested in estimating the partial derivative of a function $g:\mathbb{R}\times\mathcal{W}\rightarrow\mathbb{R}$.
+As in Klyne and Shah [2023] we start from a given arbitrary regression estimate $\widehat{g}$ of $g$ and apply the following local polynomial approach.
+Estimate, for all $i\in[n]$, polynomial coefficients $\beta\in\mathbb{R}^{3}$ such that the polynomial $p:\mathbb{R}\rightarrow\mathbb{R}$ defined for all $\ell\in\mathbb{R}$ by
+$p(\ell)=\beta^{1}+\beta^{2}(\ell-L_{i})+\beta^{3}(\ell-L_{i})^{2}$
+is a good local approximation of $\widehat{g}$ around $L_{i}$.
+Formally, for a given weight matrix $K\in\mathbb{R}^{n\times n}$, which encodes a closeness measure between sample points, we minimize for all $i\in[n]$ the weighted mean squared loss
+$\hat{\beta}(i):=\operatorname{argmin}_{\beta\in\mathbb{R}^{3}}\sum_{k=1}^{n}\left(\widehat{g}(L_{k},W_{k})-\sum_{j=1}^{3}\beta^{j}(L_{k}-L_{i})^{j-1}\right)^{2}K_{i,\ell}.$ (S32)
+Then, using the estimated coefficients $\hat{\beta}(i)$, we get a smoothed estimate $\widehat{g}_{s}$ which at each sample point $i\in[n]$ is given by
+$\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)\quad\text{and}\quad\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i).$ (S33)
+As weights $K$ we could, for example, use kernel weights, i.e., $K_{i,\ell}=k((X_{i}-X_{\ell})/\sigma)$ for a kernel function $k$ and a bandwidth $\sigma>0$.
+As we are considering a potentially high-dimensional setting, we instead propose to use random forest weights [Meinshausen, 2006, Athey et al., 2019].
+The full procedure is detailed in Algorithm S5.
+Algorithm S5 Local polynomial based smoothing with random forest weights
+1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, fitted values $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$
+2:Regress $(\widehat{g}(L_{i},W_{i}))_{i\in[n]}$ on $(L_{i},W_{i})_{i\in[n]}$ using a random forest yielding $\widehat{\mu}$.
+3:Extract weight matrix $K$ from $\widehat{\mu}$.
+4:for $i=1,\ldots,n$ do
+5:     Set $\hat{\beta}(i)$ to equal the coefficient of a degree $2$ local polynomial fit in (S32).
+6:     Set $\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{1}(i)$.
+7:     Set $\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}):=\hat{\beta}^{2}(i)$.
+8:return smoothed predictions $(\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$ and derivatives $(\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$
+S4.6 Location-scale modelling of the score function
+The following is a description of the estimator of the conditional score based on a location-scale model proposed by Klyne and Shah [2023].
+We apply this in our experiments to estimate conditional score functions and include it here for completeness.
+We are interested in the estimation of the conditional score of $L$ given $W$, where $L\in\mathbb{R}$ and $W\in\mathcal{W}$, that is, the derivative of the logarithm of the conditional density of $L$ given $W$.
+Conditional density estimation is difficult even in low dimensions.
+However, if we are willing to assume that there exist functions $m:\mathcal{W}\rightarrow\mathbb{R}$, $v:\mathcal{W}\rightarrow[0,\infty)$ and a random variable $\zeta\in\mathbb{R}$ that is independent of $W$ such that
+$L=m(W)+v(W)^{1/2}\zeta,$
+we can simplify estimation of the score function.
+The procedure is outlined in Algorithm S6 below.
+In all of our experiments, we use the univariate score estimation method given by Ng [1994] that is based on ideas from Cox [1985].
+Algorithm S6 Location-scale score modelling estimation
+1:Input: Data $(L_{i},W_{i})_{i\in[n]}$, nuisance regression methods $\widehat{\eta}=(\widehat{m}$, $\widehat{v})$, univariate score estimator $\widehat{\rho}$.
+2:Split indices $\{1,\dots,n\}$ into $2$ disjoint sets $I_{1},I_{2}$ of (roughly) equal size.
+3:Regress $L$ on $W$ using $I_{1}$ yielding regression estimate $\widehat{m}$.
+4:Regress $(L-\widehat{m}(w))^{2}$ on $W$ using $I_{1}$ yielding variance estimate $\widehat{v}$.
+5:For $i\in I_{2}$, set $\widehat{\xi}_{i}:=\frac{L_{i}-\widehat{m}(W_{i})}{\widehat{v}(W_{i})^{1/2}}$.
+6:Estimate a univariate score from $(\widehat{\xi}_{i})_{i\in I_{2}}$ yielding $\widehat{\rho}$.
+7:return $\widehat{\rho}$.
+Appendix S5 Details and additional results from numerical experiments
+Section S5.1 provides the details behind the experiments in Table S1 in Section 1.
+Section S5.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional.
+Section S5.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1.
+Section S5.4 includes figures visualizing the semi-synthetic dataset used in Section 4.1.1.
+Section S5.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2.
+Section S5.7 contains additional plots illustrating results from the microbiome data analysis in Section 4.2.
+S5.1 Simulations from Section S2.2
+In this section we describe the data-generating processes and methods used in the numerical illustrations of Example S1 and Example S2. The results are provided in Table S1.
+S5.1.1 Example S1: Effects of absence or presence of microbes
+Let $L\sim\mathrm{Bern}(1/2)$, $U^{1},U^{2}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
+$Z:=(Z^{1},W^{2}(1-Z^{1}),(1-W^{2})(1-Z^{1}))\in\Delta^{2},$
+where $W^{2}:=U^{1}(1-L)+U^{1}LB$ and $Z^{1}:=(1-L)U^{2}$. Furthermore, define $Y$ such that $Y$ is Bernoulli distributed with
+$\mathbb{P}(Y=1\,|\,Z)=\sim\mathrm{Bern}\biggl{(}\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\biggr{)}.$
+In the experiment, we generate $n=1000$ i.i.d. copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
+$\mathrm{CKE}^{1}$ is computed using the $\lambda$-estimator based on a partially linear model for $Y$ (Algorithm S3) with $10$-fold cross-fitting and random forests for all regressions.
+S5.1.2 Example S2: Effects of increased diversity
+Let $\Pi$ denote the orthogonal projection onto the orthogonal complement of $\mathrm{span}((1,1,1))$.
+Moreover, let $U\sim\mathcal{N}(0,\Pi\,\Pi^{\top})$ and $\xi,\varepsilon\sim\mathcal{N}(0,1)$ be independent random variables and define
+$Z:=-DW+z_{\mathrm{cen}}\in\Delta^{2}\quad\text{and}\quad Y:=L+4W^{\top}e_{1}+\varepsilon,$
+where $W:=\frac{U}{\|U\|_{1}}$, $D:=\mathrm{expit}(W^{\top}e_{1}+\xi)$ and $L:=-\frac{2}{9}D\sum_{j=1}^{3}\sum_{k=1}^{3}|W^{j}-W^{k}|$.
+Then, by construction $L$ is minus the Gini coefficient of $Z$.
+Hence, from the partially linear model of $Y$ it follows that the true effect of increasing the negative Gini coefficient is equal to $1$.
+In the experiment, we generate $n=1000$ i.i.d.
+copies $(Y_{1},Z_{1}),\ldots,(Y_{n},Z_{n})$ of $(Y,Z)$.
+$\mathrm{CDI}_{\mathrm{Gini}}$ is computed using the $\tau$-estimator based on a partially linear model for $Y$ (Algorithm S3) with $10$-fold cross-fitting and random forests for all regressions.
+S5.2 Semiparametric estimators
+In this section we investigate the performance of the semiparametric estimators outlined in Section 3 with $W$ as a simplex-valued random variable.
+More specifically, we compare the different approaches of estimating
+$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}\quad\text{and}\quad\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}]$
+from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
+  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (10) and (11).
+See Algorithms S1 and S2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
+  • NPM_no_x: The same estimators as NPM but without cross-fitting.
+  • NPM_oracle: The same estimators as NPM including cross-fitting but using the true nuisance functions for the one-step-correction (propensity score $\pi$ in the binary case and conditional score function $\rho$ in the directional case).
+This is an oracle as the true nuisance functions are unknown in practice.
+  • PLM: The partially least model based estimators in (13) including cross-fitting. See Algorithm S3 with $K=2$ for the detailed estimators.
+  • PLM_no_x: The same estimators as PLM but without cross-fitting.
+  • plugin: The naive plug-in versions of the NPM estimates.
+In the binary effect case, the estimator in (10) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (11) without the term involving $\widehat{\rho}$.
+  • plugin_no_x: The same estimators as plugin but without cross-fitting.
+The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (9).
+Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$.
+For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction.
+For the score estimator we use the location-scale based estimator described in Section S4.6, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
+We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$.
+In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$.
+We then consider the following four generative models
+  • Binary $L$ with partially linear $Y$:
+    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=L+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
+  • Binary $L$ with nonparametric $Y$:
+    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
+  • Continuous $L$ with partially linear $Y$:
+    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
+  • Continuous $L$ with nonparametric $Y$:
+    $L=B(W_{1})+(1+B(W_{1}))\xi\quad\text{and}\quad Y=2B(W_{1})L+B(W_{1})+\varepsilon$
+The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$).
+The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$.
+For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above.
+We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=3$ in Figure S6, for $d=15$ in Figure S7 and for $d=75$ in Figure S8.
+Figure S6: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section S5.2 with $d=3$.
+PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting.
+NPM works in both settings but can be less robust for small sample sizes.
+Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
+Figure S7: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S5.2 with $d=15$.
+The patterns observed in Figure S6 are repeated here.
+Figure S8: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S5.2 with $d=75$.
+The patterns observed in Figure S6 are repeated here.
+In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated.
+We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$.
+In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g.
+for $n=1000$ and binary $L$.
+In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
+S5.3 Description of the construction of groups in Section 4.1.1
+In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.1.1.
+On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
+$(\widehat{\pi}_{R}(\mathrm{White}),\widehat{\pi}_{R}(\mathrm{Black}),\widehat{\pi}_{R}(\mathrm{Other}))$ $=(0.855,0.096,0.049),$
+$(\widehat{\pi}_{E}(0),\widehat{\pi}_{E}(1))$ $=(0.454,0.546),$
+$(\widehat{\pi}_{C}(0),\widehat{\pi}_{C}(1))$ $=(0.761,0.239),$
+respectively.
+Now, for the two types of groupings we define target probabilities $p_{i}=(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ for each individual $i$ (see below).
+Based on these probabilities we randomly draw for each individual $i$ one of the three categories using the probabilities $p_{i}$.
+Finally, we randomly group individuals with the same categories into communities of approximately $50$ (as many communities of $50$ and the remaining with $49$) and aggregate the predictors and response as described in the main text.
+The probabilities $(p_{i}^{0},p_{i}^{1},p_{i}^{2})\in\Delta^{2}$ are selected, in each of the two settings, as follows.
+Grouping by race and education: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
+$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{E}^{1}(\texttt{education}_{i}),$
+where $\texttt{race}_{i}$ and $\texttt{education}_{i}$ denote the race and education level of the $i$th individual, $p_{R}^{1}$ and $p_{E}^{1}$ are probability vectors satisfying
+$p_{R}^{1}(\mathrm{White})\propto\frac{0.8}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{1}(\mathrm{Black})\propto\frac{0.02}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{1}(\mathrm{Other})\propto\frac{0.18}{\widehat{\pi}_{R}(\mathrm{Other})},$ (S34)
+and
+$p_{E}^{1}(0)\propto\frac{0.95}{\widehat{\pi}_{E}(0)},\quad p_{E}^{1}(1)\propto\frac{0.05}{\widehat{\pi}_{E}(1)}.$
+That is, category $1$ is about as diverse ($(0.8,0.02,0.18)$ is about as diverse as $\widehat{\pi}_{R}$) and with lower education ($0.05<\widehat{\pi}_{E}(1)$).
+The probability $p_{i}^{2}$ of observation $i$ belonging to category $2$ is set to
+$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{E}^{2}(\texttt{education}_{i}),$
+where similarly $p_{R}^{2}$ and $p_{E}^{2}$ are probability vectors but now satisfying
+$p_{R}^{2}(\mathrm{White})\propto\frac{0.65}{\widehat{\pi}_{R}(\mathrm{White})},\quad p_{R}^{2}(\mathrm{Black})\propto\frac{0.12}{\widehat{\pi}_{R}(\mathrm{Black})},\quad p_{R}^{2}(\mathrm{Other})\propto\frac{0.23}{\widehat{\pi}_{R}(\mathrm{Other})}$ (S35)
+and
+$p_{E}^{2}(0)\propto\frac{0.3}{\widehat{\pi}_{E}(0)},\quad p_{E}^{2}(1)\propto\frac{0.7}{\widehat{\pi}_{E}(1)}.$
+That is, category $2$ is more diverse ($(0.65,0.12,0.23)$ is more diverse than $\widehat{\pi}_{R}$) and with higher education ($0.7>\widehat{\pi}_{E}(1)$).
+Lastly, the probability $p_{i}^{0}$ for category 00 is set to $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
+Grouping by race and compensation: The probability $p_{i}^{1}$ of observation $i$ belonging to category $1$ is set to
+$p_{i}^{1}:=p_{R}^{1}(\texttt{race}_{i})\cdot p_{C}^{1}(\texttt{compensation}_{i}),$
+where $p_{R}^{1}$ is as in (S34) and $p_{C}^{1}$ is a probability vector satisfying
+$p_{C}^{1}(0)\propto\frac{0.9}{\widehat{\pi}_{C}(0)},\quad p_{C}^{1}(1)\propto\frac{0.1}{\widehat{\pi}_{C}(1)}.$
+That is, category $1$ is about as diverse and with lower compensation ($0.05<\widehat{\pi}_{C}(1)$). The probability $p_{i}^{2}$ for category $2$ is set to
+$p_{i}^{2}:=p_{R}^{2}(\texttt{race}_{i})\cdot p_{C}^{2}(\texttt{compensation}_{i}),$
+where $p_{R}^{2}$ is as in (S35) and $p_{C}^{2}$ is a probability vector satisfying
+$p_{C}^{2}(0)\propto\frac{0.7}{\widehat{\pi}_{C}(0)},\quad p_{C}^{2}(1)\propto\frac{0.3}{\widehat{\pi}_{C}(1)}.$
+That is, category $2$ is more diverse and with higher compensation ($0.3>\widehat{\pi}_{C}(1)$). Lastly, we set the probability $p_{i}^{0}$ for category 00 is $p_{i}^{0}:=1-p_{i}^{1}-p_{i}^{2}$.
+S5.4 Illustration of semi-synthetic data from Section 4.1.1
+Figure S9 and Figure S10 illustrate the distribution of the semi-synthetic data in Section 4.1.1.
+We see that the overall relationship between diversity and compensation is negative but conditioned on $W=\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
+Figure S9: Summary of the semi-synthetic ‘Adult’ dataset when grouping on education.
+Left: Racial compositions in a ternary plot.
+Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
+Right: Plot of average compensation per group ($Y$) against one minus the Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
+Depending on the color the dependence between $Y$ and the Gini coefficient changes.
+Figure S10: Summary of the semi-synthetic ‘Adult’ dataset when grouping on compensation.
+Left: Racial compositions in a ternary plot.
+Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
+Right: Plot of average compensation per group ($Y$) against one minus the Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle).
+Depending on the color the dependence between $Y$ and the Gini coefficient changes.
+S5.5 Semi-synthetic individual-level experiment using groups from Section 4.1.1
+In this section we consider a variant of the experiment performed in Section 4.1.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups.
+More specifically, we generate the groups as before and estimate the probability that $Y=1$ within each group and use this probability to generate new $Y$s for each observation.
+For each individual, we thus have a response $Y$, a race $R$, a ‘contextual’ variable $Z$ that gives the racial distribution of each group and additional covariates $X$.
+By construction of the groups and sampling of the $Y$, we know that there is a positive effect of diversity on the income.
+We now apply the naive_diversity and $\mathrm{CDI}_{\mathrm{Gini}}$ methods as in Section 4.1.1 except we always control for the race of inviduals and omit the unstable estimators that condition on $Z$.
+The results can be seen in Table S3.
+Method Grouping on compensation
+Estimate $95\%$ CI
+[…]
+[…]
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.459}}$ $(0.247,0.671)$
+$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.488}}$ $(0.282,0.694)$
+Table S3: Table of estimates in the individual-level ‘Adult’ data experiment.
+Bold red numbers indicate a significant negative effect of diversity at a 5% level, while bold green indicates a significant positive effect.
+The naive diversity method indicates that there is a significant negative effect of diversity on compensation, even when controlling for individual race ($R$) and additional covariates ($X$), despite the fact that the data is constructed with a positive effect of diversity.
+The CDI correctly identifies that the effect is positive.
+We see that, as in Section 4.1.1, the naive_diversity method incorrectly estimates a negative effect of diversity on income while $\mathrm{CDI}_{\mathrm{Gini}}$ correctly identifies that the effect is positive.
+S5.6 Description of additional control variables used in Section 4.1.2
+The additional background variables in the New York schools data used in Section 4.1.2 are described below. Unless otherwise mentioned, no preprocessing was used.
+School income estimate
+The school income estimates were grouped into four categories; one for missing values and three more based on quantiles of the observed values.
+School grade range (grade referring to years of school)
+The range of grades for the schools were grouped into three categories; one for ‘high grades’ where the lowest grade was greater or equal to 6th, one for ‘low grades’ where the highest grade was less than or equal to 4th and the remaining schools were put in a ‘mixed’ category.
+Student attendance rate
+This number was computed as the total number of days attended by all students / total number of days on register for all students.
+% chronically absent students
+The percentage of students missing 10% of school days or at least 18 days per year in a 180-day school year.
+Economic Need Index
+Computed as the percentage of students in temporary housing plus 0.5 times the percentage of students that are eligible for Health Reimbursement Arrangements plus 0.5 times the percentage of students eligible for free lunch.
+Community school
+An indicator of whether the school is a community school.
+% English language learners
+Percentage of students who are learning English for the first time.
+Rigorous Instruction %
+School rating for ‘how well the curriculum and instruction engage students, build critical-thinking skills, and are aligned to the Common Core’.
+Collaborative Teachers %
+School rating for ‘how well teachers participate in opportunities to develop, grow, and contribute to the continuous improvement of the school community’.
+Supportive Environment %
+School rating for ‘how well the school establishes a culture where students feel safe, challenged to grow, and supported to meet high expectations’.
+Effective School Leadership %
+School rating for ‘how well school leadership inspires the school community with a clear instructional vision and effectively distributes leadership to realize this vision’.
+Strong Family-Community Ties %
+School rating for ‘how well the school forms effective partnerships with families to improve the school’.
+Trust %
+School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
+S5.7 Additional plots from gut microbiome analysis in Section 4.2
+In this section we provide additional figures showing our results from the analysis of the gut microbiome data in Section 4.2.
+To investigate how the log-contrast approach compares to the perturbation-based approach, we plot the most significant hits using $\mathrm{CFI}_{\mathrm{mult}}$ (with random forests as regression method) and a log-contrast regression model in the left plot of Figure S11.
+These are comparable quantities since $\mathrm{CFI}_{\mathrm{mult}}^{j}$ is equal to the $j$th coefficient in a log-contrast model when it is correctly specified and no zeros are present.
+We omit plotting the $\ell^{1}$-penalized log-contrast estimates as these are broadly similar to the standard log-contrast estimates in this example.
+We also plot a marginal effect of $L=\log(Z^{j}/(1-Z^{j}))$ which corresponds to predicting $Y$ using OLS on $L$ and correcting for the presence of zeros as we mention in Remark 1.
+The results with corresponding $95\%$-confidence intervals are given in the right plot of Figure 3; top $10$ most significant species sorted according to significance of $\mathrm{CFI}_{\mathrm{mult}}$ (left) and according to log-contrast coefficients (right).
+Figure S11: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ ( right).
+The labels are colored by the size of the $p$-value.
+The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model.
+While the effects overlap for some of the features, there are also significant differences.
+In Figure S12, we plot the most significant hits for $\mathrm{CFI}{\mathrm{unit}}$ and $\mathrm{CKE}$ both estimated using random forests together with $\mathrm{CFI}_{\mathrm{mult}}$ estimates.
+We plot the estimates scaled by the estimated standard deviation of $L$ to put the effects on the same scale (we do not correct the confidence intervals for the estimation of this standard deviation).
+Furthermore, we flip the sign of $\mathrm{CKE}$ to keep the interpretation similar to $\mathrm{CFI}$, that is, that we are looking at the effect of increasing $Z^{j}$.
+Figure S12: Effect estimates scaled by the estimated standard deviation of $L$ for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{unit}}$ (left) and $\mathrm{CKE}$ (right).
+The labels are colored by the size of the $p$-value.
+For the $\mathrm{CFI}_{\mathrm{unit}}$ results, the first $9$ species remain below $0.05$ after Bonferroni-correction while this is only true for $2$ for the $\mathrm{CKE}$.
+There is little overlap between the selected species across the two methods.

### v5

@@
 We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data.
-Keywords: compositional data, effect estimation, causality, perturbation, feature importance, diversity effects, nonparametrics, mixture experiments
+Keywords: feature importance, causality, nonparametrics, semiparametrics
 1 Introduction
@@
 Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982).
-We summarize the historical development of these fields in Section S2.1 of the supplementary material.
+We summarize the historical development of these fields in Section S2.1 (Sections beginning with ‘S’ refer to the supplement throughout the paper).
 A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates.
@@
 Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity.
-However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Section 4.1).
+However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects.
 In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
 Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function of $x$.
-Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$.
+Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$.
 Then, for any $x\in\mathbb{R}^{d}$, we have
@@
 Similar observations are possible regarding other interpretable effect measures for unconstrained covariates and under more general nonlinear models.
-Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined.
-The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$.
-We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures.
-In Section S2.2 of the supplementary material we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
+In the case of compositional $Z$, we will need to use perturbations that are tailored to the simplex structure of $Z$ instead of $\psi$ given above.
+In Section S2.2 we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
 A perturbation, formally defined in Section 2, is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed.
-For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
+For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$, $\psi(z,1)^{j}=0$ and $\psi(z,1)^{-j}=\mathsf{C}(z^{-j})$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one.
 By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$.
@@
 Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities.
-To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3 of the supplementary material.
+To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3.
 While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional.
@@
 Our perturbation-based framework is formally introduced and developed in Section 2.
-We describe how to construct perturbations in Section 2.1 and provide a table of pre-defined simplex effects along with examples of their use in Section 2.2.
-In Section 3 we give an overview of the semiparametric techniques used to estimate our proposed effects.
+We describe how to construct perturbations and provide a table of pre-defined simplex effects in Section 3.
+In Section 4 we describe the semiparametric techniques used to estimate our proposed effects.
 In Section 5 we provide a semi-synthetic and two real world applications of our proposal.
@@
 Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
-All references to sections and results beginning with ‘S’ point to the supplementary material.
 2 Target parameters via perturbations
@@
 $\psi(z,\gamma)=z+\gamma s_{\psi}(z)v_{\psi}(z).$ (1)
-We construct interpretable perturbations in Section 2.1.2.
+We construct interpretable perturbations in Section 3.2.
 Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex.
@@
 To motivate the upcoming condition, we consider what happens when $Z$ takes values in $\mathbb{R}^{d}$.
-In that case, we can write
+In that case, we can write (with $\omega_{\psi}$ defined in Definition 1)
 $\partial_{\gamma}f(\psi(z,\gamma))\big{|}_{\gamma=0}=\nabla f(z)^{\top}\omega_{\psi}(z).$
@@
 From this result it follows that if we are able to find a derivative-isolating reparametrization the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993).
-It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
-$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (5)
-Details on the different estimation approaches are discussed in Section 3.
+It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that $\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ Details on the different estimation approaches are discussed in Section 4.
 Remark 1 (zero speeds).
@@
 In Section S4.5, we describe how to modify the estimation procedures introduced in Section 4 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
-2.1 Perturbations on the simplex
+3 Perturbations on the simplex
 We now turn to the construction of perturbations on the simplex.
@@
 For both approaches, we provide – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
-2.1.1 Endpoints in the simplex
+3.1 Endpoints in the simplex
 As formally defined in Definition 1, an endpoint function $E:\Delta^{d-1}\rightarrow\Delta^{d-1}$, if it exists, determines where the perturbation ends when $\gamma$ is increased as much as possible.
@@
 perturbations that push towards a vertex of the simplex.
-2.1.2 Directions, speeds and derivative-isolating reparametrizations
+3.2 Directions, speeds and derivative-isolating reparametrizations
 When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as $\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$ An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
@@
 that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint.
-Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed.
+To avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed.
 In Section S2.7 we describe how working in other geometries can lead to different notions of direction.
 Taking the direction in (5) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect.
-We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed.
-In both cases, the construction of a derivative-isolating reparametrization is similar.
 As a fundamental building block for the general construction, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
@@
 Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction in Section S2.5 to start from either a given interpretable speed or a summary statistic.
-We illustrate both options in the following two examples.
+3.3 Effects of changes in individual components and diversity
+In this section, we give examples of how to construct directional perturbations which summarize perturbation effects of practical interest.
 Example 1 (Feature influence).
@@
 $\|\partial_{c}\mathsf{C}(X\odot(1+ce_{j}))\bigm{|}_{c=0}\|_{1}=2\tfrac{X^{j}}{\|X\|_{1}}\left(1-\tfrac{X^{j}}{\|X\|_{1}}\right),$
-which is scale-invariant and therefore meaningful to consider on the simplex. On the simplex, this corresponds to the speed
-$s_{\psi}(z)=2z^{j}(1-z^{j})$
-which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
+which is scale-invariant and therefore meaningful to consider on the simplex.
+On the simplex, this corresponds to the speed $s_{\psi}(z)=2z^{j}(1-z^{j})$ which we call multiplicative speed.
+To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
 $\phi=(\phi^{L},\phi^{W}):z\mapsto\left(\log\left(\frac{z^{j}}{1-z^{j}}\right),v_{\psi}(z)\right)$
 is a derivative-isolating reparametrization for the $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation $\psi$ with multiplicative speed.
+Perturbation effects that capture changes in individual components, are, for example, useful as a variable selection tool in compositional data.
+In Section 5.2, we illustrate this for a human gut microbiome study.
+For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured.
+We then want to determine for which microbes a high abundance is associated with a high BMI.
+The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
+We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$.
+We denote by $\mathrm{CFI}^{j}_{\mathrm{unit}}$ the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ the perturbation that moves at multiplicative speed.
+$\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al.
+(2023), however without an efficient estimation procedure.
+A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$.
+As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
+In datasets with many zeros, which is common for example in microbiome data, it can also be interesting to focus only on the effect of zeros in individual components.
+This can be achieved by considering a binary perturbation $\psi$ given by $\psi(z,1)^{j}=0$ and $\psi(z,1)^{-j}=\mathsf{C}(z^{-j})$.
+We call the resulting average binary perturbation effect $\lambda_{\psi}$ the compositional knock-out effect for the $j$th feature, or $\mathrm{CKE}^{j}$ for short.
+The $\mathrm{CKE}^{j}$ allows for a straightforward analysis of zeros in compositional data, which has been a long-standing problem with many existing approaches.
 Example 2 (Diversity with Gini-speed).
@@
 Similar arguments could be applied to other popular measures of diversity such as the Shannon entropy or the Gini–Simpson index.
-Figure S1 contains an illustration of the different perturbations we have considered in this work.
-2.2 Perturbation effects on the simplex
-To make our proposal more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1.
-We motivate each perturbation effect with an example application to showcase potential use cases.
-In Section S2.6 we describe an additional class of perturbation effects based on general amalgamating perturbations.
-2.2.1 Effects of changes in individual components
-As a first example we consider perturbation effects that capture how a response is associated with changes in individual components.
-These types of effects are, for example, useful as a variable selection tool in compositional data.
-In Section 4.2, we illustrate this for a human gut microbiome study.
-For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured.
-We then want to determine for which microbes a high abundance is associated with a high BMI.
-The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
-Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$.
-We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally.
-We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$.
-We already considered these perturbations in Example 1 where we examined two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed.
-$\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al.
-(2023), however without an efficient estimation procedure.
-Huang et al.
-(2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S2.8).
-A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$.
-As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
-In datasets with many zeros, which is common for example in microbiome data, it can also be interesting to focus only on the effect of zeros in individual components.
-This can be achieved by considering a binary perturbation $\psi$ given by $\psi(z,1)^{j}=0$ and $\psi(z,1)^{-j}=\mathsf{C}(z^{-j})$.
-We call the resulting average binary perturbation effect $\lambda_{\psi}$ the compositional knock-out effect for the $j$th feature, or $\mathrm{CKE}^{j}$ for short.
-The $\mathrm{CKE}^{j}$ allows for a straightforward analysis of zeros in compositional data, which has been a long-standing problem with many existing approaches.
-2.2.2 Effects of changes in diversity
-As a second example we consider perturbation effects that capture how a response is associated with increases in diversity.
-Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions.
+The effect of increasing diversity is often relevant in social sciences and econometrics, e.g., when considering racial or income distributions.
 As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response.
-Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$).
+Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 3.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$).
 We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 2 such that $L$ is equal to one minus the Gini coefficient.
-The construction in Example 2 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics.
-In Section 4.1.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression.
-In Section 4.1.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
-We do not define a binary centering perturbation here as it requires a positive probability of observing $Z=z_{\mathrm{cen}}$, which is rarely the case.
+In Section 5.1 we give a semi-synthetic and real data example of the application of $\mathrm{CDI}_{\mathrm{Gini}}$.
+In Section S2.6 we discuss average perturbation effects that capture the effect of applying amalgamations to $Z$ on $Y$ (see Section 3.1 for a definition of a general amalgamation).
+We summarize our predefined perturbation effects in Table 1.
 Effect of changes in Target Speed $s(z)$ Endpoint $E(z)$ $\phi^{L}(z)$
@@
 $\mathrm{CDI}_{\mathrm{Gini}}$ see (7) $z_{\mathrm{cen}}$ $1-G(z)$, see (6)
+amalgamations $\mathrm{CAI}^{A\to B}_{\mathrm{unit}}$ $1$ $\mathsf{A}_{A\to B}(z)$ $-2\|z^{A}\|_{1}$
+$\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ $\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}$ $\mathsf{A}_{A\to B}(z)$ $\log\big{(}\tfrac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\big{)}$
+$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
 Table 1: List of perturbation effects considered in this work and what effects they capture.
@@
 The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest.
-For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
-$\tau_{\psi}=\mathbb{E}[\partial_{\ell}f(\ell,W)\bigm{|}_{\ell=L}],$
-where $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$.
-A naive approach, which is known as plug-in estimation, is to estimate $\tau_{\psi}$ by first learning a differentiable estimate $\widehat{f}$ of $f$ using all $n$ observations and then computing
-$\widetilde{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i}).$
-While such an estimate can be consistent, it is rarely asymptotically Gaussian, which invalidates standard asymptotic inference.
+For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of $\tau_{\psi}=\mathbb{E}[\partial_{\ell}f(\ell,W)\bigm{|}_{\ell=L}],$ where $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$.
+A naive approach, which is known as plug-in estimation, is to estimate $\tau_{\psi}$ by first learning a differentiable estimate $\widehat{f}$ of $f$ using all $n$ observations and then computing $\widetilde{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i}).$ While such an estimate can be consistent, it is rarely asymptotically Gaussian, which invalidates standard asymptotic inference.
 Using the corresponding plug-in estimate of the asymptotic variance, given by
@@
 Figure 1 contains an illustration of the coverage of confidence intervals resulting from target parameter and variance estimates from the different methods mentioned above.
-3.2 Efficient estimation of perturbation effects
+4.2 Partially linear estimation of perturbation effects
 In this section, we present our proposals for estimating perturbation effects.
-Further details and full algorithms can be found in Section S4.1.
-All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
-3.2.1 Nonparametric estimation of the binary perturbation effect
-Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
-$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}.$
-The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities).
-The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature.
-The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
-$\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i},$
-where $\widehat{f}$ is an estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\pi}$ is an estimator of the propensity score $\pi:w\mapsto\mathbb{P}(L=1\,|\,W=w)$.
-An efficient estimator for $\widehat{\lambda}_{\psi}$ is thus
-$\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (10)
-3.2.2 Nonparametric estimation of the directional perturbation effect
-Consider a directional perturbation $\psi$ with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and the corresponding perturbation effect $\tau_{\psi}$.
-By the arguments in Section 2, setting $L:=\phi^{L}(Z)$ and $W:=\phi^{W}(Z)$, we have that
-$\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}].$
-This quantity is another well-known target of inference; the average partial effect (Newey and Stoker, 1993; Chernozhukov et al., 2022; Klyne and Shah, 2023).
-The one-step corrected estimator in this setting is given by
-$\widehat{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i})-\widehat{\rho}(L_{i},W_{i})(Y_{i}-\widehat{f}(L_{i},W_{i}))$ (11)
-where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$.
-For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate).
-This excludes the direct use of many popular machine learning methods such as random forests or boosted trees.
-One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed.
-Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023).
-Here, we use an alternative smoothing method which uses random forests and local polynomials.
-Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model.
-This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1).
-Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.5 and S4.6, respectively.
-3.2.3 Partially linear estimation of perturbation effects
-While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.3 for an illustration of this sensitivity).
+It is possible to estimate average perturbation effects under mild smoothness assumptions entirely nonparametrically.
+We describe these estimators in Section S4.1.
+However, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.4 for an illustration of this sensitivity).
 One way to avoid these complicated estimations is by additionally assuming a partially linear model.
-For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
-$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (12)
-With this assumption, $\theta$ is exactly equal to the perturbation effect.
+For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that $\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ With this assumption, $\theta$ is exactly equal to the perturbation effect.
 Efficient estimation of $\theta$ in a partially linear model is a well-studied problem (e.g., Robinson, 1988; Chernozhukov et al., 2018) and one estimator (sometimes referred to as the partialling out estimator) is given by
@@
 In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S4 in Section S4.4.3 for a proof).
+Further details on both the partialling out and nonparametric estimators (including full algorithms) can be found in Section S4.2.
 5 Numerical experiments
@@
 This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption.
-The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 3.2.3) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates.
+The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 4.2) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates.
 The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect.
@@
 The resulting dataset consists of 4,581 observations of $\mathrm{BMI}$ measurements $Y\in\mathbb{R}$ and the relative abundances of $561$ microbial species $Z\in\Delta^{560}$.
-We begin by evaluating, how sensitive the proposed perturbation effect estimates are with respect to the employed nuisance function estimators.
-To this end, we consider three different regression methods: random forest (rf), neural network (mlp) and kernelized support vector regression using an RBF kernel (svr).
-Each of the regressions is combined with a grid-search cross-validation to automatically tune the hyperparameters of each method.
-For each of the three regression methods and all the species we then compute $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ using the estimator based on a partially linear model and the $\mathrm{CKE}$ using the nonparametric estimator.
-As baselines, we additionally compute a nonparametric $R^{2}$ (Williamson et al., 2021) and permutation importance (Breiman, 2001).
-Figure 2 shows the Spearman correlation between the estimates from the different regression methods.
-Figure 2: Spearman correlations of variable importance measures computed on each feature using different regressions.
-On the left, these are computed on all $561$ features and on the right they are computed on the subset of features where at least one of the two methods deems the effect significant at a $5\%$ level.
-The $R^{2}$ and permutation methods do not come with significance tests and are therefore omitted from the right plot.
-While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not.
-The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator.
-For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate.
-Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021).
-Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model.
-One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
-Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option.
-To this end we consider a standard approach for estimating effects of coordinates for unconstrained data.
-More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint.
-For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression.
-As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species.
-In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
-We now analyze different types of average perturbation effects and compare them with the commonly used log-contrast based analysis.
-We fit both a standard log-contrast regression and an $\ell^{1}$-penalized log-contrast regression and then use the coefficients as feature influences.
+In Section S5.7 we show that non-compositional variable importance measures such as permutation-based feature importance or nonparametric $R^{2}$ exhibit undesirable properties when applied to this data.
+In particular, there is little correlation between important features identified by applying different nonparametric regression methods.
+As a consequence of these findings, we focus on analyzing different types of average perturbation effects and comparing them with commonly used compositional feature importance measures, namely those based on log-contrast models.
+The linear log-contrast model assumes that there exists $\beta\in\mathbb{R}^{d}$ such that
+$\mathbb{E}[Y\,|\,Z]=\sum_{j=1}^{d}\beta^{j}\log(Z^{j})\quad\text{with}\quad\sum_{j=1}^{d}\beta^{j}=0.$
+We fit this model and an $\ell^{1}$-penalized version of the same model and seek to use the coefficients as feature influences.
 Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex.
@@
 The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model.
-The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators.
+The second row corresponds to estimates of two of the target parameters given in Table 1 with random forest (rf) as nuisance function estimators.
 Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect.
@@
 The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect.
-Details on both experiments are provided in Section S5.1 of the supplementary material.
+Details on both experiments are provided in Section S5.1.
 Example S2 (Effects of increased diversity).
@@
 S2.5 Finding derivative-isolating reparametrizations
-In this section we provide a general method for finding derivative-isolating reparametrizations based on the development in Section 2.1.2. Our main result is as follows.
+In this section we provide a general method for finding derivative-isolating reparametrizations based on the development in Section 3.2. Our main result is as follows.
 Proposition S1 (Finding derivative-isolating reparametrizations).
@@
 Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized.
-In the following two examples, we give the missing details from Examples 1 and 2 from Section 2.1.2.
+In the following example, we give the missing details from Examples 1 and 2.
 Example S3.
@@
 $s_{\psi}(z):=\frac{2d}{\sum_{j=1}^{d}\sum_{k=1}^{d}|v_{\psi}(z)^{j}-v_{\psi}(z)^{k}|}.$
-S2.6 Amalgamations
-In this section we consider general amalgamating perturbations as defined in Section 2.1.1 rather than only those which push towards a vertex of the simplex, i.e., those we describe in Section 2.2.1.
+S2.6 Amalgamation effects
+In this section we consider general amalgamating perturbations as defined in Section 3.1 rather than only those which push towards a vertex of the simplex, i.e., those we describe in Section 3.3.
 We first generalize Example 1.
@@
 Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness.
-As in Section 2.2.1, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
+As in Section 3.3, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$.
 However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally.
 For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic.
-To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports.
+To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 3.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports.
 Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$.
@@
 We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
-We summarize these target parameters in Table S2 which complements Table 1.
-amalgamations $\mathrm{CAI}^{A\to B}_{\mathrm{unit}}$ $1$ $\mathsf{A}_{A\to B}(z)$ $-2\|z^{A}\|_{1}$
-$\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ $\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}$ $\mathsf{A}_{A\to B}(z)$ $\log\big{(}\tfrac{\|z^{B}\|_{1}}{\|z^{A}\|_{1}}\big{)}$
-$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
-Table S2: List of amalgamating perturbation effects considered in this work and what effects they capture.
-For each perturbation effect the corresponding speed, endpoint and $L$-term of the reparametrization used for efficient estimation is included.
-All reparametrizations use the canonical direction $v_{E}$ based on the endpoint, see (6).
-The $W$-term $\phi^{W}(z)$ of the reparametrization is $E(z)$ for the binary effects and $(v_{E}(z),E(z))$ for the directional effects.
 S2.7 Perturbations arising from alternative simplex geometries
-In this section we describe an alternative way to construct perturbations than the one considered in Sections 2.1.1 and 2.1.2.
-When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
+In this section we describe an alternative way to construct perturbations than the one considered in Sections 3.1 and 3.2.
+When constructing the directional perturbations in Section 3.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$.
 While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions.
@@
 As a first example, assume that the SCM given in Figure S2 (left) generates the data.
-Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 2.1.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
+Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 3.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
 (a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
@@
 In this section we provide details on the theory and algorithms for the estimation of perturbation effects.
+Section S4.1 introduces the nonparametric estimators of perturbation effects.
 Section S4.2 contains a summary of the theoretical results and algorithms for the main estimation procedures.
@@
 Section S4.7 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
-S4.1 Theoretical analysis of estimators
-In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (12).
+S4.1 Nonparametric estimation of perturbation effects
+S4.1.1 Binary perturbation effect
+Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
+$\lambda_{\psi}=\frac{\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]-\mathbb{E}\left[Y\right]}{\mathbb{P}(L=0)}.$
+The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities).
+The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect [e.g., Robins et al., 1994, Chernozhukov et al., 2018, Kennedy, 2023] that has been analyzed extensively in the causal inference literature.
+The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
+$\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i},$
+where $\widehat{f}$ is an estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\pi}$ is an estimator of the propensity score $\pi:w\mapsto\mathbb{P}(L=1\,|\,W=w)$.
+An efficient estimator for $\widehat{\lambda}_{\psi}$ is thus
+$\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (S8)
+S4.1.2 Directional perturbation effect
+Consider a directional perturbation $\psi$ with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and the corresponding perturbation effect $\tau_{\psi}$.
+By the arguments in Section 2, setting $L:=\phi^{L}(Z)$ and $W:=\phi^{W}(Z)$, we have that
+$\tau_{\psi}=\mathbb{E}[\partial_{\ell}\mathbb{E}[Y\,|\,L=\ell,W]\big{|}_{L=\ell}].$
+This quantity is another well-known target of inference; the average partial effect [Newey and Stoker, 1993, Chernozhukov et al., 2022, Klyne and Shah, 2023].
+The one-step corrected estimator in this setting is given by
+$\widehat{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i})-\widehat{\rho}(L_{i},W_{i})(Y_{i}-\widehat{f}(L_{i},W_{i}))$ (S9)
+where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$.
+For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate).
+This excludes the direct use of many popular machine learning methods such as random forests or boosted trees.
+One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed.
+Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah [2023].
+Here, we use an alternative smoothing method which uses random forests and local polynomials.
+Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah [2023] and assume a location-scale model.
+This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1).
+Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.6 and S4.7, respectively.
+S4.2 Theoretical analysis of estimators
+In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 4; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model.
 These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere [e.g., Klyne and Shah, 2023, Chernozhukov et al., 2018].
@@
 When we write conditional expectations given a fitted regression function $\widehat{f}$ we mean a conditional expectation given the random variables used to fit $\widehat{f}$ and any additional randomness involved in the fitting.
-S4.1.1 Nonparametric estimation for directional effects
+S4.2.1 Nonparametric estimation for directional effects
 We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem S1 (a).
-A high-level overview of this estimation was given in Section 3.2.2.
+A high-level overview of this estimation was given in Section S4.1.2.
 More formally, assume we have a dataset of $n$ i.i.d.
@@
     $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\rho}(L_{i},W_{i})-\rho_{P}(L_{i},W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
-S4.1.2 Nonparametric estimation for binary effects
+S4.2.2 Nonparametric estimation for binary effects
 We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem S1 (b).
-A high-level overview of this estimation was provided in Section 3.2.1.
+A high-level overview of this estimation was provided in Section S4.1.1.
 More formally, assume we have a dataset of $n$ i.i.d.
@@
     $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{f}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{\pi}(W_{i})-\pi_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
-S4.1.3 Partially linear model for binary/directional effects
+S4.2.3 Partially linear model for binary/directional effects
 Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem S1 (c).
-A high-level overview of this estimation was provided in Section 3.2.3.
+A high-level overview of this estimation was provided in Section 4.2.
 More formally, assume we have a dataset of $n$ i.i.d.
@@
     $\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{m}(W_{i})-m_{P}(W_{i})\right\}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\left\{\widehat{g}(W_{i})-g_{P}(W_{i})\right\}^{2}\right)=o_{\mathcal{P}}(n^{-1})$
-S4.2 Proof of Theorem S1
+S4.3 Proof of Theorem S1
 In this section, we prove Theorem S1 and some additional auxiliary results.
-S4.2.1 Proof of (a)
+S4.3.1 Proof of (a)
 Proof.
@@
 $\geq\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
-S4.2.2 Proof of (b)
+S4.3.2 Proof of (b)
 Proof.
@@
 and $\sigma^{2}_{\lambda_{P}}p_{P}^{4}\geq c(1-c)$ by (S20), the result now follows by Lemma S4 (a) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
-S4.2.3 Proof of (c)
+S4.3.3 Proof of (c)
 Proof.
@@
 The conclusion now follows from Assumption S3 (a) and (b), Lemma S4 (b) and a uniform version of Slutsky’s theorem [Shah and Peters, 2020, Lemma 20].
-S4.2.4 Auxiliary results
+S4.3.4 Auxiliary results
 Lemma S1.
@@
 ∎
-S4.3 Robustness of nuisance function estimation for different estimators
+S4.4 Robustness of nuisance function estimation for different estimators
 A key motivation for applying the estimators based on a partially linear model on $Y$ is to avoid estimating nuisance functions that result in unstable estimators.
 We illustrate this point based on three examples below.
-S4.3.1 Estimation of nuisance functions with continuous $L$
-To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section 3.2.2) compared to $m$ and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
+S4.4.1 Estimation of nuisance functions with continuous $L$
+To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section S4.1.2) compared to $m$ and $g$ (see Section 4.2) we estimate the functions on simulated data and compare the estimates to the truth.
 We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
@@
 In contrast, the derivative and score estimates, are much worse at capturing the more complex behavior (with clear biases conditioned on $W^{1}$).
-S4.3.2 Estimation of nuisance functions with binary $L$
-To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section 3.2.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth.
+S4.4.2 Estimation of nuisance functions with binary $L$
+To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section S4.1.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 4.2) we estimate the functions on simulated data and compare the estimates to the truth.
 We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
@@
 We see that the $m$, $g$ and $f$ estimates are reasonable estimates of the true functions whereas the inverse propensity (even when cutting off extreme values at $100$) is poorly estimated for $W^{1}$ close to zero.
-S4.3.3 Partially linear model estimate when partial linearity is violated
+S4.4.3 Partially linear model estimate when partial linearity is violated
 The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified.
@@
 We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
-S4.4 Computing directional perturbation effects in the presence of zero speeds
+S4.5 Computing directional perturbation effects in the presence of zero speeds
 In Proposition 1 we only considered directional perturbations where $\omega_{\psi}(Z)\neq 0$ almost surely.
@@
 and the result thus follows by Lemma S4 (a) by the bounds on the moments and variance of $(D,S)$. ∎
-S4.5 Estimating differentiable regression functions
+S4.6 Estimating differentiable regression functions
 The following is a description of the smoothing method we apply in the nonparametric estimator of the average directional perturbation effect.
@@
 8:return smoothed predictions $(\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$ and derivatives $(\partial_{\ell}\widehat{g}_{s}(L_{i},W_{i}))_{i\in[n]}$
-S4.6 Location-scale modelling of the score function
+S4.7 Location-scale modelling of the score function
 The following is a description of the estimator of the conditional score based on a location-scale model proposed by Klyne and Shah [2023].
@@
 Section S5.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 5.1.2.
+Section S5.7 describes an additional experiment on the microbiome data analyzed in Section 5.2.
 Section S5.8 contains additional plots illustrating results from the microbiome data analysis in Section 5.2.
@@
 from i.i.d. data $(Y_{1},L_{1},W_{1}),\ldots,(Y_{n},L_{n},W_{n})$. For both types of average perturbation effects we consider the following seven types of estimators.
-  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (10) and (11).
+  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (S8) and (S9).
 See Algorithms S1 and S2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
@@
   • plugin: The naive plug-in versions of the NPM estimates.
-In the binary effect case, the estimator in (10) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (11) without the term involving $\widehat{\rho}$.
+In the binary effect case, the estimator in (S8) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (S9) without the term involving $\widehat{\rho}$.
   • plugin_no_x: The same estimators as plugin but without cross-fitting.
@@
 School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
-S5.7 Additional plots from gut microbiome analysis in Section 4.2
+S5.7 Additional experiment on microbiome data in Section 5.2
+In this section we describe an additional experiment on the microbiome data from Section 5.2.
+Our goal is to determinee how sensitive the proposed perturbation effect estimates are with respect to the employed nuisance function estimators.
+To this end, we consider three different regression methods: random forest (rf), neural network (mlp) and kernelized support vector regression using an RBF kernel (svr).
+Each of the regressions is combined with a grid-search cross-validation to automatically tune the hyperparameters of each method.
+For each of the three regression methods and all the species we then compute $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ using the estimator based on a partially linear model and the $\mathrm{CKE}$ using the nonparametric estimator.
+As baselines, we additionally compute a nonparametric $R^{2}$ [Williamson et al., 2021] and permutation importance [Breiman, 2001].
+Figure S11 shows the Spearman correlation between the estimates from the different regression methods.
+Figure S11: Spearman correlations of variable importance measures computed on each feature using different regressions.
+On the left, these are computed on all $561$ features and on the right they are computed on the subset of features where at least one of the two methods deems the effect significant at a $5\%$ level.
+The $R^{2}$ and permutation methods do not come with significance tests and are therefore omitted from the right plot.
+While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not.
+The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator.
+For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate.
+Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science [e.g., Marcos-Zambrano et al., 2021].
+Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model.
+One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
+Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option.
+To this end we consider a standard approach for estimating effects of coordinates for unconstrained data.
+More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint.
+For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression.
+As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species.
+In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
+S5.8 Additional plots from gut microbiome analysis in Section 5.2
 In this section we provide additional figures showing our results from the analysis of the gut microbiome data in Section 5.2.